A Complete Bibliography of Publications in the Journal of Fish Biology (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 May 2025 Version 1.08

Title word cross-reference

0 [1095]. 11 [1473, 294, 293, 117]. 5 [1040]. ^{+} [606, 1613, 285]. ^{109} [857, 1693]. 14 [1415, 485]. ^{p} [1380]. _{1} [1230, 487, 828]. _{2} [1578]. _{2α} [780]. _{3} [391, 531]. α [1031, 734]. β [1578, 1644, 750, 1338]. × [477, 1557, 678, 553, 1218, 1341, 260, 363, 264, 257, 248, 1621, 562, 1024, 1096].

-adrenergic [1338]. -ATPase [606, 1613, 285]. -glycine [496]. -group [1095]. -hydroxyindoleacetic [1040]. -K [606, 285]. -ketoandrostenedione [1473]. -ketotestosterone [293, 117]. -methyl-testosterone [734]. -methyltestosterone [1031]. -oestradiol [750]. -oxotestosterone [294]. -pyrene [1270].

/K [1613].

143 [292]. 14C [496]. 1939 [596].

2-year-old [1018]. 20^{o} [488].

7^{o} [1160].

A. [106]. abdominal [417]. Aberdeenshire [565]. aberrant [1477]. aberration [386]. aberrations [726]. ability [960, 1236, 812]. abiotic [1742]. Abnormal [1192, 1696, 340, 1453]. abnormalities [650, 910, 33, 1344, 1641]. above [216]. Abramis [490, 1665, 670, 278, 925, 1540, 1188, 1096]. Abscessed [12]. absorption [1616, 1259, 1621]. absorptive [195]. abu [1335]. Abundance [483, 1681, 112, 1217, 372, 182, 722]. abyssal [1069]. Acanthocephala [261, 971, 939]. acanthocephalan [1547, 1567, 716, 588]. acanthocephalans [1585]. Acanthocephalus [716, 1585]. Acanthogobius [405]. Acanthogyrus [1434]. Acanthopagms [760]. Acanthopagrus [701, 306, 389, 439, 469, 964, 1577, 1562, 12]. Accelerated [728, 1373, 465]. accessory [36]. acclimated [1364, 1619]. Acclimation [841, 199, 647, 527, 1048, 82, 15, 1659, 213, 1145, 777]. Acclimation-induced [199]. accompanying [1551]. accumulation [1726, 892, 387, 331]. accumulations [916]. accuracy [793]. aceratus [219]. Acestrorhynchinae [860]. acetate [395, 137]. acetazolamide [876]. acetazolamide-sensitive [876]. acheilognathi [686]. achromatic [668]. acid [84, 183, 1533, 746, 463, 1640, 938, 1131, 713, 1278, 1711, 1142, 653, 118, 293, 995, 1550, 277, 1431, 1342, 1655, 1672, 279, 1040, 301, 720, 812, 1118, 1639]. acid-deficient [653]. acid-tolerant [938]. acid/aluminium [1711]. acid/aluminum [1278]. acidic [730, 1275, 1282]. acidification [1221]. acidity [1142, 1189]. acids [1415, 662, 421, 1397, 91, 1543, 360, 1623, 933]. Acipenser [1720, 802, 1409, 1097]. acoustic [1478, 1069, 422]. acquired [1716]. action [557, 61]. actions [5]. Activation [1235]. active [917, 528, 330]. activities [1074, 927, 1101, 1603, 1564, 1119, 1286]. Activity [945, 915, 1070, 394, 587, 823, 395, 1109, 670, 210, 1155, 528, 940, 414, 426, 1709, 1229, 1548, 326, 606, 876, 858, 19, 178, 1521, 435, 31, 1076, 1477, 283, 556, 719, 80, 23, 350, 1268, 1236, 1635, 329]. actomyosin [329]. aculeatm [1663, 1647]. aculeatus [366, 407, 323, 1473, 976, 1551, 1263, 1669, 1452, 1136, 1155, 1471, 1552, 675, 1143, 1103, 109, 139, 461]. acute [448, 778, 1636, 876, 305, 1150, 1166, 628]. acutipennis [564]. adapt [579]. adaptation [1295, 728, 388, 548, 376, 606, 1299, 231, 285]. adaptations [1211, 1424, 715, 1589]. adapted [1322, 1606, 947, 572]. adaptive [1497, 1771]. added [203]. Addendum [1122]. additional [1625]. Address [1772, 882]. Aden [809, 1079]. adenohypophysis [832]. adenylate [572]. adenylates [604]. adipose [1490, 1025]. adjustments [1166]. administered [1031]. Administration [514, 326, 531, 657, 1141]. adopting [1359]. adoption [1607]. Adoptive [108]. adrenaline [1659]. adrenergic [1338]. adrenocortical [80, 23]. adrenosterone [360]. adult [1274, 310, 48, 1531, 195, 920, 1146, 1288, 1019, 357, 1198, 478, 1510, 494, 1284, 218, 1428]. adults [260, 150, 284]. advanced [1018]. advances [728]. Aegean [1680]. aeglefinus [1196, 1378, 322]. aeneus [821]. Aequidens [613]. aerated [562]. Aerial [1179, 940, 188, 1358]. Aerobic [740]. Aeromonas [1245, 847, 1248, 816, 1260, 657, 1167]. aestivated [528]. Aetiology [378, 1348]. affect [1717]. affecting [85, 433, 775, 332, 346, 469, 669]. afferent [324]. affinis [502, 1696, 1406, 427]. aflatoxin [1230]. Afon [1103]. Africa [844, 727, 1036, 1745, 63, 1495, 1029, 973, 1496, 1763, 314, 538, 771, 671, 1421]. African [821, 810, 528, 1356, 1297, 1328, 902, 1573]. africana [717, 1041]. africanus [1328]. after [1720, 709, 1375, 912, 125, 326, 1339, 1654, 1165, 1372, 1160, 194, 446, 451, 733, 1258, 1338]. against [1102, 625, 100, 816, 805, 108]. Agassiz [695]. Age [366, 785, 189, 249, 1581, 1422, 980, 526, 1430, 262, 671, 1073, 783, 589, 1761, 172, 710, 1385, 1406, 45, 1225, 809, 1378, 620, 1648, 357, 793, 1294, 1084, 1771, 1267, 329]. age-0 [1084]. age-determination [793]. age-growth [620]. Ageing [1469, 1746, 935, 180]. agent [1102, 43]. agents [1256, 54]. agglutinating [1102]. Agglutination [718]. aggregations [791]. aggression [1669, 1143]. aggressiveness [1551]. Agonistic [1173]. Ahl [771]. air [183, 918, 1131, 1296, 1283, 747, 438, 1216, 1317, 1150, 1166, 1399]. air-breathing [183, 918, 1131, 1296, 438, 1216, 1317, 1150, 1166, 1399]. air-supersaturated [747]. airbreathing [1017]. akaara [750]. Alabama [148]. alalunga [35]. alanine [374]. alarm [1353, 856, 1400, 303]. Alaska [1428]. albacore [35]. Alberta [985, 212]. albidus [411]. albino [1232]. albula [1531, 1351]. albumin [1261]. Alburnus [1557, 411, 1218]. alcalicus [572]. alcohol [367]. Alella [1562]. Alepisauridae [1330]. Alepisaurus [1330]. Alestes [437]. alevins [250, 520, 361]. algae [12]. algal [1170, 1377]. Algeria [721]. Algoa [671]. alimentary [851, 254, 1474, 178]. alkaline [915, 670, 572]. alkalinity [855]. allele [840]. Alligator [1023]. Allocreadium [423]. Allocyttus [1686]. allograft [256]. allotriploidy [1370]. allow [1591]. Allozyme [206, 1755]. along [1133]. Alosa [854, 1511, 1068, 233, 1316]. alpine [974]. alpinus [1458, 506, 333, 176, 1555, 1376, 681, 141, 515, 484, 474, 543, 497, 944, 596, 1633, 1623, 930, 933, 1582, 1467, 1627, 1752, 761, 991, 616, 351, 352]. Alterations [1048, 913, 1553]. altering [1186]. alternative [968, 896]. altivelis [1568]. aluminium [1380, 1616, 1711, 1662, 995, 1342, 1169, 1403, 1639]. aluminium-induced [1403]. aluminum [1278]. Alvor [1515]. Amazon [979, 416, 1299, 239, 327, 451, 733, 998]. Amazonas [440]. Ambassidae [1496]. Ambassis [1763]. ambloplitis [1643]. America [101]. American [1010, 399, 928, 905, 558, 863, 673, 208, 941, 193, 1068, 1455, 948, 1294, 748, 677, 1306, 191]. americanus [946, 709, 1723, 928, 1250, 1677, 656, 283, 191, 639]. Amia [945]. amine [934]. Amino [277, 1415, 1397, 91, 1543, 1550, 1431, 933]. aminopeptidase [414]. aminotransferase [374, 823]. ammonia [448, 1431, 933, 1200, 1201, 222]. ammonium [910]. Ammotretis [1007]. among [1735, 261, 1705, 1598, 791, 435, 1752, 74]. Amount [1705]. amphibious [1358]. Amphipnous [531]. amtmlis [1614]. Anabantid [447]. Anabas [1017, 258]. anabolic [339]. anadromous [877, 483, 800, 143, 1163, 1622, 1584]. anaerobic [1719]. anaesthesia [288, 905, 1392, 1449, 706]. anaesthetic [157, 171, 651, 1292]. analis [1465, 772]. analysed [1142]. analyses [82, 1233, 1316]. Analysis [1700, 1537, 1312, 594, 487, 1546, 644, 553, 1370, 1446, 1376, 101, 1755, 111, 504, 135, 1414, 1722, 22, 1285, 1526, 300, 38, 932, 113, 1167, 1460, 688]. Anarhichadidae [1172]. Anarhichas [1172, 1287]. anatomical [46]. anatomy [674]. Anavilhanas [979]. anchovy [810, 1741, 372, 1680]. and/or [1654]. androgen [383, 137, 806, 1486]. androgenic [178]. androgens [1426]. androstenedione [1473]. aneuploidy [40]. Angel [1348]. angelfish [1067]. angle [1293]. angler [267, 967, 865]. angler-fish [865]. anglerfish [1309]. anglers’catches [1556]. Anglesey [874, 853]. Anglian [895]. angling [134, 439, 1756]. Anguilla [785, 365, 5, 1365, 1698, 399, 14, 595, 929, 905, 558, 863, 1532, 1581, 1028, 729, 208, 193, 1046, 1045, 1340, 1614, 659, 456, 716, 198, 1173, 338, 563, 471, 42, 1294, 526, 1513, 396, 408, 369, 382, 521, 404, 1088, 38, 1165, 1333, 677, 1306, 956, 1460, 783, 849, 1158, 1204]. anguillae [1365, 1585, 563]. anguillaris [1090]. anguillarum [225, 485, 108]. anhydrase [945]. animal [460]. Anisakis [638, 629]. annectens [528]. Announcement [1093, 1108, 1124, 1138, 1151, 1575, 1712]. Announcements [57, 252, 454, 958]. Annual [1663, 1477, 592, 1182, 109]. annularis [650]. annulatus [671]. annulus [1300, 1336]. anomalies [559]. anomalous [1307]. Antarctic [219, 1039, 1315, 1262, 706]. Antarctica [119]. antarcticum [1262]. anthelmintic [320]. anthracene [1270, 451]. antibody [1230, 1241, 994, 593, 1167, 831]. antibody-producing [1241]. anticoagulants [104]. antifoam [54]. Antigen [225, 625, 1259, 1654, 593, 1255, 1258]. antigen-antibody [593]. Antigen-binding [225]. antigens [994, 1634, 1243]. antimicrobial [1387]. antioestrogens [455]. antipredator [1595]. antisera [879, 1260]. Apatemon [1104]. Apeltes [976]. Apicomplexa [1176, 1133, 563]. Aplocheilichthys [771]. Aplocheilidae [700]. Aplocheilus [316, 106, 1042]. Apo [1374]. apparent [1442]. appetite [763, 168, 349]. applicable [756]. application [111, 902, 1484]. appraisal [1525, 1556]. approach [903]. appropriate [139]. Aqaba [820, 790]. aquaculture [1580, 1173]. aquaria [932]. aquatic [158, 658, 568, 904, 55, 1055]. arcasi [553]. arcasii [1596]. archipelago [1570, 979]. Archosargus [864]. Arctic [1458, 506, 1003, 176, 1555, 1376, 681, 141, 515, 484, 474, 543, 497, 944, 135, 1633, 1623, 930, 933, 1582, 1467, 1627, 1752, 761, 991, 616, 351, 352]. arcticus [135]. area [398, 1751, 612, 69, 1294, 1326, 1768, 424]. areas [1528, 210, 12]. arenarius [205]. argenteus [738]. Argentina [1657]. Argentine [1657]. argentinensis [1683]. argentiventris [52]. Ariidae [1429]. Aristichthys [817, 260, 363, 562]. Arius [864]. Arnoglossus [124]. aromatic [1586, 327, 628]. Arripis [552, 578]. arsenic [199]. arteriosus [908]. artificial [763, 1504]. artificially [1306]. artificially-matured [1306]. artisanal [973]. Arungen [1105]. Ascaridida [638]. Ascaridoidea [1748]. Ascension [98]. Ascorbic [301, 84, 183, 279, 720, 812, 1118]. Asiatic [1383]. Askö [398]. Askö-Landsort [398]. aspartate [374]. Aspects [1092, 1033, 1220, 1318, 866, 1447, 871, 218, 1428, 859, 701, 827, 1701, 921, 551, 717, 318, 309, 215, 179, 961, 1448]. Aspius [309]. Asra [92]. assay [374, 1693]. assemblage [1157, 1703]. assessing [940]. assessment [97, 895, 354, 1501, 935, 259, 1334, 439, 719, 104, 113, 687, 571, 888]. assessments [1429]. assignment [1321]. associated [822, 286, 388, 558, 1231, 1416, 806, 60, 1562, 1286, 510, 1189, 1027]. Association [422]. associations [1210]. assortative [1063]. Astatotilapia [781]. Asturias [1444]. Atherina [518, 751, 13, 480, 1408, 627, 184, 505]. Atherinidae [1370, 1367, 1683, 1023]. atherinids [1497]. Atlantic [819, 758, 1035, 1527, 302, 1020, 1214, 1759, 607, 1215, 877, 1172, 1102, 310, 574, 883, 1030, 1705, 1580, 508, 884, 1646, 368, 224, 955, 1444, 1538, 550, 76, 1058, 1666, 711, 787, 1146, 1461, 943, 1605, 1071, 1411, 1732, 358, 940, 1420, 1479, 1347, 1685, 296, 1755, 1483, 1359, 870, 1288, 1396, 1749, 1019, 999, 91, 919, 250, 143, 1373, 954, 996, 791, 1481, 1081, 105, 606, 1668, 868, 1553, 1303, 1644, 602, 1724, 123, 481, 457, 546, 871, 1194, 1330, 1309, 532, 1692, 137, 1163, 1490, 1622, 230, 1149, 1352, 1512, 1510, 1200, 1744]. Atlantic [1337, 415, 1284, 816, 1565, 1660, 150, 1290, 419, 704, 114, 74, 266, 1485, 765, 1313, 1750, 1388, 565, 576, 1604, 1756, 1535, 1748]. atlanticus [765]. atlas [1110]. ATPase [804, 606, 1613, 213, 329, 285]. atrazine [1238]. atrial [152]. attachment [229, 1664]. attempt [1571]. attention [421]. attractants [349]. attraction [930]. atypical [1574]. aurata [1086, 1488, 792, 1361, 349]. auratus [648, 843, 2, 1619, 1264, 665, 1719, 513, 581, 570, 213, 705, 1641]. aurea [1000]. aureus [915, 492, 601, 248, 1621]. australasicus [1140]. Australia [365, 463, 174, 798, 772, 405, 502, 262, 389, 439, 1140, 1023, 757]. Australian [674, 724, 317, 552, 578, 555]. australis [1340, 578, 306, 389, 439, 469, 760, 964, 1577, 1562, 424]. auto [779]. auto-immunologically [779]. autoantibodies [718]. Automated [113, 719, 1094]. autoradioraphic [133]. autosome [1207]. autumn [310, 596, 565]. autumn-spawning [596]. autumnalis [539, 589, 599]. availability [636, 315]. available [265]. Avalon [1670, 1308]. avirulent [816]. Avoidance [747, 1682, 1586]. Awasa [1437, 1113]. axial [468, 1475]. axis [581]. Ayres [270]. ayu [1568].

B [1230, 1270]. back [594, 70, 796]. back-calculation [594, 70]. backwaters [1729]. Bacteria [1034, 1273, 144, 843, 558]. Bacterial [1367, 1255, 16, 97, 1102, 1340, 963, 755, 1298, 87, 1243]. bactericidal [549]. bacterin [225, 1258]. bacterins [1241]. Bagrus [1280]. Baird [427, 1394, 1426, 502, 393]. balance [950]. Baldwin [275]. Balsfjorden [353, 390]. Baltic [398, 1445, 728, 630, 770, 1570, 1486, 1771]. Baltic-coast [1771]. band [1211]. band-fish [1211]. banded [493, 949]. banding [470, 756, 1073]. Bangor [1051]. Bank [1734]. barbatulus [875]. barbel [1422, 1678]. barberinus [820, 790]. barbine [1379]. Barbus [172, 821, 1422, 44, 1678]. Bardawil [480]. Barents [1742]. barramundi [798]. based [13, 923, 1294, 348]. Basin [1179, 637, 1629, 922, 1090, 1351]. basking [64, 377]. Basrah [1335]. bass [739, 1273, 379, 924, 590, 1336, 1432, 1643, 1498, 699, 1413, 1447, 840, 1523, 629, 936, 914, 1372, 1715, 496, 708, 850, 961]. Basset [921]. Basset-Smith [921]. bath [1371, 1258]. Bathophilus [1290]. Bathylychnops [1321, 826]. bathypelagic [1538]. bathyphilum [1033]. batis [1628]. batrachus [1206]. bavbatulus [373]. Bay [1014, 1419, 775, 1677, 262, 306, 389, 439, 671, 639, 1733, 517, 1049, 1456, 1005, 191, 845, 848, 1531, 511, 452, 535, 72]. Baylis [1434]. Beam [1503]. Bean [302, 819]. beanii [1324]. bearing [971]. beat [190]. beds [1598]. before [1460]. Behaviour [1051, 1069, 1510, 1060, 1061, 1054, 1066, 1676, 158, 431, 924, 1446, 1531, 1405, 1071, 1173, 135, 780, 1222, 1065, 1521, 1115, 933, 1114, 951, 1059, 27, 1586, 1449, 60, 1400, 283, 162, 932, 762, 984, 1023, 677, 826, 215, 859, 344]. Behavioural [1211, 852, 1550, 1136, 1072, 864]. behaviours [1684]. Belone [1419, 1701]. below [216]. Bennet [1638]. benthic [1495]. benthopelagic [149]. benzene [629]. benzo [1270]. benzocaine [157, 171, 651]. berglax [767]. Bering [1747]. Bermudan [1367]. betamethasone [381]. better [1714]. between [1425, 1720, 1759, 1215, 411, 5, 1698, 574, 967, 1758, 680, 863, 625, 20, 800, 681, 1737, 620, 33, 737, 1608, 1481, 731, 31, 930, 1358, 984, 1584, 1565, 1680, 74, 1568, 1313, 1750, 128, 1024, 1688, 1535]. bib [921]. bicolor [1251]. bifasciatum [1319]. Big [1714, 1670]. bigemina [1176, 1133]. bighead [828, 817, 158, 260, 363, 1024]. Bight [819]. bimaculatus [1220, 975]. bimodal [945]. bimodal-breathing [945]. bimodally [1283]. binding [1578, 225]. bini [1365]. binotatus [906]. Bioassay [528, 852]. Biochemical [633, 1416, 810, 1000, 680, 1705, 1296, 353, 1646, 800, 943, 101]. biochemical-genetic [1705]. biocidal [1545]. Bioenergetic [1323]. Bioenergetics [1427, 1464, 1536]. Biological [1542, 356, 1218, 501, 1241, 38]. Biology [938, 184, 1012, 701, 1092, 1511, 1033, 703, 931, 445, 403, 81, 702, 462, 490, 170, 623, 94, 827, 1144, 1701, 124, 282, 829, 480, 371, 1029, 1220, 1318, 89, 238, 83, 868, 254, 717, 1447, 502, 364, 154, 429, 322, 771, 1080, 140, 309, 218, 47, 424, 215, 179, 845, 790, 961, 1390]. biomass [1334, 1516]. biometric [1467]. Biometrical [553, 1218]. Biometrics [1079]. biosynthetic [812]. biotelemetry [190]. biotic [1742]. Biscay [452]. bivittatum [36]. biwae [1286]. bjoerkna [1540, 295]. black [979, 603, 1686, 304, 1643, 99, 845]. blackfish [270]. blackspotted [67]. Blackwater [635, 1175]. bladder [419]. bleeding [79, 37, 209, 1242, 977]. Bleeker [467, 780]. bleekeri [1495]. Blennidae [1207]. blenniid [1702]. Blenniidae [1133]. Blennioidei [1688]. Blennius [1207, 1133, 20, 1358, 682]. blenny [1211, 62]. Blicca [1540, 295]. Bloch [183, 436, 798, 568, 918, 1131, 1296, 1109, 551, 1017, 980, 1648, 1319, 107, 133, 347, 438, 9, 717, 1041, 552, 578, 258, 1317, 932, 400, 1625, 1590]. blockade [1338]. blocking [381]. Blood [533, 632, 715, 964, 1350, 1141, 725, 1273, 219, 1219, 680, 778, 1164, 1600, 508, 1465, 1278, 225, 630, 434, 335, 1340, 1603, 598, 947, 1013, 1487, 1417, 1302, 79, 339, 347, 995, 1021, 115, 1149, 861, 1008, 187, 803, 1284, 1621, 104, 231, 1, 222, 173, 381, 1168, 706, 1150, 221]. bloom [1609]. blue [69, 243, 424]. blue-eye [243]. bluefin [1687, 1073, 1645]. bluegill [909, 167, 395, 662, 1446, 694, 604, 620, 876, 869]. bluehead [1319]. bluntnose [375, 698, 1382, 649]. boddaerti [1564]. boddarti [535]. Bodensee [1188]. bodie [1401]. bodies [893]. Body [379, 862, 1425, 1120, 595, 1427, 1357, 646, 823, 1014, 1532, 2, 1476, 1616, 153, 474, 1633, 396, 991, 285, 139]. Bogota [1179]. bohar [961]. Boleophthalmus [1564, 535]. bone [1210]. bonito [1468]. Bonnaterre [35]. bony [12]. Book [1473, 29, 56, 71, 93, 103, 129, 142, 155, 166, 192, 228, 240, 251, 272, 273, 342, 355, 412, 425, 453, 466, 489, 499, 523, 540, 585, 609, 634, 661, 676, 712, 752, 768, 795, 825, 867, 969, 978, 1037, 1121, 1137, 1177, 1205, 1226, 1291, 1310, 1346, 1386, 1402, 1435, 1454, 1493, 1592, 1610, 1631, 1652, 1671, 1699, 294, 293, 1018, 117, 1040, 1095]. Books [204, 343, 586, 1123, 1178, 1611, 1632]. borchgrevinki [1315]. borders [1729]. born [1361]. Borneo [688]. Bostrychus [1328]. both [1415]. Bothnia [1439]. Bothnian [1531, 72]. Bothrimonus [1649]. Bothriocephalus [686]. Botswana [188]. bottlenecks [1753]. bottom [202, 255]. Boulenger [636, 1262]. bovine [263, 1261, 299]. Bowdich [394, 403]. box [173]. Boyer [480]. boyeri [13, 480, 1408]. Brachydanio [652, 857, 815, 554, 1400, 214]. brachyrhynchus [1321]. brackish [321, 1277, 783]. brackish-water [1277]. brain [360, 60, 934, 64]. brain-encysted [60]. brains [1040]. brama [490, 1665, 670, 278, 925, 1540, 1188, 1096]. bramae [338]. branch [258]. branchial [854, 1553, 285]. brasiliensis [1318]. Brazil [610, 1469, 1667, 440]. Brazilian [1318]. bream [1086, 490, 670, 278, 1190, 1361, 1540, 306, 389, 439, 469, 760, 1188, 295, 349, 1733, 1096]. breathing [183, 918, 1131, 1296, 945, 1283, 438, 1216, 1317, 1150, 1166, 1399]. Breeding [1072, 636, 1603, 1472, 1127, 538, 1289, 1607, 1038]. breeding-season [1038]. brevirostrum [1409]. Brevoort [1279]. Brevoortia [1014, 1646, 296]. Brief [1203, 1586]. Bristol [1569, 1504]. Britain [1585, 801]. British [1369, 1436, 1099, 287, 931, 939, 835, 1199, 1452, 1152, 1378, 1467, 1281, 1649, 295]. Brittany [1020, 1539, 1700, 1638]. Broads [894]. Broch [437]. broken [796]. broken-back [796]. brooded [1429]. brook [1670, 483, 910, 1308, 1003, 920, 1053, 1662, 9, 748, 1040, 276]. brooksi [942]. brown [1539, 1380, 1554, 5, 144, 561, 151, 953, 1184, 1225, 541, 1726, 1636, 206, 1704, 615, 940, 1420, 1479, 1347, 870, 1707, 17, 1229, 141, 1481, 1692, 486, 271, 520, 65, 305, 514, 647, 861, 748, 1008, 1076, 1601, 660, 806, 1015, 27, 1342, 1655, 1182, 1006, 1169, 1403, 1672, 180, 1584, 766, 1025, 643, 1313, 608, 887, 1535]. browni [99]. Bruchydanio [378]. Brunnich [1207]. Brunswick [1677]. Brycon [416]. bubalis [1649]. buccal [409]. Bucephaloides [211]. Buchanan [822, 652, 815, 554, 378, 214]. budget [1717, 110]. budgets [960]. buds [289, 1384]. Buglossidium [429]. bulbus [908]. bullatamdis [631]. bullhead [1554, 561, 304, 373, 1691, 1482, 1182]. bully [512, 308, 926]. Bunodera [182]. burbot [1384]. Burchell [664, 821]. Burn [565]. burrowing [1211, 62]. Burst [1325]. Byrne [758].

C [411, 1415, 488, 949, 485]. C-banded [949]. C-labelled [485]. C. [457]. Ca [428]. cabeda [411]. Cadmium [857, 815, 1690, 361, 913, 1416, 912, 281, 675, 301, 1304]. Cadmium-induced [815]. caeca [195, 1675]. cage [891]. caged [1089]. Calcaemic [531]. calcarifer [798]. calcification [67]. calcitonin [1141]. calcium [236, 1578, 391, 319, 1342, 1655, 1006, 1403, 1672, 869]. calculation [594, 70]. California [1561, 1463]. Callorhinus [1747]. caloric [698, 986, 567]. calva [945]. Camallanus [47]. Camargue [170]. camera [1492]. Camssius [513]. Can [579, 861, 1180]. Canada [1656, 1670, 883, 1277, 581, 1622, 95]. Canadian [972, 839]. canal [851, 511, 906, 178, 829]. canicula [383, 573, 1637, 435, 115, 962, 418, 641, 685, 1304]. Cannibalism [988, 1482, 1615]. Canyon [1538]. capabilities [254]. capability [1381]. capacity [1619, 347]. Cape [394, 403, 63]. capelin [297, 476]. capensis [810]. Capillaria [99]. captive [732, 1190, 1032, 576]. captivity [1361, 312, 1072]. captivity-born [1361]. capture [25, 679, 149, 552, 578, 720, 505, 706]. capture-stress [720]. Carangidae [445, 919, 1027]. carapo [610]. Carassim [1619]. Carassius [648, 843, 2, 1264, 82, 665, 1719, 581, 570, 213, 279, 705, 1344, 1641]. carbaryl [444]. carbohydrate [1314, 265]. Carbon [986, 1301, 1725, 239]. carbonic [945]. carboxylic [426]. Carboxymethylcellulase [19]. carcase [407]. carcass [48, 263]. Carcharhinus [160, 754, 99]. cardiac [1219]. Cardio [1209]. Cardio-respiratory [1209]. cardiovascular [1150]. Carmarthen [1005]. Carnation [1369]. carnitine [914]. Carotenoids [1627, 1030]. carotid [1216]. carp [828, 817, 1034, 1480, 256, 158, 431, 678, 560, 1238, 45, 170, 1272, 1011, 851, 136, 40, 304, 1048, 983, 488, 33, 315, 737, 1459, 1397, 82, 15, 130, 260, 363, 730, 1275, 178, 293, 360, 59, 1301, 1261, 1233, 1147, 801, 1431, 788, 187, 1248, 1634, 1653, 1654, 773, 977, 593, 279, 148, 275, 495, 562, 1268, 1236, 1612, 1009, 1156, 55, 1635, 1024, 1243, 422, 344, 1615]. Carpentaria [798]. carpio [256, 560, 1238, 45, 170, 1272, 600, 1011, 157, 171, 851, 136, 40, 304, 1048, 488, 737, 730, 1275, 178, 293, 360, 1301, 1261, 801, 245, 1431, 1248, 773, 1242, 977, 593, 1268, 1236, 1009, 1156, 1635, 1243, 422, 344, 1615]. caryophyllid [1206]. caryophyllum [654]. case [1300, 1327, 1505, 498, 435]. Caspiosoma [92]. caspium [92]. castration [1486]. catabolism [1465]. Catalan [298]. catalase [350]. catch [202, 1375, 973, 1744]. catch-per-effort [1744]. catchable [901]. catches [636, 267, 967, 333, 972, 50, 328, 38, 377]. catching [1198]. catchment [612, 1391, 1758, 953]. catecholamines [5, 397]. catfish [822, 384, 1179, 664, 1257, 823, 957, 249, 1234, 1109, 551, 100, 1026, 764, 665, 796, 482, 1297, 1280, 516, 1182, 1080, 200, 642, 864, 666, 194, 1590, 87, 307, 303, 1399]. catfishes [1429, 538]. cations [242, 146]. Catostomus [663, 869]. caudal [1662]. caught [787, 277, 1453]. causation [1061]. causative [1102, 43]. cause [1320]. caused [974, 870, 1267]. causes [650, 952, 808, 1145]. causing [292]. cautionary [1329]. cavity [289, 409, 64]. Cd [857, 1693]. Cell [418, 1170, 1466, 878, 778, 157, 209, 152, 1265, 861, 1008, 1358]. cells [219, 475, 878, 560, 131, 195, 1315, 1606, 225, 632, 1614, 947, 1013, 1487, 1241, 79, 606, 908, 1553, 655, 1261, 1659, 1021, 137, 675, 1371, 856, 1634, 964, 1106, 409, 1009, 1168, 619, 116, 303, 1258, 831]. Cellular [1459, 1247, 360, 1238]. cellulose [597]. Central [1374, 870, 1617, 979, 591, 1455]. centres [232, 915, 1264, 1239, 1249]. Centrocestus [1042]. Centropomidae [1029]. cephalm [1274]. cephalus [411, 1341, 971, 246, 234, 428, 30, 769, 291, 1348, 218, 301, 720, 812, 1118, 1344, 1674, 950, 734]. Cepola [1211]. Ceratothoa [317]. Cercariae [1254]. certain [888]. Cestoda [942, 233, 1159, 936, 1583, 926]. cestode [1630, 1206]. cestodes [1199]. Cetorhinus [64, 377]. cf [416]. Chabanaud [99]. Chaenocephalus [219]. Chalcalburnus [1341, 925]. chalcogramma [986, 1747, 839, 1428]. chalcoides [1341, 925]. chalk [1492]. Challenge [631, 963]. chamaeleonticeps [819]. chamber [917]. chambo [1472]. Change [723, 735, 753, 205, 692, 804, 1160, 1390]. Changes [1551, 388, 967, 1014, 778, 50, 993, 1394, 604, 1029, 963, 198, 326, 79, 209, 1656, 339, 992, 1431, 396, 408, 521, 1008, 806, 60, 409, 1404, 222, 381, 1104, 576, 1150, 696, 1761, 1533, 1706, 1407, 1374, 1554, 286, 561, 1295, 1429, 633, 1030, 1003, 32, 247, 48, 568, 1602, 199, 376, 1416, 957, 855, 670, 1132, 447, 1360, 1579, 294, 815, 125, 605, 292, 357, 1731, 51, 764, 11, 1159, 1613, 347, 188, 1677, 1659, 229, 1767, 581, 1115, 6, 369, 656, 660, 1477, 1595, 1562, 1403, 1372, 1358, 1343, 1306, 1583, 1590, 1388]. changes [1270, 807, 965, 285]. Channa [183, 436, 1131, 1296, 133, 347, 1150, 1166]. Channapunctatus [918, 438]. Channel [997, 384, 1257, 823, 957, 100, 764, 665, 796, 516, 200, 1729, 642, 666, 194, 87, 307, 303, 1569, 1504]. channels [1355, 1305]. Channidae [1466]. Chanos [1031, 507, 797, 1332]. char [141]. Chara [363]. characid [1293]. Characidae [860]. character [1316]. Characteristics [1184, 35, 749, 600, 1276, 143, 830, 1724, 1642, 1187, 1333, 1009, 888]. Characterization [76, 1340, 860, 375, 1566, 946, 1653]. characterized [1010, 1168]. characters [1486]. charr [1458, 506, 1670, 1308, 333, 176, 1555, 24, 1376, 920, 1053, 681, 515, 484, 474, 543, 497, 944, 596, 1633, 1623, 930, 933, 1582, 1467, 1627, 1752, 761, 991, 616, 351, 352]. Chatham [1686]. Chelon [994, 1253]. Chemical [68, 1681, 600, 102, 78, 852, 471, 113, 677]. chemicals [970]. chemiluminescence [1245]. chemiluminescent [669]. chemistry [335, 339, 1284, 706]. Chemo [930]. Chemo-attraction [930]. chemosensory [1449]. chemosynthesis [1764]. chemotactic [970]. Chesapeake [511]. Chilean [638]. chiliensis [1468]. Chilodonella [547]. China [836, 96]. Chinese [33, 315, 737, 187]. Chinook [1345, 174, 1276, 963, 1212, 1450, 1160]. Chitinase [587, 927, 1119]. Chitinolytic [755, 626]. chitobiase [1119]. chloride [1380, 918, 1606, 606, 1553, 675, 1106]. Choice [1483, 1062, 1552, 1482, 1063, 1143, 1064]. Chondrostei [1409]. Chondrostoma [553]. Chone [637]. Choricotyle [1111]. chorion [1001, 624, 656]. chorions [1004]. chow [495]. chromaffin [116]. chromatography [175, 397]. chromatophores [243]. Chromosomal [725, 1722, 1027, 726]. Chromosome [1466, 470, 1207, 860, 1444, 923, 433, 458]. chromosomes [1312, 756, 949, 1193]. chronic [778, 647]. chronology [1195, 620]. Chrysichthys [249, 1297]. Chrysoblephus [937]. chum [1737]. chuss [1149]. Cichlasoma [1062, 1455, 582, 393]. cichlid [696, 1061, 1062, 568, 1455, 1063, 547, 582, 393, 1064]. Cichlidae [862, 1164, 1689, 1593, 908, 1072, 975, 1457]. cichlids [1424, 1339, 1607, 1573]. Ciguatoxin [420]. ciliary [625]. Ciliata [1401, 1132]. ciliate [748]. Cinco [1469]. circular [495]. circulating [1245, 778, 1278, 383]. Circulation [1162]. circulatory [292]. Circuli [1300, 766]. cisco [1531]. citrinellum [1062]. Clarias [664, 1090, 830, 482, 1206]. clarki [855]. class [278, 898, 1320]. classification [476]. clavata [340, 238, 503]. clavula [716]. Clear [911]. climatic [1441]. climbing [258]. Clinocottus [1465]. Clinostomum [727]. clomiphene [455]. clonal [1233]. clones [998]. Closing [1772]. clots [1]. Clupea [1770, 398, 1533, 935, 1381, 692, 1411, 1732, 635, 791, 1576, 1570, 1587, 683, 534, 58]. clupeaformis [268, 722]. clupeid [717, 1208]. Clupeidae [1318, 1041]. clutches [941]. Clyde [364, 1770, 935]. Co [1111, 1496, 1763]. co-existing [1496, 1763]. Co-occurrence [1111]. coagulation [1350]. coal [811]. coal-derived [811]. coalfish [835]. Coast [394, 403, 430, 1700, 462, 1745, 1495, 1488, 1638, 1370, 1133, 564, 450, 124, 282, 1157, 480, 1146, 1081, 1041, 1496, 1453, 1703, 1588, 1771, 845, 717, 1078]. coastal [365, 189, 394, 403, 1195, 559, 798, 1347, 405, 1165, 1523, 218]. Cobitis [300, 423, 1286]. cod [226, 1676, 462, 587, 927, 21, 247, 353, 334, 1360, 1749, 390, 91, 1754, 1599, 1548, 500, 755, 1119, 624, 1734, 1200, 1744, 555, 1565, 410, 740, 1193, 1748, 1438]. Coelorinchus [123]. coexistence [691]. Cohen [1321, 826]. cohesiveness [1382]. Coho [1759, 1195, 388, 668, 78, 51, 207, 385, 479, 1369, 1586, 494, 1180, 1572, 465, 284]. Cohort [1392]. cohorts [1735, 1479]. cold [288, 1676, 1322, 1349, 876, 212]. cold- [1349]. Colisa [1317]. collecting [741]. Collection [666]. Collett [282, 546]. Colombia [1179]. colonial [1032]. colonization [1704]. Colossoma [440]. Colour [205, 668, 382, 1595, 1343]. colouration [1289, 1607]. colours [1721]. Columbia [931, 1369]. combination [320]. combinations [1142, 968]. combined [1478]. comments [415]. commercial [636, 385, 1507]. commercially [1091]. commersoni [663, 869]. common [1099, 1327, 14, 45, 170, 1272, 512, 851, 33, 315, 737, 1354, 1081, 6, 1395, 1500, 308, 1156, 280, 926, 422, 344]. commonly [10]. Communications [1203]. communis [1156]. Communities [997, 212, 903, 1520, 1519, 984]. community [1029, 1628, 1356, 1499, 1305]. Comparative [84, 931, 1638, 1349, 304, 1732, 1558, 437, 1078, 1526, 1577, 119, 1534, 175]. compared [878, 207]. Comparison [475, 741, 947, 847, 260, 106, 255, 10, 202, 785, 161, 1322, 210, 1136, 1697, 1375, 51, 479, 1637, 364, 22, 801, 649]. comparisons [678, 625, 831]. compartment [1617]. compensation [747]. Compensatory [714, 1623]. competition [954, 996, 1568, 998]. competitive [960, 1052]. competitors [1034]. completion [1375]. complex [408, 191, 1688, 639]. complexes [593]. component [222, 381]. components [1762, 568, 630, 220, 243]. Composition [1128, 1129, 1130, 464, 1060, 1374, 929, 967, 1357, 823, 1014, 809, 353, 1476, 1437, 1116, 1375, 102, 88, 1668, 1633, 1008, 510, 781, 991, 1141, 965, 285]. compounds [1387, 1543, 1594]. concavum [1452]. concentration [1672, 1025]. concentrations [448, 10, 1706, 1185, 1186, 236, 1446, 572, 1239, 51, 263, 460, 1550, 387, 319, 1442, 629, 117, 934, 331, 222, 1304, 869]. concept [1736, 896, 1769]. conch [919]. concurrently [238]. condition [1546, 167, 709, 637, 1357, 94, 927, 862, 692, 1404, 807, 539]. conditioned [422]. conditioning [344]. conditions [288, 1742, 463, 1322, 374, 1689, 1142, 1279, 1232, 588, 1417, 1441, 1710, 1125, 1641, 1615]. confined [849]. confinement [1107]. Connecticut [1068]. Consequences [804, 134, 960, 616, 58]. consideration [917]. considerations [923, 359]. Constance [1188]. constant [590, 695, 1678]. constrained [1658]. constraints [944, 1052]. constrictions [542]. construction [50]. Consumption [495, 366, 1458, 151, 1427, 1002, 863, 1532, 376, 821, 1476, 851, 1196, 1360, 311, 241, 332, 130, 1081, 59, 1468, 1623, 1728, 981, 1674, 330, 1024, 1055, 285, 139]. containing [79, 1371]. contaminated [444]. contamination [1112]. content [1770, 380, 475, 878, 1357, 48, 265, 353, 334, 879, 986, 604, 1152, 383, 1100, 175, 263, 1414, 214, 1006, 1438]. contents [698, 111, 293, 611, 1768, 567]. continental [336, 1686]. continuous [190]. contrast [226, 1176]. contrasting [1709, 1596, 1544]. Contribution [459, 89, 1756]. control [976, 776, 188, 275, 127, 169, 150]. controlled [1710, 1615]. controlling [543]. conventional [470]. conversion [863, 1476, 982, 484, 293, 1725]. convict [582, 393]. cool [1641]. cooling [1692, 505]. cooling-water [1692]. copepod [921, 1577, 1562]. Copepoda [483]. Copper [392, 1554, 873, 778, 265, 1446, 694, 604, 1672, 1112, 331, 864, 280]. coral [1725]. coral-reef [1725]. coralligenous [1703]. coregonid [917, 1043, 1474]. Coregoninae [1192, 1491]. coregonine [1753]. Coregonus [1192, 1491, 667, 646, 1531, 268, 1351, 443, 1125, 80, 1661, 722, 539, 589, 599]. Coris [1134]. Corneal [1343]. corpuscular [548]. Correlates [1618]. Correlation [20]. Correlations [840]. Corrib [612]. corticosteroid [194]. Cortisol [1640, 861, 288, 1185, 1186, 823, 1278, 492, 1392, 992, 570, 514, 1107, 1601, 773, 381, 1246, 221]. cortisol-fed [823]. Coryphaenoides [457, 1489, 1069]. Cost [679, 893, 1032]. costs [556, 410]. cotidianus [512, 308, 926]. Cottidae [1691]. Cottus [373, 1691, 1482, 1477]. Couch [1470]. couchi [1470]. couesi [1309]. count [861]. counter [354]. counting [1094]. coupled [1516]. course [912, 1613, 1644, 305, 1372]. Courtmacsherry [1419]. covariance [594]. cranial [1224]. crappies [650]. crassum [8]. Craterocephalus [1023]. Creek [1369, 1519, 1356, 911]. creeks [1016]. Crenilabrus [1563]. Crenimugil [110, 422, 344]. Creplin [1716, 1253]. criterion [80]. critical [1736, 1355, 505]. critical-period [1736]. critique [569]. croaker [368]. Cross [1297]. Crowding [1145, 315, 647]. crucian [82, 1233, 1147]. crude [709]. Cry [443]. cryopreservation [1083, 1484]. crypsis [668]. Cryptobia [1302, 75]. Cryptocotyle [1452, 1253]. Cryptopsaras [1309]. crysoleucas [1682]. Ctenochaetus [1368]. Ctenopharyngodon [487, 477, 828, 1480, 158, 678, 1397, 15, 130, 260, 363, 59, 148, 275, 562, 55]. CTX [420]. Cuban [413]. cuchia [1589, 531]. cues [1681, 205, 1405, 1056, 78, 852, 552, 677]. culling [1158]. cultivated [617]. cultural [143]. culture [725, 433, 458, 801, 788, 891, 465, 1568, 616]. cultured [558, 208, 1326, 1372, 642, 1158]. cultures [434]. current [229, 216, 1738]. Cutaneous [1252]. cutthroat [855]. Cuv [564, 316, 15]. Cuvier [518, 751, 654, 1571, 83, 1496, 627, 96, 440, 505]. cuvieri [12]. cyanellus [1675]. cyanide [448, 199]. cycle [728, 665, 1190, 1018, 760, 686, 1182, 179, 539, 109, 294]. cycles [1638, 1394, 1426, 732, 1091, 981, 1544, 1740]. cycling [819]. cyclonic [1495]. Cyclopteridae [452]. Cyclopterus [544, 1066, 1721, 624, 1047]. Cyclothone [302]. Cydopterus [605]. cymothoid [317]. Cynoglossus [99]. Cynopotaminae [860]. Cynoscion [654]. cyprinid [1012, 1060, 1557, 1427, 1464, 1536, 1717, 821, 925, 1709, 44, 1379, 309, 34]. Cyprinidae [553, 1341, 1398, 1709, 1354, 1400, 1344]. cyprinids [1398, 1424, 1558, 896, 1379]. Cypriniformes [553]. Cyprinm [977]. Cyprinodon [118, 744]. Cyprinodontidae [771]. cyprinodontiform [1110]. Cyprinodontiformes [700]. Cyprinus [256, 560, 1238, 45, 170, 1272, 600, 1011, 157, 171, 851, 136, 40, 304, 1048, 488, 737, 730, 1275, 178, 293, 360, 1301, 1261, 801, 245, 1431, 1248, 773, 1242, 593, 1268, 1236, 1009, 1156, 1635, 1243, 422, 344]. cyproterone [137]. Cypvinus [1615]. Cystidicoloides [697, 1162]. Cytochemical [872, 1164, 131, 1409, 133, 478]. cytochemistry [1046, 1116, 908]. cytochrome [1010]. Cytogenetic [834, 789, 1461, 1312]. Cytogenetics [923]. cytological [1149]. cytometric [1370]. cytometrical [1393]. cytometry [1168]. cytophotometry [977]. Cytotoxic [1673].

D [1082, 391, 531]. D. [1716]. dab [1327, 693, 1196, 107, 6]. dace [386, 1391, 1758, 229, 346, 1095]. Daily [1332, 1425, 1214, 853, 1195, 568, 851, 223, 1196, 241, 1081, 1572, 1450, 139]. Dam [1036, 1375, 50]. damselfish [1251]. Danger [1057]. Danio [822, 378]. danios [1400]. darters [1595]. data [356, 785, 1746, 1591, 1573]. dawn [1057]. day [1451, 1002, 12]. Daytime [1579]. Death [1602, 951]. debt [1697]. decagrammus [832]. Decatur [650]. decipiens [1748]. declining [716]. declivis [317]. deep [302, 577, 1324, 1351, 784, 123, 457, 546, 871, 1309, 415]. deep-sea [302, 1324, 784, 123, 457, 546, 871, 1309, 415]. defects [1439]. deficiencies [746]. deficiency [183, 1299]. deficient [436, 653]. definitive [1547]. deformability [1487]. deformation [379]. deformities [378]. degraded [1362]. degree [1369]. dehydrogenase [326, 606]. delaying [987]. Delta [1561]. demersal [1745, 1157, 1152, 602, 611]. demography [1749]. demonstrated [1140]. demonstration [144]. dendriticum [1716]. Densities [193, 577, 988, 460, 255, 1135]. Density [148, 1739, 971, 1392, 858, 1261, 578, 1089, 1587, 1738, 1448]. density-dependent [1739, 971, 1738]. dentatus [235, 132]. denticulatm [1287]. dependence [1219, 311]. dependent [1739, 971, 1429, 1637, 1738, 1733]. Depletion [1669, 194]. deposition [550, 1633, 1570, 1342, 1655]. depression [393]. deprived [277]. depth [667, 370, 337, 996, 1472]. Derichthyidae [871]. derived [1415, 1430, 811, 330]. dermal [1223]. Descent [729]. described [1144, 99]. Description [295, 1149, 300, 179]. desert [118]. despite [1145]. detectable [1573]. Detection [1176, 879, 91, 420]. detergent [5]. Determination [172, 1294, 785, 189, 1385, 1207, 917, 345, 793, 526, 262, 1073, 783, 589, 1756]. determinations [249]. determined [1765, 306, 556, 496]. determining [1056, 591]. detritivory [731]. devario [822]. developed [863]. developing [21, 319]. Development [244, 1007, 566, 1381, 1101, 296, 1232, 653, 160, 256, 1224, 1066, 802, 924, 1030, 1536, 247, 1002, 1398, 297, 1593, 341, 925, 193, 799, 1579, 1017, 500, 605, 1275, 1026, 234, 1474, 1553, 1065, 1447, 1724, 902, 1213, 1317, 27, 536, 1251, 1023, 279, 975, 837, 182, 1249, 1674]. Developmental [1760, 931, 260, 1641]. dextrin [1621]. diameter [1215, 1225, 80, 1475]. diaphanus [493]. diatom [916]. Dicentrarchus [1273, 379, 1498, 1447, 914, 1372]. dichromate [1239]. dichrowa [1293]. Dicing [951]. Diclidophoridae [1111]. Diel [1195, 1463, 1450, 570, 283, 162]. Diemen [798]. Diet [1458, 1419, 1181, 710, 1407, 924, 1495, 1276, 920, 1113, 290, 1424, 1552, 385, 665, 416, 1242, 934, 72, 1095, 1135]. Dietary [220, 1713, 746, 1415, 913, 7, 537, 265, 88, 1100, 339, 486, 1442, 1725, 1621, 1693, 991, 1438]. diets [652, 927, 763, 495]. Differences [748, 1601, 156, 1657, 695, 800, 33, 737, 791, 1353, 65, 1702, 1680, 1688, 1748]. Different [1140, 156, 7, 662, 1689, 693, 1606, 855, 210, 698, 551, 1142, 1359, 460, 1678, 1570, 1684, 762, 1169, 410, 534, 1635, 567, 432, 1135, 887]. Differential [584, 1432, 1662]. differentials [1766]. differentiation [595, 834, 1584, 1156]. Digenea [1452, 238, 1202, 1104, 1254, 1673]. digenean [423, 1348]. digenetic [338]. digested [879]. Digestibility [568, 220]. Digestion [545, 1070, 1120, 1418, 20, 1171]. Digestive [817, 802, 1437, 799, 1768, 1661, 1597, 1086, 14, 587, 414, 426, 1424, 19, 626, 755, 1119, 254, 665, 416]. dilute [508]. diluted [472]. dimensions [1541]. dimethylbenz [451]. dimorphism [1490, 1679]. dioxide [1301].

Diphyllobothrium [681, 1716]. Diplectrum [36]. Diplodus [1190]. Diploid [640, 477, 1480, 678, 1233, 307]. Diplostomatidae [238, 1202]. Diplostomiasis [518, 751]. diplostomules [1254]. Diplostomum [1389, 238, 805, 989, 1254, 1673]. Direct [292, 1597, 941, 485]. direction [1478]. disappearance [1339]. discharge [511, 1288, 907, 1692]. Discharges [892]. discolour [1072]. Discontinuous [1585, 749, 1261]. discriminates [1481]. discrimination [1753]. discs [26]. discussion [89]. disease [1102, 1117, 861, 782, 1367, 1348, 43, 1255]. Diseases [1297, 1326]. dismutase [1635]. dispersal [1421]. dispersion [370]. displacement [230, 1165]. dissolved [448, 1035, 1345, 1561, 91, 675, 1510]. distinct [1376, 437]. distinguishing [680]. Distribution [1391, 289, 626, 1042, 1015, 733, 261, 844, 1022, 85, 483, 727, 1365, 374, 145, 654, 1705, 1324, 800, 1609, 1175, 912, 894, 337, 996, 1731, 1221, 1521, 1522, 1472, 281, 1287, 1280, 372, 1729, 1421, 1049, 239, 362, 451, 83]. distributions [1585]. District [885]. Disused [1505, 1525]. ditremum [681, 1716]. Diurnal [1115]. divergence [1565, 74]. diversity [1753, 1628, 784, 435, 1516]. diving [398]. DNA [475, 878, 1753, 167, 1765, 1755, 1707, 1285, 1565, 444, 850]. dneri [1256]. do [446]. dock [1505]. docks [1505, 1525]. docmac [1280]. Does [1766]. dogfish [383, 573, 435, 115, 418, 641, 685, 1304]. Dogfishes [1128]. Dolichopteryx [1321]. dolomieui [699, 1715, 708]. Dolops [844]. domestic [575, 690]. dominant [345]. Dormitator [637]. Dorosoma [808]. dorsal [249, 794, 1625]. Dorset [751]. Dose [1578, 750]. Dose-related [1578]. doses [1601]. dosing [1639]. Double [732, 693, 645]. Dover [559, 1514, 1154, 68]. Downstream [1355, 310, 907, 230, 520, 1604]. Dr [622]. drainage [1379, 1305, 1282]. drifting [1204]. Drinking [1331]. Drug [764]. drum [533, 1246]. dry [1152]. Drysdale [1023]. Dual [468]. due [99]. Duero [1596]. Dukan [172]. dumerili [1027]. dummy [1694]. Dunbar [1015]. duperrey [1579]. Duration [311, 1142]. during [1537, 1224, 802, 1295, 1669, 633, 1030, 646, 247, 993, 698, 1394, 1426, 692, 290, 1101, 799, 1090, 421, 605, 198, 326, 1656, 1553, 1613, 1724, 992, 1659, 1431, 1127, 521, 1008, 1076, 1153, 1317, 1477, 18, 1182, 372, 1358, 231, 797, 194, 1460, 950, 576, 1150, 628]. dusky [196]. dwelling [1181, 643]. Dyková [563]. Dynamic [61, 557]. Dynamics [663, 697, 1162, 1539, 610, 1547, 885, 857, 716, 96, 1499, 182, 194, 1032, 299, 649].

E. [946]. ear [296]. ear-lateral [296]. Early [1462, 1549, 1230, 1312, 160, 1407, 802, 646, 247, 1732, 1101, 1730, 260, 500, 1553, 1361, 346, 627, 1023, 1752, 1436]. early-life [1732]. earthen [555]. East [1029, 314, 538, 559, 846, 1705, 791, 1496, 1763, 1503, 874, 1356, 1303]. eastern [853, 1383, 1495, 1277, 224, 829, 1732, 1747, 317, 123, 481, 457, 546, 871, 1194, 1330, 1309, 1453, 1696, 1140, 415, 1565, 1440, 1290, 765, 1313, 808]. Echinorhynchidae [261]. Echinorhynchus [261]. Ecological [929, 1555, 11, 1542, 1324, 681, 1522, 808]. ecologically [1376, 437]. Ecology [1567, 467, 862, 673, 1081, 1455, 1413, 123, 481, 546, 871, 1194, 415, 826, 535, 1016, 1005]. Ecomorphological [1299, 688]. ecomorphology [1558]. ecosystem [1742]. Ectoparasite [660, 844]. ectoparasites [1642, 1577]. ectoparasitic [65, 1348]. ectopic [327, 362]. Ecuador [637]. Editor [723, 735, 753]. Edwardsiella [822]. edwardsii [483]. eel [5, 1365, 14, 929, 1532, 208, 659, 1173, 338, 563, 471, 42, 396, 408, 369, 382, 521, 404, 1088, 38, 1333, 415, 531]. eels [785, 905, 558, 863, 1602, 1581, 729, 193, 1046, 1045, 1340, 1614, 456, 716, 871, 1294, 526, 1513, 1165, 677, 1306, 956, 1460, 783, 1204]. effcts [395]. Effect [183, 407, 1230, 1406, 537, 1476, 223, 580, 1479, 391, 1487, 515, 484, 385, 858, 1031, 1564, 578, 1125, 349, 722, 280, 139, 323, 1377, 1473, 854, 1387, 379, 16, 1185, 1186, 1753, 90, 236, 527, 1427, 1464, 376, 821, 157, 670, 1666, 290, 982, 250, 730, 1026, 1250, 178, 293, 360, 312, 1441, 1382, 137, 1545, 675, 1127, 1442, 404, 1213, 214, 1526, 248, 1672, 1624, 507, 1267, 850, 628, 221, 1338, 720]. effective [893]. Effects [232, 746, 873, 1238, 1357, 662, 1256, 112, 598, 315, 907, 383, 61, 102, 500, 1417, 1239, 263, 234, 1694, 995, 1301, 1623, 856, 1678, 1449, 803, 629, 1242, 1621, 897, 704, 1292, 1246, 582, 138, 734, 1158, 327, 34, 1604, 147, 1245, 1380, 242, 1542, 1517, 367, 1681, 1120, 146, 1322, 890, 560, 1751, 1745, 1495, 7, 927, 1726, 1240, 1578, 575, 690, 1336, 1109, 1278, 1360, 892, 1432, 1662, 488, 1483, 1393, 88, 122, 153, 325, 572, 954, 996, 15, 1319, 1392, 707, 11, 360, 672, 655, 1644, 665, 1361, 1633, 1443, 514, 1594, 1586, 1655]. effects [1486, 1006, 1169, 1683, 914, 200, 779, 279, 1715, 350, 934, 683, 331, 705, 1690, 1304, 1635, 1390, 1438, 737, 1018, 1118]. efficiencies [991]. efficiency [270, 1464, 1476, 1437, 982, 1471, 968, 484, 293, 591, 1725, 1395]. efficient [987]. effluent [472, 575, 690]. effort [498, 1395, 1744]. Egg [435, 1215, 1429, 952, 500, 1217, 234, 428, 107, 624, 1482, 281, 627, 616, 1448]. Egg-case [435]. eggs [236, 633, 1030, 1007, 21, 247, 1004, 986, 692, 1152, 1378, 250, 605, 601, 617, 229, 482, 656, 1213, 1570, 99, 1674, 1094, 361]. eicosanoid [1244]. Eigenmann [1469, 118]. Eigenmannia [1010]. eight [356, 1686, 1655]. elasmobranch [1128, 1129, 1130, 1116, 962]. elasmobranchiomorph [165]. electric [1010, 1080]. electrofishing [591]. electrolyte [331, 1141]. electrolytes [59]. electromyograms [556, 330]. Electron [449, 1258, 621, 219, 648, 131, 120, 195, 1044, 1593, 1046, 659, 1001, 1599, 1154, 28, 115, 1545, 478, 1443, 656, 686, 418, 949, 666, 1168]. electron-microscope [1044]. electrophoresis [13]. Electrophoretic [487, 953, 1769, 506, 644, 1316, 1467]. Electrophoretically [1573]. Electrophysiological [910]. eleotrid [154, 1328]. Eleotridae [833]. elevated [806]. elicit [1254]. elimination [912]. ell [399]. Ellerkeldia [89]. Elopidae [765]. Elopiformes [765]. elvers [1028, 42, 1513, 849, 1158]. embedded [1560]. embryo [1429, 690]. embryonic [1593, 925, 1017]. embryos [1043, 550, 912, 250, 1724, 1193]. emerged [1492]. Emergence [1445]. emergent [1276]. emerging [1759]. emigrations [565]. Emission [933]. emphasis [1477]. employed [1722]. Emptying [693, 786, 763, 168, 241, 1100, 1637, 248, 1438]. encrasicolus [1741, 1680]. encysted [60]. encystment [399]. endings [406]. Endocrine [1639]. endosulfan [188]. Endothelial [77, 908]. endotoxin [1240, 1340, 655]. endrin [1185, 1186]. Endurance [1411, 1103]. Energetic [410]. Energetics [109, 1197, 1011, 193, 743, 1679, 1428]. Energy [1770, 909, 597, 167, 1663, 1669, 1357, 1717, 48, 353, 694, 110, 1152, 102, 88, 1100, 960, 1633, 1500, 1438]. engineering [814]. England [967, 697, 1162, 380, 710, 559, 1514, 1391, 1758, 1488, 278, 1755, 522, 1519, 86, 1556, 1503, 1512, 1189, 1320]. English [374, 1187, 845]. Engraulis [566, 810, 1741, 372, 1680]. enhanced [915, 841]. enhancement [1580, 886]. enriched [3]. enrichment [1336]. Entosphenus [1487, 460]. entrainment [1506]. environment [1195, 498, 667, 977, 1440, 1206]. Environmental [976, 1289, 87, 1425, 694, 1689, 904, 1504, 840, 891, 720, 812, 1118, 565]. Environmentally [898]. Environmentally-induced [898]. environments [1676, 270, 695, 715]. Enzymatic [13, 1074]. Enzyme [1603, 817, 872, 1700, 1046, 1116, 1709, 1285, 1182, 1286]. enzymes [1101, 1564, 626, 755, 416, 1661]. eosinophilic [131, 1614]. eosinophils [983]. eperlanus [522, 1084]. epidemiology [1658]. Epidermal [709, 559, 1366, 660, 1333]. Epidermis [1486, 1723, 1018, 1371, 856, 409, 1401, 303]. Epieimeria [563]. Epinephelus [1092, 750]. epiphytic [12]. episodes [1730]. Episodic [1495, 1616]. epithelial [653, 1634]. epithelium [1295, 274, 1154, 1306, 1399]. Epizootic [1187]. equation [70]. Eremophilus [1179]. Erie [401, 51, 479]. ERM [1256]. Errata [959]. Erratum [253]. Erythrocyte [477, 955, 307, 16, 1668, 977, 1286]. erythrocytes [475, 1176, 357, 1301, 1431, 1149, 119]. erythrophthalmus [1665, 793, 1055]. erythropoiesis [133]. Esla [1629]. esmarkii [462]. Esocidae [406]. Esox [1070, 134, 81, 658, 498, 985, 1629, 988, 679, 101, 46, 946, 745, 1252, 1141, 95, 1135, 1448]. Essex [1175]. establishment [971, 1502, 227]. esterase [426, 876]. Esterases [1075, 930]. estimate [1571, 851]. estimated [1195, 1217]. Estimates [853, 973, 1020, 846, 1741]. Estimating [1314, 577]. Estimation [739, 902, 981, 237, 1374, 1174, 241, 1468, 248, 1743]. Estuaries [1529, 445, 1528, 467, 623, 1495, 1498, 1519, 1496, 1763, 1503]. estuarine [1497, 85, 702, 1505, 1501, 1520, 1525, 1297, 1500, 1523, 1509, 1507, 1516]. Estuary [1514, 1502, 1581, 1175, 1098, 1396, 1731, 1521, 1522, 154, 1297, 469, 1512, 1510, 1526, 1524, 1035, 813, 1511, 635, 1504]. Etheostoma [695]. Ethiop [1126]. Ethiopia [1437]. Ethiopian [1113]. Ethmalosa [394, 403]. ethylestrenol [339]. Etmopterus [26]. Etroplus [568]. Eubothrium [8, 233]. Euclinostomum [727]. Eucoccidia [563]. euphausiids [552]. Euphrasen [738]. Europe [877, 1383, 101, 801]. European [379, 1223, 5, 1365, 929, 1532, 729, 46, 33, 315, 737, 456, 1558, 1173, 1026, 338, 563, 1354, 1065, 471, 42, 1513, 396, 408, 369, 382, 521, 404, 856, 1333, 897]. europium [1028]. euryhaline [1393]. eurytherm [213]. Eustronglylides [511]. Eustrongylides [399]. eutrophic [987, 81, 278, 308, 1320].

Eutrophication [1544, 901, 286, 1221]. evacuation [1458, 603, 1195, 1120, 1418, 537, 851, 88, 122, 220, 1100, 754, 1125, 1572, 704]. Evaluation [838, 1480, 420, 900, 171, 889, 1693, 1185]. events [1495, 792]. Evidence [1321, 1739, 971, 668, 402, 206, 1711, 941, 966, 1160, 1252, 1750]. evokes [750]. evolution [1497, 1466, 1039]. Evolutionary [1285]. examination [267, 953, 1154, 66, 1258]. excess [391]. excretion [1427, 1726, 153, 1200, 1201]. Exe [490]. Exercise [774, 1364, 90, 446, 221, 1338]. exercise-acclimated [1364]. exerted [1571]. exilis [826]. existing [1496, 1763]. exogenous [1604]. experience [924]. Experimental [928, 1272, 745, 1624, 966, 1160, 1460, 276, 1139, 50, 1302, 22, 1057, 1125, 1135]. experimentally [1236, 1249]. experiments [988, 1155, 311, 614, 173, 485, 1484]. explanation [952, 557]. Exploitation [348, 1745]. exploited [1091]. exposed [726, 1439, 533, 1446, 604, 319, 245, 1342, 279, 811, 222, 301]. exposure [1230, 1554, 1185, 1186, 709, 778, 694, 1278, 1616, 1362, 1443, 1586, 1169, 1358, 381, 1118, 1690, 1604, 1150, 1166, 628]. expression [506]. extensions [70]. external [1480, 707]. extracellular [1248]. extraction [58]. extreme [1441, 1320]. exudate [1009]. exudates [1614]. Eye [1224, 1300, 1389, 45, 243, 805, 1673]. eyefluke [238, 1202]. eyes [238].

F [622, 487, 828]. F. [392]. Fabricus [317]. factor [121, 692, 770]. Factorial [22]. Factors [85, 433, 884, 775, 346, 669, 485, 1282, 1425, 929, 332, 543, 944, 469, 214, 1350, 1189, 887]. faecal [1427]. fails [1717]. falcatus [919]. Falkland [424]. fallax [854, 1511, 854, 1511, 233, 1316]. families [1101]. Family [701, 1466, 715, 1194, 1577, 1562, 1363]. fario [1722, 887]. farm [1444]. farmed [842, 1030, 1146, 1284]. farms [1015]. fasciatus [1317]. fast [164]. fasted [385]. fasting [1074, 290, 1564, 237]. fat [2, 447]. fathead [777, 116]. fatigue [90]. fatty [746, 713, 421]. fauna [286, 929, 1495, 345, 106]. faunas [1538]. Fawley [518, 184]. Feasibility [805]. features [1218, 1552]. fecundity [413, 1667, 1429, 7, 1378, 839, 538, 1448, 215]. fed [1074, 823, 763, 545, 363, 37, 385, 1112, 495]. feeders [817, 987]. Feeding [394, 158, 431, 1277, 1157, 1359, 1105, 1081, 298, 154, 771, 1440, 1377, 874, 1195, 597, 1066, 1676, 167, 1686, 924, 246, 1030, 865, 1197, 467, 667, 1323, 1011, 1405, 223, 20, 772, 1087, 1029, 1483, 373, 1090, 316, 61, 181, 203, 153, 325, 515, 1329, 1222, 178, 293, 360, 573, 68, 471, 1521, 1041, 1455, 1633, 1623, 1280, 1695, 1586, 1200, 162, 300, 1450, 981, 1372, 984, 1181, 1624, 237, 1612, 410, 781, 507, 419, 519, 1412, 820, 39, 1024, 422, 344, 1005]. feeds [838]. felis [864]. female [407, 158, 678, 48, 1426, 673, 1685, 857, 260, 363, 264, 750, 1622, 1395, 1182, 126, 1607, 911, 1143, 998, 109]. females [1429, 1232]. feral [561, 513, 581]. Ferguson [1404]. Fermentative [1171]. ferox [1330]. Fertility [1096]. fertilization [500, 656]. fertilized [1459, 482]. FFA [263]. fibre [321, 82, 1621, 461]. fibres [1657, 175, 26, 128, 649, 1475]. Field [1741, 1214, 785, 311, 473, 1552, 1710, 756, 981]. fifteen [1573]. figures [434]. filaments [583, 1562, 1106]. filler [597]. filter [817, 987]. filter-feeders [987]. filters [472, 575, 690]. filtration [1487]. fimbriata [394, 403]. fin [340, 1384, 1490, 180, 1625, 1025, 1344]. fin-ray [180]. final [633, 259, 950]. findings [1361]. Fine [908, 274, 243, 258, 1702]. fine-structural [1702]. fingerling [288, 1274, 197]. fingerlings [507]. Finland [1570]. Finnmark [25]. fins [52, 1192]. fire [833]. fire-tailed [833]. First [1629, 1434, 758, 149, 1509, 926, 145, 1422, 1090, 766, 1649, 1363]. first-order [1422]. Firth [364, 429, 503, 1768]. Fish [979, 1022, 846, 1727, 990, 54, 721, 1305, 1507, 687, 1529, 232, 621, 183, 436, 336, 10, 1478, 1211, 844, 302, 380, 1061, 1312, 1466, 379, 594, 703, 1010, 1374, 923, 85, 1314, 189, 161, 470, 449, 872, 749, 901, 1546, 1385, 286, 727, 463, 310, 835, 1597, 427, 472, 1686, 1266, 1560, 895, 1505, 267, 1745, 652, 865, 1323, 1508, 512, 938, 918, 1131, 1296, 70, 354, 1630, 821, 809, 1176, 1501, 1465, 171, 651, 1534, 509, 1324, 564, 1113, 1157, 41, 1155, 320, 800, 371, 892, 789, 580, 945, 903, 1520, 1525, 893, 1087]. fish [743, 1101, 1152, 598, 974, 545, 473, 1029, 1375, 4, 316, 968, 345, 122, 241, 442, 569, 1100, 847, 211, 572, 894, 1764, 857, 904, 143, 1519, 997, 1585, 1628, 420, 614, 1241, 1283, 1354, 1731, 1221, 83, 19, 626, 868, 1356, 1747, 1159, 133, 347, 438, 1650, 591, 758, 188, 602, 611, 784, 715, 123, 481, 457, 546, 1335, 405, 1366, 1767, 1085, 348, 378, 1430, 1545, 1326, 1072, 1728, 1149, 1145, 1107, 1499, 1725, 1015, 1078, 416, 756, 1530, 1684, 556, 1500, 190, 555, 1244, 981, 1299, 1738, 80, 1729, 1181, 400, 1649]. fish [104, 231, 1350, 113, 409, 897, 1509, 1703, 669, 169, 1651, 519, 1484, 1740, 1421, 1189, 1282, 688, 571, 970, 39, 706, 77, 1524, 1267, 1237, 1094, 446, 239, 1016, 1456, 1390, 1658, 34, 757]. fished [972]. Fisheries [1436, 890, 809, 885, 1746, 1747, 889, 891, 1526, 1744, 1756]. fishery [63, 32, 973, 11, 899, 902, 439]. Fishes [1183, 84, 813, 1667, 1223, 741, 939, 321, 1751, 1110, 1199, 612, 680, 1039, 1370, 1538, 1598, 925, 1730, 1210, 181, 203, 1001, 852, 1228, 645, 212, 254, 98, 159, 1063, 1259, 201, 31, 1414, 1216, 154, 437, 468, 1091, 547, 464, 782, 1171, 984, 1702, 1343, 1, 1769, 811, 165, 1412, 824, 119, 444, 529, 362, 1516, 402, 1051]. fishing [22, 900, 1766, 1507]. fitness [584, 1762]. Fitzinger [1617]. five [726, 10, 1566, 1682, 804, 1222, 18, 1201, 362]. fixation [692]. fixed [1560]. fixing [1478]. fjord [1084]. fjords [1754]. flagellate [1015]. flatfish [1224, 559, 1222, 794, 217]. flathead [205, 196]. flavescens [775, 1574]. flaviijosephi [781]. flavimanus [405]. Fleet [751, 1282]. flesh [892]. flesus [583, 710, 376, 588, 264, 257, 1710, 592, 331]. Floodplain [50, 846, 922, 1090, 1375]. floodplains [1299]. Florida [246, 275]. flounder [583, 710, 709, 1139, 1723, 928, 1250, 264, 257, 1677, 235, 1710, 1443, 656, 592, 1526, 191, 639]. flounders [1007, 376, 588]. flow [1219, 527, 1370, 328, 1510, 1168, 888, 1604]. flow-cytometric [1370]. flowing [185, 527]. fluctuating [590, 376, 695, 870]. fluctuations [898, 1114, 812]. fluid [388]. fluke [60, 805, 1673]. fluorescence [1765]. fluorescent [1733]. flushing [1495]. fluviatilis [1626, 286, 145, 333, 498, 341, 1142, 414, 426, 441, 238, 1217, 478, 47, 182, 215, 179, 1620, 1055]. fluxes [1380]. fly [188]. following [1554, 778, 79, 876, 188, 1337, 773, 977, 631, 593, 381, 657, 1141, 163, 706]. fontinalis [1670, 483, 910, 1308, 1003, 920, 1053, 1662, 748, 276, 1040]. Food [174, 767, 1598, 1087, 316, 390, 130, 363, 573, 1041, 1084, 1728, 132, 781, 366, 407, 1458, 1070, 853, 1034, 394, 1120, 1686, 636, 69, 865, 1427, 467, 798, 863, 693, 1476, 851, 772, 1196, 1360, 787, 311, 315, 922, 88, 220, 241, 325, 515, 543, 1100, 1081, 1576, 293, 460, 1468, 960, 387, 1623, 277, 532, 1339, 1280, 18, 162, 981, 1181, 562, 761, 349, 722, 582, 1568, 820, 599, 139, 1438]. food-deprived [277]. foods [603]. Foraging [1471, 1054, 1155, 1757, 875]. fork [1429]. fork-tailed [1429]. formalin [367, 1560]. formalin-fixed [1560]. Formation [1, 1300, 386, 877, 1336]. forming [831]. formosa [239, 327, 451, 733, 998]. formosanus [1042]. forms [1376]. Forsskål [1092, 1563, 829, 772, 1031, 1280, 507, 797, 1332, 961]. fossilis [1109, 551, 1648, 932, 1625, 1590]. Fouquet [100]. four [976, 380, 13, 1277, 1355, 922, 130, 495, 761]. fourhorn [1439]. Fourleague [1014]. fractions [1170]. fragile [233]. fragilities [157]. fragility [16, 947, 1301]. France [1539, 1700, 170, 1521, 1522]. Fraser [70, 1656]. free [545, 719, 190]. free-living [545]. free-ranging [719]. free-swimming [190]. French [1368, 1081]. frequencies [840]. Frequency [726, 1377, 1602, 20, 325, 515, 1746, 930, 1200, 592]. fresh [854, 530, 321, 533, 508, 133, 1294, 897]. freshly [277]. freshwater [183, 1478, 844, 380, 475, 878, 710, 1706, 923, 727, 939, 1199, 472, 612, 633, 512, 918, 1131, 1296, 171, 651, 1534, 1606, 551, 800, 804, 903, 1045, 1101, 598, 947, 968, 1001, 997, 1354, 1413, 715, 1335, 437, 555, 1201, 984, 1584, 104, 231, 1350, 1713, 531, 1740]. freshwater-adapted [1606, 947]. Fries [364]. friesii [364]. fry [236, 151, 667, 537, 1444, 988, 1355, 315, 907, 730, 1724, 1550, 1492, 520, 1342, 1655, 914, 1672, 507, 1135]. fulvescens [802]. function [1061, 296, 492, 207, 626, 1708, 1358, 1270, 1150]. functional [913]. functions [679, 1690]. Fundamental [883]. Fundulus [392, 67, 1394, 1426, 632, 511, 312, 313, 1594, 493]. fungus [1324]. Furtado [1206]. Further [1593]. Furunculosis [208, 514, 816]. fuscus [196]. fusion [1207, 955, 1027]. Future [57, 252, 454, 891].

G [470]. G-banding [470]. gadoid [1210, 211, 18]. gadoids [1739, 162]. Gadopsis [1285]. Gadus [226, 1676, 587, 927, 21, 247, 353, 334, 1360, 1749, 390, 91, 1754, 1599, 1548, 500, 755, 1119, 624, 1200, 1744, 1565, 410, 740, 1193, 1748, 1438]. Gaimard [1170, 1377, 262]. gair [1256]. gairdneri [147, 726, 1537, 1245, 1230, 1387, 16, 97, 1364, 1219, 288, 1185, 1186, 759, 1075, 1389, 530, 841, 971, 1102, 244, 197, 548, 1415, 1184, 778, 199, 265, 714, 508, 131, 120, 274, 195, 289, 1578, 1117, 575, 690, 1476, 417, 1697, 1278, 1616, 434, 335, 391, 290, 501, 112, 982, 947, 421, 1543, 1232, 653, 815, 1608, 1417, 292, 1392, 79, 37, 209, 357, 263, 858, 707, 689, 339, 640, 684, 672, 75, 1694, 747, 1613, 73, 31, 1021, 318, 1018, 319, 66, 216, 22, 1371, 213, 1695, 1601, 1015, 803, 556, 719, 126, 718, 779, 1716, 510]. gairdneri [837, 934, 805, 1693, 989, 222, 173, 381, 359, 485, 657, 1167, 1168, 1690, 1104, 108, 128, 163, 299, 330, 567, 432, 649, 926, 1254, 1673, 397, 619, 774, 361, 628, 1338, 831]. gairdnerii [242, 597, 814, 117, 237, 221]. gaivdnevi [842]. Galaxias [504]. Galaxiidae [504, 1203]. Galeus [26]. galii [833]. gall [419]. Gambia [973]. gambusia [911, 1406, 427, 502, 1696, 911, 1148]. gamma [1167]. garfish [1419, 1701]. Garra [1545]. gas [530, 304, 995]. gasterostei [238]. Gasterosteidae [1663, 691, 1038]. Gasterostem [1669]. Gasterosteus [366, 407, 323, 1473, 976, 1551, 1263, 1663, 1452, 1136, 1155, 1647, 1471, 1552, 675, 1143, 1103, 109, 139, 461]. Gastric [786, 603, 1418, 763, 168, 88, 122, 1637, 754, 1572, 1438, 1458, 1195, 1120, 537, 274, 220, 241, 1100, 248]. gastrointestinal [1387, 1154, 665]. gayi [638]. GC [1244]. GC-mass [1244]. gene [930]. genera [811]. generalizations [1728]. generalized [1142]. generally [104]. generate [1766]. generations [1720]. genesis [621]. Genetic [1518, 1762, 33, 737, 1754, 814, 1584, 1680, 1750, 1756, 506, 1626, 1705, 1580, 810, 800, 101, 1707, 1565, 1573, 608]. Genetics [782, 789, 943]. genome [1466]. genus [302, 1321, 1223, 1110, 1567, 224, 941, 123, 481, 457, 546, 1194, 1489, 1530, 1299, 415, 1203]. Geographic [948, 1324]. geomagnetic [966]. Georges [1734]. Gerreidae [467]. Gerres [623]. Gervais [94, 437]. Ghana [394, 403]. giant [1645]. gibbosus [1363]. gigas [919]. Gilbert [415]. Gilchrist [937, 810]. Gill [270, 554, 583, 746, 333, 1534, 855, 1381, 1616, 815, 606, 1613, 1562, 1106, 811, 1258, 1309]. Gillbert [868]. Gillnet [968, 50, 1744]. gillnets [972]. gills [5, 1365, 1398, 651, 1315, 44, 653, 292, 1042, 1617, 1298, 30, 66, 1431, 1545, 675, 1443, 1594, 1317, 1577, 1201, 409]. gilthead [1488, 1361, 349]. ginbuna [1147]. giorna [123]. Girard [603, 374, 427, 1394, 1426, 502, 393, 239, 451, 733]. Girella [1170, 1377, 262]. Girnock [565]. gizzard [702]. glacial [1704]. gland [51, 858, 435, 31, 769, 1688]. glandicephalus [942]. glands [1702]. glanis [1026]. glass [956, 1460, 1204]. glass-eels [956, 1460]. glioma [835]. globulin [1167]. glomerular [1039]. Glossogobius [450]. glucocorticoid [1247]. gluconeogenic [1564]. glucose [905, 713, 807]. glucosidase [927]. Glugea [928, 1250, 191, 639]. glutamate [326]. glycine [496]. glycogen [290, 1564]. glycol [1560]. glycolytic [1101, 1564]. Gmelin [283]. goatfish [820, 790]. gobiid [564, 1702]. Gobiidae [1099, 450, 1714, 1618, 1470]. Gobio [1354, 86, 1583, 373, 1691, 1482]. Gobiogobio [1542]. Gobioidei [1223, 1395]. Gobiomorphus [512, 308, 926]. gobionid [1383]. Gobius [1470, 845]. goby [1099, 450, 282, 210, 1470, 364, 1395, 1500, 845, 405]. golden [1044, 1682]. goldfish [648, 2, 1619, 665, 1719, 513, 581, 1233, 570, 705, 1641]. Gonad [595, 186, 353, 334, 1026, 581, 1306, 33]. Gonadal [1182, 1362, 792, 1190, 1361, 797]. gonadosomatic [1018]. gonadotrophic [137]. gonadotrophin [383, 150]. gonads [1563, 637, 1132, 127]. gonionotus [780]. Gonochorism [1563]. gonopodia [1110]. gonopodium [1110]. Gonostoma [1033]. Gonostomatidae [302, 1033]. Goode [302, 1014, 819, 1367]. gorbucha [992]. gorbuscha [931, 25, 993, 223, 1656, 1153]. Gorge [50]. Gorgorhynchinae [1567]. governing [884]. gracilescens [211]. gracilis [827, 1104]. gradient [872, 749, 1437, 337, 996]. gradients [509, 983, 1261]. grahami [572]. grandis [1394, 1426]. granule [131, 1614, 619]. granules [269]. granulocyte [1045, 1410]. granulocytes [1128, 1129, 1130, 1116, 1409]. granulomas [779]. grass [828, 1480, 158, 431, 678, 946, 1397, 15, 130, 260, 363, 59, 788, 148, 275, 495, 55, 1024]. gravel [58]. gravels [138]. gravity [357]. Gray [828, 478]. grayling [674, 1603, 135]. grazing [916]. great [1720, 1184, 207]. Greece [1341, 1557, 1274, 1218, 1665]. Greeley [673]. green [1010, 1465, 112, 1675]. greenback [49]. Greenland [761]. greenling [832]. grenadier [767, 1069]. grey [287, 49, 234, 428, 994, 769, 1281, 536, 218, 950, 734, 1253]. Griffith [1463]. gross [141, 258]. ground [1515]. grounds [398, 943]. Group [875, 90, 156, 1406, 387, 18, 162, 1095]. grouper [1092]. Grouping [1717]. groups [74]. Growing [121, 567]. Growth [1364, 909, 652, 823, 809, 724, 509, 442, 836, 1208, 503, 1342, 562, 496, 114, 649, 849, 848, 739, 1761, 172, 1735, 1537, 1425, 366, 407, 323, 1214, 594, 874, 710, 746, 597, 1739, 1676, 174, 637, 472, 636, 873, 595, 1391, 324, 1427, 1464, 1536, 1717, 1323, 863, 590, 714, 53, 249, 1581, 476, 575, 24, 566, 1336, 1476, 935, 698, 278, 1666, 1360, 679, 635, 775, 1422, 1279, 620, 315, 737, 294, 1090, 259, 390, 325, 515, 484, 474, 569, 944, 954, 130, 363, 980, 1173, 1608, 1026, 263, 1746, 1694, 460, 86, 1522, 960]. growth [1623, 526, 1089, 1692, 1084, 216, 1430, 262, 389, 1725, 1317, 1596, 1679, 214, 1088, 671, 1562, 762, 1006, 1169, 1403, 914, 148, 1621, 495, 237, 761, 1660, 308, 991, 1156, 349, 722, 1073, 74, 1332, 1388, 783, 790, 351, 163, 299, 432, 1095, 1475, 1158, 444, 589, 850, 1135, 1390, 887]. Growth-rate [114, 278]. grunt [1111]. grypus [172]. Guadalquivir [1581]. guatemalensis [1307]. gudgeon [1542, 833, 1354, 86, 1583]. Guichenot [638]. Guinea [846, 1429, 827, 1144, 961]. Gulf [1657, 255, 348, 1339, 314, 1404, 1014, 1426, 1439, 1751, 952, 798, 809, 1079, 1394, 1029, 1463, 201, 127, 169, 1516]. Gunner [64]. Gunnerus [377, 457]. Günther [674, 1307, 24, 1324, 1209, 1405, 1422, 1455, 808, 502, 482, 1072, 1265, 306, 389, 439, 469, 760, 217, 964, 1664]. guppies [631]. guppy [618, 7, 280]. gut [854, 1231, 131, 120, 1405, 1437, 311, 1558, 1634, 1653, 1654, 141]. gut-associated [1231]. Gwyddon [615]. Gymnoscopelus [868]. Gymnotus [610]. gynandromorphic [1559]. gynogenesis [40, 1232, 640]. gynogenetic [1232, 1233]. Gyrodactylus [358, 999, 631].

H [1380]. habit [1317]. Habitat [889, 584, 658, 1432, 1525, 316, 1472, 1516]. habitats [741, 156, 246, 1505, 1277, 904, 517]. habits [394, 772, 373, 922, 1090, 316, 573, 1041, 298, 486, 1280, 132, 562, 781, 1440, 820, 1457]. Habituation [1647]. haddock [1196, 1378, 1734, 322]. haem [1465]. Haematocrit [513]. Haematological [436, 35, 918, 1131, 1296, 957, 171, 830, 438, 188, 200, 1286, 104, 97, 384, 963, 73, 1304, 1270, 628]. haematology [607, 1078, 706]. haematopoiesis [347]. haematopoietic [436, 918, 1131, 1296, 438]. haemoglobin [1749, 37, 209]. haemoglobins [119]. Haemogregarina [1176, 1133]. Haemolytic [1229]. haemopoietic [621]. haemosiderin [147, 621]. Hagenmuller [191, 639]. hagfish [224]. hake [1657, 638, 1149, 424]. halibut [943, 1605, 1331]. Ham [44, 1545, 1589]. Hamilton [822, 652, 815, 554, 378, 214]. Hamilton-Buchanan [822, 815, 554, 378, 214]. hamiltonii [1518]. hammerhead [1463]. Hampala [1012]. handling [957, 305, 1107, 140, 705]. hangiongensis [1477]. haplochromine [1339]. Haplosporean [43]. Haplosporean-like [43]. hard [995, 86]. hard-water [86]. hardening [624]. hardness [1726, 1636, 1715]. hardyhead [1023]. harengus [398, 692, 1570, 1770, 1533, 935, 1381, 1411, 1732, 635, 791, 1576, 1587, 683, 534, 58]. harringtonensis [1367]. Hartbeespoort [1036]. Hasselt [1012]. hatch [1215, 982]. hatchery [379, 174, 185, 527, 1705, 1762, 711, 1396, 1568, 608, 284, 1604]. hatchery-produced [174]. hatchery-reared [185, 527, 1396, 1568, 1604]. hatching [236, 690, 234, 346, 1213]. Hawaiian [1579]. head [1466]. healing [936]. health [513, 113, 687, 571]. Heard [529]. Heart [1070, 269, 1104]. hearts [910]. heat [828, 1349, 955, 1693]. heat-shock [1349]. heat-shocked [955]. heat-stable [1693]. heavy [380, 1439, 245, 350]. Heckel [1335, 502, 23, 309]. Hector [89]. held [313]. helleri [502, 23]. Helminth [1628, 1172, 1159]. Hemanthias [186]. Hemichromis [1220, 1434, 975]. Hemitripterus [283]. Henderson [900]. Henle [99]. henna [1367]. Hensel [613]. henshawi [855]. hepatic [823, 1643, 479, 858, 750, 1187, 1249]. hepatocytes [507]. hepatus [459]. herbicides [55]. herbivore [1170, 1377]. herbivorous [1113, 1397, 254, 1725, 1171]. hermaphrodite [1491, 1629, 186, 493, 36]. Hermaphroditic [459, 1134, 522]. hermaphroditism [937, 89, 1696, 402, 757]. herring [1770, 398, 1765, 935, 1381, 692, 1411, 1732, 635, 1175, 791, 1576, 1570, 1684, 1587, 534, 58]. heterochir [911, 1148]. heteroclitus [392, 632, 511, 312, 313, 1594]. heterogeneity [1516]. Heterophyidae [1042]. Heteropneustes [1109, 551, 1648, 932, 1625, 1590]. heterostomum [727]. Hexagrammidae [832]. Hexagrammm [832]. Hexamita [1015]. hexasticha [547]. hierarchy [23]. high [1288, 175, 387, 230, 397]. high-performance [175]. highly [1765]. hill [532, 1545, 1639]. Hippoglossoides [107]. Hippoglossus [943, 1605, 1331]. Histochemical [321, 375, 1566, 946, 1447, 27, 1702, 116, 461]. histochemistry [441, 1182, 619]. Histocompatibility [1233]. histogenesis [256]. Histological [156, 855, 322, 1702, 1306, 1675, 116, 297, 1447, 797, 179]. Histology [832, 769, 642, 1086, 14, 368, 67, 1132, 441, 525, 1399]. Histomorphological [1554]. Histomorphology [1590]. Histopathologic [169]. Histopathological [127, 559, 292]. Histopathology [835, 942, 796, 824]. histophysiological [1156]. historical [1221]. histories [498, 1549]. history [1483, 1359, 227, 1441, 1687, 1752, 1436, 1158]. Hoffman [1015]. holbrooki [502, 1696]. holding [173]. Hollande [563]. holostean [945]. Holt [758]. home [896]. Homeostatic [876]. homogeneity [1684].

Hormonal [736]. hormone [1364, 1640, 561, 1578, 1392, 51, 263, 1656, 339, 950, 299]. hormone-induced [950]. hormones [1578, 178, 856]. Host [1658, 1111, 1698, 1452]. Host-constrained [1658]. host-specificity [1111]. hosts [1547, 588]. hot [572, 1696]. hot-spring [572]. hours [643]. Houttuyn [566]. HRBC [1147]. Hubbs [911]. hubbsi [1657, 424]. human [1167]. humoral [1241, 1250, 245]. Hunger [1684, 1405, 223, 578, 1382, 1587]. huntii [89]. Huron [775]. Huso [1720]. hybrid [487, 1557, 828, 431, 678, 1295, 490, 1341, 260, 363, 1379, 788, 495, 562]. Hybridization [411, 1720, 1313, 1535]. Hybrids [1665, 477, 1274, 158, 553, 1218, 33, 264, 257, 248, 295, 1024, 1096]. hydrating [950]. hydrilla [275]. hydroacoustic [1351]. hydrocarbons [1586, 327, 628]. hydrochloride [157, 171, 651]. hydrocortisone [1337]. Hydrodictyon [838]. Hydrodynamics [39]. hydrostatic [640, 684, 747]. hydroxyindoleacetic [1040]. Hyne [1183]. Hyperosmotic [144]. hyperplasia [653, 51]. Hypersensitivity [1228, 516]. hypoactivity [1446]. Hypophthalmichthys [363, 487, 477, 828, 817, 1034, 678, 187, 1612]. Hypophthalmicthys [158]. hypophysial [797]. hypophysial-gonadal [797]. hypothesis [1585, 1057]. hypoxia [541, 1689, 1136, 488, 200, 222, 381, 1635]. hypoxic [270, 977, 1150, 1166]. Hypseleotris [833].

icefish [219]. Iceland [1555, 1752]. Ichthyophonus [1324]. Ichthyophonus-type [1324]. Ichthyophthirim [824]. Ichthyophthirius [1257, 1266, 625, 100]. Ichthyoplankton [1506, 1609]. ictaluri [822]. Ictalurus [1554, 822, 384, 561, 823, 100, 304, 764, 665, 516, 1182, 200, 642, 666, 194, 87, 307, 303]. idella [487, 477, 828, 1480, 158, 678, 1397, 15, 130, 260, 363, 59, 148, 275, 562, 55]. Identification [115, 1281, 1718, 517, 287, 1580, 743, 1540, 1096, 284]. Identifying [1738]. idus [1044]. II [1186, 1129, 83, 1653, 812]. III [1130, 33, 1654, 1118]. Ilisha [717, 1041]. Iljin [92]. Illinois [650]. Ilyophis [415]. imbricatus [317]. immature [353, 207, 1031, 750, 1486, 567]. immediately [1375]. immersion [225, 485, 108]. Immune [516, 1251, 1269, 1185, 1186, 1257, 1256, 17, 1271, 312, 994, 1147, 245, 1654, 773, 1242, 593, 1716, 989, 1690, 1254, 1253]. immunity [1238, 1147, 1246, 108]. immunization [100, 1654, 1242, 1258]. immunizing [805]. Immunocytochemical [1371]. immunoglobulin [1250, 1371]. immunoglobulin-containing [1371]. immunohistochemical [461]. immunoinhibitory [313]. Immunological [1634, 1653, 1654, 144, 957, 1240, 1484]. immunologically [779]. Immunomodulating [1268, 1250]. immunopotentiating [1256]. Immunostimulation [1241]. immunosuppression [1247, 1243]. Impact [1221, 1029, 891]. impar [1638]. impingement [1507]. implantation [1337]. implanted [1694]. implants [1319]. Implications [542, 1429, 1508, 885, 1324, 808, 898]. importance [1580, 1498, 1714, 1634, 1653, 1654, 1201]. important [1064]. impoundments [148]. impression [687]. improved [1724, 1366]. improving [901]. Inbreeding [393]. Incidence [602, 65]. Incidences [1308]. incident [919]. including [835]. increase [1657, 861, 1286]. Increased [511, 1288]. increasing [705]. increments [1332]. incubation [1641]. index [739, 904, 1018, 444, 141]. India [564, 450, 736, 1453]. Indian [95, 529, 481]. indicating [1735]. Indicative [1025]. indicator [1070, 751, 1225, 1232, 1648, 556]. indicators [365, 167, 1000, 1767, 513]. indices [1020, 1546, 1414]. indigenous [568]. Indirect [1468]. Individual [1723, 1279, 1406]. individually [849]. individuals [743]. Indo [829, 1363]. Induced [136, 696, 1269, 1533, 913, 199, 1446, 225, 1340, 815, 1026, 764, 640, 684, 898, 1719, 1233, 660, 1403, 779, 1343, 409, 1583, 1249, 950, 455]. Induction [915, 1641, 1243, 841, 1349, 780, 936, 677]. industrial [575, 690, 138]. infected [1273, 971, 1643, 75, 629, 1236, 1104]. infecting [238, 1159]. Infection [1328, 423, 16, 97, 1698, 928, 1266, 1272, 625, 921, 1176, 4, 1302, 514, 592, 182, 72, 87]. infections [1389, 727, 942, 60, 631, 1716, 989, 824, 1253, 1748, 547]. infectious [1272, 1117, 1271]. inferred [1285]. Infestation [317, 1542, 1262, 65, 748, 1642]. Infestations [358, 999, 430]. infiltrating [144]. infiltration [144]. inflammatory [1340, 655, 1262, 964]. Influence [265, 370, 1405, 474, 766, 991, 720, 812, 1118, 1388, 1097, 299, 970, 1060, 1753, 597, 636, 667, 1532, 925, 1378, 1174, 220, 59, 328, 229, 1085, 520, 1371, 1200, 1400, 416, 1587, 1403, 977, 837, 1040, 1693, 1660, 1055]. Influences [1100, 428, 181, 203, 1289]. influencing [929, 485]. information [1508, 1746]. infrastructure [1399]. ingestible [1069]. inhibiting [214]. inhibition [945]. Inhibitory [1473]. inhibits [1462]. injected [759, 1264, 1270]. injectijon [225]. injection [1127, 773, 593, 451]. injection* [733]. injections [1644]. inland [798, 891]. inner [296, 255, 348]. innervation [468]. insecticide [188]. Inshore [1521, 336, 853, 874, 85, 1723, 1430, 529]. inspection [1647, 951]. insulin [528, 456, 1153]. Intact [1259, 1366, 1167]. intake [1070, 853, 325, 515, 543, 363, 960, 1713, 505]. integument [382]. intensity [1058]. Intensive [788]. intensively [842]. Inter [479, 954, 996]. inter-specific [954, 996]. Interaction [1758, 1496, 762, 681, 1738]. Interactions [5, 694, 1636, 920, 484, 1747, 418]. interactive [350]. interconversions [277]. Interference [492]. intergeneric [1379]. intermediate [1547, 1411, 175]. interrenal [492, 858, 647, 1145, 1076, 116]. Interspecific [1759, 920]. intertidal [673]. intestinal [648, 388, 537, 670, 1677, 1621, 1651]. intestinalis [1542, 663, 227, 1583, 926, 1390, 1658]. intestine [913, 195, 338, 1306, 1675, 619]. intra [954, 996]. intra- [954, 996]. intrafamilial [415]. intraovarian [912]. intraperitoneal [225, 1264, 451, 733]. intraperitoneal-injected [1264]. intraperitoneally [1270]. Intraspecific [695, 960, 800]. introduced [1208, 405, 1421, 1313, 887]. introduction [1029]. inversion [1134, 760, 1077]. invertebrate [987]. invertebrates [473]. investigation [398, 650, 749, 1139, 1646, 473, 543, 944, 230, 616]. investigations [703, 347, 54]. involved [4, 552]. involvement [1367, 150]. involving [1747, 295]. iodine [183]. ion [1392, 319, 950]. Ionian [1680]. ionic [1613, 995, 31, 54, 1713]. ionoregulatory [1711]. ions [1238, 428]. IPN [1272, 1271]. Ipnopidae [1194]. Iraq [172, 1335]. Iraqi [1518]. Ireland [1542, 612, 929, 953, 1705, 1419, 1701, 206, 1585, 149, 539, 589, 599]. iridescence [1343]. Irish [1199, 865, 1470, 486, 1316, 1743]. iron [519]. irrigation [906]. isheriensis [830]. Island [1374, 1421, 98]. Islands [529, 961, 997, 424]. Isle [573, 6]. Isles [1099, 1436]. isogeneic [1147]. isolated [910, 206, 1664]. isolates [625]. Isolation [164, 1067]. isomerase [1099]. isopod [317]. isopropyl [367]. isozyme [1557]. Israel [781, 1703, 1588]. Italy [411, 304, 1696, 38]. Itchen [1015]. IV [737]. Iwamoto [457].

J [1082]. jack [207, 317]. Japan [1733]. Japanese [1651]. japonica [566]. japonicus [1487]. Jauela [942]. jellycat [1287]. Jenyns [1451, 317]. Jersey [639, 997, 191]. Jiloá [1063]. Joaquin [1561]. John [1622]. Jordan [186, 1379, 1286, 820, 790]. julis [1134]. just [1666]. Juvenile [987, 401, 216, 1425, 1020, 759, 85, 909, 644, 1195, 1676, 728, 1514, 533, 370, 884, 1736, 110, 920, 223, 1053, 763, 1479, 1347, 193, 1279, 1174, 1375, 963, 745, 954, 996, 1239, 699, 1481, 293, 360, 68, 471, 1522, 1633, 1623, 532, 1692, 1708, 994, 1513, 933, 1582, 1687, 1200, 1526, 1337, 1212, 1450, 410, 740, 419, 704, 966, 1160, 1292, 1485, 534, 1388, 517, 1055, 565]. juveniles [1539, 1759, 917, 986, 1081, 930, 1281, 1733, 491, 848, 1615].

Kafue [50]. Kainji [1022, 1326, 1080]. Kalimantan [96]. Kamburu [32]. Kanawha [1729]. Kariba [1085, 636, 328, 1208, 1334]. Karyology [1203]. Karyotype [1720]. karyotyping [470]. Katsuwonus [35]. Katz [1302]. Kavirondo [255, 348]. Kay [938]. Kedah [1206]. kelp [832]. Kenya [32, 1339, 202, 1280, 1404]. Kenyan [1430]. Kessler [92]. keta [1737, 1487]. ketoandrostenedione [1473]. ketotestosterone [293, 117]. Key [1747, 939, 1199, 1379, 1540]. Keynote [882]. Kidney [1423, 1102, 1117, 855, 1393, 360, 43, 1255]. Kielder [890]. Kiernik [547]. kill [463]. Killifish [392, 1394, 1426, 312, 313, 493]. Kinetic [1446, 1724]. Kinetics [1147, 841, 905, 1242]. King [1561]. Kinneret [781]. kisutch [1759, 1195, 388, 668, 78, 51, 207, 385, 479, 1369, 31, 1586, 494, 1180, 1572, 465, 284]. Kner [1293]. knife [1010]. known [813, 793]. koelreuteri [535]. Koi [1615]. Königssee [974]. krill [578]. Kröyer [1287, 1099, 1395, 1500]. Kuala [1012]. Kuwait [535, 1456]. Kuwaiti [1092].

L [310, 311, 294, 532, 1095, 785, 490, 497, 912, 479, 526, 248, 1188]. L-thyroxine [479]. L. [1761, 1735, 583, 448, 1035, 1425, 366, 407, 323, 1770, 1458, 506, 1473, 398, 226, 1527, 1720, 1539, 1533, 1551, 379, 610, 853, 786, 874, 1380, 1542, 386, 710, 776, 607, 1439, 392, 877, 1161, 663, 544, 1626, 5, 256, 144, 1445, 81, 286, 1120, 1418, 90, 156, 236, 971, 1192, 1491, 1640, 1676, 1700, 1102, 1327, 459, 1134, 574, 559, 835, 579, 151, 560, 883, 658, 1663, 1669, 728, 1514, 145, 246, 873, 595, 929, 462, 1074, 1391, 324, 333, 498, 45, 170, 533, 843, 865, 953, 1705, 1464, 1357, 1536, 1717, 1502, 646, 587, 927]. L. [1272, 21, 247, 1721, 1580, 1184, 48, 1002, 600, 2, 1462, 3, 1629, 1419, 1701, 541, 1619, 376, 508, 53, 370, 921, 353, 334, 1531, 368, 1726, 1636, 8, 206, 171, 1044, 693, 713, 1240, 1555, 851, 1376, 550, 935, 136, 40, 1113, 1437, 1381, 988, 1136, 1155, 76, 278, 1666, 20, 340, 763, 341, 168, 1196, 1360, 681, 1704, 711, 787, 729, 615, 630, 679, 1461, 943, 1605, 1411, 101, 358, 940, 1420, 1479, 635, 632, 1142, 1685, 1384, 414, 426, 441, 511, 799, 46, 1056, 1755, 1174, 488, 1029, 737, 659, 1647, 522, 870]. L. [1098, 1288, 1396, 1707, 373, 1552, 456, 525, 1749, 383, 141, 1019, 61, 102, 88, 122, 220, 153, 390, 325, 332, 515, 484, 474, 543, 944, 82, 91, 1754, 919, 1351, 792, 745, 250, 143, 1373, 1498, 125, 227, 588, 954, 996, 1628, 791, 596, 1599, 1548, 500, 605, 1173, 326, 730, 1275, 1239, 1026, 1105, 1481, 338, 563, 1354, 1691, 105, 1217, 606, 770, 234, 428, 1668, 1119, 178, 293, 360, 624, 573, 655, 866, 1303, 1154, 68, 471, 1637, 1644, 665, 86, 793, 1521, 1522, 1482, 1301, 1447, 164, 42, 617, 732, 1724, 1261]. L. [435, 1190, 1361, 312, 313, 229, 346, 1441, 387, 277, 1492, 1692, 115, 235, 1298, 1047, 513, 581, 1710, 137, 1163, 1490, 1513, 30, 769, 570, 291, 503, 1622, 1559, 486, 801, 245, 1339, 930, 933, 1582, 675, 6, 230, 520, 1127, 478, 396, 408, 369, 382, 521, 962, 1467, 404, 856, 65, 305, 514, 647, 861, 748, 1008, 1076, 1352, 1601, 1642, 1213, 1678, 1512, 660, 806, 1248, 1015, 1510, 27, 418, 641, 685, 1570, 1200, 1077, 1342, 1655, 592, 1679, 322, 1587, 300, 423, 38, 1165, 1337, 1125, 773, 1486, 1403, 1333, 914]. L. [1372, 1358, 1672, 1627, 977, 80, 593, 1661, 279, 816, 1624, 1743, 218, 1268, 1236, 47, 182, 766, 1025, 26, 409, 683, 410, 740, 761, 331, 1660, 1193, 1404, 705, 150, 1009, 295, 991, 419, 704, 349, 1583, 720, 812, 1252, 114, 266, 1485, 1771, 956, 1460, 1544, 215, 179, 1620, 534, 643, 1331, 1344, 845, 1635, 1388, 783, 608, 616, 351, 352, 1143, 1325, 1645, 950, 734, 807, 965, 682, 1639, 1103, 849, 1158, 1641, 1096, 109, 139, 455, 422, 344, 1135, 1320, 1390, 1658, 1005, 565, 576, 887, 1206, 1756, 1438, 461, 1615]. labelled [485]. Laboratory [1552, 460, 62, 1214, 1445, 971, 709, 1765, 566, 311, 313, 932, 444]. laboratory-infected [971]. laboratory-reared [1765]. Labrador [1738]. labrax [1273, 379, 1498, 1447, 914, 1372]. labrid [1563]. Labridae [1134, 1319]. labrosus [110, 994, 1253, 422, 344]. labyrinth [1216]. Lacépède [767, 739, 1511, 663, 654, 590, 249, 233, 699, 1708, 744, 1316, 1297, 840, 1715, 496, 708, 820, 790, 869, 285]. Lacistorhynchus [936]. Lack [376, 1442, 1565]. lacking [52]. lactate [905]. lacustrine [1308]. lacustris [47]. Laécpède [854]. Laemonema [758]. laevis [971, 1547, 588, 1585, 1710]. lagoon [189, 218, 480]. lagoons [38]. Lagos [717, 430, 1041]. Lahontan [855]. Lake [1012, 1022, 1184, 775, 212, 255, 348, 818, 1339, 900, 314, 1404, 1421, 95, 81, 286, 802, 32, 974, 268, 1351, 479, 11, 1084, 1072, 542, 47, 182, 308, 722, 1320, 261, 650, 202, 584, 174, 636, 1218, 1495, 32, 1341, 1665, 885, 1437, 401, 1029, 1105, 51, 328, 1208, 1334, 1413, 1063, 1472, 1085, 1280, 1326, 538, 542, 1188, 275, 781, 1718]. lakes [365, 901, 1389, 987, 972, 985, 862, 855, 1113, 278, 711, 731, 486, 1467, 761, 1620, 869, 51, 207, 479]. lamellae [44, 1106]. lamellar [653]. Lamniformes [1130]. Lampetra [460, 9, 478]. lampretaeformis [1211, 62]. lamprey [1487, 9, 478]. lampreys [460, 1089, 1198]. lamta [1545]. lancetfish [1330]. landlocked [607, 579, 174]. Landsort [398]. Lanka [862, 980, 218]. Lankan [731, 1457]. large [463, 846, 1697, 893, 591, 1379]. largemouth [739, 924, 590, 1336, 1432, 840, 496]. larva [1605]. larvae [1533, 1765, 1007, 667, 1043, 1323, 21, 247, 1534, 509, 566, 935, 986, 692, 1709, 1643, 605, 601, 1474, 1734, 788, 1741, 60, 1523, 372, 1683, 1589, 683, 708, 1331, 1332, 1674, 1673, 1615]. Larval [833, 1275, 1568, 1735, 741, 638, 69, 1011, 1398, 635, 1175, 1142, 1101, 511, 1731, 460, 1540, 1089, 794, 1149, 1088, 629, 1661, 722, 1641, 850]. larvivorous [316]. lascaris [1638]. Latent [864]. lateral [90, 156, 321, 512, 296, 92, 1088, 165]. lateral-line [92]. laterna [124]. Lates [776, 798, 1029, 1339]. latifrons [637, 758]. latipes [281]. latitude [1160]. Latrobe [1646, 296]. lavaretus [1192, 1491, 646, 1125, 80, 1661]. Lawrence [1751]. layer [659]. layers [77]. laying [435]. lazera [664]. Ldh [1704]. Ldh-5 [1704]. Lead [395, 913, 1416, 1672]. leading [944]. Lebistes [280]. Lebowa [727]. Lee [70]. Léger [563]. Leiopotherapon [1209, 1405]. Leiostomus [654, 1708]. Length [1429, 1215, 367, 94, 1378, 1746, 1590, 329]. Length-dependent [1429]. length-frequency [1746]. length-related [1590]. lengths [70, 1430]. lens [45, 1225]. lentic [1221]. Leone [1745]. Lepidosiren [1617]. Lepomis [909, 167, 1416, 395, 662, 1446, 694, 604, 620, 876, 1114, 869, 1675]. leptocephalus [149]. Lernaeocera [921]. lernaeopodid [1562]. LeSeuer [946]. Lesina [38]. Lesion [936]. Lesions [105, 974, 1187]. Lessepsian [1440]. lesser [1571, 83, 573]. LeSueur [1554, 905, 677, 1306, 561, 863, 193, 493, 1182]. Lesueurigobius [364]. Lethal [201, 448, 1]. lethrinid [402]. Leucine [414]. Leuciscus [386, 411, 1274, 971, 1391, 1758, 1341, 1044, 229, 346, 1344, 1095]. leucocrit [571]. leucocyte [434, 112, 866]. Leucocytes [1619, 1245, 1075, 1164, 1600, 1046, 115, 1147, 478, 962, 1653, 1244]. leucopenia [1145]. leucopoietic [1619]. leukocyte [725]. levamisole [1268]. Levant [1379]. level [1698, 88, 339, 1504, 1587, 1621]. levels [561, 7, 778, 1370, 713, 1578, 1278, 447, 1609, 294, 870, 383, 792, 1202, 326, 1392, 1250, 858, 1656, 137, 1431, 1127, 1153, 806, 1182, 1169, 1372, 1040, 410, 991, 576, 221]. lewini [1463]. lewisi [1144]. licha [298]. Life [1398, 145, 646, 1732, 1101, 1709, 1730, 1483, 1359, 1441, 627, 686, 1549, 1687, 1752]. lifestyle [1211]. ligands [1693]. Light [195, 1044, 1046, 1545, 1193, 219, 667, 1164, 131, 1058, 1599, 1293, 1492, 794, 478, 1343]. Light- [1044]. light-induced [1343]. Ligula [1542, 663, 227, 1583, 926, 1390, 1658]. like [43]. Limanda [1327, 693, 713, 168, 1196, 6, 1005]. limandoides [107]. lime [1639]. Limia [1110]. limiting [1282, 887]. limits [571]. Limnothrissa [636, 1208, 1334]. lindane [1238, 541]. line [787, 296, 1265, 92, 846]. linear [1142]. lineatus [316]. lingua [1253]. Linn [1603, 1118]. Linn. [301]. Linnaeus [35, 1274, 1488, 176, 1089, 271, 1442, 1149, 1348, 1363, 1588, 95, 1433]. Linné [58]. Liparididae [546]. Liparis [452]. Lipid [421, 48, 698, 51, 263, 214, 1372, 510, 991, 567, 965]. lipids [207, 990]. lipofuscin [621]. Lipophrys [682]. lipopolysaccharide [17]. lipoprotein [1372]. lipped [110, 994, 1253, 344]. liquid [175, 397]. List [880, 1051]. literature [335]. little [813]. little-known [813]. littoral [741]. live [41, 1650]. liver [739, 407, 197, 48, 2, 353, 368, 1578, 855, 290, 326, 263, 385, 479, 360, 1127, 811]. living [855, 545]. Liza [49, 1335, 536]. lizard [400]. Llyn [145]. loach [373, 300, 423, 1286]. loading [912]. loads [713]. Lobo [610]. localization [657]. location [875]. loch [532, 643, 1491, 1376, 266, 1282]. loci [506]. locomotor [395, 283, 719]. locus [1704, 1750]. Loire [1522, 1526]. Lom [563]. Loma [1015]. Lomond [1491]. Long [1761, 1706, 1742, 1028, 238, 1202, 1159, 1554, 1745, 787, 107, 779, 1649]. long-spined [1649]. Long-term [1761, 1706, 1742, 1028, 1202, 1159, 1554, 1745, 779]. longear [1416]. longiceps [1079]. longitudinal [1731]. Lönnberg [219]. look [1054]. Lophim [865]. Lopholatilus [819]. Lortet [781]. loss [508, 692]. Lota [1384]. Lough [1542, 953, 667, 206, 1183, 539, 589, 599]. Louisiana [1014]. Low [1725, 448, 1380, 236, 146, 1609, 1071, 1662, 699, 1550, 1492, 319, 1006, 351]. low-light [1492]. Lowe [1330]. Lower [1269, 639, 1345, 818, 191]. lowland [1305, 1095]. Lucia [1495]. luciopercae [182]. Lucioperka [670]. lucius [1070, 81, 658, 498, 985, 1629, 988, 679, 101, 46, 745, 1252, 1141, 95, 1135, 1448]. luderick [1170, 1377, 262]. Lumpenus [1211, 62]. lumpfish [1066]. lumpsucker [544, 605, 624]. lumpsuckers [1721]. Lumpur [1012]. lumpus [544, 1066, 1721, 605, 624, 1047]. Lunar [1126]. lungfish [528]. lupus [1172]. luridus [1440]. lusci [921]. luscus [921]. luteum [429]. Lutjanidae [724]. Lutjanus [52, 961]. lutkeni [942]. lutrensis [948]. luzonensis [834]. lymphocystis [592]. Lymphocyte [458, 433]. lymphocytes [161, 449, 872, 749]. Lymphocytopenia [1076, 1352]. Lymphoid [1263, 256, 560, 1231, 1264]. lysozyme [1236]. Lytic [549]. Lytocestus [1206].

m [887, 1082, 476]. maccoyii [1073]. Maccullochella [555]. Macedonia [1341, 1665]. macedonicus [1341]. mackerel [1761, 76, 1058, 1071, 1411, 1303, 317, 1149, 1325]. macrocephalus [482]. macrochims [620]. macrochirus [909, 167, 395, 662, 1446, 694, 604, 876, 1114, 869]. macrolepidota [1012]. macrophage [232, 915, 1264, 1239, 1249]. macrophages [147, 1235, 1634, 1664, 970, 1673, 1237]. macrophyte [979, 568]. macrophytes [1055]. macropomum [440]. macrotrachelus [1562]. Macrouridae [123, 481, 457]. Macrourus [767]. maculata [1150, 1166]. maculatus [1686, 502, 285]. made [32, 862, 731]. main [1729]. main-channel [1729]. mainland [801]. maintained [709, 652, 1459]. maintenance [1608, 1750]. majalis [392]. major [896, 538, 1733]. Malaŵi [1472, 1421, 1718]. malachite [112]. Malaysia [1012, 906, 1206]. Male [1143, 1406, 158, 678, 1663, 1669, 294, 260, 363, 105, 257, 1714, 1618, 137, 1163, 806, 1182, 1486, 117, 677, 1306, 127, 114, 1289, 1607, 859, 1148]. male-typical [1607]. males [1274, 1429, 1462, 1472, 1688]. Mallotus [297, 476]. Malm [364]. man [32, 862, 731, 573, 6]. man-made [32, 862, 731]. managed [148]. management [883, 885, 1727, 904, 899, 898, 889, 902]. manganese [1238, 1342]. mangrove [1356, 1016]. mangroves [1499]. Manitoba [95]. maraena [674]. marginatus [96]. mariae [1126]. marine [1170, 1377, 1195, 633, 1004, 1465, 800, 1152, 588, 1628, 1159, 28, 715, 201, 1183, 1767, 1430, 1078, 984, 177, 400, 1769, 329, 58]. Marini [424]. marinus [1089]. Mark [614]. Mark-recapture [614]. marker [1704, 1707, 519]. markers [568]. market [776]. market-size [776]. marking [284]. marmorata [1106]. marmoratus [1249]. marsh [1549, 691]. marshalli [457]. martensi [1618]. masculinized [1362, 911]. masou [1279]. masquinongy [134, 946]. mass [821, 1244]. Massachusetts [1068]. masu [1279]. Mate [1062, 1482, 1063, 1143, 1064]. materials [1504, 811]. maternal [1215, 1378, 326]. Mathematical [241]. Matias [1657]. Mating [1163, 135, 859, 998]. Matrinchã [416]. Maturation [564, 633, 1462, 737, 294, 264, 257, 689, 1361, 503, 1622, 396, 408, 369, 521, 1076, 677, 1404, 797, 215, 950]. mature [310, 1139, 353, 479, 1163, 1486]. matured [1306]. maturing [1685, 105, 806]. Maturity [476, 538, 1761, 172, 1406, 1737, 216, 1148]. Mawddach [473]. maximal [1475]. Maximum [1196, 336, 1597, 1645, 139]. maximus [746, 1120, 1074, 763, 1056, 617, 732, 1298, 1442, 64, 377]. may [1544]. McDonald [529]. McDowall [308]. Me [1750]. Me-2 [1750]. meal [1120, 704]. meals [693]. measure [10, 141]. Measurement [1412, 1546, 1597, 1578, 237]. measurements [477, 367, 496, 307]. Mebendazole [320]. Mebendazole/trichlorfon [320]. mechanical [1004, 1013]. mechanism [841, 1496, 1724, 1259, 300, 1337]. mechanisms [196, 1247, 552, 1544]. medaka [834]. Mediterranean [829, 480, 1598, 758, 1741, 1687, 1363, 1703, 1588, 1440, 848]. Meetings [57, 252, 454]. meets [1300]. megalotis [1416]. Megathylacus [942]. meiotic [814]. Melamphaidae [1324]. melanin [621]. melano [232, 147, 915, 1264, 1239, 1249]. melano-macrophage [232, 915, 1264, 1239, 1249]. melano-macrophages [147]. Melanogenesis [601]. Melanogrammus [1196, 1378, 322]. melanophore [1322]. melanops [603]. melanopterus [416]. melas [304]. melastigma [106]. melastomus [26]. Melvin [206]. membranes [1668]. membras [398, 1570]. menhaden [1014, 1646, 296]. Menidia [1214, 1215, 1370]. mento [925]. mercury [365, 1443, 1025]. meristic [1646, 791, 1316]. merits [568]. Merlangim [1642]. Merlangius [1418, 462]. merlangus [1418, 462, 1642]. Merluccius [1657, 638, 424]. Mersey [1524]. mesh [333, 972, 968, 1526]. meso [1538]. meso- [1538]. Mesodinium [1609]. Metabolic [646, 1240, 1166, 1070, 842, 1174, 61, 181, 203, 995, 750, 556, 740]. metabolism [917, 1323, 694, 391, 492, 1564, 1708, 1442]. metabolite [1392, 858]. metacercariae [1452, 1042, 211]. metacercarial [727, 1253]. metal [380, 1439, 1693]. metallic [519]. metallothionein [1127, 1442]. metals [1682, 245, 1655, 177, 350, 897]. metamorphosing [794]. metamorphosis [1224, 198, 1089]. metaphase [814]. Metazoan [95, 612, 212, 906]. Metechinorhynchus [72]. methacrylate [1560]. methacrylate-embedded [1560]. methaemoglobin [1021]. methallibure [137]. method [1597, 1765, 680, 434, 940, 1487, 1232, 1366, 1591, 756, 1140, 519, 108, 571, 1094]. methodology [895, 569]. methods [10, 878, 1374, 161, 470, 935, 620, 111, 259, 1329, 1351, 22, 1526, 1743, 1260, 657]. methyl [734, 733]. methyltestosterone [1031]. metrizamide [983]. Mexican [189]. Mexico [127, 169, 1516]. Mg [428]. mice [831]. Michigan [51, 261, 479, 1089]. micro [1664]. micro-organisms [1664]. microbands [1430]. microbial [1171]. microdensitometry [878]. microdissectional [1423]. microflora [1387, 755, 1298, 1651]. microhabitat [1710]. microlepidotus [270]. Micromesistius [69, 424]. Micropogon [368]. microps [1099, 1395, 1500]. Micropterus [739, 924, 590, 1336, 699, 840, 1715, 496, 708]. microscope [449, 1044, 28, 686, 949]. microscopic [219, 131, 120, 195, 1046, 659, 1154, 1545, 418, 1258]. microscopical [621, 1164, 478, 1193]. microscopy [648, 1176, 1593, 1001, 1599, 115, 1443, 656, 666, 1168]. microspora [30]. Microsporida [928]. microsporidan [1250, 639]. microsporidian [1015]. microsporidiosis [13]. microstructure [1735, 1294]. microtags [743, 742]. Microvascular [1216]. mid [948]. mid-west [948]. Midas [1062]. migrant [1345, 829, 1440]. migrating [176, 770, 1369, 956]. Migration [759, 25, 711, 1146, 854, 1224, 310, 638, 993, 1355, 1569, 193, 1019, 78, 1105, 1656, 866, 9, 1521, 992, 1153, 1188, 377, 1460, 565]. Migrational [1068]. Migrations [1161, 1513, 1184, 306]. Migratory [1531, 1502, 885, 8, 818, 1559, 1651, 1771, 352, 888]. Mikri [1218]. milberti [99]. milkfish [1031, 797, 1332]. milt [464]. mine [1416]. mineral [1342, 1655, 1006]. minimal [750]. mink [1479]. minnow [1357, 375, 698, 1065, 948, 1441, 856, 951, 1059, 60, 1077, 777, 649, 116, 744]. minnows [1432, 941, 1353, 1382, 1057]. minutus [462, 450, 210, 1714]. miodon [636, 1208, 1334]. Mira [1528]. Mirogrex [34]. Mitchell [261, 946, 555]. Mitchill [1670, 134, 483, 1003, 775, 268, 1574, 66, 748, 722, 276]. Mitochondrial [1755, 1707, 1565, 1753, 1285]. mitogenic [1248]. mitotic [434, 814]. mixed [1060, 1731]. mixed-species [1060]. mixtures [448]. mobilization [421]. mode [1530]. model [1536, 1323, 1053, 1142, 1100, 1504, 139]. Modelling [1748]. models [1757, 241, 1637, 902]. modifications [70]. modified [1487]. modifier [1241]. Modulation [1248]. Moenkhausia [1293]. Moisie [483]. Mole [892]. molitrix [817, 1034, 187, 1612]. mollies [859]. Mollusca [919]. molly [239, 327, 451, 733, 998]. monitored [1351, 1069]. monitoring [1503, 190]. monocytes [1270]. monocytic [1252]. monodeiodination [479]. monogamous [1062]. Monogenea [1111, 1365, 631]. monogenean [1140, 1577]. monogeneans [1698]. monogenetic [320]. Monoptems [1589]. monstrosities [645]. Montagu [452]. montagui [452]. Moore [1015]. Moorea [1368]. Moray [1768]. mordax [261]. Moreton [262, 306, 389, 439, 517, 1049]. morhua [226, 1676, 587, 927, 21, 247, 353, 334, 1360, 1749, 390, 91, 1754, 1599, 1548, 500, 755, 1119, 624, 1200, 1744, 1565, 410, 740, 1193, 1748, 1438]. Mori [1477]. Morid [758]. Mormyridae [922]. Morone [1413, 629, 936, 850]. Morphological [678, 1295, 913, 1424, 791, 1222, 1443, 1009, 1004, 1722, 962, 683]. morphologically [1606]. Morphology [943, 1617, 1399, 1480, 1086, 14, 548, 1014, 1423, 1405, 186, 1362, 1471, 525, 1558, 815, 554, 198, 833, 1319, 479, 871, 1182, 826, 359, 1249]. Morphometric [1376, 66, 1316, 1537, 1646, 82, 1591, 1594, 1718]. Morphometrics [1317, 1209]. morphometry [1381]. Mortality [1036, 870, 1217, 1620, 1735, 746, 324, 652, 809, 635, 1408, 836, 1694, 1208, 808, 1741, 1570, 1403, 1743, 308, 1733, 352, 1267]. mosaic [128, 299]. mosquito [427]. mosquitofish [1406, 1362]. mossambicus [367, 584, 1295, 1036, 537, 862, 600, 1164, 1600, 1393, 665, 731, 1115, 350, 1350, 1032, 1457]. motility [242]. motion [39]. motivation [1359, 1684]. motor [1537, 468]. mouthbrooder [1032]. mouthbrooding [1607]. Movement [1527, 1485, 907, 1734, 520, 469, 314, 888, 1604]. Movements [1514, 615, 1347, 306, 1512, 702, 185, 527, 1463, 1165, 643]. Mozambique [665]. MS [905, 1292]. mucification [660]. mucosa [648, 120, 1677, 1333]. mucosal [1653, 1654]. mucous [659, 856, 1358, 409]. mucus [1616, 1711, 501, 687]. mud [531]. mudskipper [1564, 132]. mudskippers [535]. Mugil [246, 234, 428, 30, 769, 291, 1348, 218, 301, 720, 812, 1118, 1674, 950, 734]. mugilid [1335]. Mugilidae [413, 1281]. muksun [443]. Müller [971, 1547, 297, 588, 338, 1710, 182, 99]. Mullet [916, 413, 49, 246, 110, 234, 428, 994, 769, 291, 1281, 1348, 536, 218, 301, 720, 812, 1118, 1674, 950, 734, 1253, 422, 344]. mullets [287]. multi [333, 972]. multi-mesh [333, 972]. multifiliis [1257, 1266, 625, 100, 824]. multiple [941]. Multivariate [504, 1718]. mummichog [511, 1594]. municipal [904]. Munro [702]. Murray [555]. Muscle [461, 1537, 1533, 1657, 321, 353, 334, 375, 1566, 76, 1101, 946, 82, 572, 175, 263, 360, 672, 1650, 556, 1088, 26, 1156, 165, 1645, 128, 299, 649, 1475, 329]. muscles [1398, 164, 1590]. musculature [1010, 90, 156, 512, 468, 396, 27]. muskellunge [134, 946]. Mustela [1479, 1401]. mutisii [1179]. mycobacteriosis [1303]. myctophid [868]. Mylossoma [1299]. myocardial [152]. myofibrillar [804, 213, 329]. myoglobin [175]. myotomal [375, 1566, 946, 672, 1590, 649]. myotome [461]. myotomes [406]. Myoxocephalus [1439]. Myxine [224]. myxinids [224]. Myxosoma [30]. Myxosporidia [30].

n [118, 30, 733]. N-methyl-N-nitrosourea [733]. namaycush [542]. nasus [1518]. Natal [445, 467, 623]. National [1472]. nationwide [1525]. native [1029, 1313]. Natural [1535, 1557, 463, 553, 490, 1218, 1341, 4, 316, 1644, 1379, 1018, 1743, 496]. naturally [1716, 911]. nature [68, 471, 26]. Neagh [1542, 953, 539, 589, 599, 667]. nebulosus [1554, 561, 654, 1182]. necrosis [1272, 1271, 750]. Negara [1012]. Nematalosa [1518, 702]. Nematoda [638, 511, 1748]. nematode [697, 1162, 1328, 47, 99]. nematodes [545, 629]. nemipterid [757]. Nemipterus [96]. neoplasm [1252]. nerka [1300, 1606, 225]. nesting [1032, 1457]. net [465]. net-pen [465]. Netherlands [1221]. nets [333]. Nettastomatidae [871]. neurogenic [1251]. neuromasts [217]. neurosecretory [1662]. neutrophils [1234, 1340, 1236]. nevadensis [118]. Newfoundland [1670, 1622, 579, 1308, 1313, 1204]. News [1733]. Nicaragua [1063]. Niche [1763]. nicholsi [868]. nidamental [435]. Nidelva [580, 870]. niger [1686, 845]. Nigeria [430, 1022, 922, 1220, 1090, 1375, 717, 1041, 1297, 1326, 1080, 1126]. Nigerian [336, 1016]. night [1160]. nigrodigitatus [249, 1297]. nigrofasciatum [582, 393]. Nile [1087, 1029, 1339]. niloticm [1621]. niloticus [696, 838, 776, 1295, 1002, 1689, 1593, 1113, 1437, 1029, 1339, 248, 1404, 1009, 1289]. Nilsson [462]. nine [409]. nine-spined [409]. ninespine [920]. nitrate [3]. nitrite [508]. nitrogen [986, 1725, 1358]. nitrogenous [1427, 153]. nitrosourea [733]. Nitzsch [399, 1716]. nobilis [487, 477, 828, 817, 158, 678, 260, 363, 562]. nodulosmin [974]. Noemacheilus [373, 875]. nomenclature [1433]. Non [426, 183, 877, 1139, 176, 1136, 1155, 243, 1637, 54, 562, 897, 875, 1688]. non- [562]. non-anadromous [877]. non-ionic [54]. Non-metals [897]. non-migrating [176]. non-parasitized [1136, 1155]. non-reproductive [1139]. non-salmonid [897]. non-sensory [243]. Non-specific [426]. non-surface [1637]. non-territorial [1688]. non-thyroid [183]. non-visually [875]. noradrenalin [1322]. Norfolk [1538, 894]. Normal [1284, 500, 601, 1474]. Norman [424]. normani [771]. Norris [1469]. North [853, 874, 25, 145, 1538, 772, 1700, 1705, 1732, 390, 791, 1330, 1309, 1503, 1696, 1290, 1761, 1172, 1739, 1571, 738, 767, 673, 101, 941, 345, 1764, 1576, 1303, 758, 948, 123, 481, 457, 546, 871, 1194, 721, 1642, 1679, 18, 162, 415, 765, 1748]. north-east [1705, 791, 1503, 1303]. north-eastern [1732, 1330, 1309, 1696, 1290]. north-west [1700]. northeastern [602]. northerly [1441]. Northern [1670, 1542, 953, 1023, 539, 589, 599, 1458, 398, 1161, 985, 1218, 865, 1567, 353, 787, 101, 1347, 1019, 1754, 1747, 18, 162, 1703, 1252, 1189, 688, 869, 757, 1535, 334]. Northumberland [1488]. norvegicus [282]. Norway [1458, 1161, 25, 462, 353, 334, 282, 580, 1347, 870, 1098, 1396, 1019, 390, 1754, 1105, 47, 182, 377, 783]. Norwegian [767, 787, 1146, 729, 358, 940, 999, 1084]. notacanth [149]. notatus [375, 698, 1382, 649]. Note [610, 324, 879, 278, 340, 522, 203, 1329, 338, 1574, 64, 377, 1344, 362]. Notemigonus [1682]. Notes [1144, 424, 1374, 224, 123, 481, 546, 1194]. Nothobranchius [700]. Notice [1311]. nototheniid [119]. notothenioid [1039]. Notropis [941, 948]. Nottinghamshire [267]. nov [1023]. novaezeelandiae [217]. November [372]. NRDC [292]. NRDC-143 [292]. Nuclear [80, 475, 878, 307]. nucleic [293, 360, 1623]. nucleotide [209]. number [766]. numbers [1466, 591, 1430]. nurse [437]. nurseries [1517]. nursery [1515, 1528, 401]. nutrient [1437]. nutrients [1377]. nutrition [569]. nutritional [610, 862, 713, 141]. NW [380]. Nyanza [1029, 1339, 314]. Nyasalapia [1718]. Nyctiphanes [578].

O [90, 156, 387, 18, 162, 492]. O-group [90, 156, 387, 18, 162]. O. [1295, 1621]. Obituary [524, 622]. Observations [403, 1541, 865, 1721, 1605, 1045, 1614, 441, 1152, 75, 1287, 1410, 1492, 532, 486, 271, 187, 314, 162, 1165, 400, 826, 419, 351, 1457, 380, 219, 568, 1606, 1409, 325, 596, 135, 62, 319, 322, 1080, 981, 1683, 949]. observed [105, 1443, 656]. obtained [330]. obtaining [434, 1149]. Occurrence [844, 1383, 727, 1133, 40, 211, 611, 1489, 1335, 1206, 1111, 13, 738, 1629, 1370, 1117, 941, 522, 1434, 233, 1239, 338, 906, 452, 1559, 469, 493, 1477, 191, 99, 517]. occurrences [1488]. occurring [633, 1016]. ocean [1322, 602, 481, 1069, 529]. ocellatus [1563, 533, 1246]. Ocular [1225]. Odontesthes [1683]. oesophageal [1295, 659]. oesophagus [441]. oestradiol [1578, 383, 792, 326, 1644, 750, 1127]. oestradiol-17 [1644]. oestrogens [1426]. off [430, 853, 559, 462, 1745, 717, 1041, 149, 1091, 96, 1453, 1703, 377, 1204]. offish [917, 127]. offspring [1232]. Ogilby [205]. oil [1118]. Oken [1530]. old [747, 1018]. older [431]. olfaction [521]. Olfactory [1397, 1180, 835, 1146, 764, 933, 1306, 666]. Oligocottus [673]. Oligolepis [564]. oligotrophic [47, 182]. olivaceus [67]. olly [1082]. olrikii [1649]. onald [1082]. Oncorhynchm [1606]. Oncorhynchus [1345, 1561, 1300, 931, 1759, 25, 644, 1195, 174, 388, 668, 993, 223, 225, 1737, 1279, 1487, 907, 78, 51, 207, 385, 479, 1656, 1369, 992, 31, 1153, 1586, 494, 1180, 1572, 1212, 1450, 465, 966, 1160, 1292, 284]. one [1201]. onset [1670, 1616, 369]. Ontario [51, 479, 581, 900, 542]. onto [1734]. Ontogenesis [291]. Ontogenetic [1407, 1731, 256]. Ontogenic [975].

Ontogeny [1074, 1231, 833, 1474, 613]. oocyte [1579, 1447]. oocytes [297, 476, 950]. Oogenesis [1599, 1638]. open [212]. Operant [344, 422]. opercular [556]. Ophicephalus [980]. Ophiocephalus [285]. Opisthoproctidae [1321]. optic [613]. optimal [1054, 1757]. Optimization [374]. optimize [1232]. optimum [1323, 259]. orally [1031]. orally-administered [1031]. order [1422]. Oregon [1345]. oreo [1686]. Oreochromis [696, 915, 838, 776, 1295, 1036, 862, 1002, 600, 1164, 1600, 1689, 1593, 1113, 1437, 1393, 572, 665, 1472, 1115, 1265, 1664, 1621, 1350, 1404, 1009, 1032, 1289, 1718, 1457]. Oreochromismossambicus [1231]. orfe [1044]. organ [1223, 525, 764, 666]. organically [1459]. organism [43]. organisms [1664, 1260]. organization [1216]. organs [1554, 1263, 256, 1264, 1017, 1635]. Oriental [405]. Orientation [1114, 966, 1160, 1292]. Oriented [1293]. origin [834, 1436]. Orkla [1396]. ornamentation [975]. Ornithodiplostomum [60]. oropharyngeal [289]. Orthodon [270]. Orthopristis [1111]. Oryzias [834, 281]. Osmerus [261, 522, 1084]. osmolality [533]. osmoregulation [551]. Osmotic [247, 605, 744, 947, 1301, 31, 291]. Ostariophysi [856]. Osteichthyes [1469, 553, 738, 452, 765]. Osteichthyes-Elopiformes-Elopidae [765]. Osteichthys [1330]. Ostracionidae [1363]. other [559, 1686, 224, 417, 1278, 243, 364, 856, 964]. Otolith [1735, 1481, 1719, 785, 550, 935, 1294, 1430, 1073, 1733]. otolith-tags [1733]. otoliths [1425, 1214, 1385, 1332]. ouananiche [1308]. Oued [721]. output [1219, 1358]. Ovarian [819, 1319, 536, 279, 34]. ovariectomy [383]. ovaries [1560, 48, 169]. ovary [407, 2, 1394]. over-ripening [617]. overcoming [442]. overwinter [1352]. Overwintering [352, 1408]. oviparous [1534]. ovulation [482, 279, 677]. Ovulatory [617, 732]. Owen [528]. oxidase [1010]. oxidative [1101]. oxotestosterone [294]. Oxygen [151, 21, 1002, 708, 1674, 330, 448, 1035, 1345, 1561, 1502, 863, 1532, 376, 821, 1697, 1609, 332, 59, 1301, 1468, 1510, 1299, 194, 285].

P [1270]. P. [1365]. Pacific [189, 829, 692, 1487, 1069, 1449, 1363]. Padogobius [1618]. Pagbilao [1499]. Pagothenia [1315]. Pagrus [1733]. pairing [1063]. palatability [1543]. palatine [1210]. Pallas [210, 135, 1714, 839, 443, 1428, 535]. pallasi [692]. Pampus [738]. panchax [106, 1042]. pancreas [400]. pancreatic [1272, 528, 1271]. Pangasiidae [1399]. Pangasius [1399]. pantothenic [653]. Paperna [1567]. papers [1051]. Papua [846, 1429, 827, 961, 1144]. paradoxa [1617]. Paragalaxias [504, 1203]. Paralichthys [235]. Paraliparis [546]. parameter [141]. parameters [1273, 993, 1278, 598, 1393, 963, 73, 1430, 902, 803, 1284, 1286, 1304, 628]. Parangipettai [1453]. paraquat [1635]. parasite [929, 1538, 1628, 1302]. Parasites [176, 1326, 1273, 854, 703, 1172, 612, 1630, 997, 212, 906, 1159, 1335, 95]. Parasitic [430, 939, 1199, 921, 106, 602, 30]. parasitism [1670, 836]. parasitized [1136, 1155]. parasitology [1555]. Park [1472]. Parodon [1469]. Parodontidae [1469]. Parophrys [374, 1187]. parr [310, 668, 1666, 1685, 1483, 1359, 143, 606, 1553, 137, 1163, 1352, 1486, 510, 1660, 150, 114, 1321]. parr-smolt [1553, 510, 1660]. part [879]. part-digested [879]. particle [1381, 1100, 1438]. particles [239]. particular [894, 1623, 760]. particulate [1654]. partitioning [909, 1598, 1500, 1181]. partitus [1251]. Parupeneus [820, 790]. parva [1383]. parvulus [1206]. Passa [1469]. Passage [827, 1508, 537, 311, 1431]. Passamaquoddy [1677]. Patagonian [424]. patch [875]. pathogen [847]. Pathogenic [558, 1260]. Pathogenicity [1271]. pathogens [1267]. Pathological [1562]. Pathology [822, 541, 1643, 811, 1324, 547, 1577]. paths [292, 377]. pathways [583]. patronus [1014]. pattern [1162, 817, 1465, 223, 1196, 1072, 1372, 182, 451]. patterns [506, 785, 63, 1398, 210, 737, 1100, 1633, 1687, 314, 344, 1456]. Paul [159]. Paulicea [942]. Paulo [610]. p’DDD [492]. pectoral [1648]. pectoralis [371, 447]. Peculiarities [143]. peeli [555]. Pelagic [544, 1751, 1324, 328, 1334, 602, 1085, 18, 162]. pelamis [35]. Peltorhamphus [217]. pelvic [52, 1192, 1384]. pen [465]. Penetration [239]. Peninsula [1670, 1308]. pentachlorophenate [934]. Perca [1626, 286, 145, 333, 498, 341, 775, 1142, 414, 426, 441, 238, 1217, 1574, 47, 182, 215, 179, 1620, 1055]. Percafluviatilis [873]. perch [1626, 286, 145, 873, 333, 498, 1209, 1405, 670, 341, 775, 1142, 414, 426, 441, 1029, 89, 238, 1202, 1217, 1574, 258, 1339, 47, 182, 215, 179, 1620, 1055]. Perciformes [738, 1172, 1763, 715]. Percoidei [1209, 1405]. percolating [472, 575, 690]. Percoll [872, 749]. percomorph [1530]. performance [354, 1048, 941, 175, 118, 707, 42, 465, 740, 1104, 534]. perfused [5, 910]. period [1736, 346, 389, 977]. periodicity [246, 1450, 1126, 1181]. periods [163]. Periophthalmus [535]. peripheral [1164, 1600, 225, 79, 357, 347, 115]. Peritoneal [655, 1614, 1009]. peritrich [748]. permeability [696]. Permethrin [292]. permit [919]. peroxidase [1045, 1410]. perspective [1772, 1221, 1410]. Peruvian [372]. pesticides [1238, 892]. petenense [808]. Peter [280]. Peters [52, 618, 7, 537, 600, 1393, 665, 731, 1115, 631, 350, 1032, 1457]. petroleum [709, 127, 169]. Petromyzon [1089]. Petromyzonidae [1198]. PFC [1147]. PGF [780]. pH [1533, 242, 236, 151, 1636, 1662, 572, 699, 254, 665, 319, 1006, 1169, 1715, 869]. PHA [1248]. phagocyte [1245]. phagocytes [669]. Phagocytic [1168, 1270, 1236, 1009, 1237]. Phagocytosis [1234]. pharmacological [750]. phase [987, 1176, 18]. phase-contrast [1176]. phenol [222, 381]. phenotypic [504]. phenylthiourea [601]. phenylthiourea-treated [601]. Pheromones [494]. Philippine [834]. Philippines [1374, 1499]. pholis [1133, 20, 1358, 682]. phosphatase [915, 670]. phosphatases [1075]. phosphogen [572]. Phosphoglucose [1099]. phospholipid [1668]. phospholipids [990]. photobehavioural [244]. Photoperiod [1361, 570, 728, 1360, 792, 1018, 279, 837, 1388, 34]. photoperiods [1678, 1660]. phoxinm [1536]. Phoxinus [1427, 1464, 1357, 1536, 1717, 1065, 1441, 856, 1077, 139]. pHs [146]. phylogenetic [1410, 1591]. phylogeographic [1704]. physical [600, 840, 496]. Physiological [1013, 944, 319, 739, 1211, 134, 1100, 1259, 1601, 231, 173]. physiology [1003, 437]. phytophagy [731]. pickerel [946, 406]. pigmentation [1047]. pike [1070, 81, 498, 985, 1629, 670, 988, 679, 101, 46, 745, 899, 1252, 1141, 95, 1135, 1448]. pilchardus [952, 1680]. Pimelodidae [942]. Pimephales [375, 698, 1382, 60, 777, 649, 116]. Pinang [906]. pineal [525]. Pinealectomy [536, 1109]. pink [931, 25, 993, 223, 1656, 992, 1153]. piscatorius [865]. Pisces [856, 413, 1321, 1407, 1207, 1663, 1429, 1007, 862, 1341, 1133, 942, 860, 1689, 1324, 1029, 504, 919, 1714, 1041, 1763, 1194, 502, 1367, 1683, 1023, 1344, 1027, 1203, 1038, 1457]. piscivorous [1572]. pituitary [383, 51, 207, 581, 769, 279]. pituitary- [581]. PKD [1269, 1117]. place [1515]. Plaice [1517, 1735, 1425, 1700, 48, 1240, 61, 102, 88, 122, 220, 153, 325, 332, 1239, 264, 257, 655, 866, 164, 387, 277, 1127, 1679, 1743, 807, 965]. planeri [9]. planktivorous [1323]. plankton [1523]. plant [511, 1362, 1545]. planted [53]. plants [158]. plaque [831]. plaque-forming [831]. Plasma [288, 713, 1613, 1153, 194, 397, 644, 1465, 1578, 1278, 1426, 1603, 1392, 263, 858, 1656, 59, 992, 137, 1431, 1182, 1372, 117, 331, 221]. plasticity [1497]. platessa [1735, 1425, 853, 786, 874, 1700, 48, 1240, 61, 102, 88, 122, 220, 153, 325, 332, 1239, 264, 257, 655, 866, 164, 387, 277, 1127, 1679, 1743, 807, 965, 1005]. platessoides [107]. Platichthys [583, 710, 376, 588, 264, 257, 1710, 592, 331]. Platycephalidae [196]. Platycephalus [205, 196]. play [1180]. Plecoglossm [1568]. plemobipunctatus [1274]. plerocercoid [681, 1262, 936, 1716, 1583]. plerocercoid-induced [1583]. plerocercoids [654, 974]. Pleuragramma [1262]. Pleuronectes [1735, 1425, 853, 786, 874, 1700, 48, 1240, 61, 102, 88, 122, 220, 153, 325, 332, 1239, 264, 257, 655, 866, 164, 387, 277, 1127, 1679, 1743, 807, 965, 1005]. Pleuronectidae [1007]. pleuronectiform [848]. Pleuronectiformes [1224]. Ploidy [1370, 1286]. plumbeus [160, 754]. podicipina [1202]. Poecilancistrium [654]. Poecilia [618, 7, 631, 239, 327, 451, 859, 998]. poeciliid [362]. Poeciliidae [502]. Poecilla [733]. Poeciloia [843]. poisoning [541]. Poland [591]. polarization [1587]. Pollachius [1411]. pollan [539, 589, 599, 667, 539, 589, 599]. pollock [986, 1747, 839, 1428]. pollutants [726, 897]. polluted [1502, 1503]. pollution [1439, 1501]. polychromatic [1062]. polyclonal [1260]. polycyclic [327]. Polyembryony [645]. polylepis [553]. polymorphic [1555, 1752]. polymorphism [1099, 1700, 1444, 1749, 1027]. Polynesia [1368]. Polyploidy [684]. Polyprion [1530]. polyribosomal [79]. polyunsaturated [746]. Pomacentrus [1251]. Pomatoschistus [1099, 282, 210, 1714, 1395, 1500]. Pomoxis [650]. Pomphorhynchm [1547, 1585]. Pomphorhynchus [971, 588, 1710]. Pond [1670, 94, 1622, 801]. ponds [555, 562, 1568, 1135]. pool [1537]. pools [1432, 922, 1090, 691]. Poor [1107, 462]. poor-cod [462]. Population [1020, 1539, 1753, 885, 268, 125, 1353, 868, 1499, 47, 1740, 697, 1162, 366, 1458, 518, 751, 853, 874, 703, 710, 286, 1700, 167, 895, 267, 953, 1502, 1736, 577, 1408, 1754, 227, 238, 716, 1202, 79, 11, 1522, 513, 1710, 96, 902, 1212, 1450, 496, 761, 1404, 1544, 1032, 1038, 1320]. populations [1518, 877, 13, 1389, 971, 1745, 1758, 985, 1705, 862, 1567, 206, 1727, 101, 1420, 893, 1707, 904, 770, 896, 731, 1767, 1722, 1316, 1728, 840, 1526, 1140, 977, 218, 1565, 1680, 1760, 1189, 1282]. pores [417]. portalegrensis [613]. Portugal [1528]. Portuguese [921, 1133]. position [1204]. possibilities [901]. possible [650, 892, 1489, 964, 43, 377]. Post [185, 527, 1704, 1017, 656, 1088, 1735, 1407, 709, 1139, 346, 482, 1562, 704]. Post- [527]. Post-embryonic [1017]. Post-fertilization [656]. Post-glacial [1704]. post-hatching [346]. Post-larval [1088, 1735]. post-mature [1139]. post-ovulation [482]. post-prandial [704]. post-settlement [1407, 1562]. post-spawned [709]. Post-stocking [185]. postlarvae [1521]. postlarval [469]. postprandial [917]. potassium [3, 1239]. potential [973, 836, 1251, 237, 1484, 74, 1255]. potentials [571]. Pottachius [835]. pout [462]. poutassou [69]. powan [1192]. powder [519]. powdered [279]. Power [518, 184, 1508, 511, 1526, 505, 1507]. practical [359]. prandial [704]. Pre [1497, 1061, 290, 870, 704]. pre- [704]. Pre-adaptive [1497]. pre-fasting [290]. pre-smolt [870]. pre-spawning [1061]. preceding [1375]. precocious [137]. precociously [310, 105]. Predation [1155, 1098, 1396, 770, 1053, 1479, 1483, 1359, 919]. predator [776, 1706, 658, 879, 1569, 1647, 1382, 951, 875]. predators [987, 1738]. predatory [1571, 1412]. predict [1504]. Predicting [1757, 1744]. prediction [903, 465]. Preface [881, 1494, 1050]. preference [828, 668, 933, 1064]. preferences [976, 494, 1212, 956, 1460, 1024]. preferendum [259]. preferred [1547]. pregnant [912]. Preliminary [380, 749, 641, 1244, 1683, 1334, 230, 687]. premigration [198]. Preparation [1312, 458, 756, 666]. preparations [625]. presbyter [518, 751, 627, 184, 505]. Presence [1106, 680, 1382]. present [891]. presentation [1526]. presented [1051]. preservation [443]. Prespa [1218]. prespawning [1723]. pressure [1571, 640, 684, 1141, 397]. Preston [1505]. Presumptive [1303]. Prevalence [1389, 379, 663, 511, 1159, 639]. previously [99]. Prey [1212, 1706, 1681, 679, 1569, 1056, 1552, 578, 1450]. Priacanthus [836]. primary [1167]. prior [527, 1666, 1604]. probable [952]. probatocephalus [864]. problem [1077]. problems [883, 442]. procedure [1149, 934]. process [624, 1338]. processes [799]. processing [1368, 1634]. produce [1544]. produced [910, 174, 1609]. produces [750]. producing [1232, 1241]. Production [842, 1020, 874, 776, 1427, 884, 121, 879, 401, 143, 1354, 1691, 1591, 721, 1244, 1260, 127, 169, 1620, 831]. products [842, 1248]. Profiles [1709, 384]. profitability [1552]. progeny [574, 1737, 114]. progesterone [326]. programmes [1580]. progressive [1136]. proliferation [327]. proliferative [1117, 43]. prolonged [209, 1145, 200, 1166]. promelas [60, 777, 116]. promoter [1684]. pronephric [1261]. pronephros [1583]. proof [203]. proper [398]. properties [1083, 1004, 1013, 847, 164, 549, 1150]. Proprioceptive [406]. propulsive [27]. prostaglandin [780]. protandrous [760]. protease [501, 816]. protecting [805]. Protection [100]. protective [625]. Protein [506, 698, 567, 619, 10, 838, 487, 597, 1415, 680, 7, 48, 568, 334, 1578, 447, 497, 293, 339, 360, 672, 1259, 994, 570, 237, 991, 1255]. proteins [644]. Proteocephalidae [942]. Proteocephalm [1643]. proteolytic [178]. protocol [1476]. Protogynous [937, 1134, 186, 402]. Protopterus [528]. Prototransversotrema [1348]. Prototroctes [674]. Protozoa [1176, 1133, 30]. proximate [102, 1633]. Pseudapocryptes [132]. pseudo [258]. pseudobranch [1044, 1216]. pseudobranchs [28]. Pseudocyttus [1686]. Pseudodactylogyms [1365]. Pseudodactylogyrus [1698]. Pseudomugil [243]. Pseudophyllidae [233]. Pseudophyllidea [1583]. Pseudopleuronectes [709, 1723, 928, 1250, 1677, 656, 639]. Pseudopleuronects [191]. Pseudorasbora [1383]. Pseudoterranova [1748]. Pseudotolithus [430]. Pterophyllum [1067]. ptychocheilus [60]. Publications [1227]. Puget [1187]. Pulau [906]. pulse [857, 900]. punctatus [183, 436, 822, 384, 823, 1131, 1296, 100, 764, 133, 347, 665, 516, 200, 642, 666, 194, 87, 307, 303]. Pungitius [976, 920, 409]. puniceus [937]. Puntius [906, 780]. pupfish [118]. Purification [983]. Purrumbete [174]. purseseine [63]. Putative [989, 983]. Putz [1015]. pygmaea [938]. pyloric [1675]. pyrene [1270]. pyridoxine [436]. pyriformis [625].

quadracus [976]. quadricornis [1439]. qualitative [1298, 932]. quality [1527, 537, 693, 1113, 473, 904, 500, 1724, 1107]. Quantification [1021, 1693]. Quantitative [48, 1132, 497, 1298, 932]. quantitatively [741]. Quebec [483]. queen [919]. Queensland [772, 502, 306, 517, 1049]. Quensel [1081]. Quoy [1170, 1377, 262].

R [1018, 411, 1021]. R. [1020]. rabbitfish [1567]. races [33, 315, 737]. radiation [40]. radio [1478]. radioactive [1028]. radioassay [1693]. radiographic [1329, 248, 519]. radioimmunoassay [420]. radionuclides [1504]. radiotelemetry [556, 330]. Rafinesque [650, 822, 909, 384, 167, 823, 395, 662, 1446, 694, 375, 698, 100, 620, 764, 665, 1382, 516, 1114, 60, 200, 26, 642, 666, 777, 194, 87, 649, 869, 1675, 307, 116, 303]. RAFIX [1478]. Rainbow [242, 147, 726, 261, 1245, 1230, 1387, 16, 97, 1219, 288, 1185, 1186, 1075, 1389, 841, 1054, 1681, 597, 1102, 244, 197, 1235, 842, 548, 1415, 1762, 1184, 778, 199, 265, 714, 508, 131, 120, 274, 195, 289, 1578, 1117, 575, 690, 1476, 417, 1697, 1278, 1256, 1616, 1711, 434, 335, 391, 290, 501, 112, 982, 947, 421, 1543, 1232, 653, 815, 1608, 1417, 292, 1392, 1302, 79, 37, 209, 357, 263, 858, 707, 689, 339, 640, 684, 672, 75, 1694, 747, 1613, 995, 1021, 318, 1018, 319, 66, 216, 22, 1371, 1695, 1601, 1015, 814, 803, 556, 719]. rainbow [117, 126, 43, 718, 1716, 837, 934, 237, 805, 1693, 989, 222, 173, 381, 485, 657, 1167, 1168, 1690, 1104, 108, 1270, 128, 163, 299, 330, 567, 432, 649, 926, 1254, 1673, 397, 619, 774, 361, 628, 221, 1338]. rainbowtrout [779]. rainforest [688]. raised [94, 762, 1604]. Raja [340, 1628, 503]. Rajiformes [1129]. rakers [1381]. ramada [536]. Ramsay [205]. ranarum [844]. ranching [276]. range [1478, 210, 896]. ranges [1284]. ranging [719]. Rannoch [1376]. Raphidascaroides [1328]. rapid [259, 1613, 173]. Rare [813]. rate [366, 1070, 16, 786, 636, 1464, 1602, 1476, 278, 20, 679, 1174, 315, 1288, 61, 181, 203, 1608, 995, 435, 526, 933, 190, 1403, 704, 349, 1073, 114, 1055, 1390]. rates [1458, 202, 603, 724, 693, 851, 223, 1666, 241, 332, 484, 474, 960, 1623, 1572, 1624, 237, 351, 567, 432, 1024]. ratio [444]. ration [323, 1195, 1427, 1464, 1357, 1196, 679, 1608, 1200, 1587, 1572, 1450]. rations [163]. ratios [167, 1765, 209, 850, 757]. rattail [481]. rattails [1489]. raven [283]. ray [743, 742, 1453, 180]. rays [1384, 1129]. rcvs [1082]. reaction [1353, 1694]. reactions [794, 1673]. reactivity [256]. Reading [742]. Reappraisal [878, 1224]. reared [379, 1765, 185, 527, 566, 711, 630, 1396, 532, 1018, 1678, 1568, 351, 352, 284, 1604, 1615]. rearing [1406, 1279, 1373, 858]. recapture [185, 527, 1602, 1288, 614]. Received [204, 273, 343, 586, 1123, 1178, 1227, 1611, 1632]. recently [1144, 1492]. recognition [1582]. reconstitution [1484]. record [1214, 758, 149, 1330, 1309, 1649, 1363, 765]. recorded [643]. recording [522]. Records [491, 1274, 544, 926]. recoveries [306]. Recovery [305, 163, 706, 987, 977, 1338]. recrudescence [1109, 34]. recruit [440]. recruited [874]. Recruitment [776, 1475, 1456, 1517, 775, 886]. rectum [195]. red [1211, 533, 157, 947, 1013, 1487, 175, 79, 209, 672, 1659, 1021, 1149, 1596, 1348, 1246, 1733, 961, 1567]. redbanded [89]. redds [1541, 271]. reduced [1155]. reducing [861]. Reduction [608, 1021]. reef [254, 1725, 1367]. reefs [1374]. reference [390, 894, 760, 1348]. reflectance [668]. Regan [371, 447]. regenerating [67]. Regenerative [347]. regime [1633, 1623]. regimes [792, 279]. regional [1219, 1642, 1748]. regulated [870]. regulating [770]. regulation [1736, 1408, 268, 876, 1613, 31, 291, 744, 1713]. regurgitation [1412]. reinfection [1117]. related [151, 1578, 193, 1543, 894, 357, 347, 300, 148, 1590]. relating [1508]. Relation [1737, 369, 1537, 1425, 1527, 1445, 310, 574, 1002, 1079, 713, 1579, 1471, 1552, 337, 996, 1173, 1222, 339, 1510, 27, 23, 1343, 377, 419, 505, 1189, 1674, 299, 869, 888, 329]. relations [843]. Relationship [1698, 1089, 979, 94, 863, 1569, 294, 1608, 457, 31, 283, 1358, 128, 1475, 1024]. Relationships [1215, 620, 1759, 1686, 2, 1157, 1087, 390, 504, 745, 1285, 415, 643, 1456]. Relative [432, 1681, 94, 568, 1378, 1201]. release [456, 565]. released [185, 527, 1146, 1733]. relevance [1736, 1713, 677]. remembrance [417]. removing [320]. Renal [1039, 548, 1187, 1141]. Renibacterium [1102, 1260]. repair [1337]. repeated [598]. repetitive [1644]. repopulation [1722]. Report [463, 911, 559, 687]. Reproduction [49, 498, 235, 308, 167, 1422, 326, 1678, 1679, 761, 150, 582, 1032]. Reproductive [674, 94, 673, 371, 502, 1596, 1530, 1395, 1023, 790, 1012, 701, 1092, 1033, 403, 1139, 1669, 702, 462, 623, 827, 993, 980, 118, 1714, 86, 1482, 1618, 717, 1447, 1710, 154, 1091, 760, 322, 1182, 36, 215, 179, 109, 1448]. reproductively [206]. requirements [1035, 1345, 1561, 1502, 497, 708]. requiring [173]. rerio [652, 857, 815, 554, 378, 1400, 214]. research [883]. reserve [1183]. reserves [1314, 1663, 1669]. Reservoir [610, 172, 1336, 1413, 808]. reservoirs [1457]. resident [1197, 954, 818, 1584]. resistance [828, 1117, 4, 1283, 782]. resistivity [354]. resorption [1719]. Resource [691, 1757, 1052, 1181]. resources [336]. respect [896, 1623]. Respiration [488, 1179, 848]. Respiratory [1689, 1589, 270, 196, 1209, 1150, 1399]. respiring [1283]. Response [3, 606, 1112, 1269, 1364, 288, 1186, 1257, 1266, 918, 1131, 1256, 1682, 1609, 1359, 17, 1241, 1302, 37, 1250, 747, 1271, 312, 1659, 1147, 216, 245, 1262, 1107, 1248, 773, 1242, 593, 1716, 989, 669, 381, 1167, 970, 1254, 1237, 1253, 774]. responses [1230, 286, 1640, 1322, 244, 1296, 1689, 1136, 341, 1459, 852, 438, 1550, 994, 1654, 1251, 231, 531, 1639, 422]. responsiveness [1185]. resting [1174, 397]. restricted [163]. restriction [1285]. result [1048]. resulting [463, 942]. Results [1373, 917, 852, 1509]. retention [1381]. retial [198]. reticulata [618, 843, 7, 631]. reticulatum [838]. reticulatus [280]. reticulum [164]. retina [613]. Retinal [341]. retinomotor [341]. Retropinna [818]. return [763, 168, 1524]. returns [348]. reversal [1327, 1360, 736]. Review [1473, 228, 1291, 1346, 1386, 1652, 1671, 294, 293, 1018, 117, 1040, 1095, 185, 1580, 1508, 335, 111, 4, 181, 203, 569, 1228, 1052, 140]. Reviews [29, 56, 71, 93, 103, 129, 142, 155, 166, 192, 240, 251, 272, 342, 355, 412, 425, 453, 466, 489, 499, 523, 540, 585, 609, 634, 661, 676, 712, 752, 768, 795, 825, 867, 969, 978, 1037, 1121, 1137, 1177, 1205, 1226, 1310, 1402, 1435, 1454, 1493, 1592, 1610, 1631, 1699]. Rhadinorhynchidae [1567]. Rhamdia [1307]. Rheidol [1103]. Rhiniodon [491]. Rhinobatos [671]. Rhombosolea [1007]. rhombus [1005]. rhythms [617, 1485]. RIA [420, 1515]. Ribble [1510]. Rich [195, 339, 487, 817, 335]. Rich. [726, 357]. Richardson [147, 1537, 1387, 16, 97, 1219, 288, 242, 1185, 1186, 759, 1075, 1389, 530, 841, 597, 971, 1102, 244, 637, 197, 842, 548, 1415, 1184, 778, 199, 265, 714, 508, 131, 120, 274, 289, 1578, 1117, 575, 690, 1476, 417, 1697, 1256, 391, 290, 501, 112, 947, 421, 1543, 1232, 653, 815, 1608, 1417, 292, 1392, 79, 37, 209, 263, 858, 707, 689, 640, 684, 672, 75, 1694, 747, 460, 73, 31, 318, 319, 66, 818, 22, 1371, 213, 1695, 1601, 1015, 814, 803, 556, 719, 126, 779, 1716, 510, 837, 934, 237, 805, 1693, 222, 173, 381, 359]. Richardson [485, 657, 1167, 1168, 1690, 1104, 108, 163, 299, 330, 567, 432, 649, 1254, 1673, 397, 619, 774, 361, 628, 221, 1338, 831, 117]. riffle [1181]. riffle-dwelling [1181]. Rift [1113]. Rima [1375, 922, 1090]. ring [785]. ringens [372]. rings [1425, 550]. Rio [1179]. ripening [617]. Rise [1686]. risk [1155, 1483, 1359]. Risso [69, 1638, 110, 480, 1408, 994, 429, 1027, 1253, 422, 344]. River [1345, 1469, 637, 846, 1429, 1068, 1656, 818, 1297, 1126, 1729, 1161, 1391, 1758, 1003, 1581, 473, 922, 1090, 716, 1691, 896, 591, 1482, 328, 1692, 478, 721, 1477, 1181, 1624, 138, 72, 697, 1162, 1527, 483, 267, 490, 967, 615, 580, 635, 1175, 1087, 522, 870, 1098, 1396, 1019, 233, 1379, 1710, 1513, 1556, 1015, 1510, 1596, 1095]. river-spawning [72]. riverine [973, 1212, 1450]. Rivers [1023, 365, 358, 999, 86, 902, 1750, 1756, 892]. Rivulus [1249]. RNA [167, 1765, 444, 850]. RNA-DNA [444]. RNA/DNA [1765]. roach [987, 90, 156, 490, 1391, 1758, 311, 799, 525, 227, 1105, 11, 1540, 1596, 1112, 1583, 1544, 1095, 1055, 1096, 455, 1320, 1390]. Robertsonian [1027]. Rockall [69, 602, 611]. rockfish [603]. Rocks [159]. rodlet [475, 878, 908]. roker [503]. Role [924, 1528, 1501, 1711, 1408, 1264, 1764, 1283, 1522, 933, 1180, 1251, 1255, 998]. rostrata [399, 905, 558, 863, 208, 193, 198, 1294, 677, 1306, 1204]. rostratus [1007]. rough [107]. roughhead [767]. Routine [1708, 821]. ruber [1111]. rubescens [1211]. rubilio [1557, 1218]. rubrimaris [1567]. rubrum [829, 1609]. ruckeri [1264, 1260]. Rud [423]. rudd [793, 1055]. Rudolphi [727]. rule [1054]. Run [1751, 176]. running [1198]. Ruppell [437]. ruthenus [1720]. Rutilis [227]. Rutilus [1557, 90, 156, 553, 490, 1391, 1758, 1218, 1665, 311, 799, 525, 1105, 11, 1540, 1596, 1112, 1583, 1544, 1095, 1055, 1096, 455, 1320, 1390].

s [1255]. S. [1184, 370, 1479, 230, 248, 1313]. S.E.M. [1606]. Sâo [610]. sac [236, 1342, 1655, 1683, 1672, 683, 1331]. Sacramento [270]. saddleback [1579, 1367]. sagax [1451, 372]. Saint [159]. saithe [1411]. salar [448, 1035, 1527, 1020, 1759, 877, 1445, 1102, 310, 574, 579, 883, 728, 1308, 1705, 1502, 1580, 508, 53, 370, 550, 1666, 711, 787, 630, 1461, 358, 940, 1420, 1479, 1347, 1685, 1755, 294, 1483, 870, 1098, 1288, 1396, 1019, 999, 250, 143, 1373, 125, 996, 1481, 105, 606, 770, 1553, 1644, 1724, 532, 1692, 137, 1163, 1490, 1622, 230, 1352, 1512, 1510, 1337, 1486, 1284, 816, 1624, 1660, 150, 419, 704, 114, 266, 1485, 1313, 1388, 565, 576, 1756, 1535]. salaris [999]. salines [1013]. salinities [1295, 551, 130, 534]. Salinity [427, 618, 664, 1393, 285, 976, 1002, 855, 15, 234, 59, 1496, 1708, 1710, 1114, 1180, 720, 1460, 1674]. salinity-alkalinity [855]. Salmincola [483]. Salmo [147, 726, 448, 1035, 1537, 1245, 1527, 1230, 1387, 1020, 1539, 16, 97, 1364, 1219, 288, 1380, 242, 1759, 607, 1185, 1186, 877, 1161, 759, 1075, 5, 1389, 530, 144, 841, 1445, 597, 236, 146, 971, 1640, 1102, 310, 574, 579, 244, 151, 883, 197, 842, 728, 548, 1308, 1415, 324, 953, 1705, 1502, 1580, 1184, 1462, 778, 199, 265, 714, 1225, 541, 508, 53, 370, 885, 1736, 1726, 1636, 131, 120, 274, 195, 289, 8, 206, 1578, 1117, 575, 690, 855, 1476, 550, 417, 1697, 1278, 1256, 1666, 1704, 1616, 711, 787, 615, 630, 434, 1461, 358]. Salmo [940, 1420, 1479, 1347, 335, 391, 290, 1685, 501, 112, 982, 1755, 947, 294, 1483, 870, 1098, 1288, 1396, 1707, 17, 1229, 141, 1019, 421, 999, 1543, 1232, 653, 815, 250, 143, 1373, 125, 954, 996, 1608, 1417, 292, 1481, 1392, 105, 79, 37, 209, 357, 606, 770, 263, 858, 1668, 707, 689, 339, 640, 684, 672, 75, 1553, 1694, 747, 1613, 1644, 1724, 73, 31, 1021, 1492, 1722, 532, 1692, 137, 1163, 1490, 318, 1018, 319, 66, 216, 1622, 1559, 22, 486, 245, 271, 230, 520, 1371, 213, 1695, 65, 305, 514, 647, 861, 748]. Salmo [1008, 1076, 1352, 1601, 1512, 660, 806, 1015, 1510, 27, 814, 803, 1342, 1655, 556, 719, 1337, 1486, 1006, 1403, 1284, 1672, 117, 126, 718, 779, 1716, 510, 816, 1624, 1584, 837, 766, 1025, 934, 237, 805, 1693, 989, 1660, 150, 222, 173, 381, 419, 704, 359, 485, 657, 1167, 114, 266, 1485, 1168, 1690, 1771, 1104, 643, 1313, 1388, 108, 608, 128, 163, 299, 330, 567, 432, 649, 926, 1639, 1254, 1673, 397, 619, 774, 361, 565, 576, 887, 628, 221, 1756, 1535, 1338, 831]. salmoides [739, 924, 590, 1336, 840, 496]. Salmon [886, 448, 1035, 1345, 1561, 1527, 1020, 1300, 931, 1759, 607, 877, 25, 1445, 1195, 1102, 310, 574, 579, 174, 883, 728, 388, 1030, 1705, 1502, 917, 1580, 668, 993, 508, 53, 884, 1276, 955, 1606, 1444, 550, 223, 225, 1666, 711, 787, 1146, 630, 1461, 358, 940, 1420, 1479, 1737, 1347, 1685, 1279, 1755, 1487, 294, 1483, 1359, 870, 1098, 1288, 1396, 907, 963, 1019, 78, 999, 1247, 250, 143, 1373, 125, 954, 996, 1481, 105, 606, 770, 51, 207, 385, 479, 1656, 1668, 1553, 1644, 1369, 1724, 992, 552, 578, 532, 1692]. salmon [137, 1163, 1490, 1622, 230, 1352, 1153, 1512, 1510, 1586, 494, 1449, 1180, 1337, 1572, 1486, 1212, 1450, 1284, 465, 816, 1624, 1660, 150, 419, 704, 966, 1160, 1292, 114, 74, 266, 1485, 1313, 1750, 1388, 1141, 284, 565, 576, 1604, 1756, 1535]. salmonae [1015]. salmonicida [1245, 847, 1248, 816, 1260, 657, 1167]. salmonid [1541, 417, 401, 1355, 1550, 31, 891, 1145, 1107, 549, 140, 897, 1760, 444, 1258]. Salmonidae [1192, 1491]. salmonids [725, 370, 577, 337, 960, 1713, 888]. Salmoniformes [1203]. salmoninarum [1102, 1260]. salmonis [72, 261, 1015]. salmositica [1302]. salmostica [75]. salpa [1588]. salt [321, 1549, 1713, 691, 352]. salt-marsh [1549, 691]. saltwater [606, 465]. Salvelinm [1623, 1467, 1627]. Salvelinus [1458, 506, 1670, 483, 910, 1308, 333, 1003, 176, 1555, 24, 1376, 920, 1053, 681, 1662, 141, 515, 484, 474, 543, 497, 944, 596, 1633, 930, 933, 1582, 748, 542, 1752, 1040, 761, 991, 616, 351, 352, 276]. sampler [41]. samplers [741]. samples [879]. Sampling [894, 893, 598, 1417, 1392, 1650, 803, 934, 173]. San [1561, 1657]. sanctae [34]. sand [518, 751, 205, 210, 480, 1408, 1174, 671, 184, 505]. sand-smelt [184]. sandbar [754]. sandeels [1767]. sandsmelt [627]. sap [1545]. sapidissima [1068]. Saprolegnia [514]. sarcoplasmic [164]. Sarda [1468]. Sardina [1680]. sardine [952, 1318, 328, 1334, 372, 1680]. Sardinella [1079, 1318]. Sardino [952]. Sardinops [1451, 372]. Sargocentron [829]. sargus [1190]. Saronikos [952]. Sarotherodon [367, 584, 537, 492, 601, 731, 248, 350]. Sarpa [1588]. Sars [578]. saupe [1588]. Saurida [400]. saxatalis [850]. saxatilis [1413, 629, 936]. sbp [1722]. scad [1213]. scalare [1067]. scaldfish [124]. scale [386, 408, 1738]. scales [67, 793, 611, 975, 766, 496]. scalloped [1463]. Scanning [648, 120, 659, 1001, 977, 1593, 28, 1545, 1443, 656, 686, 666]. Scardinius [1665, 793, 1055]. Scaridae [1407]. Scarus [1407]. schall [437]. schedule [1560]. scheme [890]. schemes [1508]. Schlegel [1383, 827, 405, 1568]. Schmaltz [26]. Schmidt [1567]. Schneider [552, 578, 437, 1317]. schooling [1058, 1293, 60, 1738]. Schreber [1479]. Sciaenidae [715]. Sciaenops [533, 1246]. sclateri [1422]. sclerites [1140]. Sclerocollum [1567]. Scomber [1761, 76, 1058, 1071, 1411, 1303, 1149, 1140, 1325]. Scombridae [1468]. scombrids [356]. scombrus [1761, 76, 1058, 1071, 1411, 1303, 1149, 1325]. scope [740]. Scopeliformes [1330]. Scopelogadus [1324]. Scophthalmus [746, 1120, 1074, 763, 1056, 617, 732, 1298, 1442, 1005]. Scorff [1020]. scorpaenid [1598]. Scorpaeniformes [452, 196]. scorpion [1649]. Scotland [1491, 462, 1376, 124, 282, 1157, 1461, 364, 1768, 266, 1282, 565]. Scottish [1192, 574, 53, 370, 884, 532, 1692, 1620]. SCUBA [398]. SCUBA-diving [398]. sculpin [1439, 673, 1477]. Scyliorhinus [383, 573, 1637, 435, 115, 962, 418, 641, 685, 1304]. Scymnorhinus [298]. Scyphidia [748]. SDA [557]. Sea [452, 1523, 1273, 302, 379, 1161, 530, 579, 1086, 728, 548, 1462, 176, 508, 1736, 8, 121, 1324, 1498, 1553, 1613, 784, 123, 457, 546, 871, 387, 1309, 1492, 1089, 283, 1348, 1165, 415, 914, 1372, 1649, 409, 864, 1733, 276, 1761, 1742, 1739, 1571, 865, 738, 1567, 345, 1764, 1576, 836, 1747, 96, 1642, 1679, 18, 162, 1743]. Sea-bass [1523, 1498]. sea-bream [1086]. sea-run [176]. seabirds [1767]. seagrass [1598]. seals [1747, 1768]. sealworm [1748]. seamount [1463]. seas [1680]. season [1215, 121, 713, 1603, 1483, 1090, 1127, 1008, 1038]. Seasonal [697, 1374, 1547, 167, 561, 637, 63, 1003, 952, 2, 353, 334, 1117, 1109, 630, 46, 51, 906, 1119, 1522, 1677, 581, 6, 1091, 1642, 1595, 126, 510, 182, 639, 807, 965, 1516, 261, 1563, 728, 48, 670, 1132, 1394, 1426, 1360, 1613, 1659, 721, 469, 1500, 1486, 517, 539, 294, 812, 117]. Seasonality [1472]. seasons [636, 538]. seat [1497]. seatrout [654]. seaward [1369]. seawater [388, 947, 1201]. seawater-adapted [947]. Sebaou [721]. Sebastes [603]. second [1491, 1634, 1653, 1654]. secondary [773, 1486, 1167]. Secretin [456]. Secretin-stimulated [456]. secretions [802, 682]. secretory [1109, 435]. Sediment [1368, 611, 62]. sedimentation [16, 357, 138]. segment [1634, 1653, 1654]. Segregation [1752, 1420]. segregative [1059]. selected [374, 158, 963, 1001, 1417, 222, 381, 777, 1267]. selecting [139]. selection [1551, 1546, 1681, 1056, 1684, 1766, 1450]. Selective [73, 1053, 1408, 1750]. selectivity [1277, 968, 1612]. semen [600]. semilunar [1394, 1426]. seminal [1109]. sense [1146]. sensitive [1765, 876]. sensitivity [1397, 1601, 1180]. Sensory [552, 1405, 243, 1222]. separated [872, 749, 357]. separating [176]. Separation [1261, 161, 1480, 1496, 1763, 1366]. Sepik [846, 1429]. Sequence [792]. Sequential [1249]. sera [10, 1257, 549]. Seriola [1027]. serotonin [1040]. serranid [750]. Serranus [459]. serrasalmid [1299]. Serum [1102, 1047, 869, 1185, 1186, 561, 294, 1229, 51, 207, 1119, 1261, 313, 570, 930, 1236, 807, 965, 576]. settled [575, 690]. settlement [1407, 1562]. seven [847, 491]. several [878]. severe [977, 446]. severity [65]. Severn [1556, 1511, 233, 1504]. sewage [472, 575, 690]. Sex [1134, 757, 1207, 186, 736, 760, 126, 1040, 150, 280]. Sex-inversion [1134]. sexes [584, 680]. Sexual [1657, 264, 257, 689, 1490, 65, 1679, 172, 1551, 1462, 1638, 737, 396, 408, 369, 521, 1076, 1077, 1486, 837, 797]. sexuality [459]. sexually [1685, 207, 479]. shad [1511, 394, 403, 702, 1068, 233, 808, 1316]. shallow [1519, 1221, 1703, 1016]. sham [1242]. sham-immunization [1242]. shape [1481]. sharing [1052]. shark [175, 754, 671, 64, 99, 491]. sharks [1463, 26, 377, 1130]. sheepshead [864, 744]. sheets [1366]. shelf [336, 1703, 424]. shift [533]. shifting [1595]. shiner [1682]. shoal [1382, 1400, 1684, 1587, 705]. shoaling [1065, 1057]. shoals [1060, 1734, 951, 1059]. Shock [1026, 288, 1349, 876]. Shock-induced [1026]. shocked [955]. shore [98]. Short [213, 1554, 181, 203, 569, 977]. short- [1554]. short-term [977]. shrimp [336]. Siamese [1191]. Sibling [1582, 1615, 74]. Sierra [1745]. Siganus [1567, 1430, 1440]. sign [158, 678]. signal [422]. signifer [243]. significance [584, 501, 1259, 1489, 1771, 1507]. sihama [772]. Sillaginidae [517, 1049]. Sillago [1132, 772]. Siluridae [932]. Siluriformes [1179]. siluroid [1280]. siluroidea [538]. Silurus [1026]. silver [817, 1034, 729, 187, 1333, 1612, 295]. silverfish [1262]. silvering [1369]. silverside [1215, 1370, 1367]. silversides [1214]. Similarities [984]. simple [1232, 756, 1140]. simplex [638]. simulated [1055]. simultaneous [36]. sin [175]. Sinai [480]. since [50, 1029]. single [376, 693, 733]. site [1389]. site-specific [1389]. sites [127, 169]. siting [1541]. situ [851]. six [189, 472]. Size [1276, 1408, 1059, 1733, 1148, 776, 1374, 1215, 1120, 1698, 574, 1406, 1429, 1427, 1357, 952, 1532, 1602, 334, 1053, 1666, 679, 1737, 1152, 1378, 122, 442, 1100, 1351, 250, 1217, 1714, 1063, 1382, 1200, 1400, 705, 875, 704, 280, 616, 128, 1475, 1158, 1438]. size-at-maturity [1737]. Size-dependent [1733]. Size-segregative [1059]. Size-selective [1408, 1053]. sized [922]. sizes [1466, 968]. sizing [1094]. skate [1106]. skeletal [378, 1342, 1655, 26, 329]. Skin [501, 843, 651, 1534, 1459, 990, 408, 806, 1337, 1306, 687, 77, 682]. skin-scale [408]. skipjack [35, 949]. slinger [937]. slope [1686, 149, 1091]. small [286, 1276, 1697, 1479, 922, 1220, 716, 1622, 1526, 771, 148, 1771, 619]. small-mesh [1526]. small-sized [922]. smallmouth [1432, 699, 1715, 708]. smears [687]. smelt [261, 518, 751, 480, 1408, 522, 818, 1526, 184, 505]. smelts [1084]. Smith [921, 1463]. smolt [1020, 870, 606, 770, 1553, 510, 1660]. Smoltification [465, 388, 1462, 1279]. Smolting [1771, 1666, 1644, 114, 576]. smolting-rate [114]. smolts [448, 711, 1098, 1288, 1396, 1373, 1369, 74, 266, 1604]. smooth [1686]. Snail [452]. snake [1211, 1466, 62]. snake-head [1466]. snakehead [980, 285]. snappers [724]. Snieszko [622]. Snyder [1286]. snyderi [673]. Social [23, 484, 944, 762, 1158]. Society [1436]. sockeye [1300, 1606, 225]. sodium [236, 918, 1724, 934]. soft [1221, 62, 1342, 1655, 1672]. Sokoto [1375, 922, 1090]. Sokoto-Rima [922, 1090]. solar [607, 954, 1668]. sole [559, 374, 1514, 1174, 1081, 1154, 68, 1521, 1522, 1187, 99, 1433]. Solea [559, 1514, 1495, 1638, 1174, 1081, 1154, 68, 1521, 1522, 1433]. Soleidae [1638]. solenette [429]. soluble [994, 1654, 1255, 628]. Solway [429, 503]. somatic [739, 299, 432, 1475]. Some [701, 1273, 134, 827, 1004, 600, 1701, 1630, 921, 551, 153, 325, 596, 357, 717, 318, 1006, 1169, 309, 683, 215, 179, 571, 961, 475, 878, 1554, 559, 835, 1429, 633, 417, 988, 1152, 598, 974, 968, 332, 1754, 745, 732, 871, 1728, 464, 469, 1577, 1080, 1284, 984, 359, 720, 812, 1118, 165, 1189, 329, 402, 628, 1748]. sophor [44]. Sound [1354, 1691, 1187]. sounds [1695]. source [838, 494]. South [1527, 844, 1036, 1514, 615, 405, 96, 671, 1005, 1383, 559, 1686, 1391, 1758, 1495, 1581, 564, 450, 1519, 1496, 1763, 317, 1512, 1570, 1453, 1140, 1282, 845, 1010, 63, 821, 819, 775, 836, 868, 9, 808]. south-coast [845]. south-east [559, 1496, 1763]. south-eastern [1383, 1495, 1453, 1140, 808]. south-west [1686, 1391, 1758, 1581, 564, 450, 1519, 1512, 1570, 1282]. South-western [1514]. Southampton [1609]. Southern [529, 931, 710, 81, 727, 1571, 170, 1701, 810, 580, 1422, 1087, 86, 1335, 1287, 1379, 1738, 424, 1073, 1573, 1320, 1516, 95]. Soviet [703]. sp [358, 30, 748, 1023]. spacing [1300]. spaciousness [64]. Spain [452, 1629, 1581, 1422, 1596, 1535]. spangled [1209, 1405]. Spanish [1630, 1756]. sparid [964]. Sparidae [701, 937, 1577, 1562]. Sparus [1086, 1488, 792, 1361, 349]. spathaceum [805, 989, 1254, 1673]. Spatial [1365, 1049, 1420, 1496, 1738, 643]. spawned [709, 736]. spawners [1268]. Spawning [1667, 941, 389, 555, 1770, 398, 1515, 1061, 1539, 1215, 310, 1451, 1528, 1308, 1184, 1638, 993, 564, 1394, 1426, 943, 940, 1420, 1318, 791, 596, 1548, 135, 1656, 780, 9, 839, 992, 229, 1008, 1153, 932, 1126, 1752, 1040, 677, 797, 215, 72, 455, 58]. Speakers [880]. special [1465, 390, 1348]. specializations [1222]. specialized [300]. speciation [1064]. specie [238]. Species [365, 1060, 356, 976, 302, 380, 1667, 1111, 1407, 392, 189, 939, 559, 1322, 158, 1751, 700, 1199, 472, 1686, 967, 1144, 938, 1277, 224, 651, 834, 1566, 772, 800, 1355, 1046, 1045, 1152, 1709, 922, 1434, 1471, 504, 211, 175, 1731, 152, 908, 836, 1221, 1222, 1747, 1496, 1763, 123, 481, 457, 546, 1194, 1216, 1262, 1078, 18, 1080, 415, 1201, 1181, 231, 1290, 691, 362, 1688]. specific [1389, 426, 557, 954, 996, 269, 61]. specificity [1111]. specimen [52, 1491, 1307]. spectabile [695]. spectrometry [1244]. speed [1323, 1587, 505, 1645]. speeds [1071, 1411, 1068, 1325]. sperm [1083, 143]. spermatids [1477]. spermatogenesis [1473, 1031, 1477, 542, 734]. spermatozoa [242, 443, 140, 718]. spermiation [734]. Sphaerostoma [338]. Sphyrna [1463]. spilurus [1265, 1664]. spinal [1537]. spinax [26]. spine [249, 1648]. spined [366, 323, 1473, 1551, 1663, 1669, 1136, 1155, 1647, 1471, 1552, 300, 423, 1649, 409]. spinosa [99]. spinous [1286]. spleen [347, 418, 1583]. splenectomized [347]. splenectomy [147]. splenic [1239]. Spontaneous [493, 283]. spot [1266, 654, 1708, 1348]. spots [99]. spotted [654, 573]. spp [701, 430, 1389, 1698, 417, 1472, 1285, 314, 1718, 859]. sprat [1144, 534]. Spratelloides [827, 1144]. Sprattus [534]. spraying [188]. spring [1770, 572, 207, 1644, 1696, 565]. spring-spawning [1770]. sprint [1476]. Squaliformes [1128]. squamation [620]. squirrelfish [829]. Sri [862, 980, 731, 218, 1457]. ß-glucosidase [927]. St [1751, 1495, 1622]. stability [1202]. stable [1276, 1693]. stage [1117, 681, 1685, 1331]. stages [1312, 160, 575, 260, 627, 469]. staining [1010, 872, 1385, 1560]. stamina [1624]. Standardization [1266]. Stanislas [622]. start [1030]. starvation [232, 147, 915, 873, 421, 102, 79, 209, 672, 438, 1715, 507, 704, 163, 850]. starved [699, 37, 1112]. State [610, 1389, 440, 1139, 713, 141, 578]. States [808]. station [1483, 1526, 505, 518, 184]. stations [1507]. status [588, 995, 1710, 891, 92, 1286, 301, 720, 812, 1118]. steelhead [1364, 759, 530, 73, 510]. steeri [1348]. Steindachner [915, 492, 1318, 1000, 1328, 1596, 248]. stenohaline [551]. stenotherm [213]. step [376]. stephani [928, 1250, 191, 639]. Stephanidis [1274]. sterlet [1720]. steroid [339, 856, 1182]. steroids [126, 150]. sterols [1362]. stickleback [366, 323, 1473, 1551, 1663, 1669, 1277, 1647, 675, 409, 1143, 461]. sticklebacks [976, 920, 879, 1136, 1155, 1471, 1552, 1052, 1549, 691, 1038, 1103, 109]. still [22]. stimulants [68, 471]. stimulated [456]. stimulation [750]. sting [1453]. sting-ray [1453]. Stizostedion [985, 66, 900]. stock [1753, 1626, 574, 1429, 1705, 1580, 810, 1351, 255, 1769, 886, 608]. stocked [1288, 954, 996, 1622, 486, 1135, 1756]. stocking [185, 527, 125, 1392]. stocking-density [1392]. stocks [901, 1444, 892, 1764, 364, 216, 930, 1748]. Stomach [111, 693, 41, 879, 611, 1414, 404, 1125]. stomach-content [879]. stomachless [817, 20]. stomachs [1597, 602]. Stomiatoidei [302]. Stomiidae [1290]. Stomiiformes [1290]. stone [373]. stoneloach [875]. stoneroller [1432]. stones [611]. storage [147, 167, 140]. Storer [67]. stores [646, 2]. strain [816]. strains [1762, 847, 1608, 1242]. strands [1217]. strategies [1359, 1596]. strategy [1483, 894, 980, 86, 899]. stratification [1085]. stream [574, 1197, 53, 370, 885, 1276, 401, 1432, 1355, 1422, 1347, 1220, 954, 996, 9, 1492, 1545, 1771, 1639]. stream-resident [1197]. streams [324, 884, 577, 940, 1479, 125, 337, 1089, 1189, 688, 444, 887]. strength [278, 898]. strenuous [1338]. Stress [696, 1245, 288, 134, 1640, 918, 1131, 1296, 713, 1000, 1302, 655, 1719, 404, 305, 647, 381, 720, 87, 706, 774]. Stress-induced [696]. stress-response [1302]. stressed [957]. striatus [980, 1368]. strikes [1011]. striped [246, 1413, 291, 629, 936, 1118, 1674, 850]. Stromateidae [738]. Stromateoidei [738]. Strombidae [919]. Strombus [919]. Strontium [284, 1706]. strontium-90 [1706]. Structural [685, 1333, 247, 764, 1702]. Structure [44, 382, 1518, 746, 1541, 953, 274, 1534, 810, 800, 1520, 1754, 243, 51, 207, 908, 992, 1710, 258, 271, 1499, 806, 1343, 36, 1305, 1038, 1675]. structures [772, 300]. studied [1487, 1599]. Studies [874, 1415, 843, 974, 83, 471, 31, 26, 84, 436, 725, 923, 449, 433, 1501, 195, 1044, 834, 1461, 743, 1046, 1552, 442, 569, 1329, 857, 745, 238, 1298, 1545, 1371, 641, 685, 1244, 1769, 1193, 519, 1412, 1141, 116]. study [621, 1557, 1626, 1445, 1172, 384, 1086, 638, 14, 1505, 1638, 1164, 1349, 131, 120, 297, 1534, 1593, 1423, 1113, 1732, 659, 332, 175, 133, 28, 1447, 435, 478, 962, 1467, 1149, 305, 1594, 686, 27, 418, 1624, 350, 1156, 797, 119, 361, 461]. Stuhlmann [844]. stunted [1458, 11]. stunting [944]. sturgeon [1720, 802, 1409, 1097]. sturgeons [1423]. styles [1398, 1709]. sub [1735, 1003]. sub-Arctic [1003]. sub-cohorts [1735]. subject [1692]. subjected [963, 209]. sublethal [1185, 1186, 1446, 694, 1239, 629, 200, 1112, 222, 1304]. Submarine [1538]. suborder [538]. subsequent [147, 90, 227, 507, 381]. substance [1353, 313, 856, 1400, 303]. Substances [4, 856]. substrate [668, 12]. substrates [1570]. substratum [229]. Subtleties [1730]. subviridis [49]. success [690, 135, 1714, 1482, 1618, 998, 1448]. succession [186]. successive [1720]. succinic [606]. such [143]. sucker [869]. suction [39]. Sudan [776, 700, 1087]. Sudd [1087]. Sueur [399, 558, 208]. Suez [829]. sugar [1621, 221]. suitability [935]. suitable [1546]. Sulaibikhat [1456]. summer [1132, 235]. summing [1529]. sunfish [909, 1416, 395, 662, 1446, 876, 1675]. sunlight [1160]. Superficial [217]. Superior [479]. superoxide [1635]. supersaturated [304, 580, 747]. supersaturation [530]. supplementary [1442]. supply [582]. support [1057]. supporting [1764]. suppresses [1107]. suppression [987, 814, 1237]. supra [1162]. supra-population [1162]. suratensis [568]. surf [916]. Surface [1616, 682, 843, 847, 1637, 1294, 990, 54, 975]. surface-dependent [1637]. surfactants [54]. surgeonfish [1368]. Surna [1098]. Survey [895, 13, 127, 169]. surveys [1503]. Survival [53, 118, 379, 236, 146, 971, 472, 1736, 575, 1732, 1737, 1142, 588, 954, 130, 730, 428, 229, 346, 1352, 1655, 761, 722, 1135, 1448, 887]. Susceptibility [699, 1245, 514, 861, 748, 782, 280]. sustainable [336]. sustained [1364, 318]. sutchi [1399]. sutor [1430]. Swain [186]. swamp [1016]. swamps [1087]. Swedish [9]. swelling [107]. swimbladder [193, 296, 198, 1399]. swimbladder-inner [296]. Swimming [42, 534, 1364, 1323, 1446, 76, 1071, 1411, 1048, 1068, 175, 1293, 707, 27, 1587, 190, 740, 505, 1325, 1645, 1103]. Swincombe [697, 1162]. Switches [731]. Sydenham [830]. sympatric [976, 1667, 392, 1277, 206, 920, 1420, 930, 1549, 1038]. Symphodus [1563]. symposium [1051, 1772]. Synaphobranchidae [871, 415]. synchronous [493]. syndrome [1067, 796]. Synodontis [437]. Synopses [356]. synthetic [237]. System [1023, 979, 1706, 483, 846, 1003, 1662, 296, 473, 37, 1692, 994, 818, 92, 719, 190, 1509, 173, 797, 138, 1282]. systematic [92]. Systemic [1654]. systems [487, 896].

tactics [1595]. taenia [300, 423]. tag [348, 306]. tagged [1602, 1188]. tagging [1043, 1028, 707, 1743]. tags [1666, 1733]. Tagus [1528]. tail [190]. tail-beat [190]. tailed [1429, 833]. tailings [1416]. taken [262]. Takvatn [1458]. Tamar [1514]. Tambaqui [440]. tamoxifen [455]. Tanganyika [1208]. tanks [495]. tapeworm [1643, 1649, 1390, 1658]. tapeworms [1199]. tapirina [1007]. tapwater [472]. target [1537]. tarpon [765, 765]. Tasmania [1184, 504]. Taste [1384, 289]. Taurulus [1649]. tauvina [1092]. Taw [892]. taxocenes [688]. taxonomy [923]. teaching [756]. technique [725, 1385, 1650, 591]. techniques [1349, 789, 1718]. tectum [613]. Tegid [145]. telemetry [719]. teleost [1497, 1223, 459, 1134, 1322, 1004, 1315, 1566, 20, 1393, 281, 464, 1201, 1358, 80, 177, 239, 1150, 1166]. teleostean [645, 152, 1216]. Teleostei [1223, 445, 1224, 1495, 1398, 1593, 450, 1709, 1319, 1354, 1691, 908, 1496, 1194, 1470, 406, 1072, 1281, 1395, 1688]. Teleosteir [1099, 937]. Teleosti [700]. teleosts [475, 878, 633, 804, 269, 28, 1410, 1475, 329]. Telostei [1558]. Temminck [1383, 405, 1568]. temperate [903, 1728]. Temperature [1219, 828, 259, 82, 831, 1035, 1561, 323, 1215, 1445, 1322, 1427, 1464, 1357, 1002, 1532, 376, 821, 694, 1048, 925, 290, 311, 1487, 122, 153, 474, 572, 730, 1496, 1301, 201, 1659, 1708, 213, 1213, 1678, 283, 248, 279, 1715, 766, 720, 1674, 351, 1097, 139, 34, 1018]. Temperature-dependence [1219]. temperatures [652, 590, 698, 259, 762]. Temporal [447, 1420, 643, 1049]. ten [1467, 1078]. tench [3]. tendency [1551, 759]. tension [54]. tensions [995]. tentacularis [1207]. tenuis [936]. tenuissima [697, 1162]. Teraponidae [1209, 1405]. term [1761, 1706, 1554, 1742, 1745, 1028, 238, 1202, 1159, 213, 977, 779]. Terns [1767]. terrae [34]. terrae-sanctae [34]. territorial [1115, 1038, 1688]. Territory [1023, 1032]. tertiary [44]. test [1053, 1100, 571]. testes [46, 137, 779]. testicular [1109, 1477, 542, 1702, 1688]. testing [1536]. Testis [1688, 105]. Testis-testicular [1688]. testosterone [294, 1319, 117, 734]. tests [90]. testse [188]. testudineus [1017, 258]. Tetracycline [1043]. tetracyclines [1237]. Tetrahymena [625]. tetraploids [1233]. Tetrosomus [1363]. Texas [1370]. Texoma [1413]. Thailand [201]. Thalassoma [1579, 1319]. Thames [522, 1035, 813, 1710, 1513]. their [1111, 487, 828, 1274, 90, 1295, 145, 1528, 1030, 176, 885, 1727, 892, 294, 111, 1090, 1656, 11, 229, 898, 930, 933, 1285, 902, 469, 1251, 1624, 377, 1475, 329, 1096]. theoretical [1323, 359]. theory [896]. Theragra [986, 1747, 839, 1428]. there [1300]. Thermal [1336, 777, 956, 535, 1296, 695, 1689, 804, 1000, 948, 1085, 231, 1460]. thermally [1619]. Thermoregulatory [283]. thick [110, 994, 1253, 344]. thick-lipped [110, 994, 1253]. thicklipped [422]. thickness [1486]. Thingvallavatn [1555, 1752]. thiourea [759]. Thompson [667, 937, 539, 589, 599]. Threadfin [808]. three [725, 366, 323, 1473, 1551, 1667, 1111, 1663, 1669, 972, 985, 7, 568, 651, 1136, 1155, 278, 1479, 925, 1101, 1647, 1471, 1552, 211, 175, 152, 1496, 22, 437, 1091, 1642, 1549, 1015, 762, 231, 1660, 657, 691, 1038]. three-spined [366, 323, 1473, 1551, 1663, 1669, 1136, 1155, 1647, 1471, 1552]. threespine [1143, 1103]. threshold [1058]. thresholds [226]. thrombocytes [449]. throughout [145]. Thryssa [1518]. thumb [1054]. Thumbi [1421]. Thunnus [35, 1073, 1645]. Thymallus [1603, 135]. thymectomized [1167]. thymocytes [1484]. thymus [685, 359]. thynnus [1645]. Thyroid [207, 183, 1640, 561, 1578, 1392, 51, 858, 1656, 31, 327, 362]. thyroxine [759, 479, 1644, 570, 1683, 576, 1604]. tidal [1508, 210, 1710, 1510, 691]. tide [1579]. Tigak [961]. tilapia [1083, 1191, 736, 665, 1271, 1115, 1265, 1664, 1621, 1009, 838, 94, 1000, 437, 1072, 314, 762, 1126]. tilapiae [727, 1434]. tilapias [1349]. tilefish [819]. Tilting [1071]. Time [1451, 1644, 1372, 74, 1215, 1002, 982, 912, 1613, 305, 248, 705]. Time-course [1644, 1372, 912, 305]. times [482]. timing [1751]. Tinca [3]. tissue [84, 1533, 374, 1231, 528, 1616, 604, 1643, 293, 1694, 1650, 1623, 647, 1006, 237, 657, 327, 362]. tissues [621, 183, 912, 420, 992, 277, 811, 777, 720, 812, 1118, 432]. titres [1102]. Tolerance [530, 841, 427, 618, 664, 265, 695, 304, 259, 15, 948, 55, 1064]. tolerant [938]. tongue [99]. tool [889, 1503]. toothless [1489]. top [658]. topminnow [67]. Torpedo [1106]. tortuosus [1469]. total [214]. Townsville [772]. toxic [970]. toxicant [963]. Toxicity [1682, 1588, 448, 199, 395, 1636, 1711, 54, 1672, 864, 361]. toxins [1283]. trace [1655, 177]. Trachinotus [919]. Trachinus [1571, 83]. Trachurus [317, 1213]. Tracking [266, 1509, 643]. tract [1086, 14, 587, 414, 426, 1424, 626, 755, 1119, 1474, 665, 1768]. tracts [1597, 19]. training [90, 1476, 318, 774]. transaminases [326]. transfer [1613, 1147, 108]. transferring [409]. transformation [1553, 510, 1660]. Transmission [949, 1162]. transmitter [1478]. transmitters [1694, 1069]. transmontanus [1097]. transport [288, 388, 957, 277]. Transvaal [844]. transversale [423]. transverse [1175]. trap [1198, 1744]. Trapping [1264]. trawl [202, 255, 1503]. treated [601]. treatment [780, 1486, 631]. trees [1591]. Trematoda [1042]. trematode [338]. trematodes [320]. trends [891]. Trent [267, 967]. Trewavas [696, 367, 1009]. Triaenophorus [974]. tributaries [1596]. tributary [1539]. tributyltin [1594]. trichlorfon [320]. Trichogaster [371, 447]. Trichomycteridae [1179]. tricuspidata [1170, 1377, 262]. tridentata [460]. triphosphate [209]. triphosphate/haemoglobin [209]. triphosphate/red [209]. triplets [645]. triploid [487, 477, 607, 1480, 678, 264, 257, 689, 1233, 307]. triploidy [1349, 136, 1026]. Tripterygion [1688]. Trisoptems [462]. Trisopterus [921]. Trophic [254, 1413, 784, 745, 1456]. tropical [1012, 696, 302, 1667, 1322, 32, 724, 371, 447, 1424, 457, 154, 1728, 984, 1516, 1206]. Trough [602, 611]. trout [147, 726, 1537, 1245, 1230, 1387, 1539, 16, 97, 1364, 1219, 288, 1380, 242, 1185, 1186, 1161, 759, 483, 1075, 5, 1389, 530, 144, 841, 1054, 1681, 597, 236, 146, 1640, 1102, 910, 244, 151, 197, 1235, 842, 548, 1415, 324, 913, 185, 527, 953, 1003, 1197, 1762, 1184, 1462, 778, 199, 265, 714, 1225, 541, 508, 885, 1736, 1726, 1636, 131, 120, 274, 195, 289, 8, 121, 206, 1578, 1117, 575, 690, 855, 1476, 417, 1697, 1278, 1256, 1704, 1616, 1711, 615, 434, 940, 1420, 1479, 1347, 335, 391, 290, 501, 112, 982, 1662, 947, 1013, 870, 1707]. trout [17, 1229, 141, 421, 1543, 1232, 653, 815, 125, 954, 996, 1608, 1417, 292, 1481, 1392, 1302, 79, 37, 209, 357, 263, 385, 858, 707, 689, 339, 640, 684, 672, 75, 1694, 747, 1613, 995, 899, 73, 1659, 1021, 1492, 1692, 318, 1018, 319, 66, 216, 1559, 22, 486, 271, 230, 520, 1371, 1695, 65, 305, 514, 647, 861, 748, 1008, 1076, 1601, 660, 806, 1015, 27, 814, 803, 1342, 1655, 556, 719, 542, 1006, 1169, 1403, 1672, 117, 126, 43, 718, 1716, 510, 495, 180, 1584, 837, 766, 1025, 1040, 934, 237, 805, 1693, 989]. trout [222, 173, 381, 359, 485, 657, 1167, 1168, 1690, 1771, 1104, 643, 1313, 108, 608, 1270, 128, 163, 299, 330, 567, 432, 649, 926, 1639, 276, 1254, 1673, 397, 619, 774, 361, 887, 628, 221, 1535, 1338, 831]. trunkfish [1363]. trutta [1722, 887, 1539, 1380, 1161, 5, 144, 236, 146, 1640, 151, 324, 953, 1184, 1462, 1225, 541, 370, 885, 1736, 1726, 1636, 8, 206, 1704, 615, 940, 1420, 1479, 1347, 870, 1707, 17, 1229, 141, 125, 954, 996, 1481, 1492, 552, 578, 1692, 1559, 486, 245, 271, 230, 520, 65, 305, 514, 647, 861, 748, 1008, 1076, 1601, 660, 806, 1015, 27, 1342, 1655, 1006, 1403, 1672, 1584, 766, 1025, 1771, 643, 1313, 608, 1639, 1535]. Trypanorhyncha [654, 1159]. trypanospmes [641]. tryptophan [1040, 934]. tschawytscha [966, 1160, 1292]. tshawytscha [1345, 1561, 644, 174, 907, 1212, 1450]. tubifex [399]. Tubule [1401]. tumbil [400]. tumour [1252]. tumours [835, 1251]. tuna [1687, 949, 1073, 1645]. turbid [1731]. turbines [1508]. turbot [746, 1120, 1074, 763, 1056, 617, 732, 1298, 1442]. turcomanus [92]. Turkana [1404]. turnover [672]. twaite [1511, 233, 1316]. twelve [731]. twilight [1057]. twins [1191]. twitch [1645]. Two [1488, 1414, 1290, 878, 392, 817, 741, 1389, 603, 1322, 246, 652, 1638, 1376, 1113, 1709, 792, 125, 238, 1608, 326, 1731, 836, 747, 1159, 86, 28, 31, 1722, 1216, 1316, 930, 902, 1596, 1242, 984, 1181, 811, 295, 1544, 1620, 998, 1688, 1756, 848]. two-year-old [747]. Tylodelphys [238, 1202]. type [1667, 1120, 917, 1324]. types [878, 321, 418, 887, 461, 379]. typical [1607]. typus [491]. tyrannus [1646, 296]. tyrosine [823].

U.K [1015]. U.K. [1005]. U.S.A. [1389, 819, 304, 1068, 1413]. Ukoli [727]. ultimate [1475]. ultrasonic [1666, 707, 719, 190, 1188, 643]. ultrasonic-tagged [1188]. Ultrastructural [1533, 1409, 1553, 1423, 27, 116, 461]. Ultrastructure [197, 1600, 43, 303, 368, 289, 1128, 1129, 1130, 82, 385, 152, 269]. Umbra [938]. Unchanged [1362]. unchanging [1660]. underdeveloped [863]. undergoing [901]. understanding [904]. underwater [1492]. underyearling [158, 837]. undiluted [472]. undulatus [368]. unexploited [761]. unicolor [1209, 1405]. unique [1223]. unit [357]. United [808]. unknown [1262]. Untersee [1188]. unusual [871]. upland [366, 125, 337, 954, 996]. upon [1245, 1726, 672, 683]. upper [741]. Upstream [1019, 1345, 1375, 78, 956, 1460, 888]. Uptake [508, 912, 281, 183, 392, 1415, 21, 1726, 651, 1724, 1342, 1655, 496, 485, 1258, 1634]. upwelling [1079]. urban [1622]. urea [1201]. urophthalmus [1455]. Urophycis [1149]. ursinus [1747]. Use [1329, 904, 751, 594, 658, 176, 249, 41, 879, 1432, 78, 1373, 556, 1260, 687]. used [10, 161, 569, 104]. using [739, 878, 470, 267, 1349, 1593, 1487, 1232, 175, 1746, 793, 1492, 1069, 1140, 1244, 1743, 519, 1073, 1718]. Usk [1527]. utility [1757]. Utilization [1170, 1377, 584, 597, 646, 1757, 102, 1710, 275, 991, 691, 1097].

vaccinated [108]. Vaccination [816, 915, 1371]. vaccine [1256, 485]. Val [487, 817, 1034, 1079, 564, 316, 59, 96, 1683, 132, 148, 275, 1612, 55]. Val. [130]. Vala [15]. Valencia [584]. Valenciennes [828, 1480, 49, 187, 184, 765]. Validity [180]. Valley [1113]. value [1025]. values [665, 513, 381]. Var [1156]. Varano [38]. Vardnes [1161]. variability [506, 1720, 1723, 333, 972, 1762, 1732, 1744, 608, 849]. variables [739, 1417, 840, 720, 812, 1118]. Variation [1426, 1369, 1518, 1663, 1391, 1705, 952, 1580, 800, 630, 1471, 442, 1754, 1613, 948, 960, 592, 1760, 1750, 1320]. Variations [548, 1563, 1742, 167, 637, 353, 334, 46, 570, 1510, 1486, 117, 126, 510]. varied [319]. variegatus [744]. various [288, 146, 927, 575, 943, 279]. varying [1666, 907, 482]. Vascular [583, 5]. Vascularization [165]. vasculature [1617]. Vefsna [1019]. vegetation [495]. velocities [1355, 230]. velocity [357, 520]. Venda [727]. vendace [1351]. Venezuela [584]. ventilation [270]. vermiculatus [946]. Versatility [658]. versus [498, 695, 1682]. Vertebral [650, 1439]. Vertical [372, 1204, 1175]. vesicles [1109]. vetulus [374, 1187]. Viability [482, 690]. Vibrio [485, 108]. vibriosis [1348]. Vibro [225]. Victoria [174, 1029, 1339, 314, 538, 202, 365, 255, 348, 1280]. Vilaine [1521]. villosus [297, 476, 297, 476]. Vimba [925]. vipera [1571, 83]. virens [835, 1411]. virescens [1010]. Virginia [1729]. virus [1273, 1272, 1271]. Visakhapatnam [1078]. vision [521]. vison [1479]. Visual [226, 1056, 667, 205, 1552, 369, 677]. visually [875]. vitamin [391, 531]. Vitellogenesis [177]. vitellogenin [1578]. vitreum [66, 900, 985, 66, 900]. vitro [1230, 1322, 560, 1257, 1711, 479, 350, 1104, 1237, 831]. vivanus [186]. viviparous [1534]. viviparus [912, 326]. vivo [1230, 456]. vlaminghi [702]. Voil [266]. volume [1597]. volumes [1466]. Volvi [1341, 1665]. vorax [309]. vulgaris [1081]. vulnerability [230].

W [1082]. Waikato [818]. Walb [952]. Walbaum [392, 25, 644, 1195, 709, 174, 388, 668, 993, 124, 223, 225, 1737, 907, 78, 51, 207, 385, 479, 1677, 992, 31, 62, 656, 1149, 494, 1180, 1572, 1450, 629, 936, 465, 191, 639, 966, 1160, 1292, 284]. Wales [405, 1005, 1527, 853, 874, 145, 615, 1755, 473, 1519, 1189, 1103]. Wallace [846]. Walleye [985, 986, 1747, 839, 66, 900, 1428]. warm [1373, 1173]. warm-water [1173]. Washington [1389]. waste [842, 1373]. Water [950, 696, 854, 1527, 979, 530, 463, 579, 321, 151, 728, 895, 548, 533, 1277, 3, 508, 370, 334, 1726, 1636, 577, 304, 604, 580, 1355, 893, 290, 473, 1017, 870, 1288, 1459, 1373, 1173, 730, 1275, 212, 1553, 747, 1613, 133, 995, 86, 1724, 1294, 229, 387, 1198, 1692, 1430, 22, 230, 520, 1107, 494, 1342, 1655, 1672, 1715, 409, 505, 783, 352, 567, 1016, 1604, 1150, 1166, 628, 1609]. water- [1017]. water-flow [1604]. water-soluble [628]. waterborne [265, 1283, 1693, 869]. watercourse [729]. watercourses [1584]. waterfall [216]. waters [1092, 1518, 725, 413, 202, 394, 403, 559, 835, 185, 527, 1144, 798, 724, 921, 767, 787, 1152, 1378, 118, 1221, 573, 839, 298, 317, 149, 405, 1470, 1280, 6, 1091, 1367, 1281, 1165, 1649, 897, 295]. weever [1571, 83]. Weight [692, 367, 1120, 45, 94, 2, 1578, 679, 290, 1152, 1378, 153, 474, 1708, 139]. weirs [888]. well [1731]. well-mixed [1731]. Welsh [1639]. West [1745, 973, 1729, 1421, 1700, 1686, 462, 1391, 1758, 1133, 1581, 564, 450, 124, 282, 1157, 1519, 948, 1512, 1570, 1282, 771]. western [1274, 1514, 1638, 1538, 868, 1741, 1565, 848, 63, 96]. whale [491]. wheatlandi [976]. white [650, 1657, 1266, 175, 672, 1190, 1540, 861, 1088, 1097, 649, 869]. whitefish [646, 268, 443, 722, 72]. Whiting [1569, 1418, 69, 462, 1132, 772, 1642, 424, 517, 1049]. Whitley [772]. Widespread [1313, 948]. wild [971, 1765, 1030, 1705, 1462, 1567, 711, 1605, 1396, 736, 1369, 486, 1725, 1716, 1040, 1568, 576, 1756]. wild-spawned [736]. Willamette [1345]. willughbii [24]. Windermere [1626, 324, 498, 24, 614, 596]. windowpane [1443]. winter [709, 1139, 1723, 928, 1197, 1250, 1677, 656, 1715, 191, 639]. Wisconsin [869]. within [1706, 895, 1471, 1684, 1584, 1750]. within-species [1471]. without [395, 1146, 861]. Wongratana [1144]. work [756]. works [575, 690]. worms [602]. wound [1337, 936]. wrasse [1579, 1319].

X [1315, 743, 742]. X-cells [1315]. X-ray [743, 742]. X. [502]. xanthums [1708]. xanthurus [654]. Xiphophorus [502, 23].

Y-autosome [1207]. Yamaguti [686]. yaquinae [1069]. year [1670, 1036, 145, 1014, 278, 747, 898, 1018, 1715, 766, 1320]. year-class [278, 898]. yearling [431, 385, 1613, 1006]. years [1375, 125]. yellow [1581, 775, 1574, 1165, 1333, 783]. yellowfin [306, 389, 439, 469, 760]. Yersinia [1264, 1260]. Yield [440, 1374, 846, 973]. yields [336, 809, 836, 1085]. yolk [236, 680, 982, 1342, 1655, 1683, 1672, 683, 1331, 1097]. yolk-sac [1342, 1655, 1683, 1672, 683, 1331]. York [191, 639]. Young [1729, 1670, 244, 664, 1036, 1014, 1736, 153, 1294, 230, 773, 1624, 1715]. young-of-the-year [1670, 1036, 1014, 1715]. Ysabel [827]. Yugoslavian [725].

Z [26]. Z-discs [26]. Zambia [50, 1085]. Zambia/Zimbabwe [1085]. Zealand [1686, 512, 1276, 1045, 1222, 818, 1091, 1140, 1181, 308, 926]. zebra [652, 1400, 378]. zebrafish [857, 815, 554, 214]. Zeder [697, 1162]. zilli [94, 762]. zillii [838, 437]. Zimbabwe [1085]. zinc [841, 560, 1416, 1726, 1636, 1711, 387, 1127, 1442, 629, 1672, 683, 1693]. Zoarces [912, 326]. Zoo [1012]. zoophagy [731]. zooplankton [363, 1055]. Zwai [1113].

References

Smith:1980:FLB

[1] A. C. Smith. Formation of lethal blood clots in fishes. Journal of Fish Biology, 16(1):1–4, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delahunty:1980:SRO

[2] G. Delahunty and V. L. de Vlaming. Seasonal relationships of ovary weight, liver weight and fat stores with body weight in the goldfish, Carassius auratus (L.). Journal of Fish Biology, 16(1):5–13, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Demael:1980:RTT

[3] A. Demaël, D. Garin, and G. Peres. Response of the tench (Tinca tinca L.) to potassium nitrate enriched water. Journal of Fish Biology, 16 (1):15–22, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1980:SIN

[4] G. A. Ingram. Substances involved in the natural resistance of fish to infection — a review. Journal of Fish Biology, 16(1):23–60, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolis:1980:IBV

[5] L. Bolis and J. C. Rankin. Interactions between vascular actions of detergent and catecholamines in perfused gills of European eel, Anguilla anguilla L. and brown trout, Salmo trutta L. Journal of Fish Biology, 16(1):61–73, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ortega-Salas:1980:SCC

[6] A. A. Ortega-Salas. Seasonal changes in the common dab, Limanda limanda (L.) in Isle of Man waters. Journal of Fish Biology, 16(1):75– 82, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dahlgren:1980:ETD

[7] B. T. Dahlgren. The effects of three different dietary protein levels on the fecundity in the guppy, Poecilia reticulata (Peters). Journal of Fish Biology, 16(1):83–97, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahy:1980:ECM

[8] Edward Fahy. Eubothrium crassum in migratory trout, Salmo trutta L., in the sea. Journal of Fish Biology, 16(1):99–104, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malmqvist:1980:SMB

[9] B. Malmqvist. The spawning migration of the brook lamprey, Lampetra planeri Bloch, in a South Swedish stream. Journal of Fish Biology, 16 (1):105–114, January 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexander:1980:CFM

[10] J. B. Alexander and G. A. Ingram. A comparison of five of the methods commonly used to measure protein concentrations in fish sera. Journal of Fish Biology, 16(2):115–122, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linfield:1980:ECL

[11] R. S. J. Linfield. Ecological changes in a lake fishery and their effects on a stunted roach Rutilus rutilus population. Journal of Fish Biology, 16 (2):123–144, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tareen:1980:ABA

[12] I. U. Tareen. Abscessed bony areas on Acanthopagrus cuvieri (Day) as a substrate for epiphytic algae. Journal of Fish Biology, 16(2):145–148, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berrebi:1980:ESF

[13] P. Berrebi and J. Britton-Davidian. Enzymatic survey of four populations of Atherina boyeri based on electrophoresis and the occurrence of a microsporidiosis. Journal of Fish Biology, 16(2):149–157, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:1980:SHM

[14] A. J. Clarke and D. M. Witcomb. A study of the histology and morphology of the digestive tract of the common eel (Anguilla anguilla). Journal of Fish Biology, 16(2):159–170, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1980:EAS

[15] R. V. Kilambi and A. Zdinak. The effects of acclimation on the salinity tolerance of grass carp, Ctenopharyngodon idella (Cuv. and Vala.). Journal of Fish Biology, 16(2):171–175, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barham:1980:EBI

[16] W. T. Barham, G. L. Smit, and H. J. Schoonbee. The effect of bacterial infection on erythrocyte fragility and sedimentation rate of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 16(2):177– 180, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1980:IRB

[17] G. A. Ingram and J. B. Alexander. The immune response of the brown trout Salmo trutta to lipopolysaccharide. Journal of Fish Biology, 16 (2):181–197, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robb:1980:FFG

[18] A. P. Robb and J. R. G. Hislop. The food of five gadoid species during the pelagic O-group phase in the northern North Sea. Journal of Fish Biology, 16(2):199–217, February 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindsay:1980:CAD

[19] G. J. H. Lindsay and J. E. Harris. Carboxymethylcellulase activity in the digestive tracts of fish. Journal of Fish Biology, 16(3):219–233, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grove:1980:CBD

[20] D. J. Grove and C. Crawford. Correlation between digestion rate and feeding frequency in the stomachless teleost, Blennius pholis L. Journal of Fish Biology, 16(3):235–247, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:1980:OUD

[21] J. Davenport and S. Lönning. Oxygen uptake in developing eggs and larvae of the cod, Gadus morhua L. Journal of Fish Biology, 16(3):249– 256, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OGrady:1980:FAE

[22] K. T. O’Grady and P. C. R. Hughes. Factorial analysis of an experimental comparison of three methods of fishing for rainbow trout, Salmo gairdneri Richardson, in still water. Journal of Fish Biology, 16(3):257–264, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1980:SHR

[23] D. B. C. Scott and C. E. Currie. Social hierarchy in relation to adrenocortical activity in Xiphophorus helleri Heckel. Journal of Fish Biology, 16(3):265–277, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frost:1980:GCS

[24] W. E. Frost and C. Kipling. The growth of charr, Salvelinus willughbii Günther, in Windermere. Journal of Fish Biology, 16(3):279–289, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjerknes:1980:MCP

[25] V. Bjerknes and A. B. Vaag. Migration and capture of pink salmon, Oncorhynchus gorbuscha Walbaum in Finnmark, North Norway. Journal of Fish Biology, 16(3):291–297, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slinde:1980:SNZ

[26] E. Slinde and H. Kryvi. Studies on the nature of the Z-discs in skeletal muscle fibres of sharks, Etmopterus spinax L. and Galeus melastomus Rafinesque–Schmaltz. Journal of Fish Biology, 16(3):299–308, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proctor:1980:HUS

[27] C. Proctor, P. R. L. Mosse, and R. C. L. Hudson. A histochemical and ultrastructural study of the development of the propulsive musculature of the brown trout, Salmo trutta L., in relation to its swimming behaviour. Journal of Fish Biology, 16(3):309–329, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattey:1980:SEM

[28] D. L. Mattey, M. Morgan, and D. E. Wright. A scanning electron microscope study of the pseudobranchs of two marine teleosts. Journal of Fish Biology, 16(3):331–343, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRa

[29] Anonymous. Book reviews. Journal of Fish Biology, 16(3):345–348, March 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Narasimhamurti:1980:MMS

[30] C. Narasimhamurti, C. Kalavati, and B. Saratchandra. Myxosoma microspora n. sp. (Myxosporidia: Protozoa) parasitic in the gills of Mugil cephalus L. Journal of Fish Biology, 16(4):345–348, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milne:1980:SRB

[31] R. S. Milne and J. F. Leatherland. Studies on the relationship between osmotic or ionic regulation and thyroid gland activity in two salmonid fishes, Salmo gairdneri Richardson and Oncorhynchus kisutch Walbaum. Journal of Fish Biology, 16(4):349–360, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadzie:1980:RCF

[32] S. Dadzie. Recent changes in the fishery of a new tropical man-made lake, Lake Kamburu (Kenya). Journal of Fish Biology, 16(4):361–367, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulata:1980:GDB

[33] G. Hulata, R. Moav, and G. Wohlfarth. Genetic differences between the Chinese and the European races of the common carp III. Gonad abnormalities in hybrids. Journal of Fish Biology, 16(4):369–370, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yaron:1980:ETP

[34] Z. Yaron, M. Cocos, and H. Salzer. Effects of temperature and photoperiod on ovarian recrudescence in the cyprinid fish Mirogrex terraesanctae. Journal of Fish Biology, 16(4):371–382, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexander:1980:HCA

[35] N. Alexander, R. M. Laurs, A. McIntosh, and S. W. Russell. Haematological characteristics of albacore, Thunnus alalunga (Bonnaterre), and skipjack, Katsuwonus pelamis (Linnaeus). Journal of Fish Biology, 16 (4):383–395, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Touart:1980:ARS

[36] L. W. Touart and S. A. Bortone. The accessory reproductive structure in the simultaneous hermaphrodite Diplectrum bivittatum. Journal of Fish Biology, 16(4):397–403, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1980:RHS

[37] H. C. Lane. The response of the haemoglobin system of fed and starved rainbow trout, Salmo gairdneri Richardson, to bleeding. Journal of Fish Biology, 16(4):405–411, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossi:1980:BAE

[38] Remigio Rossi and Paolo Villani. A biological analysis of eel catches, Anguilla anguilla L., from the lagoons of Lesina and Varano, Italy. Journal of Fish Biology, 16(4):413–423, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weihs:1980:HSF

[39] D. Weihs. Hydrodynamics of suction feeding of fish in motion. Journal of Fish Biology, 16(4):425–433, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gervai:1980:OAR

[40] J. Gervai, T. Mariaan, Z. Krasnai, A. Nagy, and V. Csánuyi. Occurrence of aneuploidy in radiation gynogenesis of carp, Cyprinus carpio L. Journal of Fish Biology, 16(4):435–439, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1980:SSU

[41] N. Giles. A stomach sampler for use on live fish. Journal of Fish Biology, 16(4):441–444, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCleave:1980:SPE

[42] J. D. McCleave. Swimming performance of European eel (Anguilla anguilla (L.)) elvers. Journal of Fish Biology, 16(4):445–452, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seagrave:1980:UHL

[43] C. P. Seagrave, D. Bucke, and D. J. Alderman. Ultrastructure of a haplosporean-like organism: the possible causative agent of proliferative kidney disease in rainbow trout. Journal of Fish Biology, 16(4):453– 459, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1980:SGB

[44] G. M. Hughes and A. K. Mittal. Structure of the gills of Barbus sophor (Ham), a cyprinid with tertiary lamellae. Journal of Fish Biology, 16(4): 461–467, April 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crivelli:1980:ELW

[45] A. Crivelli. The eye lens weight and age in the common carp, Cyprinus carpio L. Journal of Fish Biology, 16(5):469–473, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmann:1980:SAV

[46] R. Hoffmann, P. Wondrak, and W. Groth. Seasonal anatomical variations in the testes of European pike, Esox lucius L. Journal of Fish Biology, 16(5):475–482, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skorping:1980:PBN

[47] A. Skorping. Population biology of the nematode Camallanus lacustris in perch, Perca fluviatilis L., from an oligotrophic lake in Norway. Journal of Fish Biology, 16(5):483–492, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dawson:1980:QSC

[48] A. S. Dawson and A. S. Grimm. Quantitative seasonal changes in the protein, lipid and energy content of the carcass, ovaries and liver of adult female plaice, Pleuronectes platessa L. Journal of Fish Biology, 16(5): 493–504, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chan:1980:RGG

[49] E. H. Chan and T. E. Chua. Reproduction in the greenback grey mullet, Liza subviridis (Valenciennes, 1836). Journal of Fish Biology, 16(5):505– 519, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dudley:1980:CEG

[50] R. G. Dudley and R. J. Scully. Changes in experimental gillnet catches from the Kafue Floodplain, Zambia, since construction of the Kafue Gorge Dam. Journal of Fish Biology, 16(5):521–537, May 1980. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1980:SCT

[51] J. F. Leatherland and R. A. Sonstegard. Seasonal changes in thyroid hyperplasia, serum thyroid hormone and lipid concentrations, and pituitary gland structure in Lake Ontario Coho salmon, Oncorhynchus kisutch Walbaum and a comparison with Coho salmon from Lakes Michigan and Erie. Journal of Fish Biology, 16(5):539–562, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alvarez-Leon:1980:SLA

[52] R. Alvarez-León. A specimen of Lutjanus argentiventris (Peters) lacking pelvic fins. Journal of Fish Biology, 16(5):563–564, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1980:SGS

[53] H. J. Egglishaw and P. E. Shackley. Survival and growth of salmon, Salmo salar (L.), planted in a Scottish stream. Journal of Fish Biology, 16(5):565–584, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muller:1980:FTS

[54] R. Müller. Fish toxicity and surface tension of non-ionic surfactants: investigations of antifoam agents. Journal of Fish Biology, 16(5):585– 589, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tooby:1980:TGC

[55] T. E. Tooby, J. Lucey, and B. Stott. The tolerance of grass carp, Ctenopharyngodon idella Val., to aquatic herbicides. Journal of Fish Biology, 16(5):591–597, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRb

[56] Anonymous. Book reviews. Journal of Fish Biology, 16(5):599–602, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:FMA

[57] Anonymous. Future meetings and announcements. Journal of Fish Biology, 16(5):602, May 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deGroot:1980:CMG

[58] S. J. de Groot. The consequences of marine gravel extraction on the spawning of herring, Clupea harengus Linné. Journal of Fish Biology, 16(6):605–611, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maceina:1980:ISO

[59] M. J. Maceina, F. G. Nordlie, and J. V. Shireman. The influence of salinity on oxygen consumption and plasma electrolytes in grass carp, Ctenopharyngodon idella Val. Journal of Fish Biology, 16(6):613–619, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radabaugh:1980:CMP

[60] D. C. Radabaugh. Changes in minnow, Pimephales promelas Rafinesque, schooling behaviour associated with infections of brain-encysted larvae of the fluke, Ornithodiplostomum ptychocheilus. Journal of Fish Biology, 16(6):621–628, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1980:EFM

[61] M. Jobling and P. Spencer Davies. Effects of feeding on metabolic rate, and the Specific Dynamic Action in plaice, Pleuronectes platessa L. Journal of Fish Biology, 16(6):629–638, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nash:1980:LOB

[62] R. D. M. Nash. Laboratory observations on the burrowing of the snake blenny, Lumpenus lampretaeformis (Walbaum), in soft sediment. Journal of Fish Biology, 16(6):639–648, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crawford:1980:SPS

[63] R. J. M. Crawford. Seasonal patterns in South Africa’s Western Cape purseseine fishery. Journal of Fish Biology, 16(6):649–664, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1980:NSC

[64] F. C. Stott. A note on the spaciousness of the cavity around the brain of the basking shark, Cetorhinus maximus (Gunner). Journal of Fish Biology, 16(6):665–667, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1980:SDI

[65] A. D. Pickering and P. Christie. Sexual differences in the incidence and severity of ectoparasitic infestation of the brown trout, Salmo trutta L. Journal of Fish Biology, 16(6):669–683, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niimi:1980:MEG

[66] A. J. Niimi and S. L. Morgan. Morphometric examination of the gills of walleye, Stizostedion vitreum vitreum (Mitchill) and rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 16(6):685–692, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frietsche:1980:HCR

[67] R. A. Frietsche and C. F. Bailey. The histology and calcification of regenerating scales in the blackspotted topminnow, Fundulus olivaceus (Storer). Journal of Fish Biology, 16(6):693–700, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackie:1980:CNF

[68] A. M. Mackie, J. W. Adron, and P. T. Grant. Chemical nature of feeding stimulants for the juvenile Dover sole, Solea solea (L.). Journal of Fish Biology, 16(6):701–708, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conway:1980:FLB

[69] D. V. P. Conway. The food of larval blue whiting, Micromesistius poutassou (Risso), in the Rockall area. Journal of Fish Biology, 16(6):709–723, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duncan:1980:BCF

[70] K. W. Duncan. On the back-calculation of fish lengths; modifications and extensions to the Fraser–Lee equation. Journal of Fish Biology, 16(6):725–730, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRc

[71] Anonymous. Book reviews. Journal of Fish Biology, 16(6):731–732, June 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valtonen:1980:MSI

[72] E. T. Valtonen. Metechinorhynchus salmonis infection and diet in the river-spawning whitefish of the Bothnian Bay. Journal of Fish Biology, 17(1):1–8, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meade:1980:SHP

[73] T. L. Meade and S. J. Perrone. Selective haematological parameters in steelhead trout, Salmo gairdneri Richardson. Journal of Fish Biology, 17(1):9–12, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1980:TDG

[74] J. E. Thorpe, R. I. G. Morgan, E. M. Ottaway, and M. S. Miles. Time of divergence of growth groups between potential 1+ and 2+ smolts among sibling Atlantic salmon. Journal of Fish Biology, 17(1):13–21, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe-Jinde:1980:ORT

[75] L. Lowe-Jinde. Observations of rainbow trout, Salmo gairdneri Richardson, infected with Cryptobia salmostica. Journal of Fish Biology, 17(1): 23–30, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillerman:1980:CSM

[76] R. G. Gillerman. Characterization of swimming muscle in the Atlantic mackerel, Scomber scombrus L. Journal of Fish Biology, 17(1):31–41, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitear:1980:ELF

[77] M. Whitear, A. K. Mittal, and E. B. Lane. Endothelial layers in fish skin. Journal of Fish Biology, 17(1):43–65, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsen:1980:UCC

[78] P. B. Johnsen and A. D. Hasler. The use of chemical cues in the upstream migration of Coho salmon, Oncorhynchus kisutch Walbaum. Journal of Fish Biology, 17(1):67–73, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1980:CPP

[79] H. C. Lane and T. P. Tharp. Changes in the population of polyribosomal containing red cells of peripheral blood of rainbow trout, Salmo gairdneri Richardson, following starvation and bleeding. Journal of Fish Biology, 17(1):75–81, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1980:NDC

[80] D. B. C. Scott and S. E. Rennie. Nuclear diameter as a criterion of adrenocortical activity in a teleost fish Coregonus lavaretus (L.). Journal of Fish Biology, 17(1):83–90, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bregazzi:1980:BPE

[81] P. R. Bregazzi and C. R. Kennedy. The biology of pike, Esox lucius L., in a southern eutrophic lake. Journal of Fish Biology, 17(1):91–112, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1980:TAC

[82] I. A. Johnston and B. Maitland. Temperature acclimation in crucian carp, Carassius carassius L., morphometric analyses of muscle fibre ultrastructure. Journal of Fish Biology, 17(1):113–125, July 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1980:SBL

[83] D. B. Lewis. Studies of the biology of the lesser weever fish, Trachinus vipera Cuvier II. Distribution: II. Distribution. Journal of Fish Biology, 17(2):127–133, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agrawal:1980:CTA

[84] N. K. Agrawal and C. L. Mahajan. Comparative tissue ascorbic acid studies in fishes. Journal of Fish Biology, 17(2):135–141, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1980:FAD

[85] S. J. M. Blaber and T. G. Blaber. Factors affecting the distribution of juvenile estuarine and inshore fish. Journal of Fish Biology, 17(2):143– 162, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1980:GRS

[86] R. H. K. Mann. The growth and reproductive strategy of the gudgeon, Gobio gobio (L.), in two hard-water rivers in southern England. Journal of Fish Biology, 17(2):163–176, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walters:1980:ESB

[87] G. R. Walters and J. A. Plumb. Environmental stress and bacterial infection in channel catfish, Ictalurus punctatus Rafinesque. Journal of Fish Biology, 17(2):177–185, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1980:GEPa

[88] M. Jobling. Gastric evacuation in plaice, Pleuronectes platessa L.: effects of dietary energy level and food composition. Journal of Fish Biology, 17(2):187–196, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1980:CBR

[89] G. P. Jones. Contribution to the biology of the redbanded perch, Ellerkeldia huntii (Hector), with a discussion on hermaphroditism. Journal of Fish Biology, 17(2):197–207, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broughton:1980:ETL

[90] N. M. Broughton, G. Goldspink, and N. V. Jones. The effect of training on the lateral musculature of O-group roach, Rutilus rutilus (L.), and their fatigue in subsequent exercise tests. Journal of Fish Biology, 17 (2):209–217, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnstone:1980:DDA

[91] A. D. F. Johnstone. The detection of dissolved amino acids by the Atlantic cod, Gadus morhua L. Journal of Fish Biology, 17(2):219–230, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plnchuk:1980:LLS

[92] V. I. Plnchuk. The lateral-line system of Caspiosoma caspium (Kessler, 1877) and the systematic status of Asra turcomanus Iljin, 1941. Journal of Fish Biology, 17(2):231–235, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRd

[93] Anonymous. Book reviews. Journal of Fish Biology, 17(2):237–242, August 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dadzie:1980:RBL

[94] S. Dadzie and B. C. C. Wangila. Reproductive biology, length–weight relationship and relative condition of pond raised Tilapia zilli (Gervais). Journal of Fish Biology, 17(3):243–253, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watson:1980:MPP

[95] R. A. Watson and T. A. Dick. Metazoan parasites of pike, Esox lucius Linnaeus, from Southern Indian Lake, Manitoba, Canada. Journal of Fish Biology, 17(3):255–261, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauly:1980:PDN

[96] D. Pauly and P. Martosubroto. The population dynamics of Nemipterus marginatus (Cuvier & Val.) off Western Kalimantan, South China Sea. Journal of Fish Biology, 17(3):263–273, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barham:1980:HAB

[97] W. T. Barham, G. L. Smit, and H. J. Schoonbee. The haematological assessment of bacterial infection in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 17(3):275–281, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubbock:1980:SFA

[98] R. Lubbock. The shore fishes of Ascension Island. Journal of Fish Biology, 17(3):283–303, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanBanning:1980:OBS

[99] P. Van Banning. The occurrence of black spots in the tongue sole, Cynoglossus browni Chabanaud, due to nematode eggs (Capillaria spinosa) previously described in the shark Carcharhinus milberti Müller & Henle. Journal of Fish Biology, 17(3):305–309, September 1980. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goven:1980:PCC

[100] B. A. Goven, D. L. Dawe, and J. B. Gratzek. Protection of channel catfish, Ictalurus punctatus Rafinesque, against Ichthyophthirius multifiliis Fouquet by immunization. Journal of Fish Biology, 17(3):311–316, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Healy:1980:BGA

[101] J. A. Healy and M. F. Mulcahy. A biochemical genetic analysis of populations of the northern pike, Esox lucius L., from Europe and North America. Journal of Fish Biology, 17(3):317–324, September 1980. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1980:ESP

[102] M. Jobling. Effects of starvation on proximate chemical composition and energy utilization of plaice, Pleuronectes platessa L. Journal of Fish Biology, 17(3):325–334, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRe

[103] Anonymous. Book reviews. Journal of Fish Biology, 17(3):335–336, September 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1980:HAG

[104] G. L. Smit and J. Hattingh. Haematological assessment of generally used freshwater fish blood anticoagulants. Journal of Fish Biology, 17(4):337– 341, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laird:1980:LOT

[105] L. M. Laird, A. R. Wilson, and F. G. T. Holliday. Lesions observed in the testis of precociously maturing male Atlantic salmon, Salmo salar L. Journal of Fish Biology, 17(4):343–348, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madhavi:1980:CPF

[106] R. Madhavi. Comparison of the parasitic fauna of Aplocheilus panchax and A. melastigma. Journal of Fish Biology, 17(4):349–358, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lonning:1980:SEL

[107] S. Lønning and J. Davenport. The swelling egg of the long rough dab, Hippoglossoides platessoides limandoides (Bloch). Journal of Fish Biology, 17(4):359–378, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vlele:1980:ATI

[108] D. Vlele, T. H. Kerstetter, and J. Sullivan. Adoptive transfer of immunity against Vibrio anguillarum in rainbow trout, Salmo gairdneri Richardson, vaccinated by the immersion method. Journal of Fish Biology, 17(4):379–386, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1980:EAR

[109] R. J. Wootton, J. R. M. Allen, and S. J. Cole. Energetics of the annual reproductive cycle in female sticklebacks, Gasterosteus aculeatus L. Journal of Fish Biology, 17(4):387–394, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flowerdew:1980:EBJ

[110] M. W. Flowerdew and D. J. Grove. An energy budget for juvenile thicklipped mullet, Crenimugil labrosus (Risso). Journal of Fish Biology, 17 (4):395–410, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1980:SCA

[111] E. J. Hyslop. Stomach contents analysis — a review of methods and their application. Journal of Fish Biology, 17(4):411–429, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hlavek:1980:EMG

[112] R. R. Hlavek and R. V. Bulkley. Effects of malachite green on leucocyte abundance in rainbow trout, Salmo gairdneri (Richardson). Journal of Fish Biology, 17(4):431–444, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1980:ACA

[113] A. C. Smith and F. Ramos. Automated chemical analysis in fish health assessment. Journal of Fish Biology, 17(4):445–450, October 1980. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1980:GRS

[114] J. E. Thorpe and R. I. G. Morgan. Growth-rate and smolting-rate of progeny of male Atlantic salmon parr, Salmo salar L. Journal of Fish Biology, 17(4):451–460, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrow:1980:IPB

[115] W. J. W. Morrow and A. Pulsford. Identification of peripheral blood leucocytes of the dogfish (Scyliorhinus canicula L.) by electron microscopy. Journal of Fish Biology, 17(4):461–475, October 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoakim:1980:HHU

[116] E. G. Yoakim and J. M. Grizzle. Histological, histochemical and ultrastructural studies on the interrenal and chromaffin cells of the fathead minnow, Pimephales promelas Rafinesque. Journal of Fish Biology, 17 (5):477–494, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1980:BRS

[117] A. P. Scott, V. J. Bye, S. M. Baynes, and J. R. C. Springate. Book review: Seasonal variations in plasma concentrations of 11-ketotestosterone and testosterone in male rainbow trout, Salmo gairdnerii Richardson. Journal of Fish Biology, 17(5):495–505, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1980:SRP

[118] R. M. Lee and S. D. Gerking. Survival and reproductive performance of the desert pupfish, Cyprinodon n. nevadensis (Eigenmann and Eigenmann), in acid waters. Journal of Fish Biology, 17(5):507–515, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:1980:CSE

[119] R. M. G. Wells, M. D. Ashby, S. J. Duncan, and J. A. Macdonald. Comparative study of the erythrocytes and haemoglobins in nototheniid fishes from Antarctica. Journal of Fish Biology, 17(5):517–527, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezeasor:1980:SEM

[120] Daniel N. Ezeasor and W. M. Stokoe. Scanning electron microscopic study of the gut mucosa of the rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 17(5):529–539, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahy:1980:GSF

[121] Edward Fahy. ‘growing season’ as a factor in sea trout production. Journal of Fish Biology, 17(5):541–546, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1980:GEPb

[122] M. Jobling. Gastric evacuation in plaice, Pleuronectes platessa L.: effects of temperature and fish size. Journal of Fish Biology, 17(5):547–551, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1980:NSD

[123] N. R. Merrett. A new species of the deep-sea fish genus Coelorinchus giorna (Macrouridae) from the eastern North Atlantic, with notes on its ecology. Journal of Fish Biology, 17(5):553–564, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1980:BSA

[124] R. N. Gibson and I. A. Ezzi. The biology of the scaldfish, Arnoglossus laterna (Walbaum) on the west coast of Scotland. Journal of Fish Biology, 17(5):565–575, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1980:PCA

[125] G. J. A. Kennedy and C. D. Strange. Population changes after two years of salmon (Salmo salar L.) stocking in upland trout (Salmo trutta L.) streams. Journal of Fish Biology, 17(5):577–586, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1980:SVS

[126] A. P. Scott, V. J. Bye, and S. M. Baynes. Seasonal variations in sex steroids of female rainbow trout (Salmo gairdneri Richardson). Journal of Fish Biology, 17(5):587–592, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1980:HSM

[127] G. G. Stott, N. H. Mcarthur, R. Tarpley, R. F. Sis, and V. Jacobs. Histopathological survey of male gonads offish from petroleum production and control sites in the Gulf of Mexico. Journal of Fish Biology, 17 (5):593–602, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1980:RBM

[128] A. H. Weatherley, H. S. Gill, and S. C. Rogers. The relationship between mosaic muscle fibres and size in rainbow trout (Salmo gairdneri). Journal of Fish Biology, 17(5):603–610, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRf

[129] Anonymous. Book reviews. Journal of Fish Biology, 17(5):611–612, November 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1980:FCG

[130] Raj V. Kilambi. Food consumption, growth and survival of grass carp Ctenopharyngodon idella Val. at four salinities. Journal of Fish Biology, 17(6):613–618, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezeasor:1980:CLE

[131] D. N. Ezeasor and W. M. Stokoe. A cytochemical, light and electron microscopic study of the eosinophilic granule cells in the gut of the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 17(6): 619–634, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sarker:1980:FHM

[132] A. L. Sarker, N. K. Al-Daham, and M. N. Bhatti. Food habits of the mudskipper, Pseudapocryptes dentatus (Val.). Journal of Fish Biology, 17 (6):635–639, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1980:ACS

[133] C. L. Mahajan and J. M. S. Dheer. An autoradioraphic and cytochemical study of erythropoiesis in a fresh water fish, Channa punctatus Bloch. Journal of Fish Biology, 17(6):641–648, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beggs:1980:SPC

[134] G. L. Beggs, G. F. Holeton, and E. J. Grossman. Some physiological consequences of angling stress in muskellunge, Esox masquinongy Mitchill. Journal of Fish Biology, 17(6):649–659, December 1980. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kratt:1980:ASB

[135] L. F. Kratt and R. J. F. Smith. An analysis of the spawning behaviour of the Arctic grayling Thymallus arcticus (Pallas) with observations on mating success. Journal of Fish Biology, 17(6):661–666, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gervai:1980:ITC

[136] J. Gervai, S. Páter, A. Nagy, L. Horváth, and V. Csényi. Induced triploidy in carp, Cyprinus carpio L. Journal of Fish Biology, 17(6): 667–671, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murphy:1980:EMC

[137] T. M. Murphy. The effect of methallibure and cyproterone acetate on the gonadotrophic cells, plasma androgen levels and testes of precocious 1+ male Atlantic salmon parr Salmo salar L. Journal of Fish Biology, 17 (6):673–680, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1980:ESG

[138] A. W. H. Turnpenny and R. Williams. Effects of sedimentation on the gravels of an industrial river system. Journal of Fish Biology, 17(6): 681–693, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:1980:EBW

[139] R. J. Wootton, J. R. M. Allen, and S. J. Cole. Effect of body weight and temperature on the maximum daily food consumption of Gasterosteus aculeatus L. and Phoxinus phoxinus (L.): selecting an appropriate model. Journal of Fish Biology, 17(6):695–705, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1980:RBH

[140] A. P. Scott and S. M. Baynes. A review of the biology, handling and storage of salmonid spermatozoa. Journal of Fish Biology, 17(6):707–739, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:1980:GIN

[141] Arne Johan Jensen. The ‘Gut index’, a new parameter to measure the gross nutritional state of Arctic char, Salvelinus alpinus (L.) and brown trout, Salmo trutta L. Journal of Fish Biology, 17(6):741–747, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1980:BRg

[142] Anonymous. Book reviews. Journal of Fish Biology, 17(6):749–750, December 1980. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kazakov:1981:PSP

[143] R. V. Kazakov. Peculiarities of sperm production by anadromous and parr Atlantic salmon (Salmo salar L.) and fish cultural characteristics of such sperm. Journal of Fish Biology, 18(1):1–8, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowers:1981:HII

[144] A. Bowers and J. B. Alexander. Hyperosmotic infiltration: immunological demonstration of infiltrating bacteria in brown trout, Salmo trutta L. Journal of Fish Biology, 18(1):9–13, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coles:1981:DPP

[145] T. F. Coles. The distribution of perch, Perca fluviatilis L. throughout their first year of life in Llyn Tegid, North Wales. Journal of Fish Biology, 18(1):15–30, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1981:EVC

[146] D. J. A. Brown. The effects of various cations on the survival of brown trout, Salmo trutta at low pHs. Journal of Fish Biology, 18(1):31–40, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agius:1981:ESS

[147] C. Agius. The effects of splenectomy and subsequent starvation on the storage of haemosiderin by the melano-macrophages of rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 18(1):41–44, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shelton:1981:DRG

[148] W. L. Shelton, R. O. Smitherman, and G. L. Jensen. Density related growth of grass carp, Ctenopharyngodon idella (Val.) in managed small impoundments in Alabama. Journal of Fish Biology, 18(1):45–51, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1981:FRN

[149] N. R. Merrett. First record of a notacanth leptocephalus; a benthopelagic capture in the slope waters off Ireland. Journal of Fish Biology, 18(1):53– 57, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stuart-Kregor:1981:IGS

[150] P. A. C. Stuart-Kregor, J. P. Sumpter, and J. M. Dodd. The involvement of gonadotrophin and sex steroids in the control of reproduction in the parr and adults of Atlantic salmon, Salmo salar L. Journal of Fish Biology, 18(1):59–72, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carrick:1981:OCF

[151] T. R. Carrick. Oxygen consumption in the fry of brown trout (Salmo trutta L.) related to pH of the water. Journal of Fish Biology, 18(1):73– 80, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leknes:1981:UAM

[152] I. L. Leknes. The ultrastructure of the atrial myocardial cell in three teleostean species. Journal of Fish Biology, 18(1):81–85, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:SET

[153] M. Jobling. Some effects of temperature, feeding and body weight on nitrogenous excretion in young plaice Pleuronectes platessa L. Journal of Fish Biology, 18(1):87–96, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nordlie:1981:FRB

[154] F. G. Nordlie. Feeding and reproductive biology of eleotrid fishes in a tropical estuary. Journal of Fish Biology, 18(1):97–110, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRa

[155] Anonymous. Book reviews. Journal of Fish Biology, 18(1):111–116, January 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broughton:1981:HDL

[156] N. M. Broughton, G. Goldspink, and N. V. Jones. Histological differences in the lateral musculature of O-group roach, Rutilus rutilus (L.) from different habitats. Journal of Fish Biology, 18(2):117–122, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:1981:EAB

[157] J. T. Ferreira, G. L. Smit, and H. J. Schoonbee. The effect of the anaesthetic benzocaine hydrochloride on red cell fragilities in Cyprinus carpio. Journal of Fish Biology, 18(2):123–126, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cassani:1981:FBU

[158] J. R. Cassani. Feeding behaviour of underyearling hybrids of the grass carp, Ctenopharyngodon idella [female sign] and the bighead, Hypophthalmicthys nobilis [male sign] on selected species of aquatic plants. Journal of Fish Biology, 18(2):127–133, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubbock:1981:FSP

[159] R. Lubbock and A. Edwards. The fishes of Saint Paul’s rocks. Journal of Fish Biology, 18(2):135–157, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baranes:1981:ESD

[160] A. Baranes and J. Wendling. The early stages of development in Carcharhinus plumbeus. Journal of Fish Biology, 18(2):159–175, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1981:CMU

[161] P. C. Blaxhall. A comparison of methods used for the separation of fish lymphocytes. Journal of Fish Biology, 18(2):177–181, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robb:1981:OFD

[162] A. P. Robb. Observations on the food and diel feeding behaviour of pelagic O-group gadoids in the northern North Sea. Journal of Fish Biology, 18(2):183–194, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1981:RGF

[163] A. H. Weatherley and H. S. Gill. Recovery growth following periods of restricted rations and starvation in rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 18(2):195–208, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McArdle:1981:IPS

[164] H. J. McArdle and I. A. Johnston. Isolation and properties of sarcoplasmic reticulum from the fast muscles of plaice Pleuronectes platessa L. Journal of Fish Biology, 18(2):209–221, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Totland:1981:VLM

[165] G. K. Totland, H. Kryvi, Q. Bone, and P. R. Flood. Vascularization of the lateral muscle of some elasmobranchiomorph fishes. Journal of Fish Biology, 18(2):223–234, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRb

[166] Anonymous. Book reviews. Journal of Fish Biology, 18(2):235–236, February 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bulow:1981:SVR

[167] F. J. Bulow, M. E. Zeman, J. R. Winningham, and W. F. Hudson. Seasonal variations in RNA–DNA ratios and in indicators of feeding, reproduction, energy storage, and condition in a population of bluegill, Lepomis macrochirus Rafinesque. Journal of Fish Biology, 18(3):237– 244, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gwyther:1981:GEL

[168] D. Gwyther and D. J. Grove. Gastric emptying in Limanda limanda (L.) and the return of appetite. Journal of Fish Biology, 18(3):245–259, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1981:HSO

[169] G. G. Stott, N. H. McArthur, R. Tarpley, V. Jacobs, and R. F. Sis. Histopathologic survey of ovaries of fish from petroleum production and control sites in the Gulf of Mexico. Journal of Fish Biology, 18(3):261– 269, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crivelli:1981:BCC

[170] A. J. Crivelli. The biology of the common carp, Cyprinus carpio L. in the Camargue, southern France. Journal of Fish Biology, 18(3):271– 290, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:1981:HEA

[171] J. T. Ferreira, G. L. Smit, and H. J. Schoonbee. Haematological evaluation of the anaesthetic benzocaine hydrochloride in the freshwater fish Cyprinus carpio L. Journal of Fish Biology, 18(3):291–297, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Hakim:1981:DAG

[172] A. W. H. Al-Hakim, M. I. A. Al-Mehdi, and A. H. J. Al-Salman. Determination of age, growth and sexual maturity of Barbus grypus in the Dukan reservoir of Iraq. Journal of Fish Biology, 18(3):299–308, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swift:1981:HBS

[173] D. J. Swift. A holding box system for physiological experiments on rainbow trout (Salmo gairdneri Richardson) requiring rapid blood sampling. Journal of Fish Biology, 18(3):309–319, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cadwallader:1981:FGH

[174] P. L. Cadwallader and A. K. Eden. Food and growth of hatcheryproduced Chinook salmon, Oncorhynchus tshawytscha (Walbaum), in landlocked Lake Purrumbete, Victoria, Australia. Journal of Fish Biology, 18(3):321–330, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kryvi:1981:MCR

[175] H. Kryvi, T. Flatmark, T. Flatmark, and G. K. Totland. The myoglobin content in red, intermediate and white fibres of the swimming muscle sin three species of shark — a comparative study using high-performance liquid chromatography. Journal of Fish Biology, 18(3):331–338, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dick:1981:PAC

[176] T. A. Dick and M. Belosevic. Parasites of Arctic charr Salvelinus alpinus (Linnaeus) and their use in separating sea-run and non-migrating charr. Journal of Fish Biology, 18(3):339–347, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shackley:1981:VTM

[177] S. E. Shackley, P. E. King, and S. M. Gordon. Vitellogenesis and trace metals in a marine teleost. Journal of Fish Biology, 18(3):349–352, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lone:1981:EFA

[178] K. P. Lone and A. J. Matty. The effect of feeding androgenic hormones on the proteolytic activity of the alimentary canal of carp Cyprinus carpio L. Journal of Fish Biology, 18(3):353–358, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasurer:1981:SAR

[179] J. W. Treasurer and F. G. T. Holliday. Some aspects of the reproductive biology of perch Perca fluviatilis L. A histological description of the reproductive cycle. Journal of Fish Biology, 18(3):359–376, March 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shirvell:1981:VFR

[180] C. S. Shirvell. Validity of fin-ray ageing for brown trout. Journal of Fish Biology, 18(4):377–383, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:IFMa

[181] M. Jobling. The influences of feeding on the metabolic rate of fishes: a short review. Journal of Fish Biology, 18(4):385–400, April 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skorping:1981:SDA

[182] A. Skorping. Seasonal dynamics in abundance, development and pattern of infection of Bunodera luciopercae (Müller) in perch, Perca fluviatilis L. from an oligotrophic lake in Norway. Journal of Fish Biology, 18(4):401– 410, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agrawal:1981:EAA

[183] N. K. Agrawal and C. L. Mahajan. Effect of ascorbic acid deficiency on the uptake of iodine by thyroid and non-thyroid tissues of an airbreathing freshwater fish Channa punctatus Bloch. Journal of Fish Biology, 18(4):411–416, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1981:BSS

[184] A. W. H. Turnpenny, R. N. Bamber, and P. A. Henderson. Biology of the sand-smelt (Atherina presbyter Valenciennes) around Fawley Power Station. Journal of Fish Biology, 18(4):417–427, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cresswell:1981:PSM

[185] R. C. Cresswell. Post-stocking movements and recapture of hatcheryreared trout released into flowing waters — a review. Journal of Fish Biology, 18(4):429–442, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hastings:1981:GMS

[186] P. A. Hastings. Gonad morphology and sex succession in the protogynous hermaphrodite Hemanthias vivanus (Jordan and Swain). Journal of Fish Biology, 18(4):443–454, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pieterse:1981:OBC

[187] J. J. Pieterse, G. L. Smit, K. Van Vliet, H. J. Schoonbee, and J. Hattingh. Observations on the blood of the Chinese silver carp Hypophthalmichthys molitrix (Valenciennes). Journal of Fish Biology, 18(4):455–459, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthiessen:1981:HCF

[188] P. Matthiessen. Haematological changes in fish following aerial spraying with endosulfan insecticide for testse fly control in Botswana. Journal of Fish Biology, 18(4):461–469, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:1981:ADS

[189] C. Blake and B. F. Blake. Age determination in six species of fish from a Mexican Pacific coastal lagoon. Journal of Fish Biology, 18(4):471– 478, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:1981:UBS

[190] L. G. Ross, W. Watts, and A. H. Young. An ultrasonic biotelemetry system for the continuous monitoring of tail-beat rate from freeswimming fish. Journal of Fish Biology, 18(4):479–490, April 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takvorian:1981:OGS

[191] P. M. Takvorian and A. Cali. The occurrence of Glugea stephani (Hagenmuller, 1899) in American winter flounder, Pseudopleuronects americanus (Walbaum) from the New York–New Jersey lower bay complex. Journal of Fish Biology, 18(4):491–501, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRc

[192] Anonymous. Book reviews. Journal of Fish Biology, 18(4):503–504, April 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hickman:1981:DSD

[193] R. A. Hickman. Densities and swimbladder development of juvenile American eels, Anguilla rostrata (Lesueur) as related to energetics of migration. Journal of Fish Biology, 18(5):507–517, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomasso:1981:PCD

[194] J. R. Tomasso, K. B. Davis, and N. C. Parker. Plasma corticosteroid dynamics in channel catfish, Ictalurus punctatus (Rafinesque), during and after oxygen depletion. Journal of Fish Biology, 18(5):519–526, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezeasor:1981:LEM

[195] D. N. Ezeasor and W. M. Stokoe. Light and electron microscopic studies on the absorptive cells of the intestine, caeca and rectum of the adult rainbow trout, Salmo gairdneri, rich. Journal of Fish Biology, 18(5):527– 544, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Douglas:1981:RMD

[196] W. A. Douglas and W. J. R. Lanzing. The respiratory mechanisms of the dusky flathead, Platycephalus fuscus (Platycephalidae, Scorpaeniformes). Journal of Fish Biology, 18(5):545–552, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:1981:ULF

[197] G. B. Chapman. Ultrastructure of the liver of the fingerling rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 18(5):553– 567, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kleckner:1981:CSR

[198] R. C. Kleckner and W. H. Krueger. Changes in swimbladder retial morphology in Anguilla rostrata during premigration metamorphosis. Journal of Fish Biology, 18(5):569–577, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dixon:1981:AIC

[199] D. G. Dixon and J. B. Sprague. Acclimation-induced changes in toxicity of arsenic and cyanide to rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 18(5):579–589, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1981:HEP

[200] A. L. Scott and W. A. Rogers. Haematological effects of prolonged sublethal hypoxia on channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 18(5):591–601, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menasveta:1981:LTM

[201] P. Menasveta. Lethal temperature of marine fishes of the Gulf of Thailand. Journal of Fish Biology, 18(5):603–607, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benda:1981:CBT

[202] R. S. Benda. A comparison of bottom trawl catch rates in the Kenya waters of Lake Victoria. Journal of Fish Biology, 18(5):609–613, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:IFMb

[203] M. Jobling. The influences of feeding on the metabolic rate of fishes: a short review [note added in proof]. Journal of Fish Biology, 18(5): 615, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRd

[204] Anonymous. Books received. Journal of Fish Biology, 18(5):617–618, May 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Douglas:1981:CCV

[205] W. A. Douglas and W. J. R. Lanzing. Colour change and visual cues in the sand flathead, Platycephalus arenarius (Ramsay and Ogilby). Journal of Fish Biology, 18(6):619–628, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:1981:AER

[206] A. Ferguson and F. M. Mason. Allozyme evidence for reproductively isolated sympatric populations of brown trout Salmo trutta L. in Lough Melvin, Ireland. Journal of Fish Biology, 18(6):629–642, June 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1981:TFP

[207] J. F. Leatherland and R. A. Sonstegard. Thyroid function, pituitary structure and serum lipids in Great Lakes Coho salmon, Oncorhynchus kisutch Walbaum, ‘jacks’ compared with sexually immature spring salmon. Journal of Fish Biology, 18(6):643–653, June 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayasaka:1981:FCA

[208] S. S. Hayasaka and J. Sullivan. Furunculosis in cultured American eel Anguilla rostrata (Le Sueur). Journal of Fish Biology, 18(6):655–659, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1981:CNT

[209] H. C. Lane, A. E. Rolfe, and J. R. Nelson. Changes in the nucleotide triphosphate/haemoglobin and nucleotide triphosphate/red cell ratios of rainbow trout, Salmo gairdneri Richardson, subjected to prolonged starvation and bleeding. Journal of Fish Biology, 18(6):661–668, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1981:CAP

[210] R. N. Gibson and I. H. Hesthagen. A comparison of the activity patterns of the sand goby Pomatoschistus minutus (Pallas) from areas of different tidal range. Journal of Fish Biology, 18(6):669–684, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1981:OBG

[211] B. R. Johnston and D. W. Halton. Occurrence of Bucephaloides gracilescens metacercariae in three species of gadoid fish. Journal of Fish Biology, 18(6):685–691, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leong:1981:CMP

[212] T. S. Leong and J. C. Holmest. Communities of metazoan parasites in open water fishes of Cold Lake, Alberta. Journal of Fish Biology, 18(6): 693–713, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penney:1981:STT

[213] R. K. Penney and G. Goldspink. Short term temperature acclimation in myofibrillar ATPase of a stenotherm Salmo gairdneri Richardson and an eurytherm Carassius auratus. Journal of Fish Biology, 18(6):715– 721, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roales:1981:EGI

[214] R. R. Roales. The effect of growth inhibiting factors on the total lipid content of the zebrafish, Brachydanio rerio (Hamilton-Buchanan). Journal of Fish Biology, 18(6):723–728, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasure:1981:SAR

[215] J. W. Treasure. Some aspects of the reproductive biology of perch Perca fluviatilis L. Fecundity, maturation and spawning behaviour. Journal of Fish Biology, 18(6):729–740, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Northcote:1981:JCR

[216] T. G. Northcote. Juvenile current response, growth and maturity of above and below waterfall stocks of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 18(6):741–751, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roper:1981:SNF

[217] D. S. Roper. Superficial neuromasts of the flatfish Peltorhamphus novaezeelandiae (Günther). Journal of Fish Biology, 18(6):753–758, June 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:1981:ABG

[218] E. I. L. Silva and S. S. De Silva. Aspects of the biology of grey mullet, Mugil cephalus L., adult populations of a coastal lagoon in Sri Lanka. Journal of Fish Biology, 19(1):1–10, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1981:BCA

[219] D. L. Barber, J. E. Mills Westermann, and M. G. White. The blood cells of the Antarctic icefish Chaenocephalus aceratus Lönnberg: light and electron microscopic observations. Journal of Fish Biology, 19(1): 11–28, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:DDI

[220] M. Jobling. Dietary digestibility and the influence of food components on gastric evacuation in plaice, Pleuronectes platessa L. Journal of Fish Biology, 19(1):29–36, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zelnik:1981:EEP

[221] P. R. Zelnik and G. Goldspink. The effect of exercise on plasma cortisol and blood sugar levels in the rainbow trout, Salmo gairdnerii Richardson. Journal of Fish Biology, 19(1):37–43, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swift:1981:CSB

[222] D. J. Swift. Changes in selected blood component concentrations of rainbow trout, Salmo gairdneri Richardson, exposed to hypoxia or sublethal concentrations of phenol or ammonia. Journal of Fish Biology, 19(1): 45–61, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Godin:1981:EHD

[223] J.-G. J. Godin. Effect of hunger on the daily pattern of feeding rates in juvenile pink salmon, Oncorhynchus gorbuscha Walbaum. Journal of Fish Biology, 19(1):63–71, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernholm:1981:NSH

[224] B. Fernholm. A new species of hagfish of the genus Myxine, with notes on other eastern Atlantic myxinids. Journal of Fish Biology, 19(1):73– 82, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gosting:1981:ABC

[225] L. Gosting, D. M. Mirando, and R. W. Gould. Antigen-binding cells in the peripheral blood of sockeye salmon, Oncorhynchus nerka Walbaum, induced by immersion or intraperitoneal injectijon of Vibro anguillarum bacterin. Journal of Fish Biology, 19(1):83–86, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anthony:1981:VCT

[226] P. D. Anthony. Visual contrast thresholds in the cod Gadus morhua L. Journal of Fish Biology, 19(1):87–103, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1981:ESH

[227] C. R. Kennedy and R. J. Burrough. The establishment and subsequent history of a population of Ligula intestinalis in roach Rutilis rutilis (L.). Journal of Fish Biology, 19(1):105–126, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRe

[228] Anonymous. Book review. Journal of Fish Biology, 19(1):127, July 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:1981:ADL

[229] C. A. Mills. The attachment of dace, Leuciscus leuciscus L., eggs to the spawning substratum and the influence of changes in water current on their survival. Journal of Fish Biology, 19(2):129–134, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottaway:1981:PIV

[230] E. M. Ottaway and A. Clarke. A preliminary investigation into the vulnerability of young trout (Salmo trutta L.) and Atlantic salmon (S. salar L.) to downstream displacement by high water velocities. Journal of Fish Biology, 19(2):135–145, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1981:PRB

[231] G. L. Smit, J. Hattingh, and J. T. Ferreira. The physiological responses of blood during thermal adaptation in three freshwater fish species. Journal of Fish Biology, 19(2):147–160, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agius:1981:ESM

[232] C. Agius and R. J. Roberts. Effects of starvation on the melanomacrophage centres of fish. Journal of Fish Biology, 19(2):161–169, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1981:OEF

[233] C. R. Kennedy. The occurrence of Eubothrium fragile (Cestoda: Pseudophyllidae) in twaite shad, Alosa fallax (Lacépède) in the River Severn. Journal of Fish Biology, 19(2):171–177, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1981:ESE

[234] Cheng-Sheng Lee and Bruno Menu. Effects of salinity on egg development and hatching in grey mullet Mugil cephalus L. Journal of Fish Biology, 19(2):179–188, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morse:1981:RSF

[235] W. W. Morse. Reproduction of the summer flounder, Paralichthys dentatus (L.). Journal of Fish Biology, 19(2):189–203, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1981:ESC

[236] D. J. A. Brown and S. Lynam. The effect of sodium and calcium concentrations on the hatching of eggs and the survival of the yolk sac fry of brown trout, Salmo trutta L. at low pH. Journal of Fish Biology, 19 (2):205–211, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1981:EGP

[237] M. A. K. Smith. Estimation of growth potential by measurement of tissue protein synthetic rates in feeding and fasting rainbow trout, Salmo gairdnerii Richardson. Journal of Fish Biology, 19(2):213–220, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1981:LTS

[238] C. R. Kennedy. Long term studies on the population biology of two specie of eyefluke, Diplostomum gasterostei and Tylodelphys clavata (Digenea: Diplostomatidae), concurrently infecting the eyes of perch, Perca fluviatilis. Journal of Fish Biology, 19(2):221–236, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1981:PDC

[239] A. D. Woodhead. Penetration and distribution of carbon particles in a teleost fish, Poecilia formosa (Girard), the Amazon molly. Journal of Fish Biology, 19(2):237–242, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRf

[240] Anonymous. Book reviews. Journal of Fish Biology, 19(2):243–244, August 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:MMG

[241] M. Jobling. Mathematical models of gastric emptying and the estimation of daily rates of food consumption for fish. Journal of Fish Biology, 19 (3):245–257, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baynes:1981:RTS

[242] S. M. Baynes, A. P. Scott, and A. P. Dawson. Rainbow trout, Salmo gairdnerii Richardson, spermatozoa: effects of cations and pH on motility. Journal of Fish Biology, 19(3):259–267, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1981:FSC

[243] W. J. R. Lanzing and R. G. Wright. The fine structure of the chromatophores and other non-sensory components of the eye of the blue-eye, Pseudomugil signifer. Journal of Fish Biology, 19(3):269–284, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carey:1981:DPR

[244] W. E. Carey and D. L. G. Noakes. Development of photobehavioural responses in young rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 19(3):285–296, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:1981:HIR

[245] J. G. O’Neill. The humoral immune response of Salmo trutta L. and Cyprinus carpio L. exposed to heavy metals. Journal of Fish Biology, 19 (3):297–306, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collins:1981:FPS

[246] M. R. Collins. The feeding periodicity of striped mullet, Mugil cephalus L., in two Florida habitats. Journal of Fish Biology, 19(3):307–315, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:1981:OSC

[247] J. Davenport, S. Lønning, and E. Kjørsvik. Osmotic and structural changes during early development of eggs and larvae of the cod, Gadus morhua L. Journal of Fish Biology, 19(3):317–331, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:1981:REE

[248] B. Ross and K. Jauncey. A radiographic estimation of the effect of temperature on gastric emptying time in Sarotherodon niloticus (L) × S. aureus (Steindachner) hybrids. Journal of Fish Biology, 19(3):333–344, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezenwa:1981:AGD

[249] B. I. O. Ezenwa and K. Ikusemiju. Age and growth determinations in the catfish, Chrysichthys nigrodigitatus (Lacépède) by use of the dorsal spine. Journal of Fish Biology, 19(3):345–351, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kazakov:1981:ESA

[250] R. V. Kazakov. The effect of the size of Atlantic salmon, Salmo salar L., eggs on embryos and alevins. Journal of Fish Biology, 19(3):353–360, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRg

[251] Anonymous. Book reviews. Journal of Fish Biology, 19(3):361–363, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:FMA

[252] Anonymous. Future meetings and announcements. Journal of Fish Biology, 19(3):363, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:E

[253] Anonymous. Erratum. Journal of Fish Biology, 19(3):364, September 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lobel:1981:TBH

[254] P. S. Lobel. Trophic biology of herbivorous reef fishes: alimentary pH and digestive capabilities. Journal of Fish Biology, 19(4):365–397, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muller:1981:CBT

[255] R. G. Muller and R. S. Benda. Comparison of bottom trawl stock densities in the inner Kavirondo Gulf of Lake Victoria. Journal of Fish Biology, 19(4):399–401, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Botham:1981:HLO

[256] Jane W. Botham and Margaret J. Manning. The histogenesis of the lymphoid organs in the carp Cyprinus carpio L. and the ontogenetic development of allograft reactivity. Journal of Fish Biology, 19(4):403– 414, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lincoln:1981:SMT

[257] R. F. Lincoln. Sexual maturation in triploid male plaice (Pleuronectes platessa) and plaice × flounder (Platichthys flesus) hybrids. Journal of Fish Biology, 19(4):415–426, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munshi:1981:GFS

[258] J. S. D. Munshi and G. M. Hughes. Gross and fine structure of the pseudo branch of the climbing perch, Anabas testudineus (Bloch). Journal of Fish Biology, 19(4):427–438, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1981:TTF

[259] M. Jobling. Temperature tolerance and the final preferendum — rapid methods for the assessment of optimum growth temperatures. Journal of Fish Biology, 19(4):439–455, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1981:CED

[260] R. V. Kilambi and A. Zdinak. Comparison of early developmental stages and adults of grass carp, Ctenopharyngodon idella, and hybrid carp (female grass carp × male bighead Aristichthys nobilis). Journal of Fish Biology, 19(4):457–465, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amin:1981:SDE

[261] O. M. Amin. The seasonal distribution of Echinorhynchus salmonis (Acanthocephala: Echinorhynchidae) among rainbow smelt, Osmerus mordax Mitchell, in Lake Michigan. Journal of Fish Biology, 19(4):467– 474, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1981:ADG

[262] B. R. Pollock. Age determination and growth of luderick, Girella tricuspidata (Quoy and Gaimard), taken from Moreton Bay, Australia. Journal of Fish Biology, 19(4):475–485, October 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1981:EBG

[263] J. F. Leatherland and R. N. Nuti. Effects of bovine growth hormone on plasma FFA concentrations and liver, muscle and carcass lipid content in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 19 (5):487–498, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lincoln:1981:SMF

[264] R. F. Lincoln. Sexual maturation in female triploid plaice, Pleuronectes platessa, and plaice × flounder, Platichthys flesus, hybrids. Journal of Fish Biology, 19(5):499–508, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dixon:1981:IAD

[265] D. G. Dixon and J. W. Hilton. Influence of available dietary carbohydrate content on tolerance of waterborne copper by rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 19(5):509–518, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1981:TAS

[266] J. E. Thorpe, L. G. Ross, G. Struthers, and W. Watts. Tracking Atlantic salmon smolts, Salmo salar L., through Loch Voil, Scotland. Journal of Fish Biology, 19(5):519–537, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooper:1981:EFP

[267] M. J. Cooper and G. A. Wheatley. An examination of the fish population in the River Trent, Nottinghamshire using angler catches. Journal of Fish Biology, 19(5):539–556, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:1981:PRL

[268] A. L. Jensen. Population regulation in lake whitefish, Coregonus clupeaformis (Mitchill). Journal of Fish Biology, 19(5):557–573, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leknes:1981:USH

[269] I. L. Leknes and T. S. Sætersdal. On the ultrastructure of specific heart granules in teleosts. Journal of Fish Biology, 19(5):575–579, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campagna:1981:GVR

[270] C. G. Campagna and J. J. Cech, Jr. Gill ventilation and respiratory efficiency of Sacramento blackfish, Orthodon microlepidotus Ayres, in hypoxic environments. Journal of Fish Biology, 19(5):581–591, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottaway:1981:OSB

[271] E. M. Ottaway, P. A. Carling, A. Clarke, and N. A. Reader. Observations on the structure of brown trout, Salmo trutta Linnaeus, redds. Journal of Fish Biology, 19(5):593–607, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRh

[272] Anonymous. Book reviews. Journal of Fish Biology, 19(5):609–610, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1981:BRi

[273] Anonymous. Book received. Journal of Fish Biology, 19(5):610, November 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezeasor:1981:FSG

[274] D. N. Ezeasor. The fine structure of the gastric epithelium of the rainbow trout, Salmo gairdneri, Richardson. Journal of Fish Biology, 19(6):611– 627, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shireman:1981:UGC

[275] J. V. Shireman and M. J. Maceina. The utilization of grass carp, Ctenopharyngodon idella Val., for hydrilla control in Lake Baldwin, Florida. Journal of Fish Biology, 19(6):629–636, December 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whoriskey:1981:ESR

[276] F. G. Whoriskey, R. J. Naiman, and W. Linn Montgomery. Experimental sea ranching of brook trout, Salvelinus fontinalis Mitchill. Journal of Fish Biology, 19(6):637–651, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moon:1981:AAT

[277] T. W. Moon and I. A. Johnston. Amino acid transport and interconversions in tissues of freshly caught and food-deprived plaice, Pleuronectes platessa L. Journal of Fish Biology, 19(6):653–663, December 1981. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:1981:NGR

[278] C. R. Goldspink. A note on the growth-rate and year-class strength of bream, Abramis brama (L.), in three eutrophic lakes, England. Journal of Fish Biology, 19(6):665–673, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seymour:1981:EPC

[279] E. A. Seymour. The effects of powdered carp pituitary on ovarian development, ovarian ascorbic acid and ovulation in Carassius carassius L. exposed to various photoperiod and temperature regimes. Journal of Fish Biology, 19(6):675–682, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsai:1981:ESS

[280] Chu fa Tsai and Kwei yang Chang. Effect of sex and size on copper susceptibility of the common guppy, Lebistes reticulatus (Peter). Journal of Fish Biology, 19(6):683–689, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Michibata:1981:UDC

[281] Hitoshi Michibata. Uptake and distribution of cadmium in the egg of the teleost, Oryzias latipes. Journal of Fish Biology, 19(6):691–696, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1981:BNG

[282] R. N. Gibson and I. A. Ezzi. The biology of the Norway goby, Pomatoschistus norvegicus (Collett), on the west coast of Scotland. Journal of Fish Biology, 19(6):697–714, December 1981. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reynolds:1982:TBD

[283] W. W. Reynolds and M. E. Casterlin. Thermoregulatory behaviour, diel activity and the relationship of spontaneous locomotor activity to temperature in the sea raven, Hemitripterus americanus (Gmelin). Journal of Fish Biology, 20(1):1–4, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:1982:SMH

[284] S. Behrens Yamada and T. J. Mulligan. Strontium marking of hatchery reared Coho salmon, Oncorhynchus kisutch Walbaum, identification of adults. Journal of Fish Biology, 20(1):5–9, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woo:1982:SAS

[285] N. Y. S. Woo and W. C. M. Tong. Salinity adaptation in the snakehead, Ophiocephalus maculatus Lacépède: changes in oxygen consumption, branchial Na^{+}-K^{+}-ATPase and body composition. Journal of Fish Biology, 20(1):11–19, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bregazzi:1982:RPP

[286] P. R. Bregazzi and C. R. Kennedy. The responses of a perch, Perca fluviatilis L., population to eutrophication and associated changes in fish fauna in a small lake. Journal of Fish Biology, 20(1):21–31, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1982:IBG

[287] M. Anderson. The identification of the British grey mullets. Journal of Fish Biology, 20(1):33–38, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barton:1982:PCS

[288] B. A. Barton and R. E. Peter. Plasma cortisol stress response in fingerling rainbow trout, Salmo gairdneri Richardson, to various transport conditions, anaesthesia, and cold shock. Journal of Fish Biology, 20(1): 39–51, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ezeasor:1982:DUT

[289] D. N. Ezeasor. Distribution and ultrastructure of taste buds in the oropharyngeal cavity of the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 20(1):53–68, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hilton:1982:EPF

[290] J. W. Hilton. The effect of pre-fasting diet and water temperature on liver glycogen and liver weight in rainbow trout, Salmo gairdneri Richardson, during fasting. Journal of Fish Biology, 20(1):69–78, January 1982. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nordlie:1982:OOR

[291] F. G. Nordlie, W. A. Szelistowski, and W. C. Nordlie. Ontogenesis of osmotic regulation in the striped mullet, Mugil cephalus L. Journal of Fish Biology, 20(1):79–86, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kumaraguru:1982:DCP

[292] A. K. Kumaraguru, F. W. H. Beamish, and H. W. Ferguson. Direct and circulatory paths of permethrin (NRDC-143) causing histopathological changes in the gills of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 20(1):87–91, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lone:1982:BRE

[293] K. P. Lone and A. J. Matty. Book review: The effect of feeding 11ketotestosterone on the food conversion efficiency and tissue protein and nucleic acid contents of juvenile carp, Cyprinus carpio L. Journal of Fish Biology, 20(1):93–104, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunt:1982:BRS

[294] S. M. V. Hunt, T. H. Simpson, and R. S. Wright. Book review: Seasonal changes in the levels of 11-oxotestosterone and testosterone in the serum of male salmon, Salmo salar L., and their relationship to growth and maturation cycle. Journal of Fish Biology, 20(1):105–119, January 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swinney:1982:DTH

[295] G. N. Swinney and T. F. Coles. Description of two hybrids involving silver bream, Blicca bjoerkna (L.), from British waters. Journal of Fish Biology, 20(2):121–129, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoss:1982:DFS

[296] D. E. Hoss and J. H. S. Blaxter. Development and function of the swimbladder-inner ear-lateral line system in the Atlantic menhaden, Brevoortia tyrannus (Latrobe). Journal of Fish Biology, 20(2):131–142, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forberg:1982:HSD

[297] K. G. Forberg. A histological study of development of oocytes in capelin, Mallotus villosus villosus (Müller). Journal of Fish Biology, 20(2):143– 154, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matallanas:1982:FHS

[298] J. Matallanas. Feeding habits of Scymnorhinus licha in Catalan waters. Journal of Fish Biology, 20(2):155–163, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1982:IBG

[299] A. H. Weatherley and H. S. Gill. Influence of bovine growth hormone on the growth dynamics of mosaic muscle in relation to somatic growth of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 20 (2):165–172, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robotham:1982:ASF

[300] P. W. J. Robotham. An analysis of a specialized feeding mechanism of the spined loach, Cobitis taenia (L.), and a description of the related structures. Journal of Fish Biology, 20(2):173–181, February 1982. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:1982:AAS

[301] P. Thomas, M. Bally, and J. M. Neff. Ascorbic acid status of mullet, Mugil cephalus Linn., exposed to cadmium. Journal of Fish Biology, 20 (2):183–196, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badcock:1982:NSD

[302] J. Badcock. A new species of the deep-sea fish genus Cyclothone Goode & Bean (Stomiatoidei, Gonostomatidae) from the tropical Atlantic. Journal of Fish Biology, 20(2):197–211, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoakim:1982:UAS

[303] E. G. Yoakim and J. M. Grizzle. Ultrastructure of alarm substance cells in the epidermis of the channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 20(2):213–221, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gray:1982:CTG

[304] R. H. Gray, T. L. Page, M. G. Saroglia, and P. Bronzi. Comparative tolerance to gas supersaturated water of carp, Cyprinus carpio, and black bullhead, Ictalurus melas, from the U.S.A. and Italy. Journal of Fish Biology, 20(2):223–227, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1982:RBT

[305] A. D. Pickering, T. G. Pottinger, and P. Christie. Recovery of the brown trout, Salmo trutta L., from acute handling stress: a time-course study. Journal of Fish Biology, 20(2):229–244, February 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1982:MMY

[306] B. R. Pollock. Movements and migrations of yellowfin bream, Acanthopagrus australis (Günther), in Moreton Bay, Queensland as determined by tag recoveries. Journal of Fish Biology, 20(3):245–252, March 1982. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wolters:1982:ENM

[307] W. R. Wolters, C. L. Chrisman, and G. S. Libey. Erythrocyte nuclear measurements of diploid and triploid channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 20(3):253–258, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stephens:1982:RGM

[308] R. T. T. Stephens. Reproduction, growth and mortality of the common bully, Gobiomorphus cotidianus McDowall, in a eutrophic New Zealand lake. Journal of Fish Biology, 20(3):259–270, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shafi:1982:SAB

[309] M. Shafi and B. M. Jasim. Some aspects of the biology of a cyprinid, Aspius vorax Heckel. Journal of Fish Biology, 20(3):271–278, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buck:1982:DMP

[310] R. J. G. Buck and A. F. Youngson. The downstream migration of precociously mature Atlantic salmon, Salmo salar l, parr in autumn; its relation to the spawning migration of mature adult fish. Journal of Fish Biology, 20(3):279–288, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofer:1982:DGP

[311] R. Hofer, H. Forstner, and R. Rettenwander. Duration of gut passage and its dependence on temperature and food consumption in roach, Rutilus rutilus l: laboratory and field experiments. Journal of Fish Biology, 20 (3):289–299, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:1982:ECI

[312] N. W. Miller and M. R. Tripp. The effect of captivity on the immune response of the killifish, Fundulus heteroclitus L. Journal of Fish Biology, 20(3):301–308, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:1982:ISS

[313] N. W. Miller and M. R. Tripp. An immunoinhibitory substance in the serum of laboratory held killifish, Fundulus heteroclitus L. Journal of Fish Biology, 20(3):309–316, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rinne:1982:OMP

[314] J. N. Rinne and B. Wanjala. Observations on movement patterns of Tilapia spp. in Nyanza Gulf, Lake Victoria, East Africa. Journal of Fish Biology, 20(3):317–322, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulata:1982:ECA

[315] G. Hulata, R. Moav, and G. Wohlfarth. Effects of crowding and availability of food on growth rate of fry in the European and Chinese races of the common carp. Journal of Fish Biology, 20(3):323–327, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacob:1982:FFH

[316] Sheila Susan Jacob and N. Balakrishnan Nair. Food and feeding habits of the larvivorous fish Aplocheilus lineatus (Cuv. & Val.) in its natural habitat. Journal of Fish Biology, 20(3):329–339, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maxwell:1982:IJM

[317] J. Garrey H. Maxwell. Infestation of the jack mackerel, Trachurus declivis (Jenyns), with the cymothoid isopod, Ceratothoa imbricatus (Fabricus), in south eastern Australian waters. Journal of Fish Biology, 20(3):341– 349, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nahhas:1982:SAS

[318] R. Nahhas, N. V. Jones, and G. Goldspink. Some aspects of sustained training of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 20(3):351–358, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:1982:POD

[319] J. A. Nelson. Physiological observations on developing rainbow trout, Salmo gairdneri (Richardson), exposed to low pH and varied calcium ion concentrations. Journal of Fish Biology, 20(3):359–372, March 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goven:1982:MTC

[320] Beverly A. Goven and Donald F. Amend. Mebendazole/trichlorfon combination: a new anthelmintic for removing monogenetic trematodes from fish. Journal of Fish Biology, 20(4):373–378, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carpene:1982:HFT

[321] E. Carpenè, A. Veggetti, and F. Mascarello. Histochemical fibre types in the lateral muscle of fishes in fresh, brackish and salt water. Journal of Fish Biology, 20(4):379–396, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robb:1982:HOR

[322] A. P. Robb. Histological observations on the reproductive biology of the haddock, Melanogrammus aeglefinus (L.). Journal of Fish Biology, 20(4):397–408, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:1982:ERT

[323] J. R. M. Allen and R. J. Wootton. The effect of ration and temperature on the growth of the three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 20(4):409–422, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1982:NGM

[324] J. F. Craig. A note on growth and mortality of trout, Salmo trutta L., in afferent streams of Windermere. Journal of Fish Biology, 20(4):423– 429, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1982:SOE

[325] M. Jobling. Some observations on the effects of feeding frequency on the food intake and growth of plaice, Pleuronectes platessa L. Journal of Fish Biology, 20(4):431–444, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korsgaard:1982:CAL

[326] B. Korsgaard. Changes in activity levels of glutamate dehydrogenase and two transaminases of maternal liver during reproduction and after administration of oestradiol and progesterone in Zoarces viviparus (L.). Journal of Fish Biology, 20(4):445–454, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1982:EPA

[327] A. D. Woodhead, R. B. Setlow, and V. Pond. Effects of polycyclic aromatic hydrocarbons on the proliferation of ectopic thyroid tissue in Poecilia formosa, the Amazon molly. Journal of Fish Biology, 20(4):455–463, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:1982:IRF

[328] B. E. Marshall. The influence of river flow on pelagic sardine catches in Lake Kariba. Journal of Fish Biology, 20(4):465–469, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Witthames:1982:AMA

[329] P. R. Witthames and M. Greer Walker. The activity of myofibrillar and actomyosin ATPase in the skeletal muscle of some marine teleosts in relation to their length and age. Journal of Fish Biology, 20(4):471– 478, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1982:OCA

[330] A. H. Weatherley, S. C. Rogers, D. G. Pincock, and J. R. Patch. Oxygen consumption of active rainbow trout, Salmo gairdneri Richardson, derived from electromyograms obtained by radiotelemetry. Journal of Fish Biology, 20(4):479–489, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stagg:1982:ACP

[331] R. M. Stagg and T. J. Shuttleworth. The accumulation of copper in Platichthys flesus L. and its effects on plasma electrolyte concentrations. Journal of Fish Biology, 20(4):491–500, April 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1982:SSF

[332] M. Jobling. A study of some factors affecting rates of oxygen consumption of plaice, Pleuronectes platessa L. Journal of Fish Biology, 20(5):501– 516, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1982:VCC

[333] J. F. Craig and J. M. Fletcher. The variability in the catches of charr, Salvelinus alpinus L., and perch, Perca fluviatilis L., from multi-mesh gill nets. Journal of Fish Biology, 20(5):517–526, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliassen:1982:SVG

[334] J.-E. Eliassen and O. Vahl. Seasonal variations in the gonad size and the protein and water content of cod, Gadus morhua (L.), muscle from Northern Norway. Journal of Fish Biology, 20(5):527–533, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hille:1982:LRB

[335] S. Hille. A literature review of the blood chemistry of rainbow trout, Salmo gairdneri Rich. Journal of Fish Biology, 20(5):535–569, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ajayi:1982:MSY

[336] T. O. Ajayi. The maximum sustainable yields of the inshore fish and shrimp resources of the Nigerian continental shelf. Journal of Fish Biology, 20(5):571–577, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1982:DSU

[337] G. J. A. Kennedy and C. D. Strange. The distribution of salmonids in upland streams in relation to depth and gradient. Journal of Fish Biology, 20(5):579–591, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacey:1982:NOD

[338] S. M. Lacey, I. C. Williams, and A. C. Carpenter. A note on the occurrence of the digenetic trematode Sphaerostoma bramae (Müller) in the intestine of the European eel, Anguilla anguilla (L.). Journal of Fish Biology, 20(5):593–596, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lone:1982:CBC

[339] K. P. Lone, B. W. Inch, and A. J. Matty. Changes in the blood chemistry of rainbow trout, Salmo gairdneri rich, in relation to dietary protein level, and an anabolic steroid hormone, ethylestrenol. Journal of Fish Biology, 20(5):597–606, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grove:1982:NAF

[340] D. J. Grove and L. B. Roberts. A note on an abnormal fin in Raja clavata L. Journal of Fish Biology, 20(5):607–609, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gumaa:1982:RDR

[341] S. A. Guma’a. Retinal development and retinomotor responses in perch, Perca fluviatilis L. Journal of Fish Biology, 20(5):611–618, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:BRa

[342] Anonymous. Book reviews. Journal of Fish Biology, 20(5):619–622, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:BRb

[343] Anonymous. Books received. Journal of Fish Biology, 20(5):622–623, May 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1982:OCF

[344] D. E. Wright and A. Eastcott. Operant conditioning of feeding behaviour and patterns of feeding in thick lipped mullet, Crenimugil labrosus (Risso) and common carp, Cyprinus carpio (L.). Journal of Fish Biology, 20(6):625–634, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jiming:1982:DFF

[345] Yang Jiming. The dominant fish fauna in the North Sea and its determination. Journal of Fish Biology, 20(6):635–643, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:1982:FAS

[346] C. A. Mills. Factors affecting the survival of dace, Leuciscus leuciscus (L.), in the early post-hatching period. Journal of Fish Biology, 20(6): 645–655, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1982:RCS

[347] C. L. Mahajan and T. R. Dheer. Regenerative capacity of the spleen in a splenectomized fish, Channa punctatus Bloch, with related investigations into changes in peripheral blood and haematopoiesis. Journal of Fish Biology, 20(6):657–666, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Muller:1982:EFI

[348] R. G. Muller, R. S. Benda, and T. G. Powell. Exploitation of fish in the inner Kavirondo Gulf of Lake Victoria based on tag returns. Journal of Fish Biology, 20(6):667–671, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tandler:1982:EFA

[349] A. Tandler, B. A. Berg, G. wm. Kissil, and A. M. Mackie. Effect of food attractants on appetite and growth rate of gilthead bream, Sparus aurata L. Journal of Fish Biology, 20(6):673–681, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singh:1982:VSI

[350] Sukhdev M. Singh and P. M. Sivalingam. In vitro study on the interactive effects of heavy metals on catalase activity of Sarotherodon mossambicus (Peters). Journal of Fish Biology, 20(6):683–688, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wandsvik:1982:OGR

[351] A. Wandsvik and M. Jobling. Observations on growth rates of Arctic charr, Salvelinus alpinus (L.), reared at low temperature. Journal of Fish Biology, 20(6):689–699, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wandsvik:1982:OMM

[352] A. Wandsvik and M. Jobling. Overwintering mortality of migratory Arctic charr, Salvelinus alpinus, (L.) reared in salt water. Journal of Fish Biology, 20(6):701–706, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliassen:1982:SVB

[353] Jens-Eric Eliassen and Ola Vahl. Seasonal variations in biochemical composition and energy content of liver, gonad and muscle of mature and immature cod, Gadus morhua (L.) from balsfjorden, northern Norway. Journal of Fish Biology, 20(6):707–716, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunkley:1982:APR

[354] D. A. Dunkley and W. M. Shearer. An assessment of the performance of a resistivity fish counter. Journal of Fish Biology, 20(6):717–737, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:BRc

[355] Anonymous. Book reviews. Journal of Fish Biology, 20(6):739–740, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:SBD

[356] Anonymous. Synopses of biological data on eight species of scombrids. Journal of Fish Biology, 20(6):741, June 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lane:1982:SAR

[357] H. C. Lane, J. W. Weaver, J. A. Benson, and H. A. Nichols. Some age related changes of adult rainbow trout, Salmo gairdneri Rich., peripheral erythrocytes separated by velocity sedimentation at unit gravity. Journal of Fish Biology, 21(1):1–14, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggberget:1982:IGS

[358] T. G. Heggberget and B. O. Johnsen. Infestations by Gyrodactylus sp. of Atlantic salmon, Salmo salar L., in Norwegian rivers. Journal of Fish Biology, 21(1):15–26, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatner:1982:MTS

[359] Mary F. Tatner and Margaret J. Manning. The morphology of the trout, Salmo gairdneri Richardson, thymus: some practical and theoretical considerations. Journal of Fish Biology, 21(1):27–32, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lone:1982:CEA

[360] K. P. Lone and A. J. Matty. Cellular effects of adrenosterone feeding to juvenile carp, Cyprinus carpio L., effect on liver, kidney, brain and muscle protein and nucleic acids. Journal of Fish Biology, 21(1):33– 45, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodworth:1982:CTR

[361] J. Woodworth and D. Pascoe. Cadmium toxicity to rainbow trout, Salmo gairdneri Richardson: a study of eggs and alevins. Journal of Fish Biology, 21(1):47–57, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1982:NDE

[362] A. D. Woodhead and Paulette C. Bryant. A note on the distribution of ectopic thyroid tissue in five species of poeciliid fishes. Journal of Fish Biology, 21(1):59–62, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1982:FIG

[363] R. V. Kilambi and A. Zdinak. Food intake and growth of hybrid carp (female grass carp, Ctenopharyngodon idella × male bighead, Aristichthys (Hypophthalmichthys) nobilis) fed on zooplankton and Chara. Journal of Fish Biology, 21(1):63–67, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nash:1982:BFG

[364] R. D. M. Nash. The biology of Fries’ goby, Lesueurigobius friesii (Malm), in the Firth of Clyde, Scotland, and a comparison with other stocks. Journal of Fish Biology, 21(1):69–85, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beumer:1982:SAI

[365] J. P. Beumer and G. J. Bacher. Species of Anguilla as indicators of mercury in the coastal rivers and lakes of Victoria, Australia. Journal of Fish Biology, 21(1):87–94, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:1982:AGR

[366] J. R. M. Allen and R. J. Wootton. Age, growth and rate of food consumption in an upland population of the three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 21(1):95–105, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Billy:1982:EFI

[367] Allen J. Billy. The effects of formalin and isopropyl alcohol on length and weight measurements of Sarotherodon mossambicus Trewavas. Journal of Fish Biology, 21(1):107–112, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eurell:1982:HUL

[368] J. A. Eurell and W. E. Haensly. The histology and ultrastructure of the liver of Atlantic croaker Micropogon undulatus L. Journal of Fish Biology, 21(1):113–125, July 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1982:RVC

[369] N. W. Pankhurst. Relation of visual changes to the onset of sexual maturation in the European eel Anguilla anguilla (L.). Journal of Fish Biology, 21(2):127–140, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1982:IWD

[370] H. J. Egglishaw and P. E. Shackley. Influence of water depth on dispersion of juvenile salmonids, Salmo salar L. and S. trutta L., in a Scottish stream. Journal of Fish Biology, 21(2):141–155, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hails:1982:RBT

[371] A. J. Hails and Z. Abdullah. Reproductive biology of the tropical fish Trichogaster pectoralis (Regan). Journal of Fish Biology, 21(2):157–170, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sameoto:1982:VDA

[372] D. Sameoto. Vertical distribution and abundance of the Peruvian anchovy, Engraulis ringens, and sardine, Sardinops sagax, larvae during November 1977. Journal of Fish Biology, 21(2):171–186, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1982:FHS

[373] E. J. Hyslop. The feeding habits of 0+ stone loach, Noemacheilus bavbatulus (L.), and bullhead, Cottus gobio L. Journal of Fish Biology, 21 (2):187–196, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casillas:1982:OAC

[374] E. Casillas, J. Sundquist, and W. E. Ames. Optimization of assay conditions for, and the selected tissue distribution of, alanine aminotransferase and aspartate aminotransferase of English sole, Parophrys vetulus Girard. Journal of Fish Biology, 21(2):197–204, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:1982:HCM

[375] H. S. Gill, A. H. Weatherley, and T. Bhesania. Histochemical characterization of myotomal muscle in the bluntnose minnow, Pimephales notatus Rafinesque. Journal of Fish Biology, 21(2):205–214, August 1982. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duthie:1982:ESS

[376] G. G. Duthie and D. F. Houlihan. The effect of single step and fluctuating temperature changes on the oxygen consumption of flounders, Platichthys flesus (L.): Lack of temperature adaptation. Journal of Fish Biology, 21 (2):215–226, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1982:NCB

[377] F. C. Stott. A note on catches of basking sharks, Cetorhinus maximus (Gunnerus), off Norway and their relation to possible migration paths. Journal of Fish Biology, 21(2):227–230, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newsome:1982:ASD

[378] C. S. Newsome and R. D. Piron. Aetiology of skeletal deformities in the Zebra Danio fish (Bruchydanio rerio, Hamilton-Buchanan). Journal of Fish Biology, 21(2):231–237, August 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barahona-Fernandes:1982:BDH

[379] M. H. Barahona-Fernandes. Body deformation in hatchery reared European sea bass Dicentrarchus labrax (L.) Types, prevalence and effect on fish survival. Journal of Fish Biology, 21(3):239–249, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badsha:1982:POH

[380] K. S. Badsha and C. R. Goldspink. Preliminary observations on the heavy metal content of four species of freshwater fish in NW England. Journal of Fish Biology, 21(3):251–267, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swift:1982:CSB

[381] D. J. Swift. Changes in selected blood component values of rainbow trout, Salmo gairdneri Richardson, following the blocking of the cortisol stress response with betamethasone and subsequent exposure to phenol or hypoxia. Journal of Fish Biology, 21(3):269–277, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1982:SCI

[382] N. W. Pankhurst and J. N. Lythgoe. Structure and colour of the integument of the European eel Anguilla anguilla (L.). Journal of Fish Biology, 21(3):279–296, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jenkins:1982:EOD

[383] N. Jenkins and J. M. Dodd. Effects of ovariectomy of the dogfish Scyliorhinus canicula L. on circulating levels of androgen and oestradiol and on pituitary gonadotrophin content. Journal of Fish Biology, 21 (3):297–303, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breazile:1982:SHP

[384] J. E. Breazile, L. L. Zinn, J. C. Yauk, H. J. Mass, and J. Wollscheid. A study of haematological profiles of channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 21(3):305–309, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1982:ECT

[385] J. F. Leatherland. Effect of a commercial trout diet on liver ultrastructure of fed and fasted yearling Coho salmon, Oncorhynchus kisutch Walbaum. Journal of Fish Biology, 21(3):311–319, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beaumont:1982:ASF

[386] W. R. C. Beaumont. An aberration in the scale formation of the dace Leuciscus leuciscus (L.). Journal of Fish Biology, 21(3):321–324, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:1982:AZG

[387] N. J. Milner. The accumulation of zinc by O-group plaice, Pleuronectes platessa (L.), from high concentrations in sea water and food. Journal of Fish Biology, 21(3):325–336, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collie:1982:CIF

[388] N. L. Collie and H. A. Bern. Changes in intestinal fluid transport associated with smoltification and seawater adaptation in Coho salmon, Oncorhynchus kisutch (Walbaum). Journal of Fish Biology, 21(3):337–348, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1982:SPG

[389] B. R. Pollock. Spawning period and growth of yellowfin bream, Acanthopagrus australis (Günther), in Moreton Bay, Australia. Journal of Fish Biology, 21(3):349–355, September 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1982:FGR

[390] M. Jobling. Food and growth relationships of the cod, Gadus morhua L., with special reference to balsfjorden, north Norway. Journal of Fish Biology, 21(4):357–371, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hilton:1982:EEV

[391] J. W. Hilton and H. W. Ferguson. Effect of excess vitamin D_{3} on calcium metabolism in rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 21(4):373–379, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bennett:1982:CUT

[392] R. O. Bennett and J. K. Dooley. Copper uptake by two sympatric species of killifish Fundulus heteroclitus (L.) and F. majalis (Walbaum). Journal of Fish Biology, 21(4):381–398, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winemiller:1982:IDC

[393] K. O. Winemiller and D. H. Taylor. Inbreeding depression in the convict cichlid, Cichlasoma nigrofasciatum (Baird and Girard). Journal of Fish Biology, 21(4):399–402, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blay:1982:FAF

[394] J. Blay, Jr. and K. N. Eyeson. Feeding activity and food habits of the shad, Ethmalosa fimbriata (Bowdich), in the coastal waters of Cape Coast, Ghana. Journal of Fish Biology, 21(4):403–410, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellgaard:1982:LAT

[395] E. G. Ellgaard and T. W. Rudner. Lead acetate: toxicity without effcts on the locomotor activity of the bluegill sunfish, Lepomis macrochirus Rafinesque. Journal of Fish Biology, 21(4):411–415, October 1982. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1982:CBM

[396] N. W. Pankhurst. Changes in body musculature with sexual maturation in the European eel, Anguilla anguilla (L.). Journal of Fish Biology, 21 (4):417–428, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodward:1982:PCR

[397] J. J. Woodward. Plasma catecholamines in resting rainbow trout, Salmo gairdneri Richardson, by high pressure liquid chromatography. Journal of Fish Biology, 21(4):429–432, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aneer:1982:SDI

[398] G. Aneer and S. Nellbring. A SCUBA-diving investigation of Baltic herring (Clupea harengus membras L.) spawning grounds in the Askö-Landsort area, northern Baltic proper. Journal of Fish Biology, 21(4): 433–442, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bursey:1982:ETN

[399] C. R. Bursey. Eustrongylides tubifex (Nitzsch) encystment in an American ell Anguilla rostrata (Le Sueur). Journal of Fish Biology, 21(4):443– 447, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shyamasundari:1982:OPM

[400] K. Shyamasundari, V. J. V. Raja Kumari, and K. Hanumantha Rao. Observations on the pancreas of the marine lizard fish Saurida tumbil (Bloch). Journal of Fish Biology, 21(4):449–454, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordon:1982:JSP

[401] D. J. Gordon and H. R. MacCrimmon. Juvenile salmonid production in a Lake Erie nursery stream. Journal of Fish Biology, 21(4):455–473, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:1982:EPH

[402] P. C. Young and R. B. Martin. Evidence for protogynous hermaphroditism in some lethrinid fishes. Journal of Fish Biology, 21(4):475–484, October 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blay:1982:ORB

[403] J. Blay, Jr. and K. N. Eyeson. Observations on the reproductive biology of the shad, Ethmalosa fimbriata (Bowdich), in the coastal waters of Cape Coast, Ghana. Journal of Fish Biology, 21(5):485–496, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peters:1982:ESS

[404] G. Peters. The effect of stress on the stomach of the European eel, Anguilla anguilla L. Journal of Fish Biology, 21(5):497–512, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Middleton:1982:OGA

[405] M. J. Middleton. The Oriental Goby, Acanthogobius flavimanus (Temminck and Schlegel), an introduced fish in the coastal waters of New South Wales, Australia. Journal of Fish Biology, 21(5):513–523, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ono:1982:PEM

[406] R. D. Ono. Proprioceptive endings in the myotomes of the pickerel (Teleostei: Esocidae). Journal of Fish Biology, 21(5):525–535, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:1982:EFG

[407] J. R. M. Allen and R. J. Wootton. Effect of food on the growth of carcase, liver and ovary in female Gasterosteus aculeatus L. Journal of Fish Biology, 21(5):537–547, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1982:CSS

[408] N. W. Pankhurst. Changes in the skin-scale complex with sexual maturation in the European eel, Anguilla anguilla (L.). Journal of Fish Biology, 21(5):549–561, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solanki:1982:CMC

[409] T. G. Solanki and M. Benjamin. Changes in the mucous cells of the gills, buccal cavity and epidermis of the nine-spined stickleback, Pungitius pungitius L., induced by transferring the fish to sea water. Journal of Fish Biology, 21(5):563–575, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soofiani:1982:ECD

[410] N. M. Soofiani and A. D. Hawkins. Energetic costs at different levels of feeding in juvenile cod, Gadus morhua L. Journal of Fish Biology, 21 (5):577–592, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:1982:HBA

[411] P. G. Bianco. Hybridization between Alburnus albidus (C.) and Leuciscus cephalus cabeda R. in Italy. Journal of Fish Biology, 21(5):593–603, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:BRd

[412] Anonymous. Book reviews. Journal of Fish Biology, 21(5):605–606, November 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alvarez-Lajonchere:1982:FMP

[413] L. Alvarez-Lajonchere. The fecundity of mullet (Pisces, Mugilidae) from Cuban waters. Journal of Fish Biology, 21(6):607–613, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirji:1982:LAA

[414] K. N. Hirji and W. A. M. Courtney. Leucine aminopeptidase activity in the digestive tract of perch, Perca fluviatilis L. Journal of Fish Biology, 21(6):615–622, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saldanha:1982:NSD

[415] L. Saldanha and N. R. Merrett. A new species of the deep-sea eel genus Ilyophis Gilbert (Synaphobranchidae) from the eastern North Atlantic, with comments on its ecology and intrafamilial relationships. Journal of Fish Biology, 21(6):623–636, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reimer:1982:IDD

[416] G. Reimer. The influence of diet on the digestive enzymes of the Amazon fish Matrinchã, Brycon cf. melanopterus. Journal of Fish Biology, 21(6): 637–642, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

George:1982:RAP

[417] C. J. George, A. E. Ellis, and D. W. Bruno. On remembrance of the abdominal pores in rainbow trout, Salmo gairdneri Richardson, and some other salmonid spp. Journal of Fish Biology, 21(6):643–647, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulsford:1982:CTI

[418] A. Pulsford, R. Fänge, and W. J. W. Morrow. Cell types and interactions in the spleen of the dogfish Scyliorhinus canicula L.: an electron microscopic study. Journal of Fish Biology, 21(6):649–662, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talbot:1982:OGB

[419] C. Talbot and P. J. Higgins. Observations on the gall bladder of juvenile Atlantic salmon Salmo salar L., in relation to feeding. Journal of Fish Biology, 21(6):663–669, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kimura:1982:ERR

[420] L. H. Kimura, M. A. Abad, and Y. Hokama. Evaluation of the radioimmunoassay (RIA) for detection of ciguatoxin (CTX) in fish tissues. Journal of Fish Biology, 21(6):671–680, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jezierska:1982:LMD

[421] B. Jezierska, J. R. Hazel, and S. D. Gerking. Lipid mobilization during starvation in the rainbow trout, Salmo gairdneri Richardson, with attention to fatty acids. Journal of Fish Biology, 21(6):681–692, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1982:AAS

[422] D. E. Wright and A. Eastcott. Association of an acoustic signal with operant conditioned feeding responses in thicklipped mullet, Crenimugil labrosus (Risso) and common carp, Cyprinus carpio (L.). Journal of Fish Biology, 21(6):693–698, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robotham:1982:ISL

[423] P. W. J. Robotham and J. S. Thomas. Infection of the spined loach, Cobitis taenia (L.), by the digenean, Allocreadium transversale (Rud.). Journal of Fish Biology, 21(6):699–703, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stott:1982:NBP

[424] F. C. Stott. Notes on the biology of the Patagonian hake, Merluccius hubbsi (Marini), the southern blue whiting, Micromesistius australis (Norman) and others in the Falkland Islands shelf area. Journal of Fish Biology, 21(6):705–714, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1982:BRe

[425] Anonymous. Book reviews. Journal of Fish Biology, 21(6):715–718, December 1982. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirji:1983:NSC

[426] K. N. Hirji and W. A. M. Courtney. Non-specific carboxylic esterase activity in the digestive tract of the perch, Perca fluviatilis L. Journal of Fish Biology, 22(1):1–7, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chervinski:1983:STM

[427] J. Chervinski. Salinity tolerance of the mosquito fish, Gambusia affinis (Baird and Girard). Journal of Fish Biology, 22(1):9–11, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1983:ICM

[428] C.-S. Lee and F. Hu. Influences of Ca and Mg ions on the egg survival of grey mullet, Mugil cephalus L. Journal of Fish Biology, 22(1):13–20, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nottage:1983:BSB

[429] A. S. Nottage and E. J. Perkins. The biology of solenette, Buglossidium luteum (Risso), in the Solway Firth. Journal of Fish Biology, 22(1):21– 27, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anyanwu:1983:PIP

[430] A. O. Anyanwu. Parasitic infestations of Pseudotolithus spp. off the coast of Lagos, Nigeria. Journal of Fish Biology, 22(1):29–33, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cassani:1983:FBY

[431] J. R. Cassani and W. E. Caton. Feeding behaviour of yearling and older hybrid grass carp. Journal of Fish Biology, 22(1):35–41, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1983:RGT

[432] A. H. Weatherley and H. S. Gill. Relative growth of tissues at different somatic growth rates in rainbow trout Salmo gairdneri Richardson. Journal of Fish Biology, 22(1):43–60, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1983:FAL

[433] P. C. Blaxhall. Factors affecting lymphocyte culture for chromosome studies. Journal of Fish Biology, 22(1):61–76, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartley:1983:MOM

[434] S. E. Hartley and M. T. Horne. A method for obtaining mitotic figures from blood leucocyte cultures of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 22(1):77–82, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mellinger:1983:ECD

[435] J. Mellinger. Egg-case diversity among dogfish, Scyliorhinus canicula (L.): a study of egg laying rate and nidamental gland secretory activity. Journal of Fish Biology, 22(1):83–90, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agrawal:1983:HHS

[436] N. K. Agrawal and C. L. Mahajan. Haematological and haematopoietic studies in pyridoxine deficient fish, Channa punctatus Bloch. Journal of Fish Biology, 22(1):91–103, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oni:1983:CPT

[437] S. K. Oni, J. Y. Olayemi, and J. D. Adegboye. Comparative physiology of three ecologically distinct freshwater fishes, Alestes nurse Ruppell, Synodontis schall Broch & Schneider and Tilapia zillii Gervais. Journal of Fish Biology, 22(1):105–109, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mahajan:1983:HHR

[438] C. L. Mahajan and T. R. Dheer. Haematological and haematopoietic responses to starvation in an air-breathing fish channapunctatus Bloch. Journal of Fish Biology, 22(1):111–123, January 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1983:AAF

[439] B. R. Pollock and M. J. Williams. An assessment of the angling fishery for yellowfin bream, Acanthopagrus australis (Günther), in Moreton Bay, Australia. Journal of Fish Biology, 22(2):125–132, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petrere:1983:YPR

[440] M. Petrere, Jr. Yield per recruit of the Tambaqui, Colossoma macropomum Cuvier, in the Amazonas State, Brazil. Journal of Fish Biology, 22 (2):133–144, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirji:1983:OHH

[441] K. N. Hirji. Observations on the histology and histochemistry of the oesophagus of the perch, Perca fluviatilis L. Journal of Fish Biology, 22 (2):145–152, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:GSF

[442] M. Jobling. Growth studies with fish — overcoming the problems of size variation. Journal of Fish Biology, 22(2):153–157, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plironen:1983:CPS

[443] J. Plironen and H. Hyvärinen. Cry preservation of spermatozoa of the whitefish Coregonus muksun Pallas. Journal of Fish Biology, 22(2):159– 163, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilder:1983:RDR

[444] I. B. Wilder and J. G. Stanley. RNA-DNA ratio as an index to growth in salmonid fishes in the laboratory and in streams contaminated by carbaryl. Journal of Fish Biology, 22(2):165–172, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:1983:BCT

[445] S. J. M. Blaber and D. P. Cyrus. The biology of Carangidae (Teleostei) in Natal estuaries. Journal of Fish Biology, 22(2):173–188, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:1983:WDF

[446] C. M. Wood, J. D. Turner, and M. S. Graham. Why do fish die after severe exercise? Journal of Fish Biology, 22(2):189–201, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hails:1983:TCF

[447] A. J. Hails. Temporal changes in fat and protein levels in the tropical Anabantid Trichogaster pectoralis (Regan). Journal of Fish Biology, 22 (2):203–213, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1983:ALT

[448] J. S. Alabaster, D. G. Shurben, and M. J. Mallett. The acute lethal toxicity of mixtures of cyanide and ammonia to smolts of salmon, Salmo salar L. at low concentrations of dissolved oxygen. Journal of Fish Biology, 22(2):215–222, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1983:EMS

[449] P. C. Blaxhall. Electron microscope studies of fish lymphocytes and thrombocytes. Journal of Fish Biology, 22(2):223–229, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geevarghese:1983:NGG

[450] C. Geevarghese and P. A. John. A new goby, Glossogobius minutus (Teleostei: Gobiidae) from the south-west coast of India. Journal of Fish Biology, 22(2):231–240, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1983:PDD

[451] A. D. Woodhead and A. Bornbusch. The pattern of distribution of dimethylbenz(a)anthracene in the Amazon molly, Poecilia formosa (Girard), after intraperitoneal injection. Journal of Fish Biology, 22(2): 241–248, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matallanas:1983:OMS

[452] J. Matallanas. The occurrence of Montagu’s Sea Snail, Liparis montagui (Osteichthyes–Scorpaeniformes–Cyclopteridae) in the Bay of Biscay (Spain). Journal of Fish Biology, 22(2):249–250, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRa

[453] Anonymous. Book reviews. Journal of Fish Biology, 22(2):251–252, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:FMA

[454] Anonymous. Future meetings and announcements. Journal of Fish Biology, 22(2):252, February 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Worthington:1983:ISR

[455] A. D. Worthington, N. A. A. Macfarlane, and K. W. Easton. The induced spawning of roach, Rutilus rutilus (L.), with the antioestrogens clomiphene and tamoxifen. Journal of Fish Biology, 22(3):253–257, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ince:1983:SSI

[456] B. W. Ince. Secretin-stimulated insulin release in vivo in European eels, Anguilla anguilla L. Journal of Fish Biology, 22(3):259–263, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1983:NSDa

[457] N. R. Merrett. A new species of the deep-sea fish genus Coryphaenoides gunnerus (Macrouridae) from the tropical eastern North Atlantic and its relationship with C. (Coryphaenoides) marshalli Iwamoto, 1970. Journal of Fish Biology, 22(3):265–278, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1983:LCC

[458] P. C. Blaxhall. Lymphocyte culture for chromosome preparation. Journal of Fish Biology, 22(3):279–282, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brusle:1983:CSH

[459] S. Bruslé. Contribution to the sexuality of a hermaphroditic teleost, Serranus hepatus L. Journal of Fish Biology, 22(3):283–292, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mallatt:1983:LGL

[460] J. Mallatt. Laboratory growth of larval lampreys (Lampetra (Entosphenus) tridentata Richardson) at different food concentrations and animal densities. Journal of Fish Biology, 22(3):293–301, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

teKronnie:1983:MFT

[461] G. te Kronnie, L. Tatarczuch, W. van Raamsdonk, and W. Kilarski. Muscle fibre types in the myotome of stickleback, Gasterosteus aculeatus L.; a histochemical, immunohistochemical and ultrastructural study. Journal of Fish Biology, 22(3):303–316, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooper:1983:RBP

[462] A. Cooper. The reproductive biology of poor-cod, Trisoptems minutus L., whiting, Merlangius merlangus L., and Norway pout, Trisoptems esmarkii Nilsson, off the west coast of Scotland. Journal of Fish Biology, 22(3):317–334, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1983:RLF

[463] T. E. Brown, A. W. Morley, N. T. Sanderson, and R. D. Tait. Report of a large fish kill resulting from natural acid water conditions in Australia. Journal of Fish Biology, 22(3):335–350, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piironen:1983:CMS

[464] J. Piironen and H. Hyvärinen. Composition of the milt of some teleost fishes. Journal of Fish Biology, 22(3):351–361, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saxton:1983:SNP

[465] A. M. Saxton, R. N. Iwamoto, and W. K. Hershberger. Smoltification in the net-pen culture of accelerated Coho salmon, Oncorhynchus kisutch Walbaum: prediction of saltwater performance. Journal of Fish Biology, 22(3):363–370, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRb

[466] Anonymous. Book reviews. Journal of Fish Biology, 22(3):371–372, March 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cyrus:1983:FFE

[467] D. P. Cyrus and S. J. M. Blaber. The food and feeding ecology of Gerreidae, Bleeker 1859, in the estuaries of Natal. Journal of Fish Biology, 22(4):373–393, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ono:1983:DMI

[468] R. D. Ono. Dual motor innervation in the axial musculature of fishes. Journal of Fish Biology, 22(4):395–408, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1983:SOP

[469] B. R. Pollock, H. Weng, and R. M. Morton. The seasonal occurrence of postlarval stages of yellowfin bream, Acanthopagrus australis (Günther), and some factors affecting their movement into an estuary. Journal of Fish Biology, 22(4):409–415, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1983:CKF

[470] P. C. Blaxhall. Chromosome karyotyping of fish using conventional and G-banding methods. Journal of Fish Biology, 22(4):417–424, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackie:1983:SCN

[471] A. M. Mackie and A. I. Mitchell. Studies on the chemical nature of feeding stimulants for the juvenile European eel, Anguilla anguilla (L.). Journal of Fish Biology, 22(4):425–430, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:1983:SGS

[472] E. R. Clark and J. A. L. Eraser. The survival and growth of six species of freshwater fish, in tapwater and diluted and undiluted effluent from sewage percolating filters. Journal of Fish Biology, 22(4):431–445, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Howells:1983:FIW

[473] E. J. Howells, M. E. Howells, and J. S. Alabaster. A field investigation of water quality, fish and invertebrates in the Mawddach river system, Wales. Journal of Fish Biology, 22(4):447–469, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:IBW

[474] M. Jobling. Influence of body weight and temperature on growth rates of Arctic charr, Salvelinus alpinus (L.). Journal of Fish Biology, 22(4):471– 475, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1983:CND

[475] D. L. Barber and J. E. Mills Westermann. Comparison of nuclear DNA content of rodlet cells and erythrocytes in some freshwater teleosts. Journal of Fish Biology, 22(4):477–484, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forberg:1983:MCG

[476] K. G. Forberg. Maturity classification and growth of capelin, Mallotus villosus villosus (M), oocytes. Journal of Fish Biology, 22(4):485–496, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beck:1983:EMD

[477] M. L. Beck and C. J. Biggers. Erythrocyte measurements of diploid and triploid Ctenopharyngodon idella × Hypophthalmichthys nobilis hybrids. Journal of Fish Biology, 22(4):497–502, April 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Page:1983:CLE

[478] M. Page and A. F. Rowley. A cytochemical, light and electron microscopical study of the leucocytes of the adult river lamprey, Lampetra fluviatilis (L. Gray). Journal of Fish Biology, 22(5):503–517, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1983:ILC

[479] J. F. Leatherland and R. A. Sonstegard. Inter lake comparison of liver morphology and in vitro hepatic monodeiodination of L-thyroxine in sexually mature Coho salmon, Oncorhynchus kisutch Walbaum, from Lakes Erie, Ontario, Michigan and Superior. Journal of Fish Biology, 22(5): 519–536, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gon:1983:BBS

[480] O. Gon and A. Ben-Tuvia. The biology of Boyer’s sand smelt, Atherina boyeri Risso in the Bardawil Lagoon on the Mediterranean coast of Sinai. Journal of Fish Biology, 22(5):537–547, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1983:NGS

[481] N. R. Merrett, Y. I. Sazonov, and Y. N. Shcherbachev. A new genus and species of rattail fish (Macrouridae) from the eastern North Atlantic and eastern Indian Ocean, with notes on its ecology. Journal of Fish Biology, 22(5):549–561, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mollah:1983:VCC

[482] M. F. A. Mollah and E. S. P. Tan. Viability of catfish (Clarias macrocephalus, Günther) eggs fertilized at varying post-ovulation times. Journal of Fish Biology, 22(5):563–566, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Black:1983:ADS

[483] G. A. Black, W. L. Montgomery, and F. G. Whoriskey. Abundance and distribution of Salmincola edwardsii (Copepoda) on anadromous brook trout, Salvelinus fontinalis, (Mitchill) in the Moisie River system, Quebec. Journal of Fish Biology, 22(5):567–575, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:ESI

[484] M. Jobling and A. Wandsvik. Effect of social interactions on growth rates and conversion efficiency of Arctic charr, Salvelinus alpinus L. Journal of Fish Biology, 22(5):577–584, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatner:1983:FIU

[485] M. F. Tatner and M. T. Horne. Factors influencing the uptake of ^{14}Clabelled Vibrio anguillarum vaccine in direct immersion experiments with rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 22 (5):585–591, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OGrady:1983:ODH

[486] M. F. O’Grady. Observations on the dietary habits of wild and stocked brown trout, Salmo trutta L., in Irish lakes. Journal of Fish Biology, 22(5):593–601, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beck:1983:EAP

[487] M. L. Beck, C. J. Biggers, and H. K. Dupree. Electrophoretic analysis of protein systems of Ctenopharyngodon idella (Val.), Hypophthalmichthys nobilis (Rich.) and their F_{1} triploid hybrid. Journal of Fish Biology, 22(5):603–611, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1983:RCC

[488] G. M. Hughes, C. Albers, D. Muster, and K. H. Götz. Respiration of the carp, Cyprinus carpio L., at 10 and 20^{o}C and the effects of hypoxia. Journal of Fish Biology, 22(5):613–628, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRc

[489] Anonymous. Book reviews. Journal of Fish Biology, 22(5):629–630, May 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:1983:BBA

[490] I. G. Cowx. The biology of bream, Abramis brama (L), and its natural hybrid with roach, Rutilus rutilus (L), in the River Exe. Journal of Fish Biology, 22(6):631–646, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wolfson:1983:RSJ

[491] F. H. Wolfson. Records of seven juveniles of the whale shark, Rhiniodon typus. Journal of Fish Biology, 22(6):647–655, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ilan:1983:IPI

[492] Z. Ilan and Z. Yaron. Interference of o,p’DDD with interrenal function and cortisol metabolism in Sarotherodon aureus (Steindachner). Journal of Fish Biology, 22(6):657–669, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Porter:1983:SOS

[493] D. A. Porter and A. J. Fivizzani. Spontaneous occurrence of a synchronous hermaphrodite in the banded killifish, Fundulus diaphanus (Lesueur). Journal of Fish Biology, 22(6):671–675, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quinn:1983:PWS

[494] T. P. Quinn, E. L. Brannon, and R. P. Whitman. Pheromones and the water source preferences of adult Coho salmon Oncorhynchus kisutch Walbaum. Journal of Fish Biology, 22(6):677–684, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shireman:1983:CGH

[495] J. V. Shireman, R. W. Rottmann, and F. J. Aldridge. Consumption and growth of hybrid grass carp fed four vegetation diets and trout chow in circular tanks. Journal of Fish Biology, 22(6):685–693, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smagula:1983:GNP

[496] C. M. Smagula and I. R. Adelman. Growth in a natural population of largemouth bass, Micropterus salmoides Lacépède, as determined by physical measurements and [14C]-glycine uptake by scales. Journal of Fish Biology, 22(6):695–703, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:QPR

[497] M. Jobling and A. Wandsvik. Quantitative protein requirements of Arctic charr, Salvelinus alpinus (L). Journal of Fish Biology, 22(6):705–712, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1983:REV

[498] J. F. Craig and C. Kipling. Reproduction effort versus the environment; case histories of Windermere perch, Perca fluviatilis L., and pike, Esox lucius L. Journal of Fish Biology, 22(6):713–727, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRd

[499] Anonymous. Book reviews. Journal of Fish Biology, 22(6):729–731, June 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjorsvik:1983:EEQ

[500] E. Kjørsvik and S. Lønning. Effects of egg quality on normal fertilization and early development of the cod, Gadus morhua L. Journal of Fish Biology, 23(1):1–12, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hjelmeland:1983:SMP

[501] K. Hjelmeland, M. Christie, and J. Raa. Skin mucus protease from rainbow trout, Salmo gairdneri Richardson, and its biological significance. Journal of Fish Biology, 23(1):13–22, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milton:1983:RBG

[502] D. A. Milton and A. H. Arthington. Reproductive biology of Gambusia affinis holbrooki Baird and Girard, Xiphophorus helleri (Günther) and X. maculatus (Heckel) (Pisces; Poeciliidae) in Queensland, Australia. Journal of Fish Biology, 23(1):23–41, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nottage:1983:GMR

[503] A. S. Nottage and E. J. Perkins. Growth and maturation of roker Raja clavata L. in the Solway Firth. Journal of Fish Biology, 23(1):43–48, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:1983:MAP

[504] C. R. Johnson, D. A. Ratkowsky, and R. W. G. White. Multivariate analysis of the phenotypic relationships of the species of Paragalaxias and Galaxias (Pisces: Galaxiidae) in Tasmania. Journal of Fish Biology, 23(1):49–63, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1983:CSS

[505] A. W. H. Turnpenny and R. N. Bamber. The critical swimming speed of the sand smelt (Atherina presbyter Cuvier) in relation to capture at a power station cooling water intake. Journal of Fish Biology, 23(1): 65–73, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersson:1983:PLA

[506] L. Andersson, N. Ryman, and G. Ståhl. Protein loci in the Arctic charr, Salvelinus alpinus L.: electrophoretic expression and genetic variability patterns. Journal of Fish Biology, 23(1):75–94, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Storch:1983:ESS

[507] V. Storch and J. V. Juario. The effect of starvation and subsequent feeding on the hepatocytes of Chanos chanos (Forsskål) fingerlings and fry. Journal of Fish Biology, 23(1):95–103, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eddy:1983:ULN

[508] F. B. Eddy, P. A. Kunzlik, and R. N. Bath. Uptake and loss of nitrite from the blood of rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L. in fresh water and in dilute sea water. Journal of Fish Biology, 23(1):105–116, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuiman:1983:GGF

[509] L. A. Fuiman. Growth gradients in fish larvae. Journal of Fish Biology, 23(1):117–123, July 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sheridan:1983:SVL

[510] M. A. Sheridan, W. V. Allen, and T. H. Kerstetter. Seasonal variations in the lipid composition of the steelhead trout, Salmo gairdneri Richardson, associated with the parr-smolt transformation. Journal of Fish Biology, 23(2):125–134, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirshfield:1983:IPL

[511] M. F. Hirshfield, R. P. Morin, and D. J. Hepner. Increased prevalence of larval Eustronglylides (Nematoda) in the mummichog, Fundulus heteroclitus (L.), from the discharge canal of a power plant in the Chesapeake Bay. Journal of Fish Biology, 23(2):135–142, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davison:1983:LMC

[512] W. Davison. The lateral musculature of the common bully, Gobiomorphus cotidianus, a freshwater fish from New Zealand. Journal of Fish Biology, 23(2):143–151, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munkittrick:1983:HVF

[513] K. R. Munkittrick and J. F. Leatherland. Haematocrit values in feral goldfish, Camssius auratus L., as indicators of the health of the population. Journal of Fish Biology, 23(2):153–161, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1983:ACB

[514] A. D. Pickering and J. Duston. Administration of cortisol to brown trout, Salmo trutta L., and its effects on the susceptibility to Saprolegnia infection and furunculosis. Journal of Fish Biology, 23(2):163–175, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:EFF

[515] M. Jobling. Effect of feeding frequency on food intake and growth of Arctic charr, Salvelinus alpinus L. Journal of Fish Biology, 23(2):177– 185, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauley:1983:IHC

[516] G. B. Pauley and C. M. Heartwell. Immune hypersensitivity in the channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 23(2):187–193, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weng:1983:IHS

[517] H. T. Weng. Identification, habitats and seasonal occurrence of juvenile whiting (Sillaginidae) in Moreton Bay, Queensland. Journal of Fish Biology, 23(2):195–200, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bamber:1983:DSS

[518] R. N. Bamber, R. Glover, P. A. Henderson, and A. W. H. Turnpenny. Diplostomiasis in the sand smelt, Atherina presbyter (Cuvier), population at Fawley Power Station. Journal of Fish Biology, 23(2):201–210, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talbot:1983:RMF

[519] C. Talbot and P. J. Higgins. A radiographic method for feeding studies on fish using metallic iron powder as a marker. Journal of Fish Biology, 23(2):211–220, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ottaway:1983:IWV

[520] E. M. Ottaway and D. R. Forrest. The influence of water velocity on the downstream movement of alevins and fry of brown trout, Salmo trutta L. Journal of Fish Biology, 23(2):221–227, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1983:CVO

[521] N. W. Pankhurst and J. N. Lythgoe. Changes in vision and olfaction during sexual maturation in the European eel Anguilla anguilla (L.). Journal of Fish Biology, 23(2):229–240, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchinson:1983:NRO

[522] P. Hutchinson. A note recording the occurrence of hermaphroditic smelt, Osmerus eperlanus (L.) from the River Thames, England. Journal of Fish Biology, 23(2):241–243, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRe

[523] Anonymous. Book reviews. Journal of Fish Biology, 23(2):245–248, August 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cren:1983:O

[524] E. D. Le Cren. Obituary. Journal of Fish Biology, 23(3):249–250, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jafri:1983:MHP

[525] S. I. H. Jafri and D. M. Ensor. The morphology and histology of the pineal organ in roach, Rutilus rutilus (L.). Journal of Fish Biology, 23 (3):251–256, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moriarty:1983:ADG

[526] Christopher Moriarty. Age determination and growth rate of eels, Anguilla anguilla (L). Journal of Fish Biology, 23(3):257–264, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cresswell:1983:PSM

[527] R. C. Cresswell and R. Williams. Post- stocking movements and recapture of hatchery-reared trout released into flowing waters — effect of prior acclimation to flow. Journal of Fish Biology, 23(3):265–276, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hakeem:1983:BIA

[528] O. H. El Hakeem and M. M. Babiker. Bioassay of insulin activity of pancreatic tissue in active and aestivated African lungfish, Protopterus annectens (Owen). Journal of Fish Biology, 23(3):277–282, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:1983:IFH

[529] R. Williams. The inshore fishes of Heard and McDonald Islands, Southern Indian Ocean. Journal of Fish Biology, 23(3):283–292, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouck:1983:TGS

[530] G. R. Bouck and R. E. King. Tolerance to gas supersaturation in fresh water and sea water by steelhead trout, Salmo gairdneri Richardson. Journal of Fish Biology, 23(3):293–300, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Srivastav:1983:CRF

[531] A. K. Srivastav. Calcaemic responses in the freshwater mud eel, Amphipnous cuchia, to vitamin D_{3} administration. Journal of Fish Biology, 23 (3):301–303, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrison:1983:OFJ

[532] B. R. S. Morrison. Observations on the food of juvenile Atlantic salmon, Salmo salar l, reared in a Scottish hill loch. Journal of Fish Biology, 23 (3):305–313, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crocker:1983:BOS

[533] P. A. Crocker, C. R. Arnold, J. A. DeBoer, and G. J. Holt. Blood osmolality shift in juvenile red drum, Sciaenops ocellatus L. exposed to fresh water. Journal of Fish Biology, 23(3):315–319, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1983:SPJ

[534] A. W. H. Turnpenny. Swimming performance of juvenile sprat, Sprattus sprattus L., and herring, Clupea harengus L., at different salinities. Journal of Fish Biology, 23(3):321–325, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytler:1983:TEM

[535] P. Tytler and T. Vaughan. Thermal ecology of the mudskippers, Periophthalmus koelreuteri (Pallas) and Boleophthalmus boddarti (Pallas) of Kuwait Bay. Journal of Fish Biology, 23(3):327–337, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagi:1983:POD

[536] G. Sagi, M. Abraham, and V. Hilge. Pinealectomy and ovarian development in the grey mullet, Liza ramada. Journal of Fish Biology, 23(3): 339–345, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1983:EDQ

[537] S. S. De Silva and A. A. Owoyemi. Effect of dietary quality on the gastric evacuation and intestinal passage in Sarotherodon mossambicus (Peters) fry. Journal of Fish Biology, 23(3):347–355, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rinne:1983:MFB

[538] J. N. Rinne and B. Wanjala. Maturity, fecundity, and breeding seasons of the major catfishes (suborder: siluroidea) in Lake Victoria, East Africa. Journal of Fish Biology, 23(3):357–363, September 1983. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:1983:SCC

[539] J. P. F. Wilson and T. J. Pitcher. The seasonal cycle of condition in the pollan, Coregonus autumnalis pollan Thompson, of Lough Neagh, Northern Ireland. Journal of Fish Biology, 23(3):365–370, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1983:BRf

[540] Anonymous. Book reviews. Journal of Fish Biology, 23(3):371–372, September 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drewett:1983:PLP

[541] N. Drewett and P. D. Abel. Pathology of lindane poisoning and of hypoxia in the brown trout, Salmo trutta L. Journal of Fish Biology, 23 (4):373–384, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruby:1983:ITC

[542] S. M. Ruby and V. Cairns. Implications of testicular constrictions on spermatogenesis in Lake Ontario lake trout, Salvelinus namaycush. Journal of Fish Biology, 23(4):385–395, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:IFC

[543] M. Jobling and A. Wandsvik. An investigation of factors controlling food intake in Arctic charr, Salvelinus alpinus L. Journal of Fish Biology, 23 (4):397–404, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blacker:1983:PRL

[544] R. W. Blacker. Pelagic records of the lumpsucker, Cyclopterus lumpus L. Journal of Fish Biology, 23(4):405–417, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofsten:1983:DFL

[545] A. V. Hofsten, D. Kahan, R. Katznelson, and T. Bar-El. Digestion of free-living nematodes fed to fish. Journal of Fish Biology, 23(4):419–428, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1983:NSDb

[546] N. R. Merrett. A new species of the deep-sea fish genus Paraliparis Collett (Liparididae) from the eastern North Atlantic, with notes on its ecology. Journal of Fish Biology, 23(4):429–439, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paperna:1983:PCH

[547] I. Paperna and J. G. Van As. The pathology of Chilodonella hexasticha (Kiernik). Infections in cichlid fishes. Journal of Fish Biology, 23(4):441– 450, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colville:1983:VRC

[548] T. P. Colville, R. H. Richards, and J. W. Dobbie. Variations in renal corpuscular morphology with adaptation to sea water in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 23(4):451– 456, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakai:1983:LBP

[549] D. K. Sakai. Lytic and bactericidal properties of salmonid sera. Journal of Fish Biology, 23(4):457–466, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:1983:DOR

[550] A. J. Geffen. The deposition of otolith rings in Atlantic salmon, Salmo salar L., embryos. Journal of Fish Biology, 23(4):467–474, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goswami:1983:SAO

[551] S. V. Goswami, I. Parwez, and B. I. Sundararaj. Some aspects of osmoregulation in a stenohaline freshwater catfish, Heteropneustes fossilis (Bloch), in different salinities. Journal of Fish Biology, 23(4):475–487, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:1983:SCM

[552] W. L. Morgan and D. A. Ritz. Sensory cues and mechanisms involved in the capture of euphausiids by the Australian salmon, Arripis trutta (Bloch & Schneider). Journal of Fish Biology, 23(4):489–493, October 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collares-Pereira:1983:BAC

[553] M. J. Collares-Pereira and M. M. Coelho. Biometrical analysis of Chondrostoma polylepis × Rutilus arcasi natural hybrids (Osteichthyes — Cypriniformes — Cyprinidae). Journal of Fish Biology, 23(5):495–509, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karlsson:1983:GMZ

[554] L. Karlsson. Gill morphology in the zebrafish, Brachydanio rerio (Hamilton-Buchanan). Journal of Fish Biology, 23(5):511–524, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowland:1983:SAF

[555] S. J. Rowland. Spawning of the Australian freshwater fish Murray cod, Maccullochella peeli (Mitchell), in earthen ponds. Journal of Fish Biology, 23(5):525–534, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogers:1983:UOM

[556] S. C. Rogers and A. H. Weatherley. The use of opercular muscle electromyograms as an indicator of the metabolic costs of fish activity in rainbow trout, Salmo gairdneri Richardson, as determined by radiotelemetry. Journal of Fish Biology, 23(5):535–547, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:TES

[557] M. Jobling. Towards an explanation of specific dynamic action (SDA). Journal of Fish Biology, 23(5):549–555, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:1983:PBA

[558] J. F. Davis and S. S. Hayasaka. Pathogenic bacteria associated with cultured American eels, Anguilla rostrata Le Sueur. Journal of Fish Biology, 23(5):557–564, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucke:1983:HRS

[559] D. Bucke, S. W. Feist, M. G. Norton, and M. S. Rolfe. A histopathological report of some epidermal anomalies of Dover sole, Solea solea L., and other flatfish species in coastal waters off south-east England. Journal of Fish Biology, 23(5):565–578, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cenini:1983:VEZ

[560] P. Cenini and R. J. Turner. In vitro effects of zinc on lymphoid cells of the carp, Cyprinus carpio L. Journal of Fish Biology, 23(5):579–583, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burke:1983:SCS

[561] M. Burke and J. F. Leatherland. Seasonal changes in serum thyroid hormone levels of the feral brown bullhead, Ictalurus nebulosus Lesueur. Journal of Fish Biology, 23(5):585–593, November 1983. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shireman:1983:GFH

[562] J. V. Shireman, F. J. Aldridge, and R. W. Rottmann. Growth and food habits of hybrid carp, Ctenopharyngodon idella × Aristichthys nobilis, from aerated and non- aerated ponds. Journal of Fish Biology, 23(5): 595–604, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacey:1983:EAL

[563] S. M. Lacey and I. C. Williams. Epieimeria anguillae (Léger & Hollande, 1922) Dyková & Lom, 1981 (Apicomplexa: Eucoccidia) in the European eel, Anguilla anguilla (L.). Journal of Fish Biology, 23(5): 605–609, November 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geevarghese:1983:MSG

[564] C. Geevarghese and P. A. John. Maturation and spawning of a gobiid fish, Oligolepis acutipennis (Cuv. & Val.), from the south-west coast of India. Journal of Fish Biology, 23(6):611–624, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Youngson:1983:ASE

[565] A. F. Youngson, R. J. G. Buck, T. H. Simpson, and D. W. Hay. The autumn and spring emigrations of juvenile Atlantic salmon, Salmo salar L., from the Girnock Burn, Aberdeenshire, Scotland: environmental release of migration. Journal of Fish Biology, 23(6):625–639, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukuhara:1983:DGL

[566] O. Fukuhara. Development and growth of laboratory reared Engraulis japonica (Houttuyn) larvae. Journal of Fish Biology, 23(6):641–652, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1983:PLW

[567] A. H. Weatherley and H. S. Gill. Protein, lipid, water and caloric contents of immature rainbow trout, Salmo gairdneri Richardson, growing at different rates. Journal of Fish Biology, 23(6):653–673, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1983:DAM

[568] S. S. De Silva and M. K. Perera. Digestibility of an aquatic macrophyte by the cichlid Etroplus suratensis (Bloch) with observations on the relative merits of three indigenous components as markers and daily changes in protein digestibility. Journal of Fish Biology, 23(6):675–684, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1983:SRC

[569] M. Jobling. A short review and critique of methodology used in fish growth and nutrition studies. Journal of Fish Biology, 23(6):685–703, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Noeske:1983:PDV

[570] T. A. Noeske and R. E. Spieler. Photoperiod and diel variations of serum cortisol, thyroxine and protein in goldfish, Carassius auratus L. Journal of Fish Biology, 23(6):705–710, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wedemeyer:1983:SPL

[571] G. A. Wedemeyer, R. W. Gould, and W. T. Yasutake. Some potentials and limits of the leucocrit test as a fish health assessment method. Journal of Fish Biology, 23(6):711–716, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:1983:ETM

[572] I. A. Johnston, F. B. Eddy, and G. M. O. Maloiy. The effects of temperature on muscle pH, adenylate and phosphogen concentrations in Oreochromis alcalicus grahami, a fish adapted to an alkaline hotspring. Journal of Fish Biology, 23(6):717–724, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lyle:1983:FFH

[573] J. M. Lyle. Food and feeding habits of the lesser spotted dogfish, Scyliorhinus canicula (L.), in Isle of Man waters. Journal of Fish Biology, 23(6):725–737, December 1983. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buck:1984:RBS

[574] R. J. G. Buck and D. W. Hay. The relation between stock size and progeny of Atlantic salmon, Salmo salar L., in a Scottish stream. Journal of Fish Biology, 24(1):1–11, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:1984:ESIa

[575] J. A. L. Fraser and E. R. Clark. The effects of a settled industrial– domestic sewage works effluent from percolating filters on the survival and growth of various stages of rainbow trout, Salmo gairdneri (Richardson). Journal of Fish Biology, 24(1):13–27, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Youngson:1984:CST

[576] A. F. Youngson and T. H. Simpson. Changes in serum thyroxine levels during smolting in captive and wild Atlantic salmon, Salmo salar L. Journal of Fish Biology, 24(1):29–39, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardiner:1984:EPD

[577] W. R. Gardiner. Estimating population densities of salmonids in deep water in streams. Journal of Fish Biology, 24(1):41–49, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:1984:EPD

[578] W. L. Morgan and D. A. Ritz. Effect of prey density and hunger state on capture of krill, Nyctiphanes australis Sars, by Australian salmon, Arripis trutta (Bloch & Schneider). Journal of Fish Biology, 24(1):51– 58, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1984:CNL

[579] M. P. Burton and D. R. Idler. Can Newfoundland landlocked salmon, Salmo salar L., adapt to sea water? Journal of Fish Biology, 24(1):59– 64, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heccbercet:1984:ESW

[580] Tor G. Heccbercet. Effect of supersaturated water on fish in the River Nidelva, southern Norway. Journal of Fish Biology, 24(1):65–74, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munkittrick:1984:SCP

[581] K. R. Munkittrick and J. F. Leatherland. Seasonal changes in the pituitary- gonad axis of feral goldfish, Carassius auratus L., from Ontario, Canada. Journal of Fish Biology, 24(1):75–90, January 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townshend:1984:EFS

[582] T. J. Townshend and R. J. Wootton. Effects of food supply on the reproduction of the convict cichlid, Cichlasoma nigrofasciatum. Journal of Fish Biology, 24(1):91–104, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Kadhomiy:1984:VPG

[583] N. K. Al-Kadhomiy. Vascular pathways in the gill filaments of the flounder, Platichthys flesus L. Journal of Fish Biology, 24(1):105–114, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowen:1984:DHU

[584] S. H. Bowen. Differential habitat utilization by sexes of Sarotherodon mossambicus in Lake Valencia, Venezuela: significance for fitness. Journal of Fish Biology, 24(1):115–121, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRa

[585] Anonymous. Book reviews. Journal of Fish Biology, 24(1):123–124, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRb

[586] Anonymous. Books received. Journal of Fish Biology, 24(1):124, January 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Danulat:1984:CAD

[587] E. Danulat and H. Kausch. Chitinase activity in the digestive tract of the cod, Gadus morhua (L.). Journal of Fish Biology, 24(2):125–133, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1984:SFP

[588] C. R. Kennedy. The status of flounders, Platichthys flesus L., as hosts of the Acanthocephalan Pomphorhynchus laevis (Müller) and its survival in marine conditions. Journal of Fish Biology, 24(2):135–149, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:1984:ADG

[589] J. P. F. Wilson and T. J. Pitcher. Age determination and growth of the pollan, Coregonus autumnalis pollan Thompson, of Lough Neagh, Northern Ireland. Journal of Fish Biology, 24(2):151–163, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diana:1984:GLB

[590] J. S. Diana. The growth of largemouth bass, Micropterus salmoides (Lacépède), under constant and fluctuating temperatures. Journal of Fish Biology, 24(2):165–172, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1984:ENE

[591] R. H. K. Mann and T. Penczak. The efficiency of a new electrofishing technique in determining fish numbers in a large river in Central Poland. Journal of Fish Biology, 24(2):173–185, February 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reiersen:1984:AVL

[592] L.-O. Reiersen and K. Fugelli. Annual variation in lymphocystis infection frequency in flounder, Platichthys flesus (L.). Journal of Fish Biology, 24 (2):187–191, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Secombes:1984:IRC

[593] C. J. Secombes and J. W. Resink. The immune response of carp, Cyprinus carpio L., following injection of antigen-antibody complexes. Journal of Fish Biology, 24(2):193–200, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartlett:1984:UAC

[594] J. R. Bartlett, P. F. Randerson, R. Williams, and D. M. Ellis. The use of analysis of covariance in the back-calculation of growth in fish. Journal of Fish Biology, 24(2):201–213, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colombo:1984:GDB

[595] G. Colombo, G. Grandi, and R. Rossi. Gonad differentiation and body growth in Anguilla anguilla L. Journal of Fish Biology, 24(2):215–228, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kipling:1984:SOA

[596] C. Kipling. Some observations on autumn-spawning charr, Salvelinus alpinus L., in Windermere, 1939–1982. Journal of Fish Biology, 24(2): 229–234, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromley:1984:ICF

[597] P. J. Bromley and T. C. Adkins. The influence of cellulose filler on feeding, growth and utilization of protein and energy in rainbow trout, Salmo gairdnerii Richardson. Journal of Fish Biology, 24(2):235–244, February 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmann:1984:ERB

[598] R. Hoffmann and R. Lommel. Effects of repeated blood sampling on some blood parameters in freshwater fish. Journal of Fish Biology, 24 (3):245–251, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:1984:FPC

[599] J. P. F. Wilson. The food of the pollan, Coregonus autumnalis pollan Thompson, in Lough Neagh, Northern Ireland. Journal of Fish Biology, 24(3):253–261, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeWKruger:1984:SCP

[600] J. C. De W.Kruger, G. L. Smit, J. H. J. Van Vuren, and J. T. Ferreira. Some chemical and physical characteristics of the semen of Cyprinus carpio L. and Oreochromis mossambicus (Peters). Journal of Fish Biology, 24(3):263–272, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lanzing:1984:MNP

[601] W. J. R. Lanzing. Melanogenesis in normal and phenylthiourea-treated eggs and larvae of Sarotherodon aureus. Journal of Fish Biology, 24 (3):273–280, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mauchline:1984:IPW

[602] J. Mauchline and J. D. M. Gordon. Incidence of parasitic worms in stomachs of pelagic and demersal fish of the Rockall Trough, northeastern Atlantic Ocean. Journal of Fish Biology, 24(3):281–285, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brodeur:1984:GER

[603] R. D. Brodeur. Gastric evacuation rates for two foods in the black rockfish, Sebastes melanops Girard. Journal of Fish Biology, 24(3):287–298, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heath:1984:CTA

[604] A. G. Heath. Changes in tissue adenylates and water content of bluegill, Lepomis macrochirus, exposed to copper. Journal of Fish Biology, 24 (3):299–309, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjorsvik:1984:OCD

[605] E. Kjörsvik, J. Davenport, and S. Lönning. Osmotic changes during the development of eggs and larvae of the lumpsucker, Cydopterus lumpus L. Journal of Fish Biology, 24(3):311–321, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langdon:1984:RGK

[606] J. S. Langdon and J. E. Thorpe. Response of the gill Na^{+}-K^{+}-ATPase activity, succinic dehydrogenase activity and chloride cells to saltwater adaptation in Atlantic salmon, Salmo salar L., parr and smolt. Journal of Fish Biology, 24(3):323–331, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benfey:1984:HTL

[607] T. J. Benfey and A. M. Sutterlin. The haematology of triploid landlocked Atlantic salmon, Salmo solar L. Journal of Fish Biology, 24(3):333– 338, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vuorinen:1984:RGV

[608] J. Vuorinen. Reduction of genetic variability in a hatchery stock of brown trout, Salmo trutta L. Journal of Fish Biology, 24(3):339–348, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRc

[609] Anonymous. Book reviews. Journal of Fish Biology, 24(3):349–350, March 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbieri:1984:NND

[610] G. Barbieri and M. C. Barbieri. Note on nutritional dynamics of Gymnotus carapo (L.) from the Lobo Reservoir, Sâo Paulo State, Brazil. Journal of Fish Biology, 24(4):351–355, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mauchline:1984:OSS

[611] J. Mauchline and J. D. M. Gordon. Occurrence of stones, sediment and fish scales in stomach contents of demersal fish of the Rockall Trough. Journal of Fish Biology, 24(4):357–362, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conneely:1984:MPF

[612] J. J. Conneely and T. K. McCarthy. The metazoan parasites of freshwater fishes in the Corrib catchment area, Ireland. Journal of Fish Biology, 24 (4):363–375, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:1984:ORO

[613] A. D. Munro. The ontogeny of the retina and optic tectum in Aequidens portalegrensis (Hensel). Journal of Fish Biology, 24(4):377–393, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kipling:1984:MRE

[614] C. Kipling and E. D. Le Cren. Mark-recapture experiments on fish in Windermere, 1943–1982. Journal of Fish Biology, 24(4):395–414, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harcup:1984:MBT

[615] M. F. Harcup, R. Williams, and D. M. Ellis. Movements of brown trout, Salmo trutta L., in the River Gwyddon, South Wales. Journal of Fish Biology, 24(4):415–426, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wallace:1984:ICE

[616] J. C. Wallace and D. Aasjord. An investigation of the consequences of egg size for the culture of Arctic charr, Salvelinus alpinus (L.). Journal of Fish Biology, 24(4):427–435, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McEvov:1984:ORR

[617] L.-A. McEvov (Barton). Ovulatory rhythms and over-ripening of eggs in cultivated turbot, Scophthalmus maximus L. Journal of Fish Biology, 24(4):437–448, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chervinski:1984:STG

[618] J. Chervinski. Salinity tolerance of the guppy, Poecilia reticulata Peters. Journal of Fish Biology, 24(4):449–452, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodward:1984:PHG

[619] B. Woodward and T. Bergeron. Protein histochemistry of the granule cells in the small intestine of the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 24(4):453–458, April 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hudson:1984:RBS

[620] W. F. Hudson and F. J. Bulow. Relationships between squamation chronology of the bluegill, Lepomis macrochims Rafinesque, and agegrowth methods. Journal of Fish Biology, 24(4):459–469, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agius:1984:EMS

[621] C. Agius and S. A. Agbede. An electron microscopical study on the genesis of lipofuscin, melanin and haemosiderin in the haemopoietic tissues of fish. Journal of Fish Biology, 24(4):471–488, April 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jolly:1984:ODS

[622] D. W. Jolly. Obituary: Dr Stanislas F. Snieszko. Journal of Fish Biology, 24(5):489, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cyrus:1984:RBG

[623] D. P. Cyrus and S. J. M. Blaber. The reproductive biology of Gerres in Natal estuaries. Journal of Fish Biology, 24(5):491–504, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lonning:1984:HPE

[624] S. Lönning, E. Kjörsvik, and J. Davenport. The hardening process of the egg chorion of the cod, Gadus morhua L., and lumpsucker, Cyclopterus lumpus L. Journal of Fish Biology, 24(5):505–522, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dickerson:1984:TPP

[625] H. W. Dickerson, J. Brown, D. L. Dawe, and J. B. Gratzek. Tetrahymena pyriformis as a protective antigen against Ichthyophthirius multifiliis infection: comparisons between isolates and ciliary preparations. Journal of Fish Biology, 24(5):523–528, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindsay:1984:DFD

[626] G. J. H. Lindsay. Distribution and function of digestive tract chitinolytic enzymes in fish. Journal of Fish Biology, 24(5):529–536, May 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palmer:1984:EEL

[627] C. J. Palmer and M. B. Culley. The egg and early life stages of the sandsmelt, Atherina presbyter Cuvier. Journal of Fish Biology, 24(5): 537–544, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zbanyszek:1984:EWS

[628] R. Zbanyszek and L. S. Smith. The effect of water-soluble aromatic hydrocarbons on some haematological parameters of rainbow trout, Salmo gairdneri Richardson, during acute exposure. Journal of Fish Biology, 24(5):545–552, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakanari:1984:ESC

[629] J. A. Sakanari, M. Moser, C. A. Reilly, and T. P. Yoshino. Effects of sublethal concentrations of zinc and benzene on striped bass, Morone saxatilis (Walbaum), infected with larval Anisakis nematodes. Journal of Fish Biology, 24(5):553–563, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardig:1984:SVB

[630] J. Härdig and L. B. Höglund. Seasonal variation in blood components of reared Baltic salmon, Salmo salar L. Journal of Fish Biology, 24(5):565– 579, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1984:CIG

[631] M. E. Scott and M. A. Robinson. Challenge infections of Gyrodactylus bullatamdis (Monogenea) on guppies, Poecilia reticulata (Peters), following treatment. Journal of Fish Biology, 24(5):581–586, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hightower:1984:BCF

[632] J. A. Hightower, L. J. McCumber, M. G. Welsh, D. S. Whatley, R. E. Hartvigsen, and M. M. Sigel. Blood cells of Fundulus heteroclitus (L.). Journal of Fish Biology, 24(5):587–598, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craik:1984:BCO

[633] J. C. A. Craik and S. M. Harvey. Biochemical changes occurring during final maturation of eggs of some marine and freshwater teleosts. Journal of Fish Biology, 24(5):599–610, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRd

[634] Anonymous. Book reviews. Journal of Fish Biology, 24(5):611–612, May 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1984:GML

[635] P. A. Henderson, J. W. Whitehouse, and G. H. Cartwright. The growth and mortality of larval herring, Clupea harengus L., in the River Blackwater Estuary, 1978–1980. Journal of Fish Biology, 24(6):613–622, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cochrane:1984:IFA

[636] K. L. Cochrane. The influence of food availability, breeding seasons and growth rate on commercial catches of Limnothrissa miodon (Boulenger) in Lake Kariba. Journal of Fish Biology, 24(6):623–635, June 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:1984:SVG

[637] B. D. Chang and W. Navas. Seasonal variations in growth, condition and gonads of Dormitator latifrons (Richardson) in the Chone River Basin, Ecuador. Journal of Fish Biology, 24(6):637–648, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cattan:1984:SML

[638] P. E. Cattan and J. Carvajal. A study of the migration of larval Anisakis simplex (Nematoda: Ascaridida) in the Chilean hake, Merluccius gayi (Guichenot). Journal of Fish Biology, 24(6):649–654, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takvorian:1984:SPM

[639] P. M. Takvorian and A. Cali. Seasonal prevalence of the microsporidan, Glugea stephani (Hagenmuller), in winter flounder, Pseudopleuronectes americanus (Walbaum), from the New York–New Jersey Lower Bay Complex. Journal of Fish Biology, 24(6):655–663, June 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lou:1984:DGI

[640] Y. D. Lou and C. E. Purdom. Diploid gynogenesis induced by hydrostatic pressure in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 24(6):665–670, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulsford:1984:PST

[641] A. Pulsford. Preliminary studies on trypanospmes from the dogfish, Scyliorhinus canicula L. Journal of Fish Biology, 24(6):671–682, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soderberg:1984:HCC

[642] R. W. Soderberg, M. V. McGee, and C. E. Boyd. Histology of cultured channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 24(6):683–690, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytler:1984:TSR

[643] P. Tytler and F. G. T. Holliday. Temporal and spatial relationships in the movements of loch dwelling brown trout, Salmo trutta L., recorded by ultrasonic tracking for 24 hours. Journal of Fish Biology, 24(6):691– 702, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradley:1984:EAP

[644] T. M. Bradley and A. W. Rourke. An electrophoretic analysis of plasma proteins from juvenile Oncorhynchus tshawytscha (Walbaum). Journal of Fish Biology, 24(6):703–709, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laale:1984:PTF

[645] H. W. Laale. Polyembryony in teleostean fishes: double monstrosities and triplets. Journal of Fish Biology, 24(6):711–719, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dabrowski:1984:MUB

[646] K. Dabrowski, S. J. Kaushik, and P. Luquet. Metabolic utilization of body stores during the early life of whitefish, Coregonus lavaretus L. Journal of Fish Biology, 24(6):721–729, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1984:AIT

[647] A. D. Pickering and A. Stewart. Acclimation of the interrenal tissue of the brown trout, Salmo trutta L., to chronic crowding stress. Journal of Fish Biology, 24(6):731–740, June 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caceci:1984:SEM

[648] T. Caceci. Scanning electron microscopy of goldfish, Carassius auratus, intestinal mucosa. Journal of Fish Biology, 25(1):1–12, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1984:GDW

[649] A. H. Weatherley and H. S. Gill. Growth dynamics of white myotomal muscle fibres in the bluntnose minnow, Pimephales notatus Rafinesque, and comparison with rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(1):13–24, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumann:1984:VAW

[650] P. C. Baumann and S. J. Hamilton. Vertebral abnormalities in white crappies, Pomoxis annularis Rafinesque, from Lake Decatur, Illinois, and an investigation of possible causes. Journal of Fish Biology, 25(1):25– 33, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:1984:UAB

[651] J. T. Ferreira, H. J. Schoonbee, and G. L. Smit. The uptake of the anaesthetic benzocaine hydrochloride by the gills and the skin of three freshwater fish species. Journal of Fish Biology, 25(1):35–41, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1984:GMZ

[652] J. F. Craig and J. M. Fletcher. Growth and mortality of zebra fish, Brachydanio rerio (Hamilton Buchanan), maintained at two temperatures and on two diets. Journal of Fish Biology, 25(1):43–55, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karges:1984:DLE

[653] R. G. Karges and B. Woodward. Development of lamellar epithelial hyperplasia in gills of pantothenic acid-deficient rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(1):57–62, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collins:1984:DPC

[654] M. R. Collins, M. J. Marshall, and C. A. Lanciani. The distribution of Poecilancistrium caryophyllum (Trypanorhyncha) plerocercoids in spot, Leiostomus xanthurus Lacépède, and spotted seatrout, Cynoscion nebulosus (Cuvier). Journal of Fish Biology, 25(1):63–68, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacArthur:1984:PIC

[655] J. I. MacArthur, T. C. Fletcher, B. J. S. Pirie, R. J. L. Davidson, and A. W. Thomson. Peritoneal inflammatory cells in plaice, Pleuronectes platessa L.: effects of stress and endotoxin. Journal of Fish Biology, 25 (1):69–81, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:1984:PFC

[656] D. M. Perry. Post-fertilization changes in the chorion of winter flounder, Pseudopleuronectes americanus Walbaum, eggs observed with scanning electron microscopy. Journal of Fish Biology, 25(1):83–94, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatner:1984:TLA

[657] M. F. Tatner, C. M. Johnson, and M. T. Horne. The tissue localization of Aeromonas salmonicida in rainbow trout, Salmo gairdneri Richardson, following three methods of administration. Journal of Fish Biology, 25 (1):95–108, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:1984:VHU

[658] C. A. Chapman and W. C. Mackay. Versatility in habitat use by a top aquatic predator, Esox lucius L. Journal of Fish Biology, 25(1):109– 115, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Humbert:1984:SEM

[659] W. Humbert, R. Kirsch, and M. F. Meister. Scanning electron microscopic study of the oesophageal mucous layer in the eel, Anguilla anguilla L. Journal of Fish Biology, 25(1):117–122, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:1984:EIC

[660] T. G. Pottinger, A. D. Pickering, and N. Blackstock. Ectoparasite induced changes in epidermal mucification of the brown trout, Salmo trutta L. Journal of Fish Biology, 25(1):123–128, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRe

[661] Anonymous. Book reviews. Journal of Fish Biology, 25(1):129, July 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellgaard:1984:EDA

[662] E. G. Ellgaard and J. Y. Gilmore III. Effects of different acids on the bluegill sunfish, Lepomis macrochirus Rafinesque. Journal of Fish Biology, 25(2):133–137, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Black:1984:DPL

[663] G. A. Black and J. M. Fraser. Dynamics of prevalence of Ligula intestinalis (L.) in Catostomus commersoni (Lacépède). Journal of Fish Biology, 25(2):139–146, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chervinski:1984:STY

[664] J. Chervinski. Salinity tolerance of young catfish, Clarias lazera (Burchell). Journal of Fish Biology, 25(2):147–149, August 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maier:1984:EDD

[665] K. J. Maier and R. E. Tullis. The effects of diet and digestive cycle on the gastrointestinal tract pH values in the goldfish, Carassius auratus L., Mozambique tilapia, Oreochromis mossambicus (Peters), and channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Biology, 25(2): 151–165, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tarpley:1984:CPC

[666] R. J. Tarpley, D. H. Lewis, J. E. Marks, and R. F. Sis. Collection and preparation of the channel catfish, Ictalurus punctatus Rafinesque, olfactory organ for scanning electron microscopy. Journal of Fish Biology, 25 (2):167–171, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dabrowski:1984:ILE

[667] K. R. Dabrowski and D. H. Jewson. The influence of light environment on depth of visual feeding by larvae and fry of Coregonus pollan (Thompson) in Lough Neagh. Journal of Fish Biology, 25(2):173–181, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Donnelly:1984:ECC

[668] W. A. Donnelly and L. M. Dill. Evidence for crypsis in Coho salmon, Oncorhynchus kisutch (Walbaum), parr: substrate colour preference and achromatic reflectance. Journal of Fish Biology, 25(2):183–195, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stave:1984:FAC

[669] J. W. Stave, B. S. Roberson, and F. M. Hetrick. Factors affecting the chemiluminescent response of fish phagocytes. Journal of Fish Biology, 25(2):197–206, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gelman:1984:ESC

[670] A. Gelman, S. Mokady, and U. Cogan. The effect of seasonal changes on the activity of intestinal alkaline phosphatase of pike perch, Lucioperka lucioperka and of bream, Abramis brama. Journal of Fish Biology, 25 (2):207–212, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossouw:1984:AGS

[671] G. J. Rossouw. Age and growth of the sand shark, Rhinobatos annulatus, in Algoa Bay, South Africa. Journal of Fish Biology, 25(2):213–222, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loughna:1984:ESU

[672] P. T. Loughna and G. Goldspink. The effects of starvation upon protein turnover in red and white myotomal muscle of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(2):223–230, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grossman:1984:REF

[673] G. D. Grossman and V. DeVlaming. Reproductive ecology of female Oligocottus snyderi Greeley: a North American intertidal sculpin. Journal of Fish Biology, 25(2):231–240, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berra:1984:RAA

[674] T. M. Berra. Reproductive anatomy of the Australian grayling, Prototroctes maraena Günther. Journal of Fish Biology, 25(2):241–251, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oronsaye:1984:EDC

[675] J. A. O. Oronsaye and A. E. Brafield. The effect of dissolved cadmium on the chloride cells of the gills of the stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 25(2):253–258, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRf

[676] Anonymous. Book reviews. Journal of Fish Biology, 25(2):259–260, August 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sorensen:1984:IMO

[677] P. W. Sorensen and H. E. Winn. The induction of maturation and ovulation in American eels, Anguilla rostrata (LeSueur), and the relevance of chemical and visual cues to male spawning behaviour. Journal of Fish Biology, 25(3):261–268, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cassani:1984:MCD

[678] J. R. Cassani, W. E. Caton, and B. Clark. Morphological comparisons of diploid and triploid hybrid grass carp, Ctenopharyngodon idella [female sign]× Hypophthalmichthys nobilis [male sign]. Journal of Fish Biology, 25(3):269–278, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:1984:CPC

[679] P. J. B. Hart and B. Connellan. Cost of prey capture, growth rate and ration size in pike, Esox lucius L., as functions of prey weight. Journal of Fish Biology, 25(3):279–292, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craik:1984:BMD

[680] J. C. A. Craik and S. M. Harvey. A biochemical method for distinguishing between the sexes of fishes by the presence of yolk protein in the blood. Journal of Fish Biology, 25(3):293–303, September 1984. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halvorsen:1984:EIB

[681] O. Halvorsen and K. Andersen. The ecological interaction between Arctic charr, Salvelinus alpinus (L.), and the plerocercoid stage of Diphyllobothrium ditremum. Journal of Fish Biology, 25(3):305–316, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitear:1984:SSS

[682] M. Whitear and A. K. Mittal. Surface secretions of the skin of Blennius (Lipophrys) pholis L. Journal of Fish Biology, 25(3):317–331, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Somasundaram:1984:SME

[683] B. Somasundaram, P. E. King, and S. E. Shackley. Some morphological effects of zinc upon the yolk-sac larvae of Clupea harengus L. Journal of Fish Biology, 25(3):333–343, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lou:1984:PIH

[684] Y. D. Lou and C. E. Purdom. Polyploidy induced by hydrostatic pressure in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(3):345–351, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulsford:1984:SST

[685] A. Pulsford, W. J. W. Morrow, and R. Fänge. Structural studies on the thymus of the dogfish, Scyliorhinus canicula L. Journal of Fish Biology, 25(3):353–360, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pool:1984:SEM

[686] D. Pool. A scanning electron microscope study of the life cycle of Bothriocephalus acheilognathi Yamaguti, 1934. Journal of Fish Biology, 25(3): 361–364, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wechsler:1984:FHA

[687] S. J. Wechsler. Fish health assessment: a preliminary report on the use of impression smears of skin mucus. Journal of Fish Biology, 25(3):365–370, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watson:1984:EAF

[688] D. J. Watson and E. K. Balon. Ecomorphological analysis of fish taxocenes in rainforest streams of northern Borneo. Journal of Fish Biology, 25(3):371–384, September 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lincoln:1984:SMT

[689] R. F. Lincoln and A. P. Scott. Sexual maturation in triploid rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(4):385– 392, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:1984:ESIb

[690] J. A. L. Fraser and E. R. Clark. The effects of a settled industrial domestic sewage works effluent from percolating filters on the embryo viability and hatching success of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(4):393–403, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:1984:RUC

[691] G. Walsh and G. J. FitzGerald. Resource utilization and coexistence of three species of sticklebacks (Gasterosteidae) in tidal salt-marsh pools.

Journal of Fish Biology, 25(4):405–420, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hay:1984:WLC

[692] D. E. Hay. Weight loss and change of condition factor during fixation of Pacific herring, Clupea harengus pallasi, eggs and larvae. Journal of Fish Biology, 25(4):421–433, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fletcher:1984:ERS

[693] D. J. Fletcher, D. J. Grove, R. A. Basimi, and A. Ghaddaf. Emptying rates of single and double meals of different food quality from the stomach of the dab, Limanda limanda (L.). Journal of Fish Biology, 25(4):435– 444, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Felts:1984:ITS

[694] P. A. Felts and A. G. Heath. Interactions of temperature and sublethal environmental copper exposure on the energy metabolism of bluegill, Lepomis macrochirus Rafinesque. Journal of Fish Biology, 25(4):445– 453, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feminella:1984:IDT

[695] J. W. Feminella and W. J. Matthews. Intraspecific differences in thermal tolerance of Etheostoma spectabile (Agassiz) in constant versus fluctuating environments. Journal of Fish Biology, 25(4):455–461, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adedire:1984:SIW

[696] C. O. Adedire and S. O. Oduleye. Stress-induced water permeability changes in the tropical cichlid, Oreochromis niloticus (Trewavas). Journal of Fish Biology, 25(4):463–471, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aho:1984:SPD

[697] J. M. Aho and C. R. Kennedy. Seasonal population dynamics of the nematode Cystidicoloides tenuissima (Zeder) from the River Swincombe, England. Journal of Fish Biology, 25(4):473–489, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:1984:PLC

[698] H. S. Gill and A. H. Weatherley. Protein, lipid and caloric contents of bluntnose minnow, Pimephales notatus Rafinesque, during growth at different temperatures. Journal of Fish Biology, 25(4):491–500, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kwain:1984:SSJ

[699] W. Kwain, R. W. McCauley, and J. A. MacLean. Susceptibility of starved, juvenile smallmouth bass, Micropterus dolomieui (Lacépède) to low pH. Journal of Fish Biology, 25(4):501–504, October 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chambers:1984:NNS

[700] J. Chambers. A new Nothobranchius species (Teleosti, Cyprinodontiformes, Aplocheilidae) from the Sudan. Journal of Fish Biology, 25(5): 505–513, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abu-Hakima:1984:SAR

[701] R. Abu-Hakima. Some aspects of the reproductive biology of Acanthopagrus spp. (family: Sparidae). Journal of Fish Biology, 25(5):515–526, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chubb:1984:RBE

[702] C. F. Chubb and I. C. Potter. The reproductive biology and estuarine movements of the gizzard shad, Nematalosa vlaminghi (Munro). Journal of Fish Biology, 25(5):527–543, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:1984:SIP

[703] O. N. Bauer. Soviet investigations on the population biology of fish parasites. Journal of Fish Biology, 25(5):545–550, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talbot:1984:EPP

[704] C. Talbot, P. J. Higgins, and A. M. Shanks. Effects of pre- and postprandial starvation on meal size and evacuation rate of juvenile Atlantic salmon, Salmo salar L. Journal of Fish Biology, 25(5):551–560, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Street:1984:EIS

[705] N. E. Street, A. E. Magurran, and T. J. Pitcher. The effects of increasing shoal size on handling time in goldfish, Carassius auratus L. Journal of Fish Biology, 25(5):561–566, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:1984:RSF

[706] R. M. G. Wells, V. Tetens, and A. L. Devries. Recovery from stress following capture and anaesthesia of Antarctic fish: haematology and blood chemistry. Journal of Fish Biology, 25(5):567–576, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1984:EEU

[707] A. E. Lewis and W. R. A. Muntz. The effects of external ultrasonic tagging on the swimming performance of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(5):577–585, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spoor:1984:ORL

[708] W. A. Spoor. Oxygen requirements of larvae of the smallmouth bass, Micropterus dolomieui Lacépède. Journal of Fish Biology, 25(5):587–592, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1984:ECP

[709] D. Burton, M. P. Burton, and D. R. Idler. Epidermal condition in postspawned winter flounder, Pseudopleuronectes americanus (Walbaum), maintained in the laboratory and after exposure to crude petroleum. Journal of Fish Biology, 25(5):593–606, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beaumont:1984:AGD

[710] W. R. C. Beaumont and R. H. K. Mann. The age, growth and diet of a freshwater population of the flounder, Platichthys flesus (L.), in southern England. Journal of Fish Biology, 25(5):607–616, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:1984:MWH

[711] L. P. Hansen, B. Jonsson, and K. B. Døving. Migration of wild and hatchery reared smolts of Atlantic salmon, Salmo salar L., through lakes. Journal of Fish Biology, 25(5):617–623, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:BRg

[712] Anonymous. Book reviews. Journal of Fish Biology, 25(5):625–627, November 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fletcher:1984:PGP

[713] D. J. Fletcher. Plasma glucose and plasma fatty acid levels of Limanda limanda (L.) in relation to season, stress, glucose loads and nutritional state. Journal of Fish Biology, 25(6):629–648, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dobson:1984:CGR

[714] S. H. Dobson and R. M. Holmes. Compensatory growth in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 25(6):649– 656, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mavares:1984:BAM

[715] R. N. Mavares and J. E. Pérez. Blood adaptations to marine and freshwater environments in fish of the family Sciaenidae (Perciformes). Journal of Fish Biology, 25(6):657–663, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1984:DDP

[716] C. R. Kennedy. The dynamics of a declining population of the acanthocephalan Acanthocephalus clavula in eels Anguilla anguilla in a small river. Journal of Fish Biology, 25(6):665–677, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcus:1984:SAR

[717] O. Marcus and K. Kusemiju. Some aspects of the reproductive biology of the clupeid Ilisha africana (Bloch) off the lagos Coast, Nigeria. Journal of Fish Biology, 25(6):679–689, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Secombes:1984:ASA

[718] C. J. Secombes, A. E. Lewis, L. M. Laird, E. A. Needham, and I. G. Priede. Agglutination of spermatozoa by autoantibodies in the rainbow trout, Salmo gairdneri. Journal of Fish Biology, 25(6):691–696, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogers:1984:AUT

[719] S. C. Rogers, D. W. Church, A. H. Weatherley, and D. G. Pincock. An automated ultrasonic telemetry system for the assessment of locomotor activity in free-ranging rainbow trout, Salmo gairdneri, Richardson. Journal of Fish Biology, 25(6):697–710, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:1984:ISE

[720] P. Thomas. Influence of some environmental variables on the ascorbic acid status of mullet, Mugil cephalus L., tissues. I. Effect of salinity, capture-stress, and temperature. Journal of Fish Biology, 25(6):711–720, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penczak:1984:FPO

[721] T. Penczak and M. Moliński. Fish production in oued sebaou, a seasonal river in North Algeria. Journal of Fish Biology, 25(6):723–732, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:1984:EFA

[722] W. W. Taylor and M. H. Freeberg. Effect of food abundance on larval lake whitefish, Coregonus clupeaformis Mitchill, growth and survival. Journal of Fish Biology, 25(6):733–741, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1984:CE

[723] Anonymous. Change of Editor. Journal of Fish Biology, 25(6):743, December 1984. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1985:GRL

[724] R. R. C. Edwards. Growth rates of Lutjanidae (snappers) in tropical Australian waters. Journal of Fish Biology, 26(1):1–4, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Sabti:1985:CSB

[725] Kabil Al-Sabti. Chromosomal studies by blood leukocyte culture technique on three salmonids from Yugoslavian waters. Journal of Fish Biology, 26(1):5–12, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Sabti:1985:FCA

[726] Kabil Al-Sabti. Frequency of chromosomal aberrations in the rainbow trout, Salmo gairdneri Rich., exposed to five pollutants. Journal of Fish Biology, 26(1):13–19, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britz:1985:ODC

[727] J. Britz, J. G. Van As, and J. E. Saayman. Occurrence and distribution of Clinostomum tilapiae Ukoli, 1966 and Euclinostomum heterostomum (Rudolphi, 1809) metacercarial infections of freshwater fish in Venda and Lebowa, southern Africa. Journal of Fish Biology, 26(1):21–28, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:1985:APA

[728] W. C. Clarke, H. Lundqvist, and L.-O. Eriksson. Accelerated photoperiod advances seasonal cycle of sea water adaptation in juvenile Baltic salmon, Salmo salar L. Journal of Fish Biology, 26(1):29–35, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haraldstad:1985:DES

[729] Ø. Haraldstad, L. A. Vøllestad, and B. Jonsson. Descent of European silver eels, Anguilla anguilla L., in a Norwegian watercourse. Journal of Fish Biology, 26(1):37–41, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korwin-Kossakowski:1985:ETS

[730] M. Korwin-Kossakowski and B. Jezierska. The effect of temperature on survival of carp fry, Cyprinus carpio L., in acidic water. Journal of Fish Biology, 26(1):43–47, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maitipe:1985:SBZ

[731] P. Maitipe and S. S. De Silva. Switches between zoophagy, phytophagy and detritivory of Sarotherodon mossambicus (Peters) populations in twelve man-made Sri Lankan lakes. Journal of Fish Biology, 26(1):49– 61, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McEvoy:1985:DOC

[732] L.-A. McEvoy. Double ovulatory cycles in some captive turbot, Scophthalmus maximus L. Journal of Fish Biology, 26(1):63–66, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1985:DMN

[733] A. D. Woodhead, E.-H. Cao, and T. Melgar. Distribution of N-methyl-N-nitrosourea in the Amazon molly, Poecilla formosa (Girard), after a single intraperitoneal injection*. Journal of Fish Biology, 26(1):67–75, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:1985:EMT

[734] G. M. Weber and Cheng-Sheng Lee. Effects of 17α-methyl-testosterone on spermatogenesis and spermiation in the grey mullet, Mugil cephalus L. Journal of Fish Biology, 26(1):77–84, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:CEa

[735] Anonymous. Change of Editor. Journal of Fish Biology, 26(1):85, January 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macintosh:1985:HSR

[736] D. J. Macintosh, T. J. Varghese, and G. P. Rao Satyanarayana. Hormonal sex reversal of wild-spawned tilapia in India. Journal of Fish Biology, 26(2):87–94, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulta:1985:GDB

[737] G. Hulta, G. Wohlfarth, and R. Moav. Genetic differences between the Chinese and European races of the common carp, Cyprinus carpio L. IV. Effects of sexual maturation on growth patterns. Journal of Fish Biology, 26(2):95–103, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:1985:OPA

[738] Peter Davis and Alwyne Wheeler. The occurrence of Pampus argenteus (Euphrasen, 1788), (Osteichthyes, Perciformes, Stromateoidei, Stromateidae) in the North Sea. Journal of Fish Biology, 26(2):105–109, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:1985:ELB

[739] S. M. Adams and R. B. McLean. Estimation of largemouth bass, Micropterus salmoides Lacépède, growth using the liver somatic index and physiological variables. Journal of Fish Biology, 26(2):111–126, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soofiani:1985:AMS

[740] N. M. Soofiani and I. G. Priede. Aerobic metabolic scope and swimming performance in juvenile cod, Gadus morhua L. Journal of Fish Biology, 26(2):127–138, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolle:1985:CTS

[741] L. D. La Bolle, Jr., H. W. Li, and B. C. Mundy. Comparison of two samplers for quantitatively collecting larval fishes in upper littoral habitats. Journal of Fish Biology, 26(2):139–146, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miles:1985:RXR

[742] M. S. Miles, C. Talbot, and J. E. Thorpe. Reading X-ray microtags. Journal of Fish Biology, 26(2):147–151, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Higgins:1985:XRM

[743] P. J. Higgins. X-ray microtags for identification of individuals in studies of fish energetics. Journal of Fish Biology, 26(2):153–159, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nordlie:1985:ORS

[744] Frank G. Nordlie. Osmotic regulation in the Sheepshead Minnow Cyprinodon variegatus Lacépède. Journal of Fish Biology, 26(2):161–170, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaushik:1985:ESS

[745] S. J. Kaushik, K. Dabrowski, and P. Luquet. Experimental studies on some trophic relationships in juvenile pike, Esox lucius L. Journal of Fish Biology, 26(2):171–180, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bell:1985:EDP

[746] M. V. Bell, R. J. Henderson, B. J. S. Pirie, and J. R. Sargent. Effects of dietary polyunsaturated fatty acid deficiencies on mortality, growth and gill structure in the turbot, Scophthalmus maximus. Journal of Fish Biology, 26(2):181–191, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lund:1985:ART

[747] M. Lund and T. G. Heggberget. Avoidance response of two-year-old rainbow trout, Salmo gairdneri Richardson, to air-supersaturated water: hydrostatic compensation. Journal of Fish Biology, 26(2):193–200, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1985:DSB

[748] A. D. Pickering, A. J. Strong, and J. Pollard. Differences in the susceptibility of brown trout, Salmo trutta L., and American brook trout, Salvelinus fontinalis (Mitchill), to infestation by the peritrich ciliate, Scyphidia sp. Journal of Fish Biology, 26(2):201–208, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1985:PIC

[749] P. C. Blaxhall and P. R. Sheard. Preliminary investigation of the characteristics of fish lymphocytes separated on a Percoll discontinuous gradient. Journal of Fish Biology, 26(2):209–216, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:1985:PDO

[750] T. B. Ng, P. P. L. Tam, and N. Y. S. Woo. A pharmacological dose of 17 β-oestradiol produces minimal metabolic stimulation but evokes hepatic necrosis in the immature female serranid Epinephelus akaara. Journal of Fish Biology, 26(2):217–222, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bamber:1985:DSS

[751] R. N. Bamber and P. A. Henderson. Diplostomiasis in sand smelt, Atherina presbyter Cuvier, from the Fleet, Dorset and its use as a population indicator. Journal of Fish Biology, 26(2):223–229, February 1985. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:BRa

[752] Anonymous. Book reviews. Journal of Fish Biology, 26(2):231–235, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:CEb

[753] Anonymous. Change of Editor. Journal of Fish Biology, 26(2):237, February 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medved:1985:GES

[754] R. J. Medved. Gastric evacuation in the sandbar shark, Carcharhinus plumbeus. Journal of Fish Biology, 26(3):239–253, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindsay:1985:CEB

[755] G. J. H. Lindsay and G. W. Gooday. Chitinolytic enzymes and the bacterial microflora in the digestive tract of cod, Gadus morhua. Journal of Fish Biology, 26(3):255–265, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rivlin:1985:SMP

[756] K. Rivlin, J. W. Rachlin, and G. Dale. A simple method for the preparation of fish chromosomes applicable to field work, teaching and banding. Journal of Fish Biology, 26(3):267–272, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:1985:SRH

[757] P. C. Young and R. B. Martin. Sex ratios and hermaphroditism in nemipterid fish from northern Australia. Journal of Fish Biology, 26(3): 273–287, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matallanas:1985:FRN

[758] J. Matallanas. First record of the North Atlantic Morid fish, Laemonema latifrons Holt & Byrne, 1908, in the Mediterranean. Journal of Fish Biology, 26(3):289–290, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birks:1985:MTJ

[759] E. K. Birks, R. D. Ewing, and A. R. Hemmingsen. Migration tendency in juvenile steelhead trout, Salmo gairdneri Richardson, injected with thyroxine and thiourea. Journal of Fish Biology, 26(3):291–300, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollock:1985:RCY

[760] B. R. Pollock. The reproductive cycle of yellowfin bream, Acanthopagms australis (Günther), with particular reference to protandrous sex inversion. Journal of Fish Biology, 26(3):301–311, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sparholt:1985:PSG

[761] H. Sparholt. The population, survival, growth, reproduction and food of Arctic charr, Salvelinus alpinus (L.), in four unexploited lakes in Greenland. Journal of Fish Biology, 26(3):313–330, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saclauso:1985:IGS

[762] C. A. Saclauso. Interaction of growth with social behaviour in Tilapia zilli raised in three different temperatures. Journal of Fish Biology, 26 (3):331–337, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grove:1985:GER

[763] D. J. Grove, M. A. Moctezuma, H. R. J. Flett, J. S. Foott, T. Watson, and M. W. Flowerdew. Gastric emptying and the return of appetite in juvenile turbot, Scophthalmus maximus L., fed on artificial diets. Journal of Fish Biology, 26(3):339–354, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lewis:1985:DIS

[764] D. H. Lewis, R. J. Tarpley, J. E. Marks, and R. F. Sis. Drug induced structural changes in olfactory organ of channel catfish Ictalurus punctatus, Rafinesque. Journal of Fish Biology, 26(3):355–358, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Twomey:1985:NRT

[765] E. Twomey and P. Byrne. A new record for the tarpon, Tarpon atlanticus Valenciennes (Osteichthyes-Elopiformes-Elopidae), in the eastern North Atlantic. Journal of Fish Biology, 26(3):359–362, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skurdal:1985:ITN

[766] J. Skurdal and R. Andersen. Influence of temperature on number of circuli of first year scales of brown trout, Salmo trutta L. Journal of Fish Biology, 26(3):363–366, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliassen:1985:FRG

[767] J-E. Eliassen and M. Jobling. Food of the roughhead grenadier, Macrourus berglax, lacépède in North Norwegian waters. Journal of Fish Biology, 26(3):367–376, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:BRb

[768] Anonymous. Book reviews. Journal of Fish Biology, 26(3):377–380, March 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Narayan:1985:HPG

[769] A. Narayan, K. C. George, and A. D. Diwan. Histology of the pituitary gland of the grey mullet, Mugil cephalus L. Journal of Fish Biology, 26(4):381–390, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larsson:1985:PMS

[770] P.-O. Larsson. Predation on migrating smolt as a regulating factor in Baltic salmon, Salmo salar L., populations. Journal of Fish Biology, 26(4):391–397, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romand:1985:FBA

[771] R. Romand. Feeding biology of Aplocheilichthys normani, Ahl, a small Cyprinodontidae from West Africa. Journal of Fish Biology, 26(4):399– 410, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gunn:1985:FFH

[772] J. S. Gunn and N. E. Milward. The food, feeding habits and feeding structures of the whiting species Sillago sihama (Forsskål) and Sillago analis Whitley from Townsville, North Queensland, Australia. Journal of Fish Biology, 26(4):411–427, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruglys:1985:SIR

[773] M. P. Ruglys. The secondary immune response of young carp, Cyprinus carpio L., following injection of cortisol. Journal of Fish Biology, 26(4): 429–434, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodward:1985:ETS

[774] J. J. Woodward and L. S. Smith. Exercise training and the stress response in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 26(4):435–447, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1985:FAG

[775] B. A. Henderson. Factors affecting growth and recruitment of yellow perch, Perca flavescens Mitchill, in South Bay, Lake Huron. Journal of Fish Biology, 26(4):449–458, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bedawi:1985:RCP

[776] R. M. Bedawi. Recruitment control and production of market-size Oreochromis niloticus with the predator Lates niloticus L. in the Sudan. Journal of Fish Biology, 26(4):459–464, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatakis:1985:TAS

[777] T. A. Tatakis and R. G. Hart. Thermal acclimation in selected tissues of the fathead minnow, Pimephales promelas Rafinesque. Journal of Fish Biology, 26(4):465–473, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dick:1985:CCB

[778] P. T. Dick and D. G. Dixon. Changes in circulating blood cell levels of rainbow trout, Salmo gairdneri Richardson, following acute and chronic exposure to copper. Journal of Fish Biology, 26(4):475–481, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Secombes:1985:LTE

[779] C. J. Secombes, E. A. Needham, L. M. Laird, A. E. Lewis, and I. G. Priede. The long-term effects of auto-immunologically induced granulomas on the testes of rainbowtrout, Salmo gairdneri Richardson. Journal of Fish Biology, 26(4):483–489, April 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liley:1985:ISB

[780] N. R. Liley and E. S. P. Tan. The induction of spawning behaviour in Puntius gonionotus (Bleeker) by treatment with prostaglandin PGF_{2α}. Journal of Fish Biology, 26(5):491–502, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spataru:1985:FCF

[781] P. Spataru and M. Gophen. Food composition and feeding habits of Astatotilapia flaviijosephi (Lortet) in Lake Kinneret (Israel). Journal of Fish Biology, 26(5):503–507, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Price:1985:GSR

[782] D. J. Price. Genetics of susceptibility and resistance to disease in fishes. Journal of Fish Biology, 26(5):509–519, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vollestad:1985:ADG

[783] L. A. Vøllestad. Age determination and growth of yellow eels, Anguilla anguilla (L.), from a brackish water, Norway. Journal of Fish Biology, 26(5):521–525, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mauchline:1985:TDD

[784] J. Mauchline and J. D. M. Gordon. Trophic diversity in deep-sea fish. Journal of Fish Biology, 26(5):527–535, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berg:1985:ADE

[785] R. Berg. Age determination of eels, Anguilla anguilla (L): comparison of field data with otolith ring patterns. Journal of Fish Biology, 26(5):537– 544, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Basimi:1985:GER

[786] R. A. Basimi and D. J. Grove. Gastric emptying rate in Pleuronectes platessa L. Journal of Fish Biology, 26(5):545–552, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:1985:FAS

[787] L. P. Hansen and P. Pethon. The food of Atlantic salmon, Salmo salar L., caught by long–line in northern Norwegian waters. Journal of Fish Biology, 26(5):553–562, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Opuszynski:1985:ICG

[788] K. Opuszynski, J. V. Shireman, F. J. Aldridge, and R. Rottmann. Intensive culture of grass carp and hybrid grass carp larvae. Journal of Fish Biology, 26(5):563–573, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartley:1985:CTF

[789] S. E. Hartley and M. T. Horne. Cytogenetic techniques in fish genetics. Journal of Fish Biology, 26(5):575–582, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wahbeh:1985:RBG

[790] M. I. Wahbeh and A. Ajiad. Reproductive biology and growth of the goatfish, Parupeneus barberinus (Lacépède), in Aqaba, Jordan. Journal of Fish Biology, 26(5):583–590, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:1985:MMD

[791] D. P. F. King. Morphological and meristic differences among spawning aggregations of north-east Atlantic herring, Clupea harengus L. Journal of Fish Biology, 26(5):591–607, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kadmon:1985:SGE

[792] G. Kadmon, Z. Yaron, and H. Gordin. Sequence of gonadal events and oestradiol levels in Sparus aurata (L.) under two photoperiod regimes. Journal of Fish Biology, 26(5):609–620, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1985:AAD

[793] R. H. K. Mann and B. Steinmetz. On the accuracy of age-determination using scales from rudd, Scardinius erythrophthalmus (L.), of known age. Journal of Fish Biology, 26(5):621–628, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neave:1985:DLR

[794] D. A. Neave. The dorsal light reactions of larval and metamorphosing flatfish. Journal of Fish Biology, 26(5):629–640, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:BRc

[795] Anonymous. Book reviews. Journal of Fish Biology, 26(5):641–646, May 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miyazaki:1985:HBB

[796] T. Miyazaki, J. A. Plumb, Y. P. Li, and R. T. Lovell. Histopathology of broken-back syndrome in channel catfish. Journal of Fish Biology, 26(6):647–655, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tan:1985:HSH

[797] J. D. Tan. A histological study of the hypophysial-gonadal system during sexual maturation and spawning in the milkfish, Chanos chanos (Forsskål). Journal of Fish Biology, 26(6):657–668, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:1985:FBL

[798] T. L. O. Davis. The food of barramundi, Lates calcarifer (Bloch), in coastal and inland waters of Van Diemen Gulf and the Gulf of Carpentaria, Australia. Journal of Fish Biology, 26(6):669–682, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofer:1985:DPD

[799] R. Hofer and A. Nasir Uddin. Digestive processes during the development of the roach, Rutilus rutilus L. Journal of Fish Biology, 26(6):683–689, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gyllensten:1985:GSF

[800] U. Gyllensten. The genetic structure of fish: differences in the intraspecific distribution of biochemical genetic variation between marine, anadromous, and freshwater species. Journal of Fish Biology, 26(6):691– 699, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OGrady:1985:CPC

[801] K. T. O’Grady and P. B. Spillett. A comparison of the pond culture of carp, Cyprinus carpio L., in Britain and mainland Europe. Journal of Fish Biology, 26(6):701–714, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buddington:1985:DSL

[802] R. K. Buddington. Digestive secretions of lake sturgeon, Acipenser fulvescens, during early development. Journal of Fish Biology, 26(6):715– 723, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Railo:1985:ESB

[803] E. Railo, M. Nikinmaa, and A. Soivio. Effects of sampling on blood parameters in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 26(6):725–732, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heap:1985:CTC

[804] S. P. Heap, P. W. Watt, and G. Goldspink. Consequences of thermal change on the myofibrillar ATPase of five freshwater teleosts. Journal of Fish Biology, 26(6):733–738, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Speed:1985:FPR

[805] P. Speed and G. B. Pauley. Feasibility of protecting rainbow trout, Salmo gairdneri Richardson, by immunizing against the eye fluke, Diplostomum spathaceum. Journal of Fish Biology, 26(6):739–744, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:1985:CSS

[806] T. G. Pottinger and A. D. Pickering. Changes in skin structure associated with elevated androgen levels in maturing male brown trout, Salmo trutta L. Journal of Fish Biology, 26(6):745–753, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:1985:SCS

[807] A. White and T. C. Fletcher. Seasonal changes in serum glucose and condition of the plaice, Pleuronectes platessa L. Journal of Fish Biology, 26(6):755–764, June 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McLean:1985:TSD

[808] R. B. McLean, J. S. Griffith, and M. V. McGee. Threadfin shad, Dorosoma petenense Günther, mortality: causes and ecological implications in a South-eastern United States reservoir. Journal of Fish Biology, 27(1): 1–12, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1985:GMA

[809] R. R. C. Edwards, A. Bakhader, and S. Shaher. Growth, mortality, age composition and fisheries yields of fish from the Gulf of Aden. Journal of Fish Biology, 27(1):13–21, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:1985:BGS

[810] W. S. Grant. Biochemical genetic stock structure of the southern African anchovy, Engraulis capensis Gilchrist. Journal of Fish Biology, 27(1):

23–29, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoker:1985:PGL

[811] P. W. Stoker, J. R. Larsen, G. M. Booth, and M. L. Lee. Pathology of gill and liver tissues from two genera of fishes exposed to two coal-derived materials. Journal of Fish Biology, 27(1):31–46, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:1985:ISE

[812] P. Thomas, M. B. Bally, and J. M. Neff. Influence of some environmental variables on the ascorbic acid status of mullet, Mugil cephalus L., tissues: II. Seasonal fluctuations and biosynthetic ability. Journal of Fish Biology, 27(1):47–57, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrews:1985:RLK

[813] M. J. Andrews and Alwyne Wheeler. Rare and little-known fishes in the Thames Estuary. Journal of Fish Biology, 27(1):59–71, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purdom:1985:GER

[814] C. E. Purdom, D. Thompson, and Y. D. Lou. Genetic engineering in rainbow trout, Salmo gairdnerii Richardson, by the suppression of meiotic and mitotic metaphase. Journal of Fish Biology, 27(1):73–79, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karlsson-Norrgren:1985:CIC

[815] L. Karlsson-Norrgren, P. Runn, C. Haux, and L. Förlin. Cadmiuminduced changes in gill morphology of zebrafish, Brachydanio rerio (Hamilton-Buchanan), and rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 27(1):81–95, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shieh:1985:VAS

[816] H. S. Shieh. Vaccination of Atlantic salmon, Salmo salar L., against furunculosis with protease of an avirulent strain of Aeromonas salmonicida. Journal of Fish Biology, 27(1):97–101, July 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bitterlich:1985:DEP

[817] G. Bitterlich. Digestive enzyme pattern of two stomachless filter feeders, silver carp, Hypophthalmichthys molitrix Val., and bighead carp, Aristichthys nobilis Rich. Journal of Fish Biology, 27(2):103–112, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Northcote:1985:LRM

[818] T. G. Northcote and F. J. Ward. Lake resident and migratory smelt, Retropinna retropinna (Richardson), of the lower Waikato River system, New Zealand. Journal of Fish Biology, 27(2):113–129, August 1985. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Erickson:1985:OCT

[819] D. L. Erickson, M. J. Harris, and G. D. Grossman. Ovarian cycling of tilefish, Lopholatilus chamaeleonticeps Goode and Bean, from the South Atlantic Bight, U.S.A. Journal of Fish Biology, 27(2):131–146, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wahbeh:1985:FFH

[820] M. I. Wahbeh and A. Ajiad. The food and feeding habits of the goatfish, Parupeneus barberinus (Lacépède), from Aqaba, Jordan. Journal of Fish Biology, 27(2):147–154, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eccles:1985:ETM

[821] D. H. Eccles. The effect of temperature and mass on routine oxygen consumption in the South African cyprinid fish Barbus aeneus Burchell. Journal of Fish Biology, 27(2):155–165, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blazer:1985:PAE

[822] V. S. Blazer, E. B. Shotts, and W. D. Waltman. Pathology associated with Edwardsiella ictaluri in catfish, Ictalurus punctatus Rafinesque, and Danio devario (Hamilton-Buchanan, 1822). Journal of Fish Biology, 27 (2):167–175, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:1985:GBC

[823] K. B. Davis, P. Torrance, N. C. Parker, and M. A. Suttle. Growth, body composition and hepatic tyrosine aminotransferase activity in cortisol-fed channel catfish, Ictalurus punctatus Rafinesque. Journal of Fish Biology, 27(2):177–184, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ventura:1985:HIM

[824] M. T. Ventura and I. Paperna. Histopathology of Ichthyophthirim multifiliis infections in fishes. Journal of Fish Biology, 27(2):185–203, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:BRd

[825] Anonymous. Book reviews. Journal of Fish Biology, 27(2):205–213, August 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stein:1985:OME

[826] D. L. Stein and C. E. Bond. Observations on the morphology, ecology, and behaviour of Bathylychnops exilis Cohen. Journal of Fish Biology, 27 (3):215–228, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dalzell:1985:SAR

[827] P. Dalzell. Some aspects of the reproductive biology of Spratelloides gracilis (Schlegel) in the Ysabel Passage, Papua New Guinea. Journal of Fish Biology, 27(3):229–237, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bettoli:1985:TPH

[828] P. W. Bettoli, W. H. Neill, and S. W. Kelsch. Temperature preference and heat resistance of grass carp, Ctenopharyngodon idella (Valenciennes), bighead carp, Hypophthalmichthys nobilis (Gray), and their F_{1} hybrid. Journal of Fish Biology, 27(3):239–247, September 1985. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golani:1985:BIP

[829] D. Golani and A. Ben-Tuvia. The biology of the Indo–Pacific squirrelfish, Sargocentron rubrum (Forsskål), a Suez Canal migrant to the eastern Mediterranean. Journal of Fish Biology, 27(3):249–258, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kori-Slakpere:1985:HCC

[830] O. Kori-Slakpere. Haematological characteristics of Clarias isheriensis Sydenham. Journal of Fish Biology, 27(3):259–263, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanGlnkel:1985:TCA

[831] F. W. van Glnkel, W. B. van Muiswinkel, B. Merchant, E. F. Lizzio, O. W. Dixon, and D. P. Anderson. Temperature comparisons for antibody production in vitro by plaque-forming cells from trout, Salmo gairdneri (Richardson), and mice. Journal of Fish Biology, 27(3):265– 272, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendee:1985:HAK

[832] J. C. Hendee and G. Mueller. Histology of the adenohypophysis of the kelp greenling, Hexagrammm decagrammus (Hexagrammidae). Journal of Fish Biology, 27(3):273–276, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konagai:1985:LOM

[833] M. Konagai and M. A. Rimmer. Larval ontogeny and morphology of the fire-tailed gudgeon, Hypseleotris galii (Eleotridae). Journal of Fish Biology, 27(3):277–283, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Formacion:1985:CSO

[834] M. J. Formacion and H. Uwa. Cytogenetic studies on the origin and species differentiation of the Philippine medaka, Oryzias luzonensis. Journal of Fish Biology, 27(3):285–291, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bucke:1985:HSF

[835] D. Bucke and S. W. Feist. Histopathology of some fish tumours from British waters, including an olfactory glioma in coalfish, Pottachius virens L. Journal of Fish Biology, 27(3):293–305, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lester:1985:GMP

[836] R. J. G. Lester and R. A. Watson. Growth, mortality, parasitism, and potential yields of two Priacanthus species in the South China Sea. Journal of Fish Biology, 27(3):307–318, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skarphedinsson:1985:IPS

[837] O. Skarphedinsson, V. J. Bye, and A. P. Scott. The influence of photoperiod on sexual development in underyearling rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 27(3):319–326, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Appler:1985:EHR

[838] H. N. Appler. Evaluation of Hydrodictyon reticulatum as protein source in feeds for Oreochromis (Tilapia) niloticus and Tilapia zillii. Journal of Fish Biology, 27(3):327–334, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mason:1985:FWP

[839] J. C. Mason. The fecundity of the walleye pollock, Theragra chalcogramma (Pallas), spawning in Canadian waters. Journal of Fish Biology, 27(3):335–346, September 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Philipp:1985:CAF

[840] D. P. Philipp, W. F. Childers, and G. S. Whitt. Correlations of allele frequencies with physical and environmental variables for populations of largemouth bass, Micropterus salmoides (Lacépède). Journal of Fish Biology, 27(4):347–365, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradley:1985:ART

[841] R. W. Bradley, C. DuQuesnay, and J. B. Sprague. Acclimation of rainbow trout, Salmo gairdneri Richardson, to zinc: kinetics and mechanism of enhanced tolerance induction. Journal of Fish Biology, 27(4):367–379, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:1985:PMW

[842] E. R. Clark, J. P. Harman, and J. R. M. Forster. Production of metabolic and waste products by intensively farmed rainbow trout, Salmo gaivdnevi Richardson. Journal of Fish Biology, 27(4):381–393, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crouse-Eisnor:1985:SRB

[843] R. A. Crouse-Eisnor, D. K. Cone, and P. H. Odense. Studies on relations of bacteria with skin surface of Carassius auratus L. and Poeciloia reticulata. Journal of Fish Biology, 27(4):395–402, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Avenant:1985:ODD

[844] A. Avenant and J. G. van As. Occurrence and distribution of Dolops ranarum Stuhlmann, ectoparasite of freshwater fish in the Transvaal, South Africa. Journal of Fish Biology, 27(4):403–416, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vesey:1985:BBG

[845] G. Vesey and T. E. Langford. The biology of the black goby, Gobius niger L. in an English south-coast bay. Journal of Fish Biology, 27(4): 417–429, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coates:1985:FYE

[846] David Coates. Fish yield estimates for the Sepik River, Papua New Guinea, a large floodplain system east of ‘Wallace’s Line’. Journal of Fish Biology, 27(4):431–443, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:1985:CSP

[847] C. M. Johnson, M. F. Tatner, and M. T. Horne. Comparison of the surface properties of seven strains of a fish pathogen, Aeromonas salmonicida. Journal of Fish Biology, 27(4):445–458, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deMorais:1985:GRT

[848] L. Tito de Morais. Growth and respiration of two pleuronectiform juveniles from a western Mediterranean bay. Journal of Fish Biology, 27 (4):459–468, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wickins:1985:GVI

[849] J. F. Wickins. Growth variability in individually confined elvers, Anguilla anguilla (L.). Journal of Fish Biology, 27(4):469–478, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1985:ESR

[850] D. A. Wright and F. D. Martin. The effect of starvation on RNA: DNA ratios and growth of larval striped bass, Morone saxatalis. Journal of Fish Biology, 27(4):479–485, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:1985:SED

[851] L. M. Garcia and I. R. Adelman. An in situ estimate of daily food consumption and alimentary canal evacuation rates of common carp, Cyprinus carpio L. Journal of Fish Biology, 27(4):487–493, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1985:BRF

[852] K. A. Jones and T. J. Hara. Behavioural responses of fishes to chemical cues: results from a new bioassay. Journal of Fish Biology, 27(4):495– 504, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Basimi:1985:EDF

[853] R. A. Basimi and D. J. Grove. Estimates of daily food intake by an inshore population of Pleuronectes platessa L. off eastern Anglesey, North Wales. Journal of Fish Biology, 27(4):505–520, October 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:1985:EMA

[854] Miran W. Aprahamian. The effect of the migration of Alosa fallax fallax (Laécpède) into fresh water, on branchial and gut parasites. Journal of Fish Biology, 27(5):521–532, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galat:1985:HCG

[855] D. L. Galat, G. Post, T. J. Keefe, and G. R. Bouck. Histological changes in the gill, kidney and liver of Lahontan cutthroat trout, Salmo clarki henshawi, living in lakes of different salinity-alkalinity. Journal of Fish Biology, 27(5):533–552, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pfeiffer:1985:ESH

[856] W. Pfeiffer, U. Walz, R. Wolf, and U. Mangold-Wernado. Effects of steroid hormones and other substances on alarm substance cells and mucous cells in the epidermis of the European minnow, Phoxinus phoxinus (L.), and other ostariophysi (pisces). Journal of Fish Biology, 27(5): 553–570, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karlsson-Norrgren:1985:CDF

[857] L. Karlsson-Norrgren and P. Runn. Cadmium dynamics in fish: pulse studies with ^{109}Cd in female zebrafish, Brachydanio rerio. Journal of Fish Biology, 27(5):571–581, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1985:ERD

[858] J. F. Leatherland and C. Y. Cho. Effect of rearing density on thyroid and interrenal gland activity and plasma and hepatic metabolite levels in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 27 (5):583–592, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1985:AMB

[859] A. D. Woodhead and N. Armstrong. Aspects of the mating behaviour of male mollies (Poecilia spp.). Journal of Fish Biology, 27(5):593–601, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Falcao:1985:CCA

[860] J. das Neves Falcão and L. A. C. Bertollo. Chromosome characterization in Acestrorhynchinae and Cynopotaminae (Pisces, Characidae). Journal of Fish Biology, 27(5):603–610, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1985:CCI

[861] A. D. Pickering and T. G. Pottinger. Cortisol can increase the susceptibility of brown trout, Salmo trutta L., to disease without reducing the white blood cell count. Journal of Fish Biology, 27(5):611–619, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1985:BCN

[862] S. S. De Silva. Body condition and nutritional ecology of Oreochromis mossambicus (Pisces, Cichlidae) populations of man-made lakes in Sri Lanka. Journal of Fish Biology, 27(5):621–633, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Degani:1985:RBG

[863] G. Degani and M. Lee-Gallagher. The relationship between growth, food conversion and oxygen consumption in developed and underdeveloped American eels, Anguilla rostrata Lesueur. Journal of Fish Biology, 27 (5):635–641, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Steele:1985:LBT

[864] C. W. Steele. Latent behavioural toxicity of copper to sea catfish, Arius felis, and sheepshead, Archosargus probatocephalus. Journal of Fish Biology, 27(5):643–654, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crozier:1985:OFF

[865] W. W. Crozier. Observations on the food and feeding of the anglerfish, Lophim piscatorius L., in the northern Irish Sea. Journal of Fish Biology, 27(5):655–665, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacArthur:1985:ALM

[866] J. I. MacArthur, A. W. Thomson, and T. C. Fletcher. Aspects of leucocyte migration in the plaice, Pleuronectes platessa L. Journal of Fish Biology, 27(5):667–676, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:BRe

[867] Anonymous. Book reviews. Journal of Fish Biology, 27(5):677–681, November 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linkowski:1985:PBM

[868] T. B. Linkowski. Population biology of the myctophid fish Gymnoscopelus nicholsi (Gillbert, 1911) from the western South Atlantic. Journal of Fish Biology, 27(6):683–698, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiener:1985:SCC

[869] J. G. Wiener, R. A. Jacobson, P. S. Schmidt, and P. R. Heine. Serum calcium concentrations in white sucker, Catostomus commersoni Lacépède, and bluegill, Lepomis macrochirus Rafinesque, in northern Wisconsin lakes: relation to pH and waterborne calcium. Journal of Fish Biology, 27(6):699–709, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:1985:MPS

[870] N. A. Hvidsten. Mortality of pre-smolt Atlantic salmon, Salmo salar L., and brown trout, Salmo trutta L., caused by fluctuating water levels in the regulated River Nidelva, central Norway. Journal of Fish Biology, 27 (6):711–718, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1985:AME

[871] N. R. Merrett and L. Saldanha. Aspects of the morphology and ecology of some unusual deep-sea eels (Synaphobranchidae, Derichthyidae and Nettastomatidae) from the eastern North Atlantic. Journal of Fish Biology, 27(6):719–747, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1985:CES

[872] P. C. Blaxhall and K. Hood. Cytochemical enzyme staining of fish lymphocytes separated on a Percoll gradient. Journal of Fish Biology, 27 (6):749–755, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collvin:1985:ECG

[873] L. Collvin. Effects of copper on growth and starvation in perch, Percafluviatilis L. Journal of Fish Biology, 27(6):757–764, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Basimi:1985:SFG

[874] R. A. Basimi and D. J. Grove. Studies on feeding, growth and production of a recruited inshore population of Pleuronectes platessa (L.) at East Anglesey, North Wales. Journal of Fish Biology, 27(6):765–783, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Street:1985:GSP

[875] N. E. Street and P. J. B. Hart. Group size and patch location by the stoneloach, Noemacheilus barbatulus, a non-visually foraging predator. Journal of Fish Biology, 27(6):785–792, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larrick:1985:HRA

[876] S. R. Larrick, D. S. Cherry, and J. Cairns, Jr. Homeostatic regulation of acetazolamide-sensitive esterase activity in the bluegill sunfish, Lepomis macrochirus, following acute cold shock. Journal of Fish Biology, 27(6): 793–804, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berg:1985:FNA

[877] O. K. Berg. The formation of non-anadromous populations of Atlantic salmon, Salmo salar L., in Europe. Journal of Fish Biology, 27(6):805– 815, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1985:RND

[878] D. L. Barber and J. E. Mills Westermann. Reappraisal of nuclear DNA content of rodlet cells compared to several cell types from some freshwater teleosts, using two methods of microdensitometry. Journal of Fish Biology, 27(6):817–826, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1985:NPU

[879] N. Giles and R. S. Phillips. A note on the production and use of antisera for the detection of part-digested sticklebacks in predator stomachcontent samples. Journal of Fish Biology, 27(6):827–829, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1985:LS

[880] Anonymous. List of speakers. Journal of Fish Biology, 27(sA):v–vi, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Margetts:1985:P

[881] A. R. Margetts. Preface. Journal of Fish Biology, 27(sA):vii, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1985:KA

[882] R. W. Edwards. Keynote address. Journal of Fish Biology, 27(sA):1–7, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chadwick:1985:FRP

[883] E. M. P. Chadwick. Fundamental research problems in the management of Atlantic salmon, Salmo salar L., in Atlantic Canada. Journal of Fish Biology, 27(sA):9–25, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egglishaw:1985:FGP

[884] H. J. Egglishaw and P. E. Shackley. Factors governing the production of juvenile Atlantic salmon in Scottish streams. Journal of Fish Biology, 27 (sA):27–33, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1985:PDM

[885] J. M. Elliott. Population dynamics of migratory trout, Salmo trutta, in a Lake District stream, 1966–83, and their implications for fisheries management. Journal of Fish Biology, 27(sA):35–43, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1985:SSR

[886] D. J. Solomon. Salmon stock and recruitment, and stock enhancement. Journal of Fish Biology, 27(sA):45–57, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zalewski:1985:FLG

[887] M. Zalewski, P. Frankiewicz, and B. Brewinska. The factors limiting growth and survival of brown trout, Salmo trutta m. fario L. introduced to different types of streams. Journal of Fish Biology, 27(sA):59–73, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winstone:1985:AFC

[888] A. J. Winstone, A. S. Gee, and P. V. Varallo. The assessment of flow characteristics at certain weirs in relation to the upstream movement of migratory salmonids. Journal of Fish Biology, 27(sA):75–83, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milner:1985:HEF

[889] N. J. Milner, R. J. Hemsworth, and B. E. Jones. Habitat evaluation as a fisheries management tool. Journal of Fish Biology, 27(sA):85–108, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cave:1985:EKS

[890] J. D. Cave. The effects of the Kielder scheme on fisheries. Journal of Fish Biology, 27(sA):109–121, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:1985:EIS

[891] M. J. Phillips, M. C. M. Beveridge, and L. G. Ross. The environmental impact of salmonid cage culture on inland fisheries: present status and future trends. Journal of Fish Biology, 27(sA):123–137, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamilton:1985:DPR

[892] R. M. Hamilton. Discharges of pesticides to the Rivers Mole and Taw, their accumulation in fish flesh and possible effects on fish stocks. Journal of Fish Biology, 27(sA):139–149, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hickley:1985:CES

[893] P. Hickley and A. Starkie. Cost effective sampling of fish populations in large water bodies. Journal of Fish Biology, 27(sA):151–161, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:1985:SSR

[894] D. R. Jordan and J. S. Wortley. Sampling strategy related to fish distribution, with particular reference to the Norfolk Broads. Journal of Fish Biology, 27(sA):163–173, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coles:1985:SMF

[895] T. F. Coles, J. S. Wortley, and P. Noble. Survey methodology for fish population assessment within Anglian water. Journal of Fish Biology, 27(sA):175–186, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linfield:1985:ACH

[896] R. S. J. Linfield. An alternative concept to home range theory with respect to populations of cyprinids in major river systems. Journal of Fish Biology, 27(sA):187–196, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solbe:1985:EPF

[897] J. F. de L. G. Solbé, V. A. Cooper, C. A. Willis, and M. J. Mallett. Effects of pollutants in fresh waters on European non-salmonid fish i: Non-metals. Journal of Fish Biology, 27(sA):197–207, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:1985:EIF

[898] C. A. Mills and R. H. K. Mann. Environmentally-induced fluctuations in year-class strength and their implications for management. Journal of Fish Biology, 27(sA):209–226, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1985:PMS

[899] R. H. K. Mann. A pike management strategy for a trout fishery. Journal of Fish Biology, 27(sA):227–234, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:1985:EPF

[900] D. M. Reid and W. T. Momot. Evaluation of pulse fishing for the walleye, Stizostedion vitreum vitreum, in Henderson Lake, Ontario. Journal of Fish Biology, 27(sA):235–251, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bninska:1985:PIC

[901] M. Bnińska. The possibilities of improving catchable fish stocks in lakes undergoing eutrophication. Journal of Fish Biology, 27(sA):253–261, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:1985:EPP

[902] A. I. Payne and B. McCarton. Estimation of population parameters and their application to the development of fishery management models in two African rivers. Journal of Fish Biology, 27(sA):263–277, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1985:APT

[903] P. A. Henderson. An approach to the prediction of temperate freshwater fish communities. Journal of Fish Biology, 27(sA):279–291, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katz:1985:UME

[904] A. H. Katz and H. M. Katz. Use of a municipal environmental quality index for understanding and management of aquatic habitats and fish populations. Journal of Fish Biology, 27(sA):293–300, December 1985. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cornish:1986:GLK

[905] I. M. E. Cornish and T. W. Moon. The glucose and lactate kinetics of American eels, Anguilla rostrata (LeSueur), under MS 222 anaesthesia. Journal of Fish Biology, 28(1):1–8, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leong:1986:SOM

[906] T. S. Leong. Seasonal occurrence of metazoan parasites of Puntius binotatus in an irrigation canal, Pulau Pinang, Malaysia. Journal of Fish Biology, 28(1):9–16, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irvine:1986:EVD

[907] J. R. Irvine. Effects of varying discharge on the downstream movement of salmon fry, Oncorhynchus tshawytscha Walbaum. Journal of Fish Biology, 28(1):17–28, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leknes:1986:FSC

[908] I. L. Leknes. Fine structure and cytochemistry of the endothelial cells and rodlet cells in the bulbus arteriosus in species of Cichlidae (Teleostei). Journal of Fish Biology, 28(1):29–36, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Booth:1986:GEP

[909] D. J. Booth and J. A. Keast. Growth energy partitioning by juvenile bluegill sunfish, Lepomis macrochirus Rafinesque. Journal of Fish Biology, 28(1):37–45, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bubien:1986:EAP

[910] J. K. Bubien and T. L. Meade. Electrophysiological abnormalities produced by ammonium in isolated perfused brook trout, Salvelinus fontinalis, hearts. Journal of Fish Biology, 28(1):47–53, January 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Van:1986:RNM

[911] H.-Y. Van. Report of a naturally masculinized female of the Clear Creek gambusia, Gambusia heterochir Hubbs. Journal of Fish Biology, 28(1): 55–59, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joensen:1986:UTC

[912] J. E. Joensen and B. Korsgaard. Uptake, time-course distribution and elimination of cadmium in embryos and tissues of the pregnant Zoarces viviparus (L) after intraovarian loading. Journal of Fish Biology, 28(1): 61–68, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crespo:1986:MFA

[913] S. Crespo, G. Nonnotte, D. A. Colin, C. Leray, L. Nonnotte, and A. Aubree. Morphological and functional alterations induced in trout intestine by dietary cadmium and lead. Journal of Fish Biology, 28(1): 69–80, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santulli:1986:ECG

[914] A. Santulli and V. D’Amello. The effects of carnitine on the growth of sea bass, Dicentrarchus labrax L., fry. Journal of Fish Biology, 28(1):81– 86, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agius:1986:IEA

[915] C. Agius and W. Couchman. Induction of enhanced alkaline phosphatase activity in the melano-macrophage centres of Oreochromis aureus (Steindachner) through starvation and vaccination. Journal of Fish Biology, 28(1):87–92, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romer:1986:MGS

[916] G. S. Romer and A. McLachlan. Mullet grazing on surf diatom accumulations. Journal of Fish Biology, 28(1):93–104, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dabrowski:1986:NTM

[917] K. R. Dabrowski. A new type of metabolism chamber for the determination of active and postprandial metabolism offish, and consideration of results for coregonid and salmon juveniles. Journal of Fish Biology, 28 (1):105–117, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dheer:1986:HHR

[918] J. M. S. Dheer, T. R. Dheer, and C. L. Mahajan. Haematological and haematopoietic response to sodium chloride stress in a freshwater airbreathing fish channapunctatus Bloch. Journal of Fish Biology, 28(1): 119–128, January 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jory:1986:IPQ

[919] D. E. Jory. An incident of predation on queen conch, Strombus gigas L. (Mollusca, Strombidae), by Atlantic permit, Trachinotus falcatus L.

(Pisces, Carangidae). Journal of Fish Biology, 28(2):129–131, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaudreault:1986:IID

[920] A. Gaudreault, T. Miller, W. L. Montgomery, and G. J. FitzGerald. Interspecific interactions and diet of sympatric juvenile brook charr, Salvelinus fontinalis, and adult ninespine sticklebacks, Pungitius pungitius. Journal of Fish Biology, 28(2):133–140, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eiras:1986:SAI

[921] J. C. Eiras. Some aspects of the infection of bib, Trisopterus luscus (L.), by the parasitic copepod Lernaeocera lusci (Basset-Smith, 1896) in Portuguese waters. Journal of Fish Biology, 28(2):141–145, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1986:FHF

[922] E. J. Hyslop. The food habits of four small-sized species of Mormyridae from the floodplain pools of the Sokoto-Rima river basin, Nigeria. Journal of Fish Biology, 28(2):147–151, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertollo:1986:CTC

[923] L. A. C. Bertollo, O. Moreira-Filho, and P. M. Galetti, Jr. Cytogenetics and taxonomy: considerations based on chromosome studies of freshwater fish. Journal of Fish Biology, 28(2):153–159, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colgan:1986:RDE

[924] P. W. Colgan, J. A. Brown, and S. D. Orsatti. Role of diet and experience in the development of feeding behaviour in largemouth bass, Micropterus salmoides. Journal of Fish Biology, 28(2):161–170, February 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herzig:1986:ITE

[925] A. Herzig and H. Winkler. The influence of temperature on the embryonic development of three cyprinid fishes, Abramis brama, Chalcalburnus chalcoides mento and Vimba vimba. Journal of Fish Biology, 28 (2):171–181, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weekes:1986:FRL

[926] P. J. Weekes and B. Penlington. First records of Ligula intestinalis (Cestoda) in rainbow trout, Salmo gairdneri, and common bully, Gobiomorphus cotidianus, in New Zealand. Journal of Fish Biology, 28(2):183–190, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Danulat:1986:EVD

[927] E. Danulat. The effects of various diets on chitinase and ß-glucosidase activities and the condition of cod, Gadus morhua (L.). Journal of Fish Biology, 28(2):191–197, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cali:1986:EIA

[928] A. Cali, P. M. Takvorian, J. J. Ziskowski, and T. K. Sawyer. Experimental infection of American winter flounder (Pseudopleuronectes americanus) with Glugea stephani (Microsporida). Journal of Fish Biology, 28 (2):199–206, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conneely:1986:EFI

[929] J. J. Conneely and T. K. McCarthy. Ecological factors influencing the composition of the parasite fauna of the European eel, Anguilla anguilla (L.), in Ireland. Journal of Fish Biology, 28(2):207–219, February 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:1986:CAB

[930] K. H. Olsén. Chemo-attraction between juveniles of two sympatric stocks of Arctic charr, Salvelinus alpinus (L.), and their gene frequency of serum esterases. Journal of Fish Biology, 28(2):221–231, February 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beacham:1986:CDB

[931] T. D. Beacham and C. B. Murray. Comparative developmental biology of pink salmon, Oncorhynchus gorbuscha, in southern British Columbia. Journal of Fish Biology, 28(2):233–246, February 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roy:1986:QQA

[932] S. Roy and B. C. Pal. Quantitative and qualitative analysis of spawning behaviour of Heteropneustes fossilis (Bloch.) (Siluridae) in laboratory aquaria. Journal of Fish Biology, 28(3):247–254, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:1986:ERA

[933] K. H. Olsén. Emission rate of amino acids and ammonia and their role in olfactory preference behaviour of juvenile Arctic charr, Salvelinus alpinus (L.). Journal of Fish Biology, 28(3):255–265, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloley:1986:ESP

[934] B. D. Sloley, B. E. Hickie, D. G. Dixon, R. G. H. Downer, and R. J. Martin. The effects of sodium pentachlorophenate, diet and sampling procedure on amine and tryptophan concentrations in the brain of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 28(3): 267–277, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:1986:GHL

[935] A. J. Geffen. The growth of herring larvae, Clupea harengus L., in the Clyde: an assessment of the suitability of otolith ageing methods. Journal of Fish Biology, 28(3):279–288, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakanari:1986:LIP

[936] J. A. Sakanari and M. Moserj. Lesion induction by the plerocercoid Lacistorhynchus tenuis (Cestoda) and wound healing in the striped bass Morone saxatilis (Walbaum). Journal of Fish Biology, 28(3):289–296, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garratt:1986:PHS

[937] P. A. Garratt. Protogynous hermaphroditism in the slinger, Chrysoblephus puniceus (Gilchrist & Thompson, 1908) (Teleosteir: Sparidae). Journal of Fish Biology, 28(3):297–306, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dederen:1986:BAT

[938] L. H. T. Dederen, R. S. E. W. Leuven, S. E. Wendelaar, and F. G. F. Oyen. Biology of the acid-tolerant fish species Umbra pygmaea (De Kay, 1842). Journal of Fish Biology, 28(3):307–326, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1986:KSA

[939] A. F. Brown, J. C. Chubb, and C. J. Veltkamp. A key to the species of Acanthocephala parasitic in British freshwater fishes. Journal of Fish Biology, 28(3):327–334, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggberget:1986:AMA

[940] T. G. Heggberget, T. Haukebø, and B. Veie-Rosvoll. An aerial method of assessing spawning activity of Atlantic salmon, Salmo salar L., and brown trout, Salmo trutta L., in Norwegian streams. Journal of Fish Biology, 28(3):335–342, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heins:1986:SPN

[941] D. C. Heins and F. G. Rabito, Jr. Spawning performance in North American minnows: direct evidence of the occurrence of multiple clutches in the genus Notropis. Journal of Fish Biology, 28(3):343–357, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elras:1986:HPL

[942] J. C. Elras, A. A. Rego, and G. C. Pavanelll. Histopathology in Paulicea lutkeni (Pisces: Pimelodidae) resulting from infections with Megathylacus brooksi and Jauela glandicephalus (Cestoda: Proteocephalidae). Journal of Fish Biology, 28(3):359–365, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haug:1986:MBG

[943] T. Haug and S. E. Fevolden. Morphology and biochemical genetics of Atlantic halibut, Hippoglossus hippoglossus (L.), from various spawning grounds. Journal of Fish Biology, 28(3):367–378, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1986:PSC

[944] M. Jobling and T.-G. Reinsnes. Physiological and social constraints on growth of Arctic charr, Salvelinus alpinus L.: an investigation of factors leading to stunting. Journal of Fish Biology, 28(3):379–384, March 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heming:1986:AIC

[945] T. A. Heming and T. A. Watson. Activity and inhibition of carbonic anhydrase in Amia calva, a bimodal-breathing holostean fish. Journal of Fish Biology, 28(4):385–392, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoyle:1986:HCM

[946] J. Hoyle, H. S. Gill, and A. H. Weatherley. Histochemical characterization of myotomal muscle in the grass pickerel, Esox americanus vermiculatus (LeSeuer), and the muskellunge, E. masquinongy (Mitchell). Journal of Fish Biology, 28(4):393–401, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1986:COF

[947] G. M. Hughes, I. Martinez, and G. Boeuf. Comparison of osmotic fragility of red blood cells of freshwater-adapted and seawater-adapted rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 28(4):403– 406, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthews:1986:GVT

[948] W. J. Matthews. Geographic variation in thermal tolerance of a widespread minnow Notropis lutrensis of the North American mid-west. Journal of Fish Biology, 28(4):407–417, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Song:1986:TEM

[949] Y. C. Song, F. J. Ratty, and R. M. Laurs. Transmission electron microscope observations of C-banded skipjack tuna chromosomes. Journal of Fish Biology, 28(4):419–424, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:1986:WIB

[950] W. O. Watanabe and C-M. Kuo. Water and ion balance in hydrating oocytes of the grey mullet, Mugil cephalus (L.), during hormone-induced final maturation. Journal of Fish Biology, 28(4):425–437, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:1986:DDP

[951] T. J. Pitcher, D. A. Green, and A. E. Magurran. Dicing with death: predator inspection behaviour in minnow shoals. Journal of Fish Biology, 28(4):439–448, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daoulas:1986:SVE

[952] Ch. Daoulas and A. N. Economou. Seasonal variation of egg size in the sardine, Sardino pilchardus Walb., of the Saronikos Gulf: causes and a probable explanation. Journal of Fish Biology, 28(4):449–457, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crozier:1986:EEP

[953] W. W. Crozier and A. Ferguson. Electrophoretic examination of the population structure of brown trout, Salmo trutta L., from the Lough Neagh catchment, Northern Ireland. Journal of Fish Biology, 28(4):459– 477, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1986:EIIa

[954] G. J. A. Kennedy and C. D. Strange. The effects of intra- and interspecific competition on the survival and growth of stocked juvenile Atlantic salmon, Salmo solar L., and resident trout, Salmo trutta L., in an upland stream. Journal of Fish Biology, 28(4):479–489, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fox:1986:EFH

[955] D. P. Fox, R. Johnstone, and E. Durward. Erythrocyte fusion in heatshocked Atlantic salmon. Journal of Fish Biology, 28(4):491–499, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tongiorgi:1986:TPU

[956] P. Tongiorgi, L. Tosi, and M. Balsamo. Thermal preferences in upstream migrating glass-eels of Anguilla anguilla (L.). Journal of Fish Biology, 28(4):501–510, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellsaesser:1986:HIC

[957] C. F. Ellsaesser and L. W. Clem. Haematological and immunological changes in channel catfish stressed by handling and transport. Journal of Fish Biology, 28(4):511–521, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:A

[958] Anonymous. Announcements. Journal of Fish Biology, 28(4):523, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:E

[959] Anonymous. Errata. Journal of Fish Biology, 28(4):524, April 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:1986:IVC

[960] N. B. Metcalfe. Intraspecific variation in competitive ability and food intake in salmonids: consequences for energy budgets and growth rates. Journal of Fish Biology, 28(5):525–531, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1986:SAB

[961] A. Wright, P. J. Dalzell, and A. H. Richards. Some aspects of the biology of the red bass, Lutjanus bohar (Forsskål), from the Tigak Islands, Papua New Guinea. Journal of Fish Biology, 28(5):533–544, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parish:1986:LES

[962] N. Parish, A. Wrathmell, S. Hart, and J. E. Harris. The leucocytes of the elasmobranch Scyliorhinus canicula L. — a morphological study. Journal of Fish Biology, 28(5):545–561, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Iwama:1986:CSH

[963] G. K. Iwama, G. L. Greek, and D. J. Randall. Changes in selected haematological parameters in juvenile Chinook salmon subjected to a bacterial challenge and a toxicant. Journal of Fish Biology, 28(5):563– 572, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roubal:1986:BOP

[964] Frank R. Roubal. Blood and other possible inflammatory cells in the sparid Acanthopagrus australis (Günther). Journal of Fish Biology, 28 (5):573–593, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:1986:SCS

[965] A. White, T. C. Fletcher, and J. A. Pope. Seasonal changes in serum lipid composition of the plaice, Pleuronectes platessa L. Journal of Fish Biology, 28(5):595–606, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:1986:EEG

[966] P. B. Taylor. Experimental evidence for geomagnetic orientation in juvenile salmon, Oncorhynchus tschawytscha Walbaum. Journal of Fish Biology, 28(5):607–623, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:1986:CSC

[967] I. G. Cowx and N. M. Broughton. Changes in the species composition of anglers’ catches in the River Trent (England) between 1969 and 1984. Journal of Fish Biology, 28(5):625–636, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:1986:GSE

[968] J. W. Jensen. Gillnet selectivity and the efficiency of alternative combinations of mesh sizes for some freshwater fish. Journal of Fish Biology, 28(5):637–646, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:BRa

[969] Anonymous. Book reviews. Journal of Fish Biology, 28(5):647–652, May 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weeks:1986:ITC

[970] B. A. Weeks, J. E. Warinner, P. L. Mason, and D. S. McGinnis. Influence of toxic chemicals on the chemotactic response of fish macrophages. Journal of Fish Biology, 28(6):653–658, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1986:EDD

[971] A. F. Brown. Evidence for density-dependent establishment and survival of Pomphorhynchus laevis (Müller, 1776) (Acanthocephala) in laboratory-infected Salmo gairdneri Richardson and its bearing on wild populations in Leuciscus cephalus (L.). Journal of Fish Biology, 28(6): 659–669, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1986:VCM

[972] J. F. Craig, A. Sharma, and K. Smile. The variability in catches from multi-mesh gillnets fished in three Canadian lakes. Journal of Fish Biology, 28(6):671–678, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lesack:1986:ECP

[973] L. F. W. Lesack. Estimates of catch and potential yield for the riverine artisanal fishery in The Gambia, West Africa. Journal of Fish Biology, 28(6):679–700, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmann:1986:SLC

[974] R. W. Hoffmann, J. Meder, M. Klein, K. Osterkornj, and R. D. Negele. Studies on lesions caused by plerocercoids of Triaenophorus nodulosmin some fish of an alpine lake, the Königssee. Journal of Fish Biology, 28(6):701–712, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sire:1986:ODS

[975] J.-Y. Sire. Ontogenic development of surface ornamentation in the scales of Hemichromis bimaculatus (Cichlidae). Journal of Fish Biology, 28(6): 713–724, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Audet:1986:ECS

[976] C. Audet, G. J. Fitzgerald, and H. Guderley. Environmental control of salinity preferences in four sympatric species of sticklebacks: Gasterosteus aculeatus, Gasterosteus wheatlandi, Pungitius pungitius and Apeltes quadracus. Journal of Fish Biology, 28(6):725–739, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schindler:1986:SCC

[977] J. F. Schindler and U. de Vries. Scanning cytophotometry of carp, Cyprinm carpio L., erythrocyte populations: the influence of short-term hypoxic environment and the recovery period following severe bleeding. Journal of Fish Biology, 28(6):741–752, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:BRb

[978] Anonymous. Book reviews. Journal of Fish Biology, 28(6):753–756, June 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araujo-Lima:1986:FMR

[979] C. A. R. M. Araujo-Lima, L. P. S. Portugal, and E. G. Ferreira. Fish– macrophyte relationship in the Anavilhanas Archipelago, a black water system in the Central Amazon. Journal of Fish Biology, 29(1):1–11, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kilambi:1986:AGR

[980] R. V. Kilambi. Age, growth and reproductive strategy of the snakehead, Ophicephalus striatus Bloch, from Sri Lanka. Journal of Fish Biology, 29 (1):13–22, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sainsbury:1986:EFC

[981] K. J. Sainsbury. Estimation of food consumption from field observations of fish feeding cycles. Journal of Fish Biology, 29(1):23–36, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hodson:1986:ETH

[982] P. V. Hodson and B. R. Blunt. The effect of time from hatch on the yolk conversion efficiency of rainbow trout, Salmo gairdneri. Journal of Fish Biology, 29(1):37–46, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendrick:1986:PCP

[983] M. Hendrick, A. DiNapoli, P. Cammarata, and S. Pincus. Purification of carp putative eosinophils on metrizamide gradients. Journal of Fish Biology, 29(1):47–51, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sazima:1986:SFB

[984] I. Sazima. Similarities in feeding behaviour between some marine and freshwater fishes in two tropical communities. Journal of Fish Biology, 29(1):53–65, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:1986:WSV

[985] J. F. Craig and K. Smiley. Walleye, Stizostedion vitreum, and northern pike, Esox lucius, populations in three Alberta lakes. Journal of Fish Biology, 29(1):67–85, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:1986:CNC

[986] R. K. Harris, T. Nishiyama, and A. J. Paul. Carbon, nitrogen and caloric content of eggs, larvae, and juveniles of the walleye pollock, Theragra chalcogramma. Journal of Fish Biology, 29(1):87–98, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brabrand:1986:JRI

[987] Å. Brabrand, B. Faafeng, and J. P. M. Nilssen. Juvenile roach and invertebrate predators: delaying the recovery phase of eutrophic lakes by suppression of efficient filter-feeders. Journal of Fish Biology, 29(1): 99–106, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1986:CPF

[988] N. Giles, R. M. Wright, and M. E. Nord. Cannibalism in pike fry, Esox lucius L.: some experiments with fry densities. Journal of Fish Biology, 29(1):107–113, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stables:1986:PIR

[989] J. N. Stables and L. H. Chappell. Putative immune response of rainbow trout, Salmo gairdneri, to Diplostomum spathaceum infections. Journal of Fish Biology, 29(1):115–122, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mittal:1986:FSS

[990] A. K. Mittal and G. D. Nigam. Fish skin surface lipids: phospholipids. Journal of Fish Biology, 29(1):123–138, July 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tabachek:1986:IDP

[991] J. L. Tabachek. Influence of dietary protein and lipid levels on growth, body composition and utilization efficiencies of Arctic charr, Salvelinus alpinus L. Journal of Fish Biology, 29(2):139–151, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mcbride:1986:CST

[992] J. R. Mcbride, U. H. M. Fagerlund, H. M. Dye, and J. Bagshaw. Changes in structure of tissues and in plasma cortisol during the spawning migration of pink salmon, Oncorhynchus gorbucha (Walbaum). Journal of Fish Biology, 29(2):153–166, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dye:1986:CRP

[993] H. M. Dye, J. P. Sumpter, U. H. M. Fagerlund, and E. M. Donaldson. Changes in reproductive parameters during the spawning migration of pink salmon, Oncorhynchus gorbuscha (Walbaum). Journal of Fish Biology, 29(2):167–176, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mughal:1986:ISJ

[994] M. S. Mughal and M. J. Manning. The immune system of juvenile thicklipped grey mullet, Chelon labrosus Risso: antibody responses to soluble protein antigens. Journal of Fish Biology, 29(2):177–186, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malte:1986:EAH

[995] H. Malte. Effects of aluminium in hard, acid water on metabolic rate, blood gas tensions and ionic status in the rainbow trout. Journal of Fish Biology, 29(2):187–198, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1986:EIIb

[996] G. J. A. Kennedy and C. D. Strange. The effects of intra- and interspecific competition on the distribution of stocked juvenile Atlantic salmon, Salmo salar L., in relation to depth and gradient in an upland trout, Salmo trutta L., stream. Journal of Fish Biology, 29(2):199–214, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1986:CPF

[997] C. R. Kennedy, D. d’A. Laffoley, G. Bishop, P. Jones, and M. Taylor. Communities of parasites of freshwater fish of Jersey, Channel Islands. Journal of Fish Biology, 29(2):215–226, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodhead:1986:RFC

[998] A. D. Woodhead and C. Carlson. The role of female competition in the mating success of two clones of the Amazon molly, Poecilia formosa. Journal of Fish Biology, 29(2):227–232, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsen:1986:IAS

[999] B. O. Johnsen and A. J. Jensen. Infestations of Atlantic salmon, Salmo salar, by Gyrodactylus salaris in Norwegian rivers. Journal of Fish Biology, 29(2):233–241, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kindle:1986:BIT

[1000] K. R. Kindle and D. H. Whitmore. Biochemical indicators of thermal stress in Tilapia aurea (Steindachner). Journal of Fish Biology, 29(2): 243–255, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:1986:SEM

[1001] E. Z. Johnson and R. G. Werner. Scanning electron microscopy of the chorion of selected freshwater fishes. Journal of Fish Biology, 29(2):257– 265, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:1986:OCO

[1002] C. D. De Silva, S. Premawansa, and C. N. Keembiyahetty. Oxygen consumption in Oreochromis niloticus (L.) in relation to development, salinity, temperature and time of day. Journal of Fish Biology, 29(2): 267–277, August 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunjak:1986:SCP

[1003] R. A. Cunjak and G. Power. Seasonal changes in the physiology of brook trout, Salvelinus fontinalis (Mitchill), in a sub-Arctic river system. Journal of Fish Biology, 29(3):279–288, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:1986:SMM

[1004] J. Davenport, S. Lønning, and E. Kjørsvik. Some mechanical and morphological properties of the chorions of marine teleost eggs. Journal of Fish Biology, 29(3):289–301, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wyche:1986:FEP

[1005] C. J. Wyche and S. E. Shackley. The feeding ecology of Pleuronectes platessa L., Limanda limanda (L.) and Scophthalmus rhombus (L.) in Carmarthen Bay, South Wales, U.K. Journal of Fish Biology, 29(3): 303–311, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadler:1986:SEL

[1006] K. Sadler and S. Lynam. Some effects of low pH and calcium on the growth and tissue mineral content of yearling brown trout, Salmo trutta. Journal of Fish Biology, 29(3):313–324, September 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crawford:1986:DEL

[1007] C. M. Crawford. Development of eggs and larvae of the flounders Rhombosolea tapirina and Ammotretis rostratus (Pisces: Pleuronectidae). Journal of Fish Biology, 29(3):325–334, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1986:CBC

[1008] A. D. Pickering. Changes in blood cell composition of the brown trout, Salmo trutta L., during the spawning season. Journal of Fish Biology, 29 (3):335–347, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suzuki:1986:MPC

[1009] K. Suzuki. Morphological and phagocytic characteristics of peritoneal exudate cells in tilapia, Oreochromis niloticus (Trewavas), and carp, Cyprinus carpio L. Journal of Fish Biology, 29(3):349–364, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Behrend:1986:MEF

[1010] K. Behrend. The musculature of the electric fish Eigenmannia virescens (South American green knife fish) characterized with cytochrome oxidase staining. Journal of Fish Biology, 29(3):365–369, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drost:1986:EFS

[1011] M. R. Drost and J. G. M. van den Boogaart. The energetics of feeding strikes in larval carp, Cyprinus carpio. Journal of Fish Biology, 29(3): 371–379, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abidin:1986:RBT

[1012] A. Z. Abidin. The reproductive biology of a tropical cyprinid, Hampala macrolepidota (Van Hasselt), from Zoo Negara Lake, Kuala Lumpur, Malaysia. Journal of Fish Biology, 29(3):381–391, September 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1986:PSM

[1013] G. M. Hughes, Y. Kikuchi, and J. Barrington. Physiological salines and the mechanical properties of trout red blood cells. Journal of Fish Biology, 29(3):393–402, September 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deegan:1986:CBC

[1014] L. A. Deegan. Changes in body composition and morphology of youngof-the-year gulf menhaden, Brevoortia patronus Goode, in Fourleague Bay, Louisiana. Journal of Fish Biology, 29(4):403–415, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poynton:1986:DFH

[1015] S. L. Poynton. Distribution of the flagellate Hexamita salmonis Moore, 1922 and the microsporidian Loma salmonae Putz, Hoffman and Dunbar, 1965 in brown trout, Salmo trutta L., and rainbow trout, Salmo gairdneri Richardson, in the River Itchen (U.K.) and three of its fish farms. Journal of Fish Biology, 29(4):417–429, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1986:EFO

[1016] J. M. Wright. The ecology of fish occurring in shallow water creeks of a Nigerian mangrove swamp. Journal of Fish Biology, 29(4):431–441, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1986:PED

[1017] G. M. Hughes, J. S. Datta Munshi, and J. Ojha. Post-embryonic development of water- and airbreathing organs of Anabas testudineus (Bloch). Journal of Fish Biology, 29(4):443–450, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakari:1986:BRE

[1018] T. Nakari, A. Soivio, and S. Pesonen. Book review: Effects of advanced photoperiod cycle on the epidermis and gonadosomatic index of 2-year-old rainbow trout, Salmo gairdneri R., reared at natural temperature. Journal of Fish Biology, 29(4):451–457, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:1986:UMA

[1019] A. J. Jensen, T. G. Heggberget, and B. O. Johnsen. Upstream migration of adult Atlantic salmon, Salmo salar L., in the River Vefsna, northern Norway. Journal of Fish Biology, 29(4):459–465, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagliniere:1986:PEJ

[1020] J.-L. Baglinière and A. Champigneulle. Population estimates of juvenile Atlantic salmon, Salmo salar, as indices of smolt production in the R. Scorff, Brittany. Journal of Fish Biology, 29(4):467–482, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohr:1986:QMR

[1021] A. Mohr, W. Wolf, M. Bohl, and R. Hoffmann. Quantification of methaemoglobin reduction in red blood cells of the rainbow trout Salmo gairdneri R. Journal of Fish Biology, 29(4):483–487, October 1986. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balogun:1986:FDK

[1022] J. K. Balogun. Fish distribution in Kainji Lake, Nigeria. Journal of Fish Biology, 29(4):489–498, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Semple:1986:RBE

[1023] G. P. Semple. Reproductive behaviour and early development of the Drysdale hardyhead, Craterocephalus sp. nov. (Pisces: Atherinidae), from the Alligator Rivers System, Northern Territory, Australia. Journal of Fish Biology, 29(4):499–506, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiley:1986:RBF

[1024] M. J. Wiley, S. M. Pescitelli, and L. D. Wike. The relationship between feeding preferences and consumption rates in grass carp and grass carp × bighead carp hybrids. Journal of Fish Biology, 29(4):507–514, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skurdal:1986:IVM

[1025] J. Skurdal, O. K. Skogheim, and T. Qvenild. Indicative value of mercury concentration in adipose fin of brown trout, Salmo trutta L. Journal of Fish Biology, 29(4):515–517, October 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krasznai:1986:SIT

[1026] Z. Krasznai and T. Márían. Shock-induced triploidy and its effect on growth and gonad development of the European catfish, Silurus glanis L. Journal of Fish Biology, 29(5):519–527, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vitturi:1986:CPA

[1027] R. Vitturi, A. Mazzola, M. Macaluso, and E. Catalano. Chromosomal polymorphism associated with Robertsonian fusion in Seriola dumerili (Risso, 1810) (Pisces: Carangidae). Journal of Fish Biology, 29(5):529– 534, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:1986:LTT

[1028] H. J. M. Hansen and A. T. A. Fattah. Long-term tagging of elvers, Anguilla anguilla, with radioactive europium. Journal of Fish Biology, 29 (5):535–540, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1986:CFB

[1029] N. F. Hughes. Changes in the feeding biology of the Nile perch, Lates niloticus (L.) (Pisces: Centropomidae), in Lake Victoria, East Africa since its introduction in 1960, and its impact on the native fish community of the Nyanza Gulf. Journal of Fish Biology, 29(5):541–548, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craik:1986:CEW

[1030] J. C. A. Craik and S. M. Harvey. The carotenoids of eggs of wild and farmed Atlantic salmon, and their changes during development to the start of feeding. Journal of Fish Biology, 29(5):549–565, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1986:EOA

[1031] C. S. Lee, G. M. Weber, and C. S. Tamaru. Effect of orallyadministered 17α-methyltestosterone on spermatogenesis in immature milkfish, Chanos chanos Forsskål. Journal of Fish Biology, 29(5):567– 572, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:1986:TDC

[1032] G. F. Turner. Territory dynamics and cost of reproduction in a captive population of the colonial nesting mouthbrooder Oreochromis mossambicus (Peters). Journal of Fish Biology, 29(5):573–587, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badcock:1986:ARB

[1033] J. Badcock. Aspects of the reproductive biology of Gonostoma bathyphilum (Gonostomatidae). Journal of Fish Biology, 29(5):589–603, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bitterlich:1986:BFF

[1034] G. Bitterlich and E. Schaber. Bacteria — food or food competitors of silver carp, Hypophthalmichthys molitrix Val.? Journal of Fish Biology, 29 (5):605–612, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1986:DOT

[1035] J. S. Alabaster and P. J. Gough. The dissolved oxygen and temperature requirements of Atlantic salmon, Salmo salar L., in the Thames Estuary. Journal of Fish Biology, 29(5):613–621, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cochrane:1986:MYY

[1036] K. L. Cochrane. Mortality of young-of-the-year Oreochromis mossambicus in Hartbeespoort Dam, South Africa. Journal of Fish Biology, 29 (5):623–630, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:BRc

[1037] Anonymous. Book reviews. Journal of Fish Biology, 29(5):631–633, November 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whoriskey:1986:BSP

[1038] F. G. Whoriskey, G. J. Fitzgerald, and S. G. Reebst. The breeding-season population structure of three sympatric, territorial sticklebacks (Pisces: Gasterosteidae). Journal of Fish Biology, 29(6):635–648, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eastman:1986:RGE

[1039] J. T. Eastman and A. L. Devries. Renal glomerular evolution in Antarctic notothenioid fishes. Journal of Fish Biology, 29(6):649–662, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloley:1986:BRI

[1040] B. D. Sloley, R. A. Cunjak, G. Power, and R. G. H. Downer. Book review: The influence of sex and spawning on levels of tryptophan, serotonin and 5-hydroxyindoleacetic acid in the brains of wild brook trout, Salvelinus fontinalis. Journal of Fish Biology, 29(6):663–669, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcus:1986:FFH

[1041] O. Marcus. Food and feeding habits of Ilisha africana (Bloch) (Pisces: Clupeidae) off the Lagos coast, Nigeria. Journal of Fish Biology, 29(6): 671–683, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madhavi:1986:DMC

[1042] R. Madhavi. Distribution of metacercariae of Centrocestus formosanus (Trematoda: Heterophyidae) on the gills of Aplocheilus panchax. Journal of Fish Biology, 29(6):685–690, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dabrowski:1986:TTC

[1043] K. Dabrowski and K. Tsukamoto. Tetracycline tagging in coregonid embryos and larvae. Journal of Fish Biology, 29(6):691–698, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fischer-scherl:1986:LEM

[1044] Th. Fischer-scherl and R. Hoffmann. Light- and electron-microscope studies on the pseudobranch of the golden orfe, Leuciscus idus L. Journal of Fish Biology, 29(6):699–709, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1986:OGP

[1045] P. M. Hine, J. M. Wain, and D. M. Dunlop. Observations on granulocyte peroxidase in New Zealand freshwater eels, Anguilla species. Journal of Fish Biology, 29(6):711–720, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1986:LEM

[1046] P. M. Hine, J. M. Wain, N. C. Boustead, and D. M. Dunlop. Light and electron microscopic studies on the enzyme cytochemistry of leucocytes of eels, Anguilla species. Journal of Fish Biology, 29(6):721–735, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mudge:1986:SPC

[1047] S. M. Mudge and J. Davenport. Serum pigmentation in Cyclopterus lumpus L. Journal of Fish Biology, 29(6):737–745, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heap:1986:ASP

[1048] S. P. Heap and G. Goldspink. Alterations to the swimming performance of carp, Cyprinus carpio, as a result of temperature acclimation. Journal of Fish Biology, 29(6):747–753, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weng:1986:STD

[1049] H. T. Weng. Spatial and temporal distribution of whiting (Sillaginidae) in Moreton Bay, Queensland. Journal of Fish Biology, 29(6):755–764, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:1986:P

[1050] Tony Pitcher, Anne Magurran, and Dick Margetts. Preface. Journal of Fish Biology, 29(sA):v, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1986:LPP

[1051] Anonymous. List of papers presented at the Bangor symposium on the Behaviour of Fishes. Journal of Fish Biology, 29(sA):vii–x, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milinski:1986:RCR

[1052] M. Milinski. A review of competitive resource sharing under constraints in sticklebacks. Journal of Fish Biology, 29(sA):1–14, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:1986:TSS

[1053] J. W. A. Grant and D. L. G. Noakes. A test of a size-selective predation model with juvenile brook charr, Salvelinus fontinalis. Journal of Fish Biology, 29(sA):15–23, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bres:1986:NLO

[1054] M. Bres. A new look at optimal foraging behaviour; rule of thumb in the rainbow trout. Journal of Fish Biology, 29(sA):25–36, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:1986:ISA

[1055] I. J. Winfield. The influence of simulated aquatic macrophytes on the zooplankton consumption rate of juvenile roach, Rutilus rutilus, rudd, Scardinius erythrophthalmus, and perch, Perca fluviatilis. Journal of Fish Biology, 29(sA):37–48, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmes:1986:VCD

[1056] R. A. Holmes and R. N. Gibson. Visual cues determining prey selection by the turbot, Scophthalmus maximus L. Journal of Fish Biology, 29 (sA):49–58, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:1986:DDE

[1057] T. J. Pitcher and J. R. Turner. Danger at dawn: experimental support for the twilight hypothesis in shoaling minnows. Journal of Fish Biology, 29 (sA):59–70, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glass:1986:LIT

[1058] C. W. Glass, C. S. Wardle, and W. R. Mojsiewicz. A light intensity threshold for schooling in the Atlantic mackerel, Scomber scombrus. Journal of Fish Biology, 29(sA):71–81, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:1986:SSB

[1059] T. J. Pitcher, A. E. Magurran, and J. R. Allan. Size-segregative behaviour in minnow shoals. Journal of Fish Biology, 29(sA):83–95, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allan:1986:ISC

[1060] J. R. Allan. The influence of species composition on behaviour in mixedspecies cyprinid shoals. Journal of Fish Biology, 29(sA):97–106, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baerends:1986:CFP

[1061] G. P. Baerends. On causation and function of the pre-spawning behaviour of cichlid fish. Journal of Fish Biology, 29(sA):107–121, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barlow:1986:MCM

[1062] G. W. Barlow. Mate choice in the monogamous and polychromatic Midas cichlid, Cichlasoma citrinellum. Journal of Fish Biology, 29(sA):123– 133, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKaye:1986:MCS

[1063] K. R. McKaye. Mate choice and size assortative pairing by the cichlid fishes of Lake Jiloá, Nicaragua. Journal of Fish Biology, 29(sA):135–150, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deCaprona:1986:PTC

[1064] M.-D. Crapon de Caprona. Are ‘preferences’ and ‘tolerances’ in cichlid mate choice important for speciation? Journal of Fish Biology, 29(sA): 151–158, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MaGurran:1986:DSB

[1065] A. E. MaGurran. The development of shoaling behaviour in the European minnow, Phoxinus phoxinus. Journal of Fish Biology, 29(sA):159–169, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1986:DFB

[1066] J. A. Brown. The development of feeding behaviour in the lumpfish, Cyclopterus lumpus. Journal of Fish Biology, 29(sA):171–178, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomez-Laplaza:1986:TIS

[1067] L. M. Gómez-Laplaza and E. Morgan. Towards an isolation syndrome for the angelfish, Pterophyllum scalare. Journal of Fish Biology, 29(sA): 179–187, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katz:1986:MSS

[1068] H. M. Katz. Migrational swimming speeds of the American shad, Alosa sapidissima, in the Connecticut River, Massachusetts, U.S.A. Journal of Fish Biology, 29(sA):189–197, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:1986:BAG

[1069] I. G. Priede and K. L. Smith, Jr. Behaviour of the abyssal grenadier, Coryphaenoides yaquinae, monitored using ingestible acoustic transmitters in the Pacific Ocean. Journal of Fish Biology, 29(sA):199–206, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:1986:HRI

[1070] J. D. Armstrong. Heart rate as an indicator of activity, metabolic rate, food intake and digestion in pike, Esox lucius. Journal of Fish Biology, 29(sA):207–221, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:1986:TBA

[1071] P. He and C. S. Wardle. Tilting behaviour of the Atlantic mackerel, Scomber scombrus, at low swimming speeds. Journal of Fish Biology, 29(sA):223–232, December 1986. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Owusu-Frimpong:1987:BBP

[1072] M. Owusu-Frimpong. Breeding behavioural pattern of the lake fish Tilapia discolour Günther (Teleostei, Cichlidae) in captivity. Journal of Fish Biology, 30(1):1–5, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorogood:1987:AGR

[1073] J. Thorogood. Age and growth rate determination of southern bluefin tuna, Thunnus maccoyii, using otolith banding. Journal of Fish Biology, 30(1):7–14, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cousin:1987:OEA

[1074] J. C. B. Cousin, F. Baudin-Laurencin, and J. Gabaudan. Ontogeny of enzymatic activities in fed and fasting turbot, Scophthalmus maximus L. Journal of Fish Biology, 30(1):15–33, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaxhall:1987:EPL

[1075] P. C. Blaxhall and T. Doggett. Esterases and phosphatases in the leucocytes of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 30(1):35–40, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1987:LIA

[1076] A. D. Pickering and T. G. Pottinger. Lymphocytopenia and interrenal activity during sexual maturation in the brown trout, Salmo trutta L. Journal of Fish Biology, 30(1):41–50, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rasotto:1987:PSI

[1077] M. B. Rasotto, P. Cardellini, and E. Marconato. The problem of sexual inversion in the minnow, Phoxinus phoxinus (L.). Journal of Fish Biology, 30(1):51–57, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rambhaskar:1987:CHT

[1078] B. Rambhaskar and K. Srinivasa Rao. Comparative haematology of ten species of marine fish from visakhapatnam Coast. Journal of Fish Biology, 30(1):59–66, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:1987:BSL

[1079] R. R. C. Edwards and S. Shaher. Biometrics of Sardinella longiceps Val. in relation to upwelling in the Gulf of Aden. Journal of Fish Biology, 30(1):67–73, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagua:1987:NSE

[1080] V. O. Sagua. On a new species of electric catfish from Kainji, Nigeria, with some observations on its biology. Journal of Fish Biology, 30(1):75– 89, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lagardere:1987:FED

[1081] J. P. Lagardère. Feeding ecology and daily food consumption of common sole, Solea vulgaris Quensel, juveniles on the French Atlantic coast. Journal of Fish Biology, 30(1):91–104, January 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:DOW

[1082] Anonymous. D onald w. J olly M rcvs. Journal of Fish Biology, 30(2): 105, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chao:1987:PTS

[1083] Nai-Hsien Chao, Wei-Chung Chao, Kuo-Chun Liu, and I-Chiu Liao. The properties of tilapia sperm and its cryopreservation. Journal of Fish Biology, 30(2):107–118, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naesje:1987:FGA

[1084] T. F. Næsje, B. Jonsson, L. Klyve, and O. T. Sandlund. Food and growth of age-0 smelts, Osmerus eperlanus, in a Norwegian fjord lake. Journal of Fish Biology, 30(2):119–126, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mtada:1987:ITS

[1085] O. S. M. Mtada. The influence of thermal stratification on pelagic fish yields in Lake Kariba, Zambia/Zimbabwe. Journal of Fish Biology, 30 (2):127–133, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cataldi:1987:SHM

[1086] E. Cataldi, S. Cataudella, G. Monaco, A. Rossi, and L. Tancioni. A study of the histology and morphology of the digestive tract of the sea-bream, Sparus aurata. Journal of Fish Biology, 30(2):135–145, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hickley:1987:FFR

[1087] P. Hickley and R. G. Bailey. Food and feeding relationships of fish in the Sudd swamps (River Nile, southern Sudan). Journal of Fish Biology, 30 (2):147–159, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romanello:1987:PLG

[1088] M. G. Romanello, P. A. Scapolo, S. Luprano, and F. Mascarello. Post-larval growth in the lateral white muscle of the eel, Anguilla anguilla. Journal of Fish Biology, 30(2):161–172, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morman:1987:RDG

[1089] R. H. Morman. Relationship of density to growth and metamorphosis of caged larval sea lampreys, Petromyzon marinus Linnaeus, in Michigan streams. Journal of Fish Biology, 30(2):173–181, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1987:GFH

[1090] E. J. Hyslop. The growth and feeding habits of Clarias anguillaris during their first season in the floodplain pools of the Sokoto-Rima river basin, Nigeria. Journal of Fish Biology, 30(2):183–192, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:1987:SRC

[1091] N. W. Pankhurst, P. J. McMillan, and D. M. Tracey. Seasonal reproductive cycles in three commercially exploited fishes from the slope waters off New Zealand. Journal of Fish Biology, 30(2):193–211, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abu-Hakima:1987:ARB

[1092] R. Abu-Hakima. Aspects of the reproductive biology of the grouper, Epinephelus tauvina (Forsskål), in Kuwaiti waters. Journal of Fish Biology, 30(2):213–222, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Aa

[1093] Anonymous. Announcement. Journal of Fish Biology, 30(2):223, February 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Witthames:1987:AMC

[1094] P. R. Witthames and M. Greer Walker. An automated method for counting and sizing fish eggs. Journal of Fish Biology, 30(3):225–235, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1987:BRD

[1095] N. S. Weatherley. Book review: The diet and growth of 0-group dace, Leuciscus leuciscus (L.), and roach, Rutilus rutilus (L.), in a lowland river. Journal of Fish Biology, 30(3):237–247, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:1987:FRB

[1096] A. B. Wood and D. R. Jordan. Fertility of roach × bream hybrids, Rutilus rutilus (L.) × Abramis brama (L.), and their identification. Journal of Fish Biology, 30(3):249–261, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:1987:ITY

[1097] Y. L. Wang, R. K. Buodington, and S. I. Doroshov. Influence of temperature on yolk utilization by the white sturgeon, Acipenser transmontanus. Journal of Fish Biology, 30(3):263–271, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:1987:PSS

[1098] N. A. Hvidsten and P. I. Møkkelgjerd. Predation on salmon smolts, Salmo salar L., in the estuary of the River Surna, Norway. Journal of Fish Biology, 30(3):273–280, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Hassan:1987:PIP

[1099] L. A. J. Al-Hassan, C. J. Webb, M. Giama, and P. J. Miller. Phosphoglucose isomerase polymorphism in the common goby, Pomatoschistus microps (Krøyer) (Teleosteir Gobiidae), around the British Isles. Journal of Fish Biology, 30(3):281–298, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:1987:IFP

[1100] M. Jobling. Influences of food particle size and dietary energy content on patterns of gastric evacuation in fish: test of a physiological model of gastric emptying. Journal of Fish Biology, 30(3):299–314, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinterleitner:1987:DAO

[1101] S. Hinterleitner, U. Platzer, and W. Wieser. Development of the activities of oxidative, glycolytic and muscle enzymes during early larval life in three families of freshwater fish. Journal of Fish Biology, 30(3):315– 326, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruno:1987:SAT

[1102] D. W. Bruno. Serum agglutinating titres against Renibacterium salmoninarum the causative agent of bacterial kidney disease, in rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L. Journal of Fish Biology, 30(3):327–334, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whoriskey:1987:SET

[1103] F. G. Whoriskey and R. J. Wootton. The swimming endurance of threespine sticklebacks, Gasterosteus aculeatus L., from the Afon Rheidol, Wales. Journal of Fish Biology, 30(3):335–339, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tort:1987:CVH

[1104] L. Tort, J. J. Watson, and I. G. Priede. Changes in in vitro heart performance in rainbow trout, Salmo gairdneri Richardson, infected with Apatemon gracilis (Digenea). Journal of Fish Biology, 30(3):341–347, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LAbee-Lund:1987:FMR

[1105] J. H. L’Abée-Lund and L. A. Vøllestad. Feeding migration of roach, Rutilus rutilus (L.), in Lake Arungen, Norway. Journal of Fish Biology, 30(3):349–355, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sala:1987:PCC

[1106] R. Sala, S. Crespo, V. Martin, and O. Castell. Presence of chloride cells in gill filaments and lamellae of the skate Torpedo marmorata. Journal of Fish Biology, 30(3):357–361, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1987:PWQ

[1107] A. D. Pickering and T. G. Pottinger. Poor water quality suppresses the cortisol response of salmonid fish to handling and confinement. Journal of Fish Biology, 30(3):363–374, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Ab

[1108] Anonymous. Announcement. Journal of Fish Biology, 30(3):375, March 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garg:1987:SEP

[1109] S. K. Garg. Seasonal effects of pinealectomy on testicular recrudescence and secretory activity of seminal vesicles in the catfish Heteropneustes fossilis (Bloch). Journal of Fish Biology, 30(4):377–387, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chambers:1987:CGA

[1110] J. Chambers. The cyprinodontiform gonopodium, with an atlas of the gonopodia of the fishes of the genus Limia. Journal of Fish Biology, 30(4):389–418, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bashirullah:1987:COT

[1111] A. K. M. Bashirullah and N. E. Rado. Co-occurrence of three species of Choricotyle (Monogenea: Diclidophoridae) in the grunt Orthopristis ruber and their host-specificity. Journal of Fish Biology, 30(4):419–422, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Segner:1987:RFS

[1112] H. Segner. Response of fed and starved roach, Rutilus rutilus, to sublethal copper contamination. Journal of Fish Biology, 30(4):423–437, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Getachew:1987:SHF

[1113] T. Getachew. A study on an herbivorous fish, Oreochromis niloticus L., diet and its quality in two Ethiopian Rift Valley lakes, Awasa and Zwai. Journal of Fish Biology, 30(4):439–449, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peterson:1987:OBL

[1114] M. S. Peterson, D. E. Gustafson, Jr., and F. R. Moore. Orientation behaviour of Lepomis macrochirus Rafinesque to salinity fluctuations. Journal of Fish Biology, 30(4):451–458, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:1987:DCT

[1115] A. D. Munro and I. Singh. Diurnal changes in the territorial behaviour of the tilapia Oreochromis mossambicus (Peters). Journal of Fish Biology, 30(4):459–464, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1987:ECC

[1116] P. M. Hine and J. M. Wain. The enzyme cytochemistry and composition of elasmobranch granulocytes. Journal of Fish Biology, 30(4):465–475, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foott:1987:SOI

[1117] J. S. Foott and R. P. Hedrick. Seasonal occurrence of the infectious stage of proliferative kidney disease (PKD) and resistance of rainbow trout, Salmo gairdneri Richardson, to reinfection. Journal of Fish Biology, 30(4):477–483, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:1987:ISE

[1118] P. Thomas. Influence of some environmental variables on the ascorbic acid status of striped mullet, Mugil cephalus Linn., tissues. III. Effects of exposure to oil. Journal of Fish Biology, 30(4):485–494, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindsay:1987:SAC

[1119] G. J. H. Lindsay. Seasonal activities of chitinase and chitobiase in the digestive tract and serum of cod, Gadus morhua (L.). Journal of Fish Biology, 30(4):495–500, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromley:1987:EFT

[1120] P. J. Bromley. The effects of food type, meal size and body weight on digestion and gastric evacuation in turbot, Scophthalmus maximus L. Journal of Fish Biology, 30(4):501–512, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRa

[1121] Anonymous. Book reviews. Journal of Fish Biology, 30(4):513–517, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Ac

[1122] Anonymous. Addendum. Journal of Fish Biology, 30(4):517, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRb

[1123] Anonymous. Books received. Journal of Fish Biology, 30(4):517, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Ad

[1124] Anonymous. Announcement. Journal of Fish Biology, 30(4):519, April 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rousch:1987:EEC

[1125] R. Röusch. Effect of experimental conditions on the stomach evacuation of Coregonus lavaretus L. Journal of Fish Biology, 30(5):521–531, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schwanck:1987:LPS

[1126] E. Schwanck. Lunar periodicity in the spawning of Tilapia mariae in the Ethiop River, Nigeria. Journal of Fish Biology, 30(5):533–537, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Overnell:1987:LLM

[1127] J. Overnell, R. McIntosh, and T. C. Fletcher. The levels of liver metallothionein and zinc in plaice, Pleuronectes platessa L., during the breeding season, and the effect of oestradiol injection. Journal of Fish Biology, 30(5):539–546, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1987:CUEa

[1128] P. M. Hine and J. M. Wain. Composition and ultrastructure of elasmobranch granulocytes. I. Dogfishes (Squaliformes). Journal of Fish Biology, 30(5):547–556, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1987:CUEb

[1129] P. M. Hine and J. M. Wain. Composition and ultrastructure of elasmobranch granulocytes. II. Rays (Rajiformes). Journal of Fish Biology, 30(5):557–565, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1987:CUEc

[1130] P. M. Hine and J. M. Wain. Composition and ultrastructure of elasmobranch granulocytes. III. Sharks (Lamniformes). Journal of Fish Biology, 30(5):567–576, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dheer:1987:HHR

[1131] J. M. S. Dheer, T. R. Dheer, and C. L. Mahajan. Haematological and haematopoietic response to acid stress in an air-breathing freshwater fish, Channa punctatus Bloch. Journal of Fish Biology, 30(5):577–588, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodall:1987:QHS

[1132] J. A. Goodall, M. F. Capra, and A. W. Blackshaw. Quantitative histology of the seasonal changes in the gonads of summer whiting, Sillago ciliata. Journal of Fish Biology, 30(5):589–595, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eiras:1987:OHB

[1133] J. C. Eiras. Occurrence of Haemogregarina bigemina (Protozoa: Apicomplexa) in Blennius pholis (Pisces: Blenniidae) along the Portuguese west coast. Journal of Fish Biology, 30(5):597–603, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruslea:1987:SIH

[1134] S. Brusléa. Sex-inversion of the hermaphroditic, protogynous teleost Coris julis L. (Labridae). Journal of Fish Biology, 30(5):605–616, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1987:SGD

[1135] R. M. Wright and N. Giles. The survival, growth and diet of pike fry, Esox lucius L., stocked at different densities in experimental ponds. Journal of Fish Biology, 30(5):617–629, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1987:CBR

[1136] N. Giles. A comparison of the behavioural responses of parasitized and non-parasitized three-spined sticklebacks, Gasterosteus aculeatus L., to progressive hypoxia. Journal of Fish Biology, 30(5):631–638, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRc

[1137] Anonymous. Book reviews. Journal of Fish Biology, 30(5):639–641, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Ae

[1138] Anonymous. Announcement. Journal of Fish Biology, 30(5):642, May 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1987:EIN

[1139] M. P. Burton and D. R. Idler. An experimental investigation of the non-reproductive, post-mature state in winter flounder. Journal of Fish Biology, 30(6):643–650, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rohde:1987:DPS

[1140] K. Rohde. Different populations of Scomber australasicus in New Zealand and south-eastern Australia, demonstrated by a simple method using monogenean sclerites. Journal of Fish Biology, 30(6):651–657, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wales:1987:BPE

[1141] N. A. M. Wales and T. Gaunt. Blood pressure, electrolyte composition and renal studies of pike, Esox lucius, following administration of salmon calcitonin. Journal of Fish Biology, 30(6):659–665, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hilden:1987:SLP

[1142] M. Hildén and J.-P. Hirvi. The survival of larval perch, Perca fluviatilis L., under different combinations of acidity and duration of acid conditions, analysed with a generalized linear model. Journal of Fish Biology, 30(6):667–677, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:1987:MAF

[1143] G. Ward and G. J. FitzGerald. Male aggression and female mate choice in the threespine stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 30(6):679–690, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dalzell:1987:NBS

[1144] P. Dalzell. Notes on the biology of Spratelloides lewisi (Wongratana, 1983), a recently described species of sprat from Papua New Guinea waters. Journal of Fish Biology, 30(6):691–700, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1987:CCP

[1145] A. D. Pickering and T. G. Pottinger. Crowding causes prolonged leucopenia in salmonid fish, despite interrenal acclimation. Journal of Fish Biology, 30(6):701–712, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:1987:MFA

[1146] L. P. Hansen, K. B. Døsving, and B. Jonsson. Migration of farmed adult Atlantic salmon with and without olfactory sense, released on the Norwegian coast. Journal of Fish Biology, 30(6):713–721, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakanishi:1987:KTI

[1147] T. Nakanishi. Kinetics of transfer of immunity by immune leucocytes and PFC response to HRBC in isogeneic ginbuna crucian carp. Journal of Fish Biology, 30(6):723–729, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yan:1987:SMM

[1148] H. Y. Yan. Size at maturity in male Gambusia heterochir. Journal of Fish Biology, 30(6):731–741, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perry:1987:POE

[1149] D. M. Perry. A procedure for obtaining erythrocytes from larval fish for cytological study and a description of larval blood of red hake, Urophycis chuss (Walbaum) and Atlantic mackerel, Scomber scombrus (Linnaeus). Journal of Fish Biology, 30(6):743–748, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yu:1987:CBR

[1150] K. L. Yu and N. Y. S. Woo. Changes in blood respiratory properties and cardiovascular function during acute exposure to hypoxic water in an air-breathing teleost, Channa maculata. Journal of Fish Biology, 30 (6):749–760, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:Af

[1151] Anonymous. Announcement. Journal of Fish Biology, 30(6):761, June 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hislop:1987:OSD

[1152] J. R. G. Hislop and M. A. Bell. Observations on the size, dry weight and energy content of the eggs of some demersal fish species from British marine waters. Journal of Fish Biology, 31(1):1–20, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plisetskaya:1987:PIL

[1153] E. M. Plisetskaya, E. M. Donaldson, and H. M. Dye. Plasma insulin levels during the spawning migration of the pink salmon, Oncorhynchus gorbuscha. Journal of Fish Biology, 31(1):21–26, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macdonald:1987:EME

[1154] N. L. Macdonald. An electron microscopic examination of the gastrointestinal epithelium in Dover sole, Solea solea (L.). Journal of Fish Biology, 31(1):27–36, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giles:1987:PRR

[1155] N. Giles. Predation risk and reduced foraging activity in fish: experiments with parasitized and non-parasitized three-spined sticklebacks, Gasterosteus aculeatus L. Journal of Fish Biology, 31(1):37–44, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Talesara:1987:HSM

[1156] C. L. Talesara and A. J. Urfi. A histophysiological study of muscle differentiation and growth in the common carp, Cyprinus carpio var. communis. Journal of Fish Biology, 31(1):45–54, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1987:FRD

[1157] R. N. Gibson and I. A. Ezzi. Feeding relationships of a demersal fish assemblage on the west coast of Scotland. Journal of Fish Biology, 31 (1):55–69, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wickins:1987:ESC

[1158] J. F. Wickins. Effects of size, culling and social history on growth of cultured elvers, Anguilla anguilla (L.). Journal of Fish Biology, 31(1): 71–82, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackenzie:1987:LTC

[1159] K. Mackenzie. Long-term changes in the prevalence of two helminth parasites (Cestoda: Trypanorhyncha) infecting marine fish. Journal of Fish Biology, 31(1):83–87, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:1987:EEJ

[1160] P. B. Taylor. Experimental evidence for juvenile Chinook salmon, Oncorhynchus tschawytscha Walbaum, orientation at night and in sunlight after a 7^{o}change in latitude. Journal of Fish Biology, 31(1):89–111, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berg:1987:MST

[1161] O. K. Berg and M. Berg. Migrations of sea trout, Salmo trutta L., from the Vardnes river in northern Norway. Journal of Fish Biology, 31(1): 113–121, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aho:1987:CPT

[1162] J. M. Aho and C. R. Kennedy. Circulation pattern and transmission dynamics of the supra-population of the nematode Cystidicoloides tenuissima (Zeder) in the River Swincombe, England. Journal of Fish Biology, 31(1):123–141, July 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Myers:1987:MAA

[1163] R. A. Myers and J. A. Hutchings. Mating of anadromous Atlantic salmon, Salmo salar L., with mature male parr. Journal of Fish Biology, 31(2): 143–146, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doggett:1987:CLM

[1164] T. A. Doggett, A. B. Wrathmell, and J. E. Harris. A cytochemical and light microscopical study of the peripheral blood leucocytes of Oreochromis mossambicus, Cichlidae. Journal of Fish Biology, 31(2):147– 153, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rossi:1987:OMY

[1165] R. Rossi, M. Bianchini, A. Carrieri, and P. Franzoi. Observations on movements of yellow eels, Anguilla anguilla L., after displacement from coastal waters to sea. Journal of Fish Biology, 31(2):155–164, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yu:1987:MAA

[1166] K. L. Yu and N. Y. S. Woo. Metabolic adjustments of an air-breathing teleost, Channa maculata, to acute and prolonged exposure to hypoxic water. Journal of Fish Biology, 31(2):165–175, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatner:1987:APS

[1167] M. F. Tatner, A. Adams, and W. Leschen. An analysis of the primary and secondary antibody response in intact and thymectomized rainbow trout, Salmo gairdneri Richardson, to human gamma globulin and Aeromonas salmonicida. Journal of Fish Biology, 31(2):177–195, August 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thuvander:1987:PCB

[1168] A. Thuvander, L. Norrgren, and C. Fossum. Phagocytic cells in blood from rainbow trout, Salmo gairdneri (Richardson), characterized by flow cytometry and electron microscopy. Journal of Fish Biology, 31(2):197– 208, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadler:1987:SEG

[1169] K. Sadler and S. Lynam. Some effects on the growth of brown trout from exposure to aluminium at different pH levels. Journal of Fish Biology, 31(2):209–219, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1987:UAC

[1170] T. A. Anderson. Utilization of algal cell fractions by the marine herbivore the luderick, Girella tricuspidata (Quoy and Gaimard). Journal of Fish Biology, 31(2):221–228, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rimmer:1987:FMD

[1171] D. W. Rimmer and W. J. Wiebe. Fermentative microbial digestion in herbivorous fishes. Journal of Fish Biology, 31(2):229–236, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bray:1987:SHP

[1172] R. A. Bray. A study of the helminth parasites of Anarhichas lupus (Perciformes: Anarhichadidae) in the North Atlantic. Journal of Fish Biology, 31(2):237–264, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knights:1987:ABG

[1173] B. Knights. Agonistic behaviour and growth in the European eel, Anguilla anguilla L., in relation to warm-water aquaculture. Journal of Fish Biology, 31(2):265–276, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Howell:1987:ISE

[1174] B. R. Howell and A. V. M. Canario. The influence of sand on the estimation of resting metabolic rate of juvenile sole, Solea solea (L.). Journal of Fish Biology, 31(2):277–280, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1987:VTD

[1175] P. A. Henderson. The vertical and transverse distribution of larval herring in the River Blackwater estuary, Essex. Journal of Fish Biology, 31 (2):281–290, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eiras:1987:DHB

[1176] J. C. Eiras. Detection of Haemogregarina bigemina (Protozoa: Apicomplexa) infection in fish erythrocytes by phase-contrast microscopy. Journal of Fish Biology, 31(2):291–292, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRd

[1177] Anonymous. Book reviews. Journal of Fish Biology, 31(2):293–300, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRe

[1178] Anonymous. Books received. Journal of Fish Biology, 31(2):300, August 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cala:1987:ARC

[1179] P. Cala. Aerial respiration in the catfish, Eremophilus mutisii (Trichomycteridae, Siluriformes), in the Rio Bogota Basin, Colombia. Journal of Fish Biology, 31(3):301–303, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rehnberg:1987:OSC

[1180] B. G. Rehnberg and C. B. Schreck. Olfactory sensitivity of Coho salmon, Oncorhynchus kisutch (Walbaum): can salinity play a role? Journal of Fish Biology, 31(3):305–307, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scrimgeour:1987:DFR

[1181] G. J. Scrimgeour and M. J. Winterbourn. Diet, food resource partitioning and feeding periodicity of two riffle-dwelling fish species in a New Zealand river. Journal of Fish Biology, 31(3):309–324, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosenblum:1987:GME

[1182] P. M. Rosenblum, J. Pudney, and I. P. Callard. Gonadal morphology, enzyme histochemistry and plasma steroid levels during the annual reproductive cycle of male and female brown bullhead catfish, Ictalurus nebulosus Lesueur. Journal of Fish Biology, 31(3):325–341, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchin:1987:FLH

[1183] D. Minchin. Fishes of the Lough Hyne marine reserve. Journal of Fish Biology, 31(3):343–352, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davies:1987:CSM

[1184] P. E. Davies and R. D. Sloane. Characteristics of the spawning migrations of brown trout, Salmo trutta L., and rainbow trout, S. gairdneri Richardson, in Great Lake, Tasmania. Journal of Fish Biology, 31(3): 353–373, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bennett:1987:ESEa

[1185] R. O. Bennett and R. E. Wolke. The effect of sublethal endrin exposure on rainbow trout, Salmo gairdneri Richardson. I. Evaluation of serum cortisol concentrations and immune responsiveness. Journal of Fish Biology, 31(3):375–385, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bennett:1987:ESEb

[1186] R. O. Bennett and R. E. Wolke. The effect of sublethal endrin exposure on rainbow trout, Salmo gairdneri Richardson. II. The effect of altering serum cortisol concentrations on the immune response. Journal of Fish Biology, 31(3):387–394, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rhodes:1987:ECH

[1187] L. D. Rhodes, M. S. Myers, W. D. Gronlund, and B. B. Mccain. Epizootic characteristics of hepatic and renal lesions in English sole, Parophrys vetulus, from Puget Sound. Journal of Fish Biology, 31(3):395–407, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schulz:1987:MUT

[1188] U. Schulz and R. Berg. The migration of ultrasonic-tagged bream, Abramis brama (L), in Lake Constance (Bodensee–Untersee). Journal of Fish Biology, 31(3):409–414, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1987:FPS

[1189] A. W. H. Turnpenny, K. Sadler, R. J. Aston, A. G. P. Milner, and S. Lynam. The fish populations of some streams in Wales and northern England in relation to acidity and associated factors. Journal of Fish Biology, 31(3):415–434, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Micale:1987:GCC

[1190] V. Micale, F. Perdichizzi, and G. Santangelo. The gonadal cycle of captive white bream, Diplodus sargus (L.). Journal of Fish Biology, 31(3): 435–440, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:1987:STT

[1191] C. M. Huang, H. J. Cheng, S. L. Chang, N. H. Chao, and I. C. Liao. Siamese twins in tilapia. Journal of Fish Biology, 31(3):441–442, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1987:APF

[1192] E. A. R. Brown and D. B. C. Scott. Abnormal pelvic fins in Scottish powan, Coregonus lavaretus (L.) (Salmonidae, Coregoninae). Journal of Fish Biology, 31(3):443–444, September 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stene:1987:LMS

[1193] A. Stene. Light microscopical studies of chromosomes in embryos of cod, Gadus morhua L. Journal of Fish Biology, 31(4):445–450, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:1987:NGS

[1194] N. R. Merrett and J. G. Nielsen. A new genus and species of the family Ipnopidae (Pisces, Teleostei) from the eastern North Atlantic, with notes on its ecology. Journal of Fish Biology, 31(4):451–464, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brodeur:1987:DFC

[1195] R. D. Brodeur and W. G. Pearcy. Diel feeding chronology, gastric evacuation and estimated daily ration of juvenile Coho salmon, Oncorhynchus kisutch (Walbaum), in the coastal marine environment. Journal of Fish Biology, 31(4):465–477, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:1987:MDR

[1196] S. J. Hall. Maximum daily ration and the pattern of food consumption in haddock, Melanogrammus aeglefinus (L.), and dab, Limanda limanda (L.). Journal of Fish Biology, 31(4):479–491, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunjak:1987:FES

[1197] R. A. Cunjak and G. Power. The feeding and energetics of streamresident trout in winter. Journal of Fish Biology, 31(4):493–511, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morris:1987:TCA

[1198] K. H. Morris and P. S. Maitland. A trap for catching adult lampreys (Petromyzonidae) in running water. Journal of Fish Biology, 31(4):513– 516, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chubb:1987:KSC

[1199] J. C. Chubb, D. W. Pool, and C. J. Veltkamp. A key to the species of cestodes (tapeworms) parasitic in British and Irish freshwater fishes. Journal of Fish Biology, 31(4):517–543, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramnarine:1987:IRS

[1200] I. W. Ramnarine, J. M. Pirie, A. D. F. Johnstone, and G. W. Smith. The influence of ration size and feeding frequency on ammonia excretion by juvenile Atlantic cod, Gadus morhua L. Journal of Fish Biology, 31 (4):545–559, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sayer:1987:RIG

[1201] M. D. J. Sayer and J. Davenport. The relative importance of the gills to ammonia and urea excretion in five seawater and one freshwater teleost species. Journal of Fish Biology, 31(4):561–570, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1987:LTS

[1202] C. R. Kennedy. Long-term stability in the population levels of the eyefluke Tylodelphys podicipina (Digenea: Diplostomatidae) in perch. Journal of Fish Biology, 31(4):571–581, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:1987:BCK

[1203] R. W. G. White, Y. A. E. Bick, and Fu Wenqing. Brief communications: Karyology of the genus Paragalaxias (Pisces, Salmoniformes, Galaxiidae). Journal of Fish Biology, 31(4):583–585, October 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williamson:1987:VDP

[1204] G. R. Williamson. Vertical drifting position of glass eels, Anguilla rostrata, off Newfoundland. Journal of Fish Biology, 31(4):587–588, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRf

[1205] Anonymous. Book reviews. Journal of Fish Biology, 31(4):589–590, October 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zaman:1987:OCC

[1206] Z. Zaman and T. S. Leong. Occurrence of the caryophyllid cestode Lytocestus parvulus Furtado, 1963 in Clarias batrachus (L.) in a tropical environment, Kedah, Malaysia. Journal of Fish Biology, 31(5):591–596, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carbone:1987:CSD

[1207] P. Carbone, R. Vitturi, E. Catalano, and M. Macaluso. Chromosome sex determination and Y-autosome fusion in Blennius tentacularis Brunnich, 1765 (Pisces, Blennidae). Journal of Fish Biology, 31(5):597–602, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:1987:GMI

[1208] B. E. Marshall. Growth and mortality of the introduced Lake Tanganyika clupeid, Limnothrissa miodon, in Lake Kariba. Journal of Fish Biology, 31(5):603–615, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gehrke:1987:CRM

[1209] P. C. Gehrke. Cardio-respiratory morphometrics of spangled perch, Leiopotherapon unicolor (Günther, 1859), (Percoidei, Teraponidae). Journal of Fish Biology, 31(5):617–623, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Howes:1987:PBA

[1210] G. J. Howes. The palatine bone and its associations in gadoid fishes. Journal of Fish Biology, 31(5):625–637, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atkinson:1987:BPA

[1211] R. J. A. Atkinson, B. Pelster, C. R. Bridges, A. C. Taylor, and S. Morris. Behavioural and physiological adaptations to a burrowing lifestyle in the snake blenny, Lumpenus lampretaeformis, and the red band-fish, Cepola rubescens. Journal of Fish Biology, 31(5):639–659, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagar:1987:PPR

[1212] P. M. Sagar and G. J. Glova. Prey preferences of a riverine population of juvenile Chinook salmon, Oncorhynchus tshawytscha. Journal of Fish Biology, 31(5):661–673, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pipe:1987:ETD

[1213] R. K. Pipe and P. Walker. The effect of temperature on development and hatching of scad, Trachurus trachurus L., eggs. Journal of Fish Biology, 31(5):675–682, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barkman:1987:RDG

[1214] R. C. Barkman and D. A. Bengtson. The record of daily growth in otoliths of Atlantic silversides, Menidia menidia, from field and laboratory. Journal of Fish Biology, 31(5):683–695, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bengtson:1987:RBM

[1215] D. A. Bengtson, R. C. Barkman, and W. J. Berry. Relationships between maternal size, egg diameter, time of spawning season, temperature, and length at hatch of Atlantic silverside, Menidia menidia. Journal of Fish Biology, 31(5):697–704, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munshi:1987:MOP

[1216] J. S. D. Munshi and G. M. Hughes. Microvascular organization of the pseudobranch and carotid labyrinth of two species of air-breathing teleostean fishes. Journal of Fish Biology, 31(5):705–714, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lang:1987:MPP

[1217] C. Lang. Mortality of perch, Perca fluviatilis L., estimated from the size and abundance of egg strands. Journal of Fish Biology, 31(5):715–720, November 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crivelli:1987:BBF

[1218] A. J. Crivelli and F. Dupont. Biometrical and biological features of Alburnus alburnus × Rutilus rubilio natural hybrids from Lake Mikri Prespa, northern Greece. Journal of Fish Biology, 31(6):721–733, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barron:1987:TDC

[1219] M. G. Barron, B. D. Tarr, and W. L. Hayton. Temperature-dependence of cardiac output and regional blood flow in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 31(6):735–744, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1987:ABH

[1220] E. J. Hyslop. Aspects of the biology of Hemichromis bimaculatus from a small stream in Nigeria. Journal of Fish Biology, 31(6):745–751, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leuven:1987:IAE

[1221] R. S. E. W. Leuven and F. G. F. Oyen. Impact of acidification and eutrophication on the distribution of fish species in shallow and lentic soft waters of The Netherlands: an historical perspective. Journal of Fish Biology, 31(6):753–774, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Livingston:1987:MSS

[1222] M. E. Livingston. Morphological and sensory specializations of five New Zealand flatfish species, in relation to feeding behaviour. Journal of Fish Biology, 31(6):775–795, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:1987:UTD

[1223] P. G. Bianco, A. M. Bullock, P. J. Miller, and F. R. Roubal. A unique teleost dermal organ in a new European genus of fishes (Teleostei: Gobioidei). Journal of Fish Biology, 31(6):797–803, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brewster:1987:EMC

[1224] B. Brewster. Eye migration and cranial development during flatfish metamorphosis: a reappraisal (Teleostei: Pleuronectiformes). Journal of Fish Biology, 31(6):805–833, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Douglas:1987:OLD

[1225] R. H. Douglas. Ocular lens diameter as an indicator of age in brown trout, Salmo trutta. Journal of Fish Biology, 31(6):835–836, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:BRg

[1226] Anonymous. Book reviews. Journal of Fish Biology, 31(6):837–841, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1987:PR

[1227] Anonymous. Publications received. Journal of Fish Biology, 31(6):841, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jurd:1987:HFR

[1228] R. D. Jurd. Hypersensitivity in fishes: a review. Journal of Fish Biology, 31(sA):1–7, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingram:1987:HAS

[1229] G. A. Ingram. Haemolytic activity in the serum of brown trout, Salmo trutta. Journal of Fish Biology, 31(sA):9–17, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkoosh:1987:EEA

[1230] M. R. Arkoosh and S. L. Kaattari. Effect of early aflatoxin B_{1} exposure on in vivo and in vitro antibody responses in rainbow trout, Salmo gairdneri. Journal of Fish Biology, 31(sA):19–22, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doggett:1987:OGA

[1231] T. A. Doggett and J. E. Harris. The ontogeny of gut-associated lymphoid tissue in Oreochromismossambicus. Journal of Fish Biology, 31(sA):23– 27, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaastrup:1987:DSM

[1232] P. Kaastrup and V. Hørlyck. Development of a simple method to optimize the conditions for producing gynogenetic offspring, using albino rainbow trout, Salmo gairdneri Richardson, females as an indicator for gynogenesis. Journal of Fish Biology, 31(sA):29–33, December 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakanishi:1987:HAT

[1233] T. Nakanishi. Histocompatibility analyses in tetraploids induced from clonal triploid crucian carp and in gynogenetic diploid goldfish. Journal of Fish Biology, 31(sA):35–40, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finco-Kent:1987:PCN

[1234] D. Finco-Kent and R. L. Thune. Phagocytosis by catfish neutrophils. Journal of Fish Biology, 31(sA):41–49, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chung:1987:ART

[1235] S. Chung and C. J. Secombes. Activation of rainbow trout macrophages. Journal of Fish Biology, 31(sA):51–56, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siwicki:1987:PAN

[1236] A. Siwicki and M. Studnicka. The phagocytic ability of neutrophils and serum lysozyme activity in experimentally infected carp, Cyprinus carpio L. Journal of Fish Biology, 31(sA):57–60, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wishkovsky:1987:VSP

[1237] A. Wishkovsky, B. S. Roberson, and F. M. Hetrick. In vitro suppression of the phagocytic response of fish macrophages by tetracyclines. Journal of Fish Biology, 31(sA):61–65, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cossarini-Dunier:1987:EPA

[1238] M. Cossarini-Dunier. Effects of the pesticides atrazine and lindane and of manganese ions on cellular immunity of carp, Cyprinus carpio. Journal of Fish Biology, 31(sA):67–73, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kranz:1987:ESC

[1239] H. Kranz and J. Gercken. Effects of sublethal concentrations of potassium dichromate on the occurrence of splenic melano-macrophage centres in juvenile plaice, Pleuronectes platessa, L. Journal of Fish Biology, 31 (sA):75–80, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fletcher:1987:MIE

[1240] T. C. Fletcher and A. White. Metabolic and immunological effects of endotoxin in the plaice, Pleuronectes platessa L. Journal of Fish Biology, 31(sA):81–85, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitao:1987:IAP

[1241] T. Kitao, T. Yoshida, D. P. Anderson, O. W. Dixon, and A. Blanch. Immunostimulation of antibody-producing cells and humoral antibody to fish bacterins by a biological response modifier. Journal of Fish Biology, 31(sA):87–91, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salomoni:1987:EDS

[1242] C. Salomoni, M. Fiorentino, and D. L. Palenzona. Effects of diet, shamimmunization and bleeding on the immune response kinetics for two strains of Cyprinus carpio. Journal of Fish Biology, 31(sA):93–99, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wishkovsky:1987:IIB

[1243] A. Wishkovsky and R. R. Avtalion. Induction of immunosuppression to bacterial antigens in carp, Cyprinus carpio. Journal of Fish Biology, 31(sA):101–106, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowley:1987:PSE

[1244] A. F. Rowley, S. E. Barrow, and T. C. Hunt. Preliminary studies on eicosanoid production by fish leucocytes, using GC-mass spectrometry. Journal of Fish Biology, 31(sA):107–111, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Angelidis:1987:SRT

[1245] P. Angelidis, F. Baudin-Laurencin, and P. Youinou. Stress in rainbow trout, Salmo gairdneri: effects upon phagocyte chemiluminescence, circulating leucocytes and susceptibility to Aeromonas salmonicida. Journal of Fish Biology, 31(sA):113–122, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thomas:1987:ECI

[1246] P. Thomas and D. H. Lewis. Effects of cortisol on immunity in red drum, Sciaenops ocellatus. Journal of Fish Biology, 31(sA):123–127, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaattari:1987:CMG

[1247] S. L. Kaattari and R. A. Tripp. Cellular mechanisms of glucocorticoid immunosuppression in salmon. Journal of Fish Biology, 31(sA):129–132, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pourreau:1987:MMP

[1248] C. N. Pourreau, M. B. H. Koopman, G. F. R. Hendriks, D. Evenberg, and W. B. van Muiswinkel. Modulation of the mitogenic PHA response of carp, Cyprinus carpio L., by extracellular products of Aeromonas salmonicida. Journal of Fish Biology, 31(sA):133–143, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vogelbein:1987:SDM

[1249] W. K. Vogelbein, J. W. Fournie, and R. M. Overstreet. Sequential development and morphology of experimentally induced hepatic melanomacrophage centres in Rivulus marmoratus. Journal of Fish Biology, 31(sA):145–153, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laudan:1987:IEM

[1250] R. Laudan, J. S. Stolen, and A. Cali. The immunomodulating effect of the microsporidan Glugea stephani on the humoral response and immunoglobulin levels in winter flounder, Pseudopleuronectes americanus. Journal of Fish Biology, 31(sA):155–160, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schmale:1987:IRB

[1251] M. C. Schmale and E. C. Mckinney. Immune responses in the bicolor damselfish, Pomacentrus partitus, and their potential role in the development of neurogenic tumours. Journal of Fish Biology, 31(sA):161–166, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thompson:1987:CTN

[1252] J. S. Thompson, A. A. I. Kostiala, and M. Miettinen. Cutaneous tumour of northern pike Esox lucius L.: evidence for a monocytic neoplasm. Journal of Fish Biology, 31(sA):167–173, December 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:1987:IRT

[1253] B. P. Wood and R. A. Matthews. The immune response of the thicklipped grey mullet, Chelon labrosus (Risso, 1826), to metacercarial infections of Cryptocotyle lingua (Creplin, 1825). Journal of Fish Biology, 31(sA):175–183, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whyte:1987:CDD

[1254] S. K. Whyte, J. C. Allan, C. J. Secombes, and L. H. Chappell. Cercariae and diplostomules of Diplostomum spathaceum (Digenea) elicit an immune response in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 31(sA):185–190, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turaga:1987:BKD

[1255] P. Turaga, G. Wiens, and S. Kaattari. Bacterial kidney disease: the potential role of soluble protein antigen(s). Journal of Fish Biology, 31(sA):191–194, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grayson:1987:EIA

[1256] T. H. Grayson, R. J. Williams, A. B. Wrathmell, C. B. Munn, and J. E. Harris. Effects of immunopotentiating agents on the immune response of rainbow trout, Salmo gair dneri Richardson, to ERM vaccine. Journal of Fish Biology, 31(sA):195–202, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:1987:VRI

[1257] T. G. Clark, H. W. Dickerson, J. B. Gratzek, and R. C. Findly. In vitro response of Ichthyophthirius multifiliis to sera from immune channel catfish. Journal of Fish Biology, 31(sA):203–208, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zapata:1987:EME

[1258] A. G. Zapata, M. Torroba, F. Alvarez, D. P. Anderson, O. W. Dixon, and M. Wisniewski. Electron microscopic examination of antigen uptake by salmonid gill cells after bath immunization with a bacterin. Journal of Fish Biology, 31(sA):209–217, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mclean:1987:IPA

[1259] E. Mclean and R. Ash. Intact protein (antigen) absorption in fishes: mechanism and physiological significance. Journal of Fish Biology, 31(sA):219–223, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1987:PMU

[1260] A. M. Smith, O. L. Goldring, and G. Dear. The production and methods of use of polyclonal antisera to the pathogenic organisms Aeromonas salmonicida, Yersinia ruckeri and Renibacterium salmoninarum. Journal of Fish Biology, 31(sA):225–226, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mcdowall:1987:SCC

[1261] A. A. Mcdowall, L. J. Healey, and M. J. Manning. Separation of carp, Cyprinus carpio L., pronephric cells on discontinuous bovine serum albumin density gradients. Journal of Fish Biology, 31(sA):227–229, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

ONeill:1987:IRA

[1262] J. G. O’Neill, M. G. White, T. A. Sims, and D. L. Barber. The inflammatory response of the Antarctic silverfish, Pleuragramma antarcticum Boulenger, 1902, to a plerocercoid (species unknown) infestation. Journal of Fish Biology, 31(sA):231–232, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bigaj:1987:LOG

[1263] J. Bigaj, J. Dulak, and B. Plytycz. Lymphoid organs of Gasterosteus aculeatus. Journal of Fish Biology, 31(sA):233–234, December 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herraez:1987:TII

[1264] M. P. Herraez and A. Zapata. Trapping of intraperitoneal-injected Yersinia ruckeri in the lymphoid organs of Carassius auratus: the role of melano-macrophage centres. Journal of Fish Biology, 31(sA):235–237, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parkinson:1987:CLT

[1265] C. Parkinson and C. Agius. A cell line from the tilapia Oreochromis spilurus Günther. Journal of Fish Biology, 31(sA):239–240, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clayton:1987:SIR

[1266] G. M. Clayton and D. J. Price. Standardization of infection and response to white spot, Ichthyophthirius multifiliis, in fish. Journal of Fish Biology, 31(sA):241–242, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wishkovsky:1987:EFA

[1267] A. Wishkovsky, N. Garber, and R. R. Avtalion. The effect of fish age on the mortality caused by selected fish pathogens. Journal of Fish Biology, 31(sA):243–244, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siwicki:1987:IAL

[1268] A. Siwicki. Immunomodulating activity of levamisole in carp spawners, Cyprinus carpio L. Journal of Fish Biology, 31(sA):245–246, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Angelidis:1987:LIR

[1269] P. Angelidis, F. Baudin-Laurencin, C. Quentel, and P. Youinou. Lower immune response induced by PKD. Journal of Fish Biology, 31(sA): 247–250, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walczak:1987:PFM

[1270] B. Z. Walczak, B. R. Blunt, and P. V. Hodson. Phagocytic function of monocytes and haematological changes in rainbow trout injected intraperitoneally with benzo(a)-pyrene (B(a)P) and benzo(a)anthracene (B(a)A). Journal of Fish Biology, 31(sA):251–253, December 1987. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mangunwiryo:1987:PIP

[1271] H. Mangunwiryo and C. Agius. Pathogenicity of infectious pancreatic necrosis (IPN) virus to tilapia and its immune response. Journal of Fish Biology, 31(sA):255–256, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daud:1987:EIC

[1272] H. M. Daud and C. Agius. Experimental infection of common carp, Cyprinus carpio L., by infectious pancreatic necrosis (IPN) virus. Journal of Fish Biology, 31(sA):257–258, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alvarez-Pellitero:1987:SBP

[1273] P. Alvarez-Pellitero and R. M. Pintó. Some blood parameters in sea bass, Dicentrarchus labrax, infected by bacteria, virus and parasites. Journal of Fish Biology, 31(sA):259–261, December 1987. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:1988:LCL

[1274] P. G. Bianco. Leuciscus cephalm (Linnaeus), with records of fingerling adult males, Leuciscus plemobipunctatus (Stephanidis) and their hybrids from western Greece. Journal of Fish Biology, 32(1):1–16, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korwin-Kossakowski:1988:LDC

[1275] M. Korwin-Kossakowski. Larval development of carp, Cyprinus carpio L., in acidic water. Journal of Fish Biology, 32(1):17–26, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Field-Dodgson:1988:SCD

[1276] M. S. Field-Dodgson. Size characteristics and diet of emergent Chinook salmon in a small, stable, New Zealand stream. Journal of Fish Biology, 32(1):27–40, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delbeek:1988:FSF

[1277] J. C. Delbeek and D. D. Williams. Feeding selectivity of four species of sympatric stickleback in brackish-water habitats in eastern Canada. Journal of Fish Biology, 32(1):41–62, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goss:1988:EAA

[1278] G. G. Goss and C. M. Wood. The effects of acid and acid/aluminum exposure on circulating plasma cortisol levels and other blood parameters in the rainbow trout, Salmo gairdneri. Journal of Fish Biology, 32(1):63– 76, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirata:1988:IGS

[1279] T. Hirata, A. Goto, and F. Yamazaki. Individual growth and smoltification of juvenile masu salmon, Oncorhynchus masou Brevoort, under rearing conditions. Journal of Fish Biology, 32(1):77–84, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okach:1988:FFH

[1280] J. O. Okach and S. Dadzie. The food, feeding habits and distribution of a siluroid catfish, Bagrus docmac (Forsskål), in the Kenya waters of Lake Victoria. Journal of Fish Biology, 32(1):85–94, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reay:1988:IGM

[1281] P. J. Reay and V. Cornell. Identification of grey mullet (Teleostei: Mugilidae) juveniles from British waters. Journal of Fish Biology, 32(1):95–99, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1988:FLF

[1282] A. W. H. Turnpenny, C. H. Dempsey, M. H. Davis, and J. M. Fleming. Factors limiting fish populations in the Loch Fleet system, an acidic drainage system in south-west Scotland. Journal of Fish Biology, 32 (1):101–118, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kulakkattolickal:1988:RAB

[1283] A. T. Kulakkattolickal and D. L. Kramer. The role of air breathing in the resistance of bimodally respiring fish to waterborne toxins. Journal of Fish Biology, 32(1):119–127, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sandnes:1988:NRS

[1284] K. Sandnes, Ø. Lie, and R. Waagbø. Normal ranges of some blood chemistry parameters in adult farmed Atlantic salmon, Salmo salar. Journal of Fish Biology, 32(1):129–136, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ovenden:1988:ERG

[1285] J. R. Ovenden, R. W. G. White, and A. C. Sanger. Evolutionary relationships of Gadopsis spp. inferred from restriction enzyme analysis of their mitochondrial DNA. Journal of Fish Biology, 32(1):137–148, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sezaki:1988:HPE

[1286] K. Sezaki, S. Watabe, and K. Hashimoto. Haematological parameters and erythrocyte enzyme activities associated with increase in ploidy status of the spinous loach, Cobitis biwae Jordan and Snyder. Journal of Fish Biology, 32(1):149–150, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchin:1988:OSD

[1287] D. Minchin. Observations on the southern distribution of the jellycat, Anarhichas denticulatm Kröyer. Journal of Fish Biology, 32(1):151–152, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:1988:IRR

[1288] N. A. Hvidsten and L. P. Hansen. Increased recapture rate of adult Atlantic salmon, Salmo salar L., stocked as smolts at high water discharge. Journal of Fish Biology, 32(1):153–154, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:1988:EIM

[1289] G. F. Turner. Environmental influences on male breeding colouration in Oreochromis niloticus. Journal of Fish Biology, 32(1):155–156, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swinney:1988:TSB

[1290] G. N. Swinney. Two species of Bathophilus (Stomiiformes, Stomiidae) new to the north-eastern Atlantic. Journal of Fish Biology, 32(1):157– 158, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRa

[1291] Anonymous. Book review. Journal of Fish Biology, 32(1):159–160, January 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:1988:EAM

[1292] P. B. Taylor. Effects of anaesthetic MS 222 on the orientation of juvenile salmon, Oncorhynchus tschawytscha Walbaum. Journal of Fish Biology, 32(2):161–168, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Levin:1988:OSA

[1293] L. E. Levin and P. Belmonte. Oriented swimming at an angle to light in a schooling characid, Moenkhausia dichrowa Kner. Journal of Fish Biology, 32(2):169–177, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Michaud:1988:DAY

[1294] M. Michaud, J.-D. Dutil, and J. J. Dodson. Determination of the age of young American eels, Anguilla rostrata, in fresh water, based on otolith surface area and microstructure. Journal of Fish Biology, 32(2):179–189, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cataldi:1988:MCO

[1295] E. Cataldi, D. Crosetti, G. Conte, D. D’Ovidio, and S. Cataudella. Morphological changes in the oesophageal epithelium during adaptation to salinities in Oreochromis mossambicus, O. niloticus and their hybrid. Journal of Fish Biology, 32(2):191–196, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dheer:1988:HHB

[1296] J. M. S. Dheer. Haematological, haematopoietic and biochemical responses to thermal stress in an air-breathing freshwater fish, Channa punctatus Bloch. Journal of Fish Biology, 32(2):197–206, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Obiekezie:1988:DAE

[1297] A. I. Obiekezie, H. Möer, and K. Anders. Diseases of the African estuarine catfish Chrysichthys nigrodigitatus (Lacépède) from the Cross River estuary, Nigeria. Journal of Fish Biology, 32(2):207–221, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mudarris:1988:QQS

[1298] M. Mudarris and B. Austin. Quantitative and qualitative studies of the bacterial microflora of turbot, Scophthalmus maximus L., gills. Journal of Fish Biology, 32(2):223–229, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saint-Paul:1988:EAO

[1299] U. Saint-Paul and G. M. Soares. Ecomorphological adaptation to oxygen deficiency in Amazon floodplains by serrasalmid fish of the genus Mylossoma. Journal of Fish Biology, 32(2):231–236, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:1988:CSA

[1300] W. E. Barber and R. J. Walker. Circuli spacing and annulus formation: is there more than meets the eye? The case for sockeye salmon, Oncorhynchus nerka. Journal of Fish Biology, 32(2):237–245, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez:1988:ETO

[1301] I. Martinez, G. Viscor, and J. Palomeque. Effects of temperature, oxygen and carbon dioxide on osmotic fragility of carp, Cyprinus carpio L., erythrocytes. Journal of Fish Biology, 32(2):247–252, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laidley:1988:SRR

[1302] C. W. Laidley, P. T. K. Woo, and J. F. Leatherland. The stress-response of rainbow trout to experimental infection with the blood parasite Cryptobia salmositica Katz, 1951. Journal of Fish Biology, 32(2):253–261, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacKenzie:1988:PMN

[1303] K. MacKenzie. Presumptive mycobacteriosis in North-east Atlantic mackerel, Scomber scombrus L. Journal of Fish Biology, 32(2):263–275, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tort:1988:ESC

[1304] L. Tort and P. Torres. The effects of sublethal concentrations of cadmium on haematological parameters in the dogfish, Scyliorhinus canicula. Journal of Fish Biology, 32(2):277–282, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townsend:1988:FCS

[1305] C. R. Townsend and G. Peirson. Fish community structure in lowland drainage channels. Journal of Fish Biology, 32(2):283–295, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sorensen:1988:HCG

[1306] P. W. Sorensen and N. W. Pankhurst. Histological changes in the gonad, skin, intestine and olfactory epithelium of artificially-matured male American eels, Anguilla rostrata (LeSueur). Journal of Fish Biology, 32 (2):297–307, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cabrera:1988:ASR

[1307] J. Cabrera, G. Martinez, and A. Alforo. An anomalous specimen of Rhamdia guatemalensis (Günther, 1864). Journal of Fish Biology, 32 (2):309–310, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowan:1988:ILS

[1308] G. I. McT. Cowan and E. M. Baggs. Incidences of lacustrine spawning of the ouananiche, Salmo salar, and the brook charr, Salvelinus fontinalis, on the Avalon Peninsula, Newfoundland. Journal of Fish Biology, 32 (2):311–312, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchin:1988:RDS

[1309] D. Minchin. A record of the deep-sea anglerfish, Cryptopsaras couesi Gill, from the north-eastern Atlantic. Journal of Fish Biology, 32(2): 313, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRb

[1310] Anonymous. Book reviews. Journal of Fish Biology, 32(2):315–317, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:N

[1311] Anonymous. Notice. Journal of Fish Biology, 32(2):319, February 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baksi:1988:PCE

[1312] S. M. Baksi and J. C. Means. Preparation of chromosomes from early stages of fish for cytogenetic analysis. Journal of Fish Biology, 32(3):321– 325, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:1988:WHB

[1313] E. Verspoor. Widespread hybridization between native Atlantic salmon, Salmo salar, and introduced brown trout, S. trutta, in eastern Newfoundland. Journal of Fish Biology, 32(3):327–334, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Black:1988:ECR

[1314] D. Black and R. Malcolm Love. Estimating the carbohydrate reserves in fish. Journal of Fish Biology, 32(3):335–340, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franklin:1988:XCG

[1315] C. E. Franklin and W. Davison. X-cells in the gills of an Antarctic teleost, Pagothenia borchgrevinki. Journal of Fish Biology, 32(3):341–353, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OMaoileidigh:1988:MMC

[1316] N. O’Maoileidigh, S. Cawdery, J. J. Bracken, and A. Ferguson. Morphometric, meristic character and electrophoretic analyses of two Irish populations of twaite shad, Alosa fallax (Lacépède). Journal of Fish Biology, 32(3):355–366, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prasad:1988:MGD

[1317] M. S. Prasad. Morphometrics of gills during growth and development of the air-breathing habit in Colisa fasciatus (Bloch and Schneider). Journal of Fish Biology, 32(3):367–381, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Isaac-Nahum:1988:ASB

[1318] V. J. Isaac-Nahum, R. de D. Cardoso, G. Servo, and C. L. del B. Rossi-Wongtschowski. Aspects of the spawning biology of the Brazilian sardine, Sardinella brasiliensis (Steindachner, 1879), (Clupeidae). Journal of Fish Biology, 32(3):383–396, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kramer:1988:ETI

[1319] C. R. Kramer, S. Koulish, and P. L. Bertacchi. The effects of testosterone implants on ovarian morphology in the bluehead wrasse, Thalassoma bifasciatum (Bloch) (Teleostei: Labridae). Journal of Fish Biology, 32 (3):397–407, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wyatt:1988:CEY

[1320] R. J. Wyatt. The cause of extreme year class variation in a population of roach, Rutilus rutilus L., from a eutrophic lake in southern England. Journal of Fish Biology, 32(3):409–421, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Badcock:1988:EAD

[1321] J. Badcock. Evidence for the assignment of Dolichopteryx brachyrhynchus Parr to the genus Bathylychnops Cohen (Pisces, Opisthoproctidae). Journal of Fish Biology, 32(3):423–432, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1988:CET

[1322] D. Burton. A comparison of the effects of temperature on melanophore in vitro responses to noradrenalin in two teleost species adapted to cold ocean and tropical conditions. Journal of Fish Biology, 32(3):433–442, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Darowski:1988:BMP

[1323] K. Darowski, F. Takashima, and Y. K. Law. Bioenergetic model of planktivorous fish feeding, growth and metabolism: theoretical optimum swimming speed of fish larvae. Journal of Fish Biology, 32(3):443–458, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gartner:1988:ITF

[1324] J. V. Gartner and D. E. Zwerner. An Ichthyophonus-type fungus in the deep-sea pelagic fish Scopelogadus beanii (Günther) (Pisces: Melamphaidae) pathology, geographic distribution and ecological implications. Journal of Fish Biology, 32(3):459–470, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wardle:1988:BSS

[1325] C. S. Wardle and P. He. Burst swimming speeds of mackerel, Scomber scombrus L. Journal of Fish Biology, 32(3):471–478, March 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okaeme:1988:PDC

[1326] A. N. Okaeme, A. I. Obiekezie, J. Lehman, E. E. Antai, and C. T. Madu. Parasites and diseases of cultured fish of Lake Kainji area, Nigeria. Journal of Fish Biology, 32(3):479–481, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruno:1988:CRC

[1327] D. W. Bruno and C. O. Fraser. A case of reversal in the common dab Limanda limanda (L.). Journal of Fish Biology, 32(3):483–484, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oyetayo:1988:IAE

[1328] A. S. Oyetayo. Infection of the African eleotrid Bostrychus africanus (Steindachner 1980) by the nematode Raphidascaroides africanus. Journal of Fish Biology, 32(3):485–486, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jorgensen:1988:URM

[1329] E. Jørgensen and M. Jobling. Use of radiographic methods in feeding studies: a cautionary note. Journal of Fish Biology, 32(3):487–488, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchin:1988:NRL

[1330] D. Minchin and Alwyne Wheeler. A new record for the lancetfish, Alepisaurus ferox Lowe (Osteichthys: Scopeliformes: Alepisauridae), in the north-eastern Atlantic. Journal of Fish Biology, 32(3):489–491, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytler:1988:DYS

[1331] P. Tytler and J. H. S. Blaxter. Drinking in yolk-sac stage larvae of the halibut, Hippoglossus hippoglossus (L.). Journal of Fish Biology, 32(3): 493–494, March 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tzeng:1988:DGI

[1332] W.-N. Tzeng and S.-Y. Yu. Daily growth increments in otoliths of milkfish, Chanos chanos (Forsskål), larvae. Journal of Fish Biology, 32(4): 495–504, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saglio:1988:SCE

[1333] Ph. Saglio, A. M. Escaffre, and J. M. Blanc. Structural characteristics of the epidermal mucosa in yellow and silver European eel, Anguilla anguilla (L.). Journal of Fish Biology, 32(4):505–514, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:1988:PAB

[1334] B. E. Marshall. A preliminary assessment of the biomass of the pelagic sardine Limnothrissa miodon in Lake Kariba. Journal of Fish Biology, 32(4):515–524, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mhaisen:1988:OPF

[1335] F. T. Mhaisen, N. K. Al-Salim, and N. R. Khamees. Occurrence of parasites of the freshwater mugilid fish Liza abu (Heckel) from Basrah, southern Iraq. Journal of Fish Biology, 32(4):525–532, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galloway:1988:TER

[1336] M. L. Galloway and R. V. Kilambi. Thermal enrichment of a reservoir and the effects on annulus formation and growth of largemouth bass, Micropterus salmoides. Journal of Fish Biology, 32(4):533–543, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roubal:1988:MWR

[1337] F. R. Roubal and A. M. Bullock. The mechanism of wound repair in the skin of juvenile Atlantic salmon, Salmo salar L., following hydrocortisone implantation. Journal of Fish Biology, 32(4):545–555, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanDijk:1988:EAB

[1338] P. L. M. van Dijk and C. M. Wood. The effect of β-adrenergic blockade on the recovery process after strenuous exercise in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 32(4):557–570, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ogari:1988:FNP

[1339] J. Ogari and S. Dadzie. The food of the Nile perch, Lates niloticus (L.), after the disappearance of the haplochromine cichlids in the Nyanza Gulf of Lake Victoria (Kenya). Journal of Fish Biology, 32(4):571–577, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1988:CIN

[1340] P. M. Hine and J. M. Wain. Characterization of inflammatory neutrophils induced by bacterial endotoxin in the blood of eels, Anguilla australis. Journal of Fish Biology, 32(4):579–592, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Economidis:1988:NHL

[1341] P. S. Economidis and A. I. Sinis. A natural hybrid of Leuciscus cephalus macedonicus × Chalcalburnus chalcoides macedonicus (Pisces, Cyprinidae) from Lake Volvi (Macedonia, Greece). Journal of Fish Biology, 32(4):593–605, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reader:1988:GMU

[1342] J. P. Reader, T. R. K. Dalziel, and R. Morris. Growth, mineral uptake and skeletal calcium deposition in brown trout, Salmo trutta L., yolk-sac fry exposed to aluminium and manganese in soft acid water. Journal of Fish Biology, 32(4):607–624, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shand:1988:CIF

[1343] J. Shand. Corneal iridescence in fishes: light-induced colour changes in relation to structure. Journal of Fish Biology, 32(4):625–632, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valente:1988:NFA

[1344] A. C. N. Valente. A note on fin abnormalities in Leuciscus cephalus L. and Carassius carassius L. (Pisces: Cyprinidae). Journal of Fish Biology, 32(4):633–634, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1988:DOR

[1345] J. S. Alabaster. The dissolved oxygen requirements of upstream migrant Chinook salmon, Oncorhynchus tshawytscha, in the lower Willamette River, Oregon. Journal of Fish Biology, 32(4):635–636, April 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRc

[1346] Anonymous. Book review. Journal of Fish Biology, 32(4):637–638, April 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:1988:MBT

[1347] T. Hesthagen. Movements of brown trout, Salmo trutta, and juvenile Atlantic salmon, Salmo salar, in a coastal stream in northern Norway. Journal of Fish Biology, 32(5):639–653, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodgers:1988:ARS

[1348] L. J. Rodgers and J. B. Burke. Aetiology of ‘red spot’ disease (vibriosis) with special reference to the ectoparasitic digenean Prototransversotrema steeri (Angel) and the sea mullet, Mugil cephalus (Linnaeus). Journal of Fish Biology, 32(5):655–663, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Don:1988:CSI

[1349] J. Don and R. R. Avtalion. Comparative study on the induction of triploidy in tilapias, using cold- and heat-shock techniques. Journal of Fish Biology, 32(5):665–672, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smit:1988:BCF

[1350] G. L. Smit and H. J. Schoonbee. Blood coagulation factors in the freshwater fish Oreochromis mossambicus. Journal of Fish Biology, 32(5): 673–677, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jurvelius:1988:SVC

[1351] J. Jurvelius, T. Lindem, and T. Heikkinen. The size of a vendace, Coregonus albula L., stock in a deep lake basin monitored by hydroacoustic methods. Journal of Fish Biology, 32(5):679–687, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1988:LOS

[1352] A. D. Pickering and T. G. Pottinger. Lymphocytopenia and the overwinter survival of Atlantic salmon parr, Salmo salar L. Journal of Fish Biology, 32(5):689–697, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Levesley:1988:PDR

[1353] P. B. Levesley and A. E. Magurran. Population differences in the reaction of minnows to alarm substance. Journal of Fish Biology, 32(5):699– 706, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ladich:1988:SPG

[1354] F. Ladich. Sound production by the gudgeon, Gobio gobio L., a common European freshwater fish (Cyprinidae, Teleostei). Journal of Fish Biology, 32(5):707–715, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggenes:1988:DMC

[1355] J. Heggenes and T. Traaen. Downstream migration and critical water velocities in stream channels for fry of four salmonid species. Journal of Fish Biology, 32(5):717–727, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Little:1988:FCE

[1356] M. C. Little, P. J. Reay, and S. J. Grove. The fish community of an East African mangrove creek. Journal of Fish Biology, 32(5):729–747, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:1988:ERT

[1357] Y. Cui and R. J. Wootton. Effects of ration, temperature and body size on the body composition, energy content and condition of the minnow, Phoxinus phoxinus (L.). Journal of Fish Biology, 32(5):749–764, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sayer:1988:RBN

[1358] M. D. J. Sayer and N. Blackstock. The relationship between nitrogen output and changes in mucous cell function in the amphibious teleost Blennius pholis L. during aerial exposure. Journal of Fish Biology, 32 (5):765–776, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:1988:FMR

[1359] F. A. Huntingford, N. B. Metcalfe, and J. E. Thorpe. Feeding motivation and response to predation risk in Atlantic salmon parr adopting different life history strategies. Journal of Fish Biology, 32(5):777–782, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:1988:ERS

[1360] S. J. Hall. The effects of reversal of seasonal changes in photoperiod on the growth and food consumption of cod, Gadus morhua L. Journal of Fish Biology, 32(5):783–792, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Micale:1988:PEG

[1361] V. Micale and F. Perdichizzi. Photoperiod effects on gonadal maturation in captivity-born gilthead bream, Sparus aurata (L.): early findings. Journal of Fish Biology, 32(5):793–794, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunsinger:1988:UGM

[1362] R. N. Hunsinger, B. R. Byram, and W. M. Howell. Unchanged gonadal morphology of mosquitofish masculinized by exposure to degraded plant sterols. Journal of Fish Biology, 32(5):795–796, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spanier:1988:IPT

[1363] E. Spanier and M. Goren. An Indo–Pacific trunkfish Tetrosomus gibbosus (Linnaeus): first record of the family Ostracionidae in the Mediterranean. Journal of Fish Biology, 32(5):797–798, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrett:1988:GHR

[1364] B. A. Barrett and B. A. McKeown. Growth hormone response to sustained swimming in exercise-acclimated steelhead trout, Salmo gairdneri. Journal of Fish Biology, 32(5):799–800, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchmann:1988:SDP

[1365] K. Buchmann. Spatial distribution of Pseudodactylogyms anguillae and P. bini (Monogenea) on the gills of the European eel, Anguilla anguilla. Journal of Fish Biology, 32(5):801–802, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mittal:1988:IMS

[1366] A. K. Mittal and T. K. Garg. An improved method of separation of intact epidermal sheets in fish. Journal of Fish Biology, 32(5):803–804, May 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rand:1988:BIS

[1367] T. G. Rand and M. Wiles. Bacterial involvement in a saddleback disease of the reef silverside, At henna harringtonensis Goode (Pisces: Atherinidae) from Bermudan waters. Journal of Fish Biology, 32(6): 805–816, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:1988:SPS

[1368] S. G. Nelson and S. DeC. Wilkins. Sediment processing by the surgeonfish Ctenochaetus striatus at Moorea, French Polynesia. Journal of Fish Biology, 32(6):817–824, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMahon:1988:VDS

[1369] T. E. McMahon and G. F. Hartman. Variation in the degree of silvering of wild Coho salmon, Oncorhynchus kisutch, smolts migrating seaward from Carnation Creek, British Columbia. Journal of Fish Biology, 32(6): 825–833, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Echelle:1988:PLS

[1370] A. A. Echelle, A. F. Echelle, L. E. DeBault, and D. W. Durham. Ploidy levels in silverside fishes (Atherinidae, Menidia) on the Texas coast: flowcytometric analysis of the occurrence of allotriploidy. Journal of Fish Biology, 32(6):835–844, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peleteiro:1988:ISI

[1371] M. C. Peleteiro and R. H. Richards. Immunocytochemical studies on immunoglobulin-containing cells in the epidermis of rainbow trout Salmo gairdneri Richardson: influence of bath vaccination. Journal of Fish Biology, 32(6):845–858, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santulli:1988:TCC

[1372] A. Santulli, A. Curatolo, A. Modica, and V. D’Amelio. Time-course changes of plasma lipid levels and lipoprotein pattern after feeding in cultured sea bass, Dicentrarchus labrax L. Journal of Fish Biology, 32 (6):859–867, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kazakov:1988:RAR

[1373] R. V. Kazakov, O. L. Christoforov, I. G. Murza, S. A. Ilyenkova, and S. F. Titov. Results of accelerated rearing of Atlantic salmon, Salmo salar L., smolts by use of warm waste water. Journal of Fish Biology, 32(6):869–880, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bellwood:1988:SCS

[1374] D. R. Bellwood. Seasonal changes in the size and composition of the fish yield from reefs around Apo Island, Central Philippines, with notes on methods of yield estimation. Journal of Fish Biology, 32(6):881– 893, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1988:CCJ

[1375] E. J. Hyslop. A comparison of the composition of the juvenile fish catch from the sokoto–rima floodplain, Nigeria in years preceding and immediately after upstream dam completion. Journal of Fish Biology, 32(6):

895–899, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gardner:1988:MAT

[1376] A. S. Gardner, A. F. Walker, and R. B. Greer. Morphometric analysis of two ecologically distinct forms of Arctic charr, Salvelinus alpinus (L.), in Loch Rannoch, Scotland. Journal of Fish Biology, 32(6):901–910, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:1988:EFF

[1377] T. A. Anderson. The effect of feeding frequency on utilization of algal nutrients by the marine herbivore, the luderick, Girella tricuspidata (Quoy and Gaimard). Journal of Fish Biology, 32(6):911–921, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hislop:1988:IML

[1378] J. R. G. Hislop. The influence of maternal length and age on the size and weight of the eggs and the relative fecundity of the haddock, Melanogrammus aeglefinus, in British waters. Journal of Fish Biology, 32(6):923–930, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mir:1988:NNI

[1379] S. Mir, A. Al-Absy, and F. Krupp. A new natural intergeneric cyprinid hybrid from the Jordan River drainage, with a key to the large barbine cyprinids of the southern Levant. Journal of Fish Biology, 32(6):931– 936, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Battram:1988:EAL

[1380] J. C. Battram. The effects of aluminium and low^{p} H on chloride fluxes in the brown trout, Salmo trutta L. Journal of Fish Biology, 32(6):937– 947, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:1988:DMP

[1381] R. N. Gibson. Development, morphometry and particle retention capability of the gill rakers in the herring, Clupea harengus L. Journal of Fish Biology, 32(6):949–962, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:1988:EHS

[1382] M. J. Morgan. The effect of hunger, shoal size and the presence of a predator on shoal cohesiveness in bluntnose minnows, Pimephales notatus Rafinesque. Journal of Fish Biology, 32(6):963–971, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:1988:OAG

[1383] P. G. Bianco. Occurrence of the Asiatic gobionid Pseudorasbora parva (Temminck and Schlegel) in south-eastern Europe. Journal of Fish Biology, 32(6):973–974, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinkens:1988:TBP

[1384] E. Hinkens and P. A. Cochran. Taste buds on pelvic fin rays of the burbot, Lota lota (L.). Journal of Fish Biology, 32(6):975, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bouain:1988:NTS

[1385] A. Bouain and Y. Siau. A new technique for staining fish otoliths for age determination. Journal of Fish Biology, 32(6):977–978, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRd

[1386] Anonymous. Book review. Journal of Fish Biology, 32(6):979, June 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Austin:1988:EAC

[1387] B. Austin and A. M. J. Al-Zahrani. The effect of antimicrobial compounds on the gastrointestinal microflora of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 33(1):1–14, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villarreal:1988:IPG

[1388] C. A. Villarreal, J. E. Thorp, and M. S. Miles. Influence of photoperiod on growth changes in juvenile Atlantic salmon, Salmo salar L. Journal of Fish Biology, 33(1):15–30, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bortz:1988:PTS

[1389] B. M. Bortz, G. E. Kenny, G. B. Pauley, and A. H. Bunt-Milam. Prevalence of two site-specific populations of Diplostomum spp. in eye infections of rainbow trout, Salmo gairdneri Richardson, from lakes in Washington State, U.S.A. Journal of Fish Biology, 33(1):31–43, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wyatt:1988:ECG

[1390] R. J. Wyatt and C. R. Kennedy. The effects of a change in the growth rate of roach, Rutilus rutilus (L.), on the biology of the fish tapeworm Ligula intestinalis (L.). Journal of Fish Biology, 33(1):45–57, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:1988:DVG

[1391] I. G. Cowx. Distribution and variation in the growth of roach, Rutilus rutilus (L.), and dace, Leuciscus leuciscus (L.), in a river catchment in south-west England. Journal of Fish Biology, 33(1):59–72, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laidley:1988:CSA

[1392] C. W. Laidley and J. F. Leatherland. Cohort sampling, anaesthesia and stocking-density effects on plasma cortisol, thyroid hormone, metabolite and ion levels in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 33(1):73–88, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hwang:1988:SEC

[1393] P. P. Hwang and S. M. Wu. Salinity effects on cytometrical parameters of the kidney in the euryhaline teleost Oreochromis mossambicus Peters. Journal of Fish Biology, 33(1):89–95, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greeley:1988:COG

[1394] M. S. Greeley, Jr., R. Macgregor III, and K. R. Marion. Changes in the ovary of the Gulf killifish, Fundulus grandis (Baird and Girard), during seasonal and semilunar spawning cycles. Journal of Fish Biology, 33(1): 97–107, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogers:1988:REE

[1395] S. I. Rogers. Reproductive effort and efficiency in the female common goby, Pomatoschistus microps (Krøyer) (Teleostei: Gobioidei). Journal of Fish Biology, 33(1):109–119, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:1988:PHR

[1396] N. A. Hvidsten and R. A. Lund. Predation on hatchery-reared and wild smolts of Atlantic salmon, Salmo salar L., in the estuary of River Orkla, Norway. Journal of Fish Biology, 33(1):121–126, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsen:1988:OSH

[1397] P. B. Johnsen, H. Zhou, and M. A. Adams. Olfactory sensitivity of the herbivorous grass carp, Ctenopharyngodon idella, to amino acids. Journal of Fish Biology, 33(1):127–134, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

El-Fiky:1988:LSP

[1398] N. El-Fiky and W. Wieser. Life styles and patterns of development of gills and muscles in larval cyprinids (Cyprinidae; Teleostei). Journal of Fish Biology, 33(1):135–145, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zheng:1988:MHS

[1399] W. B. Zheng and W. S. Liu. Morphology and histology of the swimbladder and infrastructure of respiratory epithelium in the air-breathing catfish, Pangasius sutchi (Pangasiidae). Journal of Fish Biology, 33(1): 147–154, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rehnberg:1988:IAS

[1400] B. G. Rehnberg and R. J. F. Smith. The influence of alarm substance and shoal size on the behaviour of zebra danios, Brachydanio rerio (Cyprinidae). Journal of Fish Biology, 33(1):155–163, July 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitear:1988:TBE

[1401] M. Whitear. ‘Tubule bodies’ in the epidermis of Ciliata mustela. Journal of Fish Biology, 33(1):165–166, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRe

[1402] Anonymous. Book reviews. Journal of Fish Biology, 33(1):167–170, July 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadler:1988:ICA

[1403] K. Sadler and S. Lynam. The influence of calcium on aluminium-induced changes in the growth rate and mortality of brown trout, Salmo trutta L. Journal of Fish Biology, 33(2):171–179, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart:1988:CCM

[1404] K. M. Stewart. Changes in condition and maturation of the Oreochromis niloticus L. population of Ferguson’s Gulf, Lake Turkana, Kenya. Journal of Fish Biology, 33(2):181–188, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gehrke:1988:IGM

[1405] P. C. Gehrke. Influence of gut morphology, sensory cues and hunger on feeding behaviour of spangled perch, Leiopotherapon unicolor (Günther, 1859), (Percoidei, Teraponidae). Journal of Fish Biology, 33(2):189– 201, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campton:1988:EIG

[1406] D. E. Campton and G. A. E. Gall. Effect of individual and group rearing on age and size at maturity of male mosquitofish, Gambusia affinis. Journal of Fish Biology, 33(2):203–212, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bellwood:1988:OCD

[1407] D. R. Bellwood. Ontogenetic changes in the diet of early post-settlement Scarus species (Pisces: Scaridae). Journal of Fish Biology, 33(2):213– 219, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1988:SSO

[1408] P. A. Henderson, R. H. A. Holmes, and R. N. Bamber. Size-selective overwintering mortality in the sand smelt, Atherina boyeri Risso, and its role in population regulation. Journal of Fish Biology, 33(2):221– 233, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1988:UCO

[1409] P. M. Hine and J. M. Wain. Ultrastructural and cytochemical observations on the granulocytes of the sturgeon, Acipenser brevirostrum (Chondrostei). Journal of Fish Biology, 33(2):235–245, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mine:1988:OGP

[1410] P. M. Mine and J. M. Wain. Observations on the granulocyte peroxidase of teleosts: a phylogenetic perspective. Journal of Fish Biology, 33(2): 247–254, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:1988:EIS

[1411] P. He and C. S. Wardle. Endurance at intermediate swimming speeds of Atlantic mackerel, Scomber scombrus L., herring, Clupea harengus L., and saithe, Pollachius virens L. Journal of Fish Biology, 33(2):255– 266, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasurer:1988:MRF

[1412] J. W. Treasurer. Measurement of regurgitation in feeding studies of predatory fishes. Journal of Fish Biology, 33(2):267–271, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthews:1988:TES

[1413] W. J. Matthews, L. G. Hill, D. R. Edds, J. J. Hoover, and T. G. Heger. Trophic ecology of striped bass, Morone saxatilis, in a freshwater reservoir (Lake Texoma, U.S.A.). Journal of Fish Biology, 33(2):273–288, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mohan:1988:TNI

[1414] M. V. Mohan and T. M. Sankaran. Two new indices for stomach content analysis of fishes. Journal of Fish Biology, 33(2):289–292, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowey:1988:SUA

[1415] C. B. Cowey and M. J. Walton. Studies on the uptake of (^{14}C) amino acids derived from both dietary (^{14}C) protein and dietary (^{14}C) amino acids by rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 33(2):293–305, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dwyer:1988:BCL

[1416] F. J. Dwyer, C. J. Schmitt, S. E. Finger, and P. M. Mehrle. Biochemical changes in longear sunfish, Lepomis megalotis, associated with lead, cadmium and zinc from mine tailings. Journal of Fish Biology, 33(2): 307–317, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korcock:1988:ESC

[1417] D. E. Korcock, A. H. Houston, and J. D. Gray. Effects of sampling conditions on selected blood variables of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 33(2):319–330, August 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromley:1988:GDE

[1418] P. J. Bromley. Gastric digestion and evacuation in whiting, Merlangius merlangus (L.). Journal of Fish Biology, 33(2):331–338, August 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorman:1988:DGB

[1419] J. A. Dorman. Diet of the garfish, Belone belone (L.), from Courtmacsherry Bay, Ireland. Journal of Fish Biology, 33(3):339–346, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggberget:1988:TSS

[1420] T. G. Heggberget, T. Haukebø, J. Mork, and G. Ståhl. Temporal and spatial segregation of spawning in sympatric populations of Atlantic salmon, Salmo salar L., and brown trout, Salmo trutta L. Journal of Fish Biology, 33(3):347–356, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trendall:1988:DDI

[1421] J. Trendall. The distribution and dispersal of introduced fish at Thumbi West Island in Lake Malaŵi, Africa. Journal of Fish Biology, 33(3): 357–369, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herrera:1988:AGR

[1422] M. Herrera, J. A. Hernando, C. Fernandez-Delgado, and M. Bellido. Age, growth and reproduction of the barbel, Barbus sclateri (Günther, 1868), in a first-order stream in southern Spain. Journal of Fish Biology, 33(3): 371–381, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gambaryan:1988:KMS

[1423] S. P. Gambaryan. Kidney morphology in sturgeons: a microdissectional and ultrastructural study. Journal of Fish Biology, 33(3):383–398, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofer:1988:MAD

[1424] R. Hofer. Morphological adaptations of the digestive tract of tropical cyprinids and cichlids to diet. Journal of Fish Biology, 33(3):399–408, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alhossaini:1988:RBD

[1425] M. Alhossaini and T. J. Pitcher. The relation between daily rings, body growth and environmental factors in plaice, Pleuronectes platessa L., juvenile otoliths. Journal of Fish Biology, 33(3):409–418, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greeley:1988:VPO

[1426] M. S. Greeley, Jr., R. Macgregor III, and K. R. Marion. Variation in plasma oestrogens and androgens during the seasonal and semilunar spawning cycles of female gulf killifish, Fundulus grandis (Baird and Girard). Journal of Fish Biology, 33(3):419–429, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:1988:BGCa

[1427] Y. Cui and R. J. Wootton. Bioenergetics of growth of a cyprinid, Phoxinus phoxinus: the effect of ration, temperature and body size on food consumption, faecal production and nitrogenous excretion. Journal of Fish Biology, 33(3):431–443, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1988:AEA

[1428] R. L. Smith, A. J. Paul, and J. M. Paul. Aspects of energetics of adult walleye pollock, Theragra chalcogramma (Pallas), from Alaska. Journal of Fish Biology, 33(3):445–454, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coates:1988:LDC

[1429] D. Coates. Length-dependent changes in egg size and fecundity in females, and brooded embryo size in males, of fork-tailed catfishes (Pisces: Ariidae) from the Sepik River, Papua New Guinea, with some implications for stock assessments. Journal of Fish Biology, 33(3):455–464, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ntiba:1988:AGP

[1430] M. J. Ntiba and V. Jaccarini. Age and growth parameters of Siganus sutor in Kenyan marine inshore water, derived from numbers of otolith microbands and fish lengths. Journal of Fish Biology, 33(3):465–470, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ogata:1988:CAA

[1431] H. Ogata and T. Murai. Changes in ammonia and amino acid levels in the erythrocytes and plasma of carp, Cyprinus carpio, during passage through the gills. Journal of Fish Biology, 33(3):471–479, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harvey:1988:DEL

[1432] B. C. Harvey, R. C. Cashner, and W. J. Matthews. Differential effects of largemouth and smallmouth bass on habitat use by stoneroller minnows in stream pools. Journal of Fish Biology, 33(3):481–487, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1988:NSS

[1433] Alwyne Wheeler. The nomenclature of the sole, Solea solea (Linnaeus, 1758). Journal of Fish Biology, 33(3):489–490, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyslop:1988:FOA

[1434] E. J. Hyslop. First occurrence of Acanthogyrus tilapiae (Baylis 1948) in Hemichromis species. Journal of Fish Biology, 33(3):491–492, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRf

[1435] Anonymous. Book reviews. Journal of Fish Biology, 33(3):493–496, September 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1988:FSB

[1436] Alwyne Wheeler. The Fisheries Society of the British Isles: its origin and early history. Journal of Fish Biology, 33(4):497–499, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Getachew:1988:DEN

[1437] T. Getachew. Digestive efficiency and nutrient composition gradient in the gut of Oreochromis niloticus L. in Lake Awasa, Ethiopia. Journal of Fish Biology, 33(4):501–509, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dosSantos:1988:GEC

[1438] J. dos Santos and M. Jobling. Gastric emptying in cod, Gadus morhua L.: effects of food particle size and dietary energy content. Journal of Fish Biology, 33(4):511–516, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bengtsson:1988:VDF

[1439] Å. Bengtsson, B. E. Bengtsson, and G. Lithner. Vertebral defects in fourhorn sculpin, Myoxocephalus quadricornis L., exposed to heavy metal pollution in the Gulf of Bothnia. Journal of Fish Biology, 33(4):517–529, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stergiou:1988:FHL

[1440] K. I. Stergiou. Feeding habits of the Lessepsian migrant Siganus luridus in the eastern Mediterranean, its new environment. Journal of Fish Biology, 33(4):531–543, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:1988:EEN

[1441] C. A. Mills. The effect of extreme northerly climatic conditions on the life history of the minnow, Phoxinus phoxinus (L.). Journal of Fish Biology, 33(4):545–561, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Overnell:1988:ALE

[1442] J. Overnell, T. C. Fletcher, and R. McIntosh. The apparent lack of effect of supplementary dietary zinc on zinc metabolism and metallothionein concentrations in the turbot, Scophthalmus maximus (Linnaeus). Journal of Fish Biology, 33(4):563–570, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:1988:MEM

[1443] J. J. Pereira. Morphological effects of mercury exposure on windowpane flounder gills as observed by scanning electron microscopy. Journal of Fish Biology, 33(4):571–580, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Vazquez:1988:CPF

[1444] E. García-Vazquez, A. R. Linde, G. Blanco, J. A. Sanchez, E. Vázquez, and J. Rubio. Chromosome polymorphism in farm fry stocks of Atlantic salmon from asturias. Journal of Fish Biology, 33(4):581–587, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brannas:1988:EBS

[1445] E. Brännäs. Emergence of Baltic salmon, Salmo salar L., in relation to temperature: a laboratory study. Journal of Fish Biology, 33(4):589– 600, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellgaard:1988:KAS

[1446] E. G. Ellgaard and J. L. Guillot. Kinetic analysis of the swimming behaviour of bluegill sunfish, Lepomis macrochirus Rafinesque, exposed to copper: hypoactivity induced by sublethal concentrations. Journal of Fish Biology, 33(4):601–608, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mayer:1988:ARB

[1447] I. Mayer, S. E. Shackley, and J. S. Ryland. Aspects of the reproductive biology of the bass, Dicentrarchus labrax L. I. An histological and histochemical study of oocyte development. Journal of Fish Biology, 33 (4):609–622, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1988:RSP

[1448] R. M. Wright and E. A. Shoesmith. The reproductive success of pike, Esox lucius: aspects of fecundity, egg density and survival. Journal of Fish Biology, 33(4):623–636, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quinn:1988:EAC

[1449] T. P. Quinn, A. F. Olson, and J. T. Konecki. Effects of anaesthesia on the chemosensory behaviour of Pacific salmon. Journal of Fish Biology, 33(4):637–641, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagar:1988:DFP

[1450] P. M. Sagar and G. J. Glova. Diel feeding periodicity, daily ration and prey selection of a riverine population of juvenile Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Biology, 33(4):643– 653, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coello:1988:TDS

[1451] S. Coello. Time of day of spawning in Sardinops sagax (Jenyns). Journal of Fish Biology, 33(4):655–656, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Donoghue:1988:GAN

[1452] S. Donoghue. Gasterosteus aculeatus, a new British host for the metacercariae of Cryptocotyle concavum (Digenea). Journal of Fish Biology, 33(4):657–658, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramaiyan:1988:ASR

[1453] V. Ramaiyan and R. Sivakumar. An abnormal sting-ray caught off Parangipettai on the south-eastern coast of India. Journal of Fish Biology, 33(4):659–660, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRg

[1454] Anonymous. Book reviews. Journal of Fish Biology, 33(4):661–663, October 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez-Palacios:1988:FEC

[1455] C. A. Martinez-Palacios and L. G. Ross. The feeding ecology of the Central American cichlid Cichlasoma urophthalmus (Günther). Journal of Fish Biology, 33(5):665–670, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:1988:RPT

[1456] J. M. Wright. Recruitment patterns and trophic relationships of fish in Sulaibikhat Bay, Kuwait. Journal of Fish Biology, 33(5):671–687, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deSilva:1988:ONH

[1457] S. S. de Silva and H. K. G. Sirisena. Observations on the nesting habits of Oreochromis mossambicus (Peters) (Pisces: Cichlidae) in Sri Lankan reservoirs. Journal of Fish Biology, 33(5):689–696, November 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amundsen:1988:DGE

[1458] P. A. Amundsen and A. Klemetsen. Diet, gastric evacuation rates and food consumption in a stunted population of Arctic charr, Salvelinus alpinus L., in Takvatn, northern Norway. Journal of Fish Biology, 33 (5):697–709, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Iger:1988:CRS

[1459] Y. Iger, M. Abraham, A. Dotan, B. Fattal, and E. Rahamim. Cellular responses in the skin of carp maintained in organically fertilized water. Journal of Fish Biology, 33(5):711–720, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tosi:1988:EAT

[1460] L. Tosi, L. Sala, C. Sola, A. Spampanato, and P. Tongiorgi. Experimental analysis of the thermal and salinity preferences of glass-eels, Anguilla anguilla (L.), before and during the upstream migration. Journal of Fish Biology, 33(5):721–733, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartley:1988:CSA

[1461] S. E. Hartley. Cytogenetic studies of Atlantic salmon, Salmo salar L., in Scotland. Journal of Fish Biology, 33(5):735–740, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dellefors:1988:ESM

[1462] C. Dellefors and U. Faremo. Early sexual maturation in males of wild sea trout, Salmo trutta L., inhibits smoltification. Journal of Fish Biology, 33 (5):741–749, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klimley:1988:DMS

[1463] A. P. Klimley, S. B. Butler, D. R. Nelson, and A. T. Stull. Diel movements of scalloped hammerhead sharks, Sphyrna lewini Griffith and Smith, to and from a seamount in the Gulf of California. Journal of Fish Biology, 33 (5):751–761, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:1988:BGCb

[1464] Y. Cui and R. J. Wootton. Bioenergetics of growth of a cyprinid, Phoxinus phoxinus (L.): the effect of ration and temperature on growth rate and efficiency. Journal of Fish Biology, 33(5):763–773, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fang:1988:SPH

[1465] L.-S. Fang and J. L. Bada. A special pattern of haem catabolism in a marine fish, Clinocottus analis, with green blood plasma. Journal of Fish Biology, 33(5):775–780, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banerjee:1988:CNG

[1466] S. K. Banerjee, K. K. Misra, S. Banerjee, and S. P. Ray-Chaudhuri. Chromosome numbers, genome sizes, cell volumes and evolution of snakehead fish (family Channidae). Journal of Fish Biology, 33(5):781–789, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Partington:1988:EBS

[1467] J. D. Partington and C. A. Mills. An electrophoretic and biometric study of Arctic charr, Salvelinm alpinus (L.), from ten British lakes. Journal of Fish Biology, 33(5):791–814, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendo:1988:IEO

[1468] J. Mendo and D. Pauly. Indirect estimation of oxygen and food consumption in bonito, Sarda chiliensis (Scombridae). Journal of Fish Biology, 33 (5):815–817, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbieri:1988:APT

[1469] G. Barbieri and M. C. Barbieri. Ageing of Parodon tortuosus Eigenmann and Norris, 1900 (Osteichthyes, Parodontidae) from the Passa Cinco River, Brazil. Journal of Fish Biology, 33(5):819, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Minchin:1988:CGG

[1470] D. Minchin. Couch’s goby, Gobius couchi (Teleostei: Gobiidae), from Irish waters. Journal of Fish Biology, 33(5):821–822, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibrahim:1988:FER

[1471] A. A. Ibrahim and F. A. Huntingford. Foraging efficiency in relation to within-species variation in morphology in three-spined sticklebacks, Gasterosteus aculeatus. Journal of Fish Biology, 33(5):823–824, November 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKaye:1988:SDH

[1472] K. R. McKaye and J. R. Stauffer, Jr. Seasonality, depth and habitat distribution of breeding males of Oreochromis spp., ‘chambo’, in Lake Malaŵi National Park. Journal of Fish Biology, 33(6):825–834, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersson:1988:BRI

[1473] E. Andersson, I. Mayer, and B. Borg. Book review: Inhibitory effect of 11-ketoandrostenedione and androstenedione on spermatogenesis in the three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 33(6):835–840, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loewe:1988:OAT

[1474] H. Loewe and R. Eckmann. The ontogeny of the alimentary tract of coregonid larvae: normal development. Journal of Fish Biology, 33(6): 841–850, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weatherley:1988:RMD

[1475] A. H. Weatherley, H. S. Gill, and A. F. Lobo. Recruitment and maximal diameter of axial muscle fibres in teleosts and their relationship to somatic growth and ultimate size. Journal of Fish Biology, 33(6):851–859, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gamperl:1988:EST

[1476] A. K. Gamperl, J. Bryant, and E. D. Stevens. Effect of a sprint training protocol on growth rate, conversion efficiency, food consumption and body composition of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 33(6):861–870, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quinitio:1988:ACT

[1477] G. F. Quinitio, H. Takahashi, and A. Goto. Annual changes in the testicular activity of the river sculpin, Cottus hangiongensis Mori, with emphasis on the occurrence of aberrant spermatids during spermatogenesis. Journal of Fish Biology, 33(6):871–878, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:1988:CRA

[1478] J. D. Armstrong, M. Lucas, J. French, L. Vera, and I. G. Priede. A combined radio and acoustic transmitter for fixing direction and range of freshwater fish (RAFIX). Journal of Fish Biology, 33(6):879–884, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggenes:1988:EMM

[1479] J. Heggenes and R. Borgstrøm. Effect of mink, Mustela vison Schreber, predation on cohorts of juvenile Atlantic salmon, Salmo salar L., and brown trout, S. trutta L., in three small streams. Journal of Fish Biology, 33(6):885–894, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bonar:1988:EST

[1480] S. A. Bonar, G. L. Thomas, and G. B. Pauley. Evaluation of the separation of triploid and diploid grass carp, Ctenopharyngodon idella (Valenciennes), by external morphology. Journal of Fish Biology, 33(6):895–898, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

LAbee-Lund:1988:OSD

[1481] J. H. L’Abée-Lund. Otolith shape discriminates between juvenile Atlantic salmon, Salmo salar L., and brown trout, Salmo trutta L. Journal of Fish Biology, 33(6):899–903, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marconato:1988:MCE

[1482] A. Marconato and A. Bisazza. Mate choice, egg cannibalism and reproductive success in the river bullhead, Cottus gobio L. Journal of Fish Biology, 33(6):905–916, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:1988:CFS

[1483] F. A. Huntingford, N. B. Metcalfe, and J. E. Thorpe. Choice of feeding station in Atlantic salmon, Salmo salar, parr: effects of predation risk, season and life history strategy. Journal of Fish Biology, 33(6):917–924, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tatner:1988:CFT

[1484] M. F. Tatner. The cryopreservation of fish thymocytes and its potential application for immunological reconstitution experiments. Journal of Fish Biology, 33(6):925–929, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1988:MRJ

[1485] J. E. Thorpe, R. I. G. Morgan, D. Pretswell, and P. J. Higgins. Movement rhythms in juvenile Atlantic salmon, Salmo salar L. Journal of Fish Biology, 33(6):931–940, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rydevik:1988:ETS

[1486] M. Rydevik. Epidermis thickness and secondary sexual characters in mature male and immature Baltic salmon, Salmo salar L., parr: seasonal variations and effects of castration and androgen treatment. Journal of Fish Biology, 33(6):941–944, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:1988:ETD

[1487] G. M. Hughes and T. Koyama. Effect of temperature on the deformability of a lamprey, Entosphenus japonicus, and Pacific salmon, Oncorhynchus keta, red blood cells, studied using a modified filtration method. Journal of Fish Biology, 33(6):945–950, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:1988:TOG

[1488] P. S. Davis. Two occurrences of the gilthead, Sparus aurata Linnaeus 1758, on the coast of Northumberland, England. Journal of Fish Biology, 33(6):951, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrettm:1988:OTR

[1489] N. R. Merrettm and C. Karrer. Occurrence of toothless rattails of the genus Coryphaenoides and its possible significance. Journal of Fish Biology, 33(6):953–954, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naesje:1988:SDA

[1490] T. F. Næsje, L. P. Hansen, and T. Järvi. Sexual dimorphism in the adipose fin of Atlantic salmon, Salmo salar L. Journal of Fish Biology, 33 (6):955–956, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1988:SHS

[1491] E. A. R. Brown and D. B. C. Scott. A second hermaphrodite specimen of Coregonus lavaretus (L.) (Salmonidae, Coregoninae) from Loch Lomond, Scotland. Journal of Fish Biology, 33(6):957–958, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:1988:ORE

[1492] A. Moore and A. Scott. Observations of recently emerged sea trout, Salmo trutta L., fry in a chalk stream, using a low-light underwater camera. Journal of Fish Biology, 33(6):959–960, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1988:BRh

[1493] Anonymous. Book reviews. Journal of Fish Biology, 33(6):961–967, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Margetts:1988:P

[1494] A. R. Margetts. Preface. Journal of Fish Biology, 33(sA):iii, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cyrus:1988:EEE

[1495] D. P. Cyrus. Episodic events and estuaries: effects of cyclonic flushing on the benthic fauna and diet of Solea bleekeri (Teleostei) in Lake St Lucia on the south-eastern coast of Africa. Journal of Fish Biology, 33 (sA):1–7, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:1988:IST

[1496] T. J. Martin. Interaction of salinity and temperature as a mechanism for spatial separation of three co-existing species of ambassidae (Cuvier) (Teleostei) in estuaries on the south-east coast of Africa. Journal of Fish Biology, 33(sA):9–15, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bamber:1988:PAP

[1497] R. N. Bamber and P. A. Henderson. Pre-adaptive plasticity in atherinids and the estuarine seat of teleost evolution. Journal of Fish Biology, 33 (sA):17–23, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelley:1988:IES

[1498] D. F. Kelley. The importance of estuaries for sea-bass, Dicentrarchus labrax (L.). Journal of Fish Biology, 33(sA):25–33, December 1988. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinto:1988:PDC

[1499] L. Pinto. Population dynamics and community structure of fish in the mangroves of pagbilao, Philippines. Journal of Fish Biology, 33(sA):35– 43, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogers:1988:SPE

[1500] S. I. Rogers. The seasonal partitioning of energy in an estuarine fish, the common goby, Pomatoschistus microps Krøyer. Journal of Fish Biology, 33(sA):45–50, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1988:RFS

[1501] M. Elliott, A. H. Griffiths, and C. J. L. Taylor. The role of fish studies in estuarine pollution assessment. Journal of Fish Biology, 33(sA):51–61, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curran:1988:ORP

[1502] J. C. Curran and A. R. Henderson. The oxygen requirements of a polluted estuary for the establishment of a migratory salmon, Salmo salar L., population. Journal of Fish Biology, 33(sA):63–69, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pomfret:1988:BTS

[1503] J. R. Pomfret, G. S. Turner, and S. Phillips. Beam trawl surveys as a monitoring tool in polluted estuaries in north-east England. Journal of Fish Biology, 33(sA):71–77, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McColl:1988:MPL

[1504] N. P. McColl. A model to predict the level of artificial radionuclides in environmental materials in the Severn Estuary and the Bristol Channel. Journal of Fish Biology, 33(sA):79–84, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conlan:1988:DDH

[1505] K. Conlan, K. Hendry, K. N. White, and S. J. Hawkins. Disused docks as habitats for estuarine fish: a case study of Preston dock. Journal of Fish Biology, 33(sA):85–91, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dempsey:1988:IE

[1506] C. H. Dempsey. Ichthyoplankton entrainment. Journal of Fish Biology, 33(sA):93–102, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turnpenny:1988:FIE

[1507] A. W. H. Turnpenny. Fish impingement at estuarine power stations and its significance to commercial fishing. Journal of Fish Biology, 33(sA): 103–110, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davies:1988:RIR

[1508] J. K. Davies. A review of information relating to fish passage through turbines: implications to tidal power schemes. Journal of Fish Biology, 33(sA):111–126, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solomon:1988:FRN

[1509] D. J. Solomon and E. C. E. Potter. First results with a new estuarine fish tracking system. Journal of Fish Biology, 33(sA):127–132, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:1988:BAA

[1510] I. G. Priede, J. F. de L. G. Solbé, J. E. Nott, K. T. O’Grady, and D. Cragg-Hine. Behaviour of adult Atlantic salmon, Salmo salar L., in the estuary of the River Ribble in relation to variations in dissolved oxygen and tidal flow. Journal of Fish Biology, 33(sA):133–139, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:1988:BTS

[1511] M. W. Aprahamian. The biology of the twaite shad, Alosa fallax fallax (Lacépède), in the Severn Estuary. Journal of Fish Biology, 33(sA): 141–152, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Potter:1988:MAS

[1512] E. C. E. Potter. Movements of Atlantic salmon, Salmo salar L., in an estuary in south-west England. Journal of Fish Biology, 33(sA):153–159, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naismith:1988:MEJ

[1513] I. A. Naismith and B. Knights. Migrations of elvers and juvenile European eels, Anguilla anguilla L., in the River Thames. Journal of Fish Biology, 33(sA):161–175, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coggan:1988:MJD

[1514] R. A. Coggan and P. R. Dando. Movements of juvenile Dover sole, Solea solea (L.), in the Tamar Estuary, South-western England. Journal of Fish Biology, 33(sA):177–184, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Antunes:1988:RAS

[1515] M. M. Antunes, P. L. da Cunha, A. P. Duarte, and E. P. Mendonça. Ria de Alvor as a spawning place and a nursery ground. Journal of Fish Biology, 33(sA):185–190, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yaanez-Arancibia:1988:SBD

[1516] A. Yåñez-Arancibia, A. L. Lara-Dominguez, J. L. Rojas-Galaviz, P. Sånchez-Gil, J. W. Day, Jr., and C. J. Madden. Seasonal biomass and diversity of estuarine fishes coupled with tropical habitat heterogeneity (southern Gulf of Mexico). Journal of Fish Biology, 33(sA):191–200, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergman:1988:PNE

[1517] M. J. N. Bergman, H. W. van der Veer, and J. J. Zulstra. Plaice nurseries: effects on recruitment. Journal of Fish Biology, 33(sA):201–218, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Hassan:1988:GVP

[1518] L. A. J. Al-Hassan. Genetic variation and populations structure in Thryssa hamiltonii and Nematalosa nasus in Iraqi waters. Journal of Fish Biology, 33(sA):219–220, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelley:1988:SCF

[1519] D. F. Kelley and P. J. Reay. The shallow creek fish communities of southwest England and west Wales estuaries. Journal of Fish Biology, 33(sA): 221–222, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1988:SEF

[1520] P. A. Henderson. The structure of estuarine fish communities. Journal of Fish Biology, 33(sA):223–225, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchand:1988:IMD

[1521] J. Marchand and G. Masson. Inshore migration, distribution, behaviour and feeding activity of sole, Solea solea (L.), postlarvae in the Vilaine estuary, France. Journal of Fish Biology, 33(sA):227–228, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchand:1988:SDG

[1522] J. Marchand. Seasonal distribution, growth and ecological role of the juvenile sole, Solea solea L., population in the Loire estuary, France. Journal of Fish Biology, 33(sA):229, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sabriye:1988:SBL

[1523] A. S. Sabriye, P. J. Reay, and S. H. Coombs. Sea-bass larvae in coastal and estuarine plankton. Journal of Fish Biology, 33(sA):231–233, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:1988:RFM

[1524] K. W. Wilson, B. J. D’Arcy, and S. Taylor. The return of fish to the Mersey estuary. Journal of Fish Biology, 33(sA):235–238, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendry:1988:DDH

[1525] K. Hendry, K. Conlan, K. N. White, A. Bewsher, and S. J. Hawkins. Disused docks as a habitat for estuarine fish: a nationwide appraisal. Journal of Fish Biology, 33(sA):239–241, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robin:1988:CAE

[1526] J.-P. Robin. Comparative analysis of the effect of a power station and of small-mesh fisheries on the juvenile populations of smelt and flounder in the Loire estuary: presentation of methods. Journal of Fish Biology, 33(sA):243–244, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:1988:MAS

[1527] M. W. Aprahamian, C. D. Strange, and C. Dimond. Movement of Atlantic salmon, Salmo salar L., into the River Usk, South Wales, in relation to water quality. Journal of Fish Biology, 33(sA):245–247, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:1988:TME

[1528] M. J. Costa. The Tagus and Mira estuaries (Portugal) and their role as spawning and nursery areas. Journal of Fish Biology, 33(sA):249–250, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1988:FES

[1529] Alwyne Wheeler. Fish in estuaries: a summing up. Journal of Fish Biology, 33(sA):251–254, December 1988. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roberts:1989:RMP

[1530] C. D. Roberts. Reproductive mode in the percomorph fish genus Polyprion Oken. Journal of Fish Biology, 34(1):1–9, January 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Enderlein:1989:MBA

[1531] O. Enderlein. Migratory behaviour of adult cisco, Coregonus albula L., in the Bothnian Bay. Journal of Fish Biology, 34(1):11–18, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Degani:1989:IBS

[1532] G. Degani, M. L. Gallagher, and A. Meltzer. The influence of body size and temperature on oxygen consumption of the European eel, Anguilla anguilla. Journal of Fish Biology, 34(1):19–24, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bahgat:1989:UCM

[1533] F. J. Bahgat, P. E. King, and S. E. Shackley. Ultrastructural changes in the muscle tissue of Clupea harengus L. larvae induced by acid pH. Journal of Fish Biology, 34(1):25–30, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foscarini:1989:CSS

[1534] R. Foscarini. A comparative study of the skin and gill structure in oviparous and viviparous freshwater fish larvae. Journal of Fish Biology, 34(1):31–40, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deLeaniz:1989:NHB

[1535] C. Garcia de Leaniz and E. Verspoor. Natural hybridization between Atlantic salmon, Salmo salar, and brown trout, Salmo trutta, in northern Spain. Journal of Fish Biology, 34(1):41–46, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:1989:BGC

[1536] Y. Cui and R. J. Wootton. Bioenergetics of growth of a cyprinid, Phoxinus phoxinm (L.): development and testing of a growth model. Journal of Fish Biology, 34(1):47–64, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alfei:1989:MAS

[1537] L. Alfei, G. Bertoncello, E. Gelosi, and F. Parvopassu. A morphometric analysis of the spinal motor pool in relation to its target muscle during growth in the trout, Salmo gairdneri Richardson. Journal of Fish Biology, 34(1):65–78, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gartner:1989:PFM

[1538] J. V. Gartner, Jr. and D. E. Zwerner. The parasite faunas of mesoand bathypelagic fishes of Norfolk Submarine Canyon, western North Atlantic. Journal of Fish Biology, 34(1):79–95, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagliniere:1989:PDB

[1539] J. L. Bagliniére, G. Maisse, P. Y. Lebail, and A. Nihouarn. Population dynamics of brown trout, Salmo trutta L., in a tributary in Brittany (France): spawning and juveniles. Journal of Fish Biology, 34(1):97– 110, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mooij:1989:KIL

[1540] W. M. Mooij. A key to the identification of larval bream, Abramis brama, white bream, Blicca bjoerkna, and roach, Rutilus rutilus. Journal of Fish Biology, 34(1):111–118, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crisp:1989:OSD

[1541] D. T. Crisp and P. A. Carling. Observations on siting, dimensions and structure of salmonid redds. Journal of Fish Biology, 34(1):119–134, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bean:1989:BEE

[1542] C. W. Bean and I. J. Winfield. Biological and ecological effects of a Ligula intestinalis (L.) infestation of the gudgeon, Gobiogobio (L.), in Lough Neagh, Northern Ireland. Journal of Fish Biology, 34(1):135–147, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:1989:PAA

[1543] K. A. Jones. The palatability of amino acids and related compounds to rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 34 (1):149–160, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townsend:1989:EMP

[1544] C. R. Townsend and M. R. Perrow. Eutrophication may produce population cycles in roach, Rutilus rutilus (L.), by two contrasting mechanisms. Journal of Fish Biology, 34(1):161–164, January 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojha:1989:LSE

[1545] J. Ojha, N. C. Rooj, A. K. Mittal, and J. S. Datta Munshi. Light and scanning electron microscopic studies on the effect of biocidal plant sap on the gills of a hill stream fish, Garra lamta (Ham.). Journal of Fish Biology, 34(2):165–170, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolger:1989:SSI

[1546] T. Bolger and P. L. Connolly. The selection of suitable indices for the measurement and analysis of fish condition. Journal of Fish Biology, 34 (2):171–182, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1989:SDA

[1547] A. F. Brown. Seasonal dynamics of the acanthocephalan Pomphorhynchm laevis (Müller, 1776) in its intermediate and preferred definitive hosts. Journal of Fish Biology, 34(2):183–194, February 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjesbu:1989:SAC

[1548] O. S. Kjesbu. The spawning activity of cod, Gadus morhua L. Journal of Fish Biology, 34(2):195–206, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulin:1989:ELH

[1549] R. Poulin and G. J. Fitzgerald. Early life histories of three sympatric sticklebacks in a salt-marsh. Journal of Fish Biology, 34(2):207–221, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mearns:1989:BRS

[1550] K. J. Mearns. Behavioural responses of salmonid fry to low amino acid concentrations. Journal of Fish Biology, 34(2):223–232, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bakker:1989:CST

[1551] Th. C. M. Bakker and P. Sevenster. Changes in the sexual tendency accompanying selection for aggressiveness in the three-spined stickleback, Gasterosteus aculeatus L. Journal of Fish Biology, 34(2):233–243, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibrahim:1989:LFS

[1552] A. A. Ibrahim and F. A. Huntingford. Laboratory and field studies on diet choice in three-spined sticklebacks, Gasterosteus aculeatus L., in relation to profitability and visual features of prey. Journal of Fish Biology, 34(2):245–257, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lubin:1989:UAB

[1553] R. T. Lubin, A. W. Rourke, and T. M. Bradley. Ultrastructural alterations in branchial chloride cells of Atlantic salmon, Salmo salar, during parr-smolt transformation and early development in sea water. Journal of Fish Biology, 34(2):259–272, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benedetti:1989:HCS

[1554] I. Benedetti, A. G. Albano, and L. Mola. Histomorphological changes in some organs of the brown bullhead, Ictalurus nebulosus LeSueur, following short- and long-term exposure to copper. Journal of Fish Biology, 34 (2):273–280, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frandsen:1989:EPP

[1555] F. Frandsen, H. J. Malmquist, and S. S. Snorrason. Ecological parasitology of polymorphic Arctic charr, Salvelinus alpinus (L.), in Thingvallavatn, Iceland. Journal of Fish Biology, 34(2):281–297, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

North:1989:AAR

[1556] E. North and P. Hickley. An appraisal of anglers’catches in the River Severn, England. Journal of Fish Biology, 34(2):299–306, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berrebi:1989:ISN

[1557] P. Berrebi, F. Dupont, G. Cattaneo-Berrebi, and A. J. Crivelli. An isozyme study of the natural cyprinid hybrid Alburnus alburnus × Rutilus rubilio in Greece. Journal of Fish Biology, 34(2):307–313, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Junger:1989:CME

[1558] H. Junger, K. Kotrschal, and A. Goldschmid. Comparative morphology and ecomorphology of the gut in European cyprinids (Telostei). Journal of Fish Biology, 34(2):315–326, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OFarrell:1989:OGM

[1559] M. M. O’Farrell and R. E. Peirce. The occurrence of a gynandromorphic migratory trout, Salmo trutta L. Journal of Fish Biology, 34(2):327, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coello:1989:NSS

[1560] S. Coello. A new staining schedule for formalin-fixed, glycol methacrylate-embedded fish ovaries. Journal of Fish Biology, 34(2): 329–330, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alabaster:1989:DOT

[1561] J. S. Alabaster. The dissolved oxygen and temperature requirements of king salmon, Oncorhynchus tshawytscha, in the San Joaquin Delta, California. Journal of Fish Biology, 34(2):331–332, February 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roubal:1989:PCG

[1562] F. R. Roubal. Pathological changes in the gill filaments of Acanthopagrus australis (family Sparidae) associated with the post-settlement growth of a lernaeopodid copepod, Alella macrotrachelus. Journal of Fish Biology, 34(3):333–342, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bentivegna:1989:GSV

[1563] F. Bentivegna and F. Benedetto. Gonochorism and seasonal variations in the gonads of the labrid Symphodus (Crenilabrus) ocellatus (Forsskål). Journal of Fish Biology, 34(3):343–348, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lim:1989:EFG

[1564] A. L. L. Lim and Y. K. Ip. Effect of fasting on glycogen metabolism and activities of glycolytic and gluconeogenic enzymes in the mudskipper Boleophthalmus boddaerti. Journal of Fish Biology, 34(3):349–367, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1989:MDA

[1565] P. J. Smith, A. J. Birley, A. Jamieson, and C. A. Bishop. Mitochondrial DNA in the Atlantic cod, Gadus morhua: lack of genetic divergence between eastern and western populations. Journal of Fish Biology, 34 (3):369–373, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:1989:HCM

[1566] H. S. Gill, A. H. Weatherley, R. Lee, and D. Legere. Histochemical characterization of myotomal muscle of five teleost species. Journal of Fish Biology, 34(3):375–386, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diamant:1989:EAS

[1567] A. Diamant. Ecology of the acanthocephalan Sclerocollum rubrimaris Schmidt and Paperna, 1978 (Rhadinorhynchidae: Gorgorhynchinae) from wild populations of rabbitfish (genus Siganus) in the northern Red Sea. Journal of Fish Biology, 34(3):387–397, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uchida:1989:LCF

[1568] K. Uchida, K. Tsukamoto, S. Ishii, R. Ishida, and T. Kajihara. Larval competition for food between wild and hatchery-reared ayu, Plecoglossm altivelis Temminck et Schlegel, in culture ponds. Journal of Fish Biology, 34(3):399–407, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:1989:WMB

[1569] P. A. Henderson and R. H. A. Holmes. Whiting migration in the Bristol Channel: a predator–prey relationship. Journal of Fish Biology, 34(3): 409–416, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rajasilta:1989:DME

[1570] M. Rajasilta, J. Eklund, J. Kääriä, and K. Ranta-Aho. The deposition and mortality of the eggs of the Baltic herring Clupea harengus membras L., on different substrates in the south-west archipelago of Finland. Journal of Fish Biology, 34(3):417–427, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Creutzberg:1989:AEP

[1571] F. Creutzberg and J. Ij. Witte. An attempt to estimate the predatory pressure exerted by the lesser weever, Trachinus vipera Cuvier, in the southern North Sea. Journal of Fish Biology, 34(3):429–449, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruggerone:1989:GER

[1572] G. T. Ruggerone. Gastric evacuation rates and daily ration of piscivorous Coho salmon, Oncorhynchus kisutch Walbaum. Journal of Fish Biology, 34(3):451–463, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanderBank:1989:EDG

[1573] F. H. Van der Bank, W. S. Grant, and J. T. Ferreira. Electrophoretically detectable genetic data for fifteen southern African cichlids. Journal of Fish Biology, 34(3):465–483, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacCrimmon:1989:NAY

[1574] H. R. MacCrimmon, B. L. Gots, and R. B. MacGregor. A note on an atypical yellow perch, Perca flavescens Mitchill. Journal of Fish Biology, 34(3):485–486, March 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:Aa

[1575] Anonymous. Announcement. Journal of Fish Biology, 34(4):487, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Last:1989:FHC

[1576] J. M. Last. The food of herring, Clupea harengus, in the North Sea, 1983–1986. Journal of Fish Biology, 34(4):489–501, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roubal:1989:CPS

[1577] F. R. Roubal. Comparative pathology of some monogenean and copepod ectoparasites on the gills of Acanthopagrus australis (family Sparidae). Journal of Fish Biology, 34(4):503–514, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flett:1989:DRE

[1578] P. A. Flett and J. F. Leatherland. Dose-related effects of 17β oestradiol (E_{2}) on liver weight, plasma E_{2}, protein, calcium and thyroid hormone levels, and measurement of the binding of thyroid hormones to vitellogenin in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 34(4):515–527, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffman:1989:DCO

[1579] K. S. Hoffman and E. G. Grau. Daytime changes in oocyte development with relation to the tide for the Hawaiian saddleback wrasse, Thalassoma duperrey. Journal of Fish Biology, 34(4):529–546, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidson:1989:RGV

[1580] W. S. Davidson, T. P. Birt, and J. M. Green. A review of genetic variation in Atlantic salmon, Salmo salar L., and its importance for stock identification, enhancement programmes and aquaculture. Journal of Fish Biology, 34(4):547–560, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernendez-Delgado:1989:AGY

[1581] C. Fernéndez-Delgado, J. A. Hernando, M. Herrera, and M. Bellido. Age and growth of yellow eels, Anguilla anguilla, in the estuary of the Guadalquivir river (south-west Spain). Journal of Fish Biology, 34(4): 561–570, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:1989:SRJ

[1582] K. H. Olsén. Sibling recognition in juvenile Arctic charr, Salvelinus alpinus (L.). Journal of Fish Biology, 34(4):571–581, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:1989:LIC

[1583] M. Taylor and D. Hoole. Ligula intestinalis (L.) (Cestoda: Pseudophyllidea): plerocercoid-induced changes in the spleen and pronephros of roach, Rutilus rutilus (L.), and gudgeon, Gobio gobio (L.). Journal of Fish Biology, 34(4):583–596, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skaala:1989:GDB

[1584] Ø. Skaala and G. Nævdal. Genetic differentiation between freshwater resident and anadromous brown trout, Salmo trutta, within watercourses. Journal of Fish Biology, 34(4):597–605, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1989:DDF

[1585] C. R. Kennedy, R. M. Bates, and A. F. Brown. Discontinuous distributions of the fish acanthocephalans Pomphorhynchm laevis and Acanthocephalus anguillae in Britain and Ireland: an hypothesis. Journal of Fish Biology, 34(4):607–619, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purdy:1989:EBE

[1586] J. E. Purdy. The effects of brief exposure to aromatic hydrocarbons on feeding and avoidance behaviour in Coho salmon, Oncorhynchus kisutch.

Journal of Fish Biology, 34(4):621–629, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:1989:IHR

[1587] C. J. Robinson and T. J. Pitcher. The influence of hunger and ration level on shoal density, polarization and swimming speed of herring, Clupea harengus L. Journal of Fish Biology, 34(4):631–633, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spanier:1989:TSS

[1588] E. Spanier, Y. Finkelstein, and B. Raikhlin-Eisenkraft. Toxicity of the saupe, Sarpa salpa (Linnaeus, 1758), on the Mediterranean coast of Israel. Journal of Fish Biology, 34(4):635–636, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singh:1989:RAL

[1589] B. N. Singh, M. A. Towheed, and J. S. D. Munshi. Respiratory adaptations in the larvae of Monoptems cuchia (Ham.). Journal of Fish Biology, 34(4):637–638, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urfi:1989:HLR

[1590] A. J. Urfi and C. L. Talesara. Histomorphology of length-related changes in myotomal muscles of catfish, Heteropneustes fossilis (Bloch). Journal of Fish Biology, 34(4):639–641, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Partington:1989:MAP

[1591] J. D. Partington. A method to allow the production of phylogenetic trees from morphometric data. Journal of Fish Biology, 34(4):643–644, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRa

[1592] Anonymous. Book reviews. Journal of Fish Biology, 34(4):645–652, April 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galman:1989:FSE

[1593] O. R. Galman and R. R. Avtalion. Further study of the embryonic development of Oreochromis niloticus (Cichlidae, Teleostei) using scanning electron microscopy. Journal of Fish Biology, 34(5):653–664, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinkney:1989:MSE

[1594] A. E. Pinkney, D. A. Wright, and G. M. Hughes. A morphometric study of the effects of tributyltin compounds on the gills of the mummichog, Fundulus heteroclitus. Journal of Fish Biology, 34(5):665–677, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radabaugh:1989:SCC

[1595] D. C. Radabaugh. Seasonal colour changes and shifting antipredator tactics in darters. Journal of Fish Biology, 34(5):679–685, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rincon:1989:RGS

[1596] P. A. Rincon and J. Lobon-Cervia. Reproductive and growth strategies of the red roach, Rutilus arcasii (Steindachner, 1866), in two contrasting tributaries of the River Duero, Spain. Journal of Fish Biology, 34(5):687– 705, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burley:1989:MDM

[1597] C. C. Burley and S. Vigg. A method for direct measurement of the maximum volume of fish stomachs or digestive tracts. Journal of Fish Biology, 34(5):707–714, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harmelin-Vivien:1989:FPA

[1598] M. L. Harmelin-Vivien, R. A. Kaim-Malka, M. Ledoyer, and S. S. Jacob-Abraham. Food partitioning among scorpaenid fishes in Mediterranean seagrass beds. Journal of Fish Biology, 34(5):715–734, May 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjesbu:1989:OCG

[1599] O. S. Kjesbu and H. Kryvi. Oogenesis in cod, Gadus morhua L., studied by light and electron microscopy. Journal of Fish Biology, 34(5):735– 746, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doggett:1989:UPB

[1600] T. A. Doggett and J. E. Harris. Ultrastructure of the peripheral blood leucocytes of Oreochromis mossambicus. Journal of Fish Biology, 34(5): 747–756, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pickering:1989:DSB

[1601] A. D. Pickering, T. G. Pottinger, and J. F. Carragher. Differences in the sensitivity of brown trout, Salmo trutta L., and rainbow trout, Salmo gairdneri Richardson, to physiological doses of cortisol. Journal of Fish Biology, 34(5):757–768, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dekker:1989:DRR

[1602] W. Dekker. Death rate, recapture frequency and changes in size of tagged eels. Journal of Fish Biology, 34(5):769–777, May 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hlavova:1989:EAB

[1603] V. Hlavová. Enzyme activities in the blood plasma of grayling, Thymallus thymallus (Linn.), in the breeding season. Journal of Fish Biology, 34 (5):779–789, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Youngson:1989:EET

[1604] A. F. Youngson, L. P. Hansen, B. Jonsson, and T. F. Naesje. Effects of exogenous thyroxine or prior exposure to raised water-flow on the downstream movement of hatchery-reared Atlantic salmon smolts. Journal of Fish Biology, 34(5):791–797, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haug:1989:OWA

[1605] T. Haug, E. Kjøsrsvik, and K. Pittman. Observations on a wild Atlantic halibut larva, Hippoglossus hippoglossus (L.). Journal of Fish Biology, 34(5):799–801, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franklin:1989:MOM

[1606] C. E. Franklin and W. Davison. S.E.M. observations of morphologically different chloride cells in freshwater-adapted sockeye salmon, Oncorhynchm nerka. Journal of Fish Biology, 34(5):803–804, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:1989:AMT

[1607] G. F. Turner and U. Falter. On the adoption of ‘male-typical’ breeding colouration by female mouthbrooding cichlids. Journal of Fish Biology, 34(5):805–806, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolok:1989:RBM

[1608] A. S. Kolok. The relationship between maintenance ration and growth rate in two strains of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 34(5):807–809, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayes:1989:DIS

[1609] G. C. Hayes, D. A. Purdie, and J. A. Williams. The distribution of ichthyoplankton in Southampton Water in response to low oxygen levels produced by a Mesodinium rubrum bloom. Journal of Fish Biology, 34 (5):811–813, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRb

[1610] Anonymous. Book reviews. Journal of Fish Biology, 34(5):815–818, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRc

[1611] Anonymous. Books received. Journal of Fish Biology, 34(5):818, May 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1989:FSS

[1612] D. W. Smith. The feeding selectivity of silver carp, Hypophthalmichthys molitrix Val. Journal of Fish Biology, 34(6):819–828, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madsen:1989:PIR

[1613] S. S. Madsen and E. T. Naamansen. Plasma ionic regulation and gill Na^{+} /K^{+}-ATPase changes during rapid transfer to sea water of yearling rainbow trout, Salmo gairdneri: time course and seasonal variation. Journal of Fish Biology, 34(6):829–840, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hine:1989:OEG

[1614] P. M. Hine and J. M. Wain. Observations on eosinophilic granule cells in peritoneal exudates of eels, Anguilla amtmlis. Journal of Fish Biology, 34(6):841–853, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanDamme:1989:SCK

[1615] P. van Damme, S. Appelbaum, and T. Hecht. Sibling cannibalism in Koi carp, Cypvinus carpio L., larvae and juveniles reared under controlled conditions. Journal of Fish Biology, 34(6):855–863, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Handy:1989:SAA

[1616] R. D. Handy and F. B. Eddy. Surface absorption of aluminium by gill tissue and body mucus of rainbow trout, Salmo gairdneri, at the onset of episodic exposure. Journal of Fish Biology, 34(6):865–874, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:1989:MCC

[1617] M. Morgan and D. E. Wright. Morphology of central compartment and vasculature of the gills of Lepidosiren paradoxa (Fitzinger). Journal of Fish Biology, 34(6):875–888, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marconato:1989:CMR

[1618] A. Marconato, A. Bisazza, and G. Marin. Correlates of male reproductive success in Padogobius martensi (Gobiidae). Journal of Fish Biology, 34 (6):889–899, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunn:1989:LLC

[1619] S. E. Dunn, A. Murad, and A. H. Houston. Leucocytes and leucopoietic capacity in thermally acclimated goldfish, Carassim auratus L. Journal of Fish Biology, 34(6):901–911, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasurer:1989:MPP

[1620] J. W. Treasurer. Mortality and production of 0+ perch, Perca fluviatilis L., in two Scottish lakes. Journal of Fish Biology, 34(6):913–928, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shiau:1989:EDF

[1621] S. Y. Shiau, C. C. Kwok, C. J. Chen, H. T. Hong, and H. B. Hsieh. Effects of dietary fibre on the intestinal absorption of dextrin, blood sugar level and growth of tilapia, Oreochromis niloticm × O. aureus. Journal of Fish Biology, 34(6):929–935, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnell:1989:MAF

[1622] M. F. O’Connell and R. J. Gibson. The maturation of anadromous female Atlantic salmon, Salmo salar L., stocked in a small pond in urban St John’s, Newfoundland, Canada. Journal of Fish Biology, 34(6):937– 946, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mlglavs:1989:EFR

[1623] I. Mlglavs and M. Jobling. Effects of feeding regime on food consumption, growth rates and tissue nucleic acids in juvenile Arctic charr, Salvelinm alpinus, with particular respect to compensatory growth. Journal of Fish Biology, 34(6):947–957, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shustov:1989:ESE

[1624] Yu. A. Shustov and I. L. Shchurov. Experimental study of the effect of young salmon, Salmo salar L., stamina on their feeding rates in a river. Journal of Fish Biology, 34(6):959–961, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Singh:1989:ADF

[1625] H. S. Singh and T. V. Reddy. An additional dorsal fin in Heteropneustes fossilis (Bloch). Journal of Fish Biology, 34(6):963–964, June 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bodaly:1989:GSS

[1626] R. A. Bodaly, R. D. Ward, and C. A. Mills. A genetic stock study of perch, Perca fluviatilis L., in Windermere. Journal of Fish Biology, 34(6):965–967, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scalia:1989:CAC

[1627] S. Scalia, M. Isaksen, and G. W. Francis. Carotenoids of the Arctic charr, Salvelinm alpinus (L.). Journal of Fish Biology, 34(6):969–970, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1989:HPC

[1628] C. R. Kennedy and H. H. Williams. Helminth parasite community diversity in a marine fish, Raja batis L. Journal of Fish Biology, 34(6): 971–972, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dominguez:1989:FOH

[1629] J. Dominguez, J. C. Pena, and G. González. First occurrence of hermaphrodite pike, Esox lucius L., in the Esla basin (Spain). Journal of Fish Biology, 34(6):973–975, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duran:1989:SSC

[1630] M. L. Sanmartín Durán, P. Quinteiro Alonso, A. Rodriguez, and J. A. Fernándeez. Some Spanish cestode fish parasites. Journal of Fish Biology, 34(6):977–978, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRd

[1631] Anonymous. Book reviews. Journal of Fish Biology, 34(6):979–983, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRe

[1632] Anonymous. Books received. Journal of Fish Biology, 34(6):983–984, June 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miglavs:1989:EFR

[1633] I. Miglavs and M. Jobling. The effects of feeding regime on proximate body composition and patterns of energy deposition in juvenile Arctic charr, Salvelinus alpinus. Journal of Fish Biology, 35(1):1–11, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rombout:1989:IISa

[1634] J. H. W. M. Rombout and A. A. van den Berg. Immunological importance of the second gut segment of carp. I. Uptake and processing of antigens by epithelial cells and macrophages. Journal of Fish Biology, 35(1):13–22, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vig:1989:EHP

[1635] É. Víg and J. Nemcsók. The effects of hypoxia and paraquat on the superoxide dismutase activity in different organs of carp, Cyprinus carpio L. Journal of Fish Biology, 35(1):23–25, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everall:1989:IWH

[1636] N. C. Everall, N. A. A. Macfarlane, and R. W. Sedgwick. The interactions of water hardness and pH with the acute toxicity of zinc to the brown trout, Salmo trutta L. Journal of Fish Biology, 35(1):27–36, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macpherson:1989:GES

[1637] E. Macpherson, J. Lleonart, and P. Sanchez. Gastric emptying in Scyliorhinus canicula (L.): a comparison of surface-dependent and nonsurface dependent models. Journal of Fish Biology, 35(1):37–48, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deniel:1989:CSS

[1638] C. Deniel, C. Le Blanc, and A. Rodriguez. Comparative study of sexual cycles, oogenesis and spawning of two Soleidae, Solea lascaris (Risso, 1810) and Solea impar (Bennet, 1831), on the western coast of Brittany. Journal of Fish Biology, 35(1):49–58, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:1989:ERB

[1639] C. Whitehead and J. A. Brown. Endocrine responses of brown trout, Salmo trutta L., to acid, aluminium and lime dosing in a Welsh hill stream. Journal of Fish Biology, 35(1):59–71, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1989:CTH

[1640] J. A. Brown, D. Edwards, and C. Whitehead. Cortisol and thyroid hormone responses to acid stress in the brown trout, Salmo trutta L. Journal of Fish Biology, 35(1):73–84, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiegand:1989:IDA

[1641] M. D. Wiegand, J. M. Hataley, C. L. Kitchen, and L. G. Buchanan. Induction of developmental abnormalities in larval goldfish, Carassius auratus L., under cool incubation conditions. Journal of Fish Biology, 35(1):85–95, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pilcher:1989:SRI

[1642] M. W. Pilcher, P. J. Whitfield, and J. D. Riley. Seasonal and regional infestation characteristics of three ectoparasites of whiting, Merlangim merlangus L., in the North Sea. Journal of Fish Biology, 35(1):97–110, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joy:1989:PBB

[1643] J. E. Joy and E. Madan. Pathology of black bass hepatic tissue infected with larvae of the tapeworm Proteocephalm ambloplitis. Journal of Fish Biology, 35(1):111–118, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madsen:1989:TCE

[1644] S. S. Madsen and B. Korsgaard. Time-course effects of repetitive oestradiol-17β and thyroxine injections on the natural spring smolting of Atlantic salmon, Salmo salar L. Journal of Fish Biology, 35(1):119– 128, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wardle:1989:MTM

[1645] C. S. Wardle, J. J. Videler, T. Arimoto, J. M. Franco, and P. He. The muscle twitch and the maximum swimming speed of giant bluefin tuna, Thunnus thynnus L. Journal of Fish Biology, 35(1):129–137, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Epperly:1989:MMB

[1646] S. P. Epperly. A meristic, morphometric and biochemical investigation of Atlantic menhaden, Brevoortia tyrannus (Latrobe). Journal of Fish Biology, 35(1):139–152, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:1989:HPI

[1647] F. A. Huntingford and R. M. Coulter. Habituation of predator inspection in the three-spined stickleback, Gasterosteus aculeatm L. Journal of Fish Biology, 35(1):153–154, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kohli:1989:PSI

[1648] M. P. Singh Kohli. The pectoral spine as an indicator of age in Heteropneustes fossilis (Bloch). Journal of Fish Biology, 35(1):155–156, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smew:1989:FRT

[1649] I. A. Smew and I. C. Williams. The first record of the tapeworm Bothrimonus olrikii from a fish, the long-spined sea scorpion, Taurulus bubalis, in British waters. Journal of Fish Biology, 35(1):157–158, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mair:1989:TSM

[1650] G. C. Mair. A technique for sampling muscle tissue from live fish. Journal of Fish Biology, 35(1):159–160, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sugita:1989:IMJ

[1651] H. Sugita, C. Mlyajima, C. Sugiura, M. Iwata, S. Kato, and Y. Deguchi. The intestinal microflora of Japanese migratory fish. Journal of Fish Biology, 35(1):161–163, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRf

[1652] Anonymous. Book review. Journal of Fish Biology, 35(1):165–166, July 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rombout:1989:IISb

[1653] J. H. W. M. Rombout, H. E. Bot, and J. J. Taverne-Thiele. Immunological importance of the second gut segment of carp. II. Characterization of mucosal leucocytes. Journal of Fish Biology, 35(2):167–178, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rombout:1989:IISc

[1654] J. H. W. M. Rombout, A. A. van den Berg, C. T. G. A. van den Berg, P. Witte, and E. Egberts. Immunological importance of the second gut segment of carp. III. Systemic and/or mucosal immune responses after immunization with soluble or particulate antigen. Journal of Fish Biology, 35(2):179–186, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reader:1989:EET

[1655] J. P. Reader, N. C. Everall, M. D. J. Sayer, and R. Morris. The effects of eight trace metals in acid soft water on survival, mineral uptake and skeletal calcium deposition in yolk-sac fry of brown trout, Salmo trutta L. Journal of Fish Biology, 35(2):187–198, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:1989:CPT

[1656] J. F. Leatherland, N. E. Down, E. M. Donaldson, and H. M. Dye. Changes in plasma thyroid hormone levels in pink salmon, Oncorhynchus gorbuscha, during their spawning migration in the Fraser River (Canada). Journal of Fish Biology, 35(2):199–205, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Calvo:1989:SDI

[1657] J. Calvo. Sexual differences in the increase of white muscle fibres in Argentine hake, Merluccius hubbsi, from the San Matias Gulf (Argentina).

Journal of Fish Biology, 35(2):207–214, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wyatt:1989:HCE

[1658] R. J. Wyatt and C. R. Kennedy. Host-constrained epidemiology of the fish tapeworm Ligula intestinalis (L.). Journal of Fish Biology, 35(2): 215–227, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milligan:1989:RTR

[1659] C. L. Milligan, M. S. Graham, and A. P. Farrell. The response of trout red cells to adrenaline during seasonal acclimation and changes in temperature. Journal of Fish Biology, 35(2):229–236, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stefansson:1989:ITU

[1660] S. O. Stefansson, G. Naevdal, and T. Hansen. The influence of three unchanging photoperiods on growth and parr-smolt transformation in Atlantic salmon, Salmo salar L. Journal of Fish Biology, 35(2):237– 247, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Segner:1989:DEL

[1661] H. Segner, R. Rösch, H. Schmidt, and K. J. von Poeppinghausen. Digestive enzymes in larval Coregonus lavaretus L. Journal of Fish Biology, 35(2):249–263, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hontela:1989:DEL

[1662] A. Hontela, Y. Roy, R. van Collie, K. Lederis, and G. Chevalier. Differential effects of low pH and aluminium on the caudal neurosecretory system of the brook trout, Salvelinus fontinalis. Journal of Fish Biology, 35(2):265–273, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chellappa:1989:AVE

[1663] S. Chellappa, F. A. Huntingford, R. H. C. Strang, and R. Y. Thomson. Annual variation in energy reserves in male three-spined stickleback, Gasterosteus aculeatm L. (Pisces, Gasterosteidae). Journal of Fish Biology, 35(2):275–286, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saggers:1989:AMO

[1664] B. A. Saggers and M. L. Gould. The attachment of micro-organisms to macrophages isolated from the tilapia Oreochromis spilurus Günther. Journal of Fish Biology, 35(2):287–294, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Economidis:1989:HAB

[1665] P. S. Economidis and Alwyne Wheeler. Hybrids of Abramis brama with Scardinius erythrophthalmus and Rutilus rutilus from Lake Volvi, Macedonia, Greece. Journal of Fish Biology, 35(2):295–299, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greenstreet:1989:EUT

[1666] S. P. R. Greenstreet and R. I. G. Morgan. The effect of ultrasonic tags on the growth rates of Atlantic salmon, Salmo salar L., parr of varying size just prior to smolting. Journal of Fish Biology, 35(2):301–309, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbieri:1989:STF

[1667] G. Barbieri. Spawning type and fecundity of three sympatric species of tropical fishes in Brazil. Journal of Fish Biology, 35(2):311–312, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:1989:PCA

[1668] K. K. Lee, R. S. Raynard, and A. E. Ellis. The phospholipid composition of Atlantic salmon, Salmo solar L., erythrocyte membranes. Journal of Fish Biology, 35(2):313–314, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chellappa:1989:DER

[1669] S. Chellappa and F. A. Huntingford. Depletion of energy reserves during reproductive aggression in male three-spined stickleback, Gasterostem aculeatus L. Journal of Fish Biology, 35(2):315–316, August 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baggs:1989:OPY

[1670] E. M. Baggs and G. I. McT. Cowan. The onset of parasitism of youngof-the-year brook charr, Salvelinus fontinalis (Mitchill), in Big Northern Pond, Avalon Peninsula, Newfoundland, Canada. Journal of Fish Biology, 35(2):317–319, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRg

[1671] Anonymous. Book review. Journal of Fish Biology, 35(2):321–322, August 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sayer:1989:ECC

[1672] M. D. J. Sayer, J. P. Reader, and R. Morris. The effect of calcium concentration on the toxicity of copper, lead and zinc to yolk-sac fry of brown trout, Salmo trutta L., in soft, acid water. Journal of Fish Biology, 35(3):323–332, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whyte:1989:CRR

[1673] S. K. Whyte, L. H. Chappell, and C. J. Secombes. Cytotoxic reactions of rainbow trout, Salmo gairdneri Richardson, macrophages for larvae of the eye fluke Diplostomum spathaceum (Digenea). Journal of Fish Biology, 35(3):333–345, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:1989:OCE

[1674] W. A. Walsh, C. Swanson, C.-S. Lee, J. E. Banno, and H. Eda. Oxygen consumption by eggs and larvae of striped mullet, Mugil cephalus, in relation to development, salinity and temperature. Journal of Fish Biology, 35(3):347–358, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:1989:HSI

[1675] J. A. Williams and B. B. Nickol. Histological structure of the intestine and pyloric caeca of the green sunfish, Lepomis cyanellus Rafinesque. Journal of Fish Biology, 35(3):359–372, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:1989:FGB

[1676] J. A. Brown, P. Pepin, D. A. Methven, and D. C. Somerton. The feeding, growth and behaviour of juvenile cod, Gadus morhua L., in cold environments. Journal of Fish Biology, 35(3):373–380, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McLeese:1989:SCI

[1677] J. M. McLeese and T. W. Moon. Seasonal changes in the intestinal mucosa of winter flounder, Pseudopleuronectes americanus (Walbaum), from Passamaquoddy Bay, New Brunswick. Journal of Fish Biology, 35 (3):381–393, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poncin:1989:EDP

[1678] P. Poncin. Effects of different photoperiods on the reproduction of the barbel, Barbus barbus (L.), reared at constant temperature. Journal of Fish Biology, 35(3):395–400, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rljnsdorp:1989:SDE

[1679] A. D. Rljnsdorp and B. Ibelings. Sexual dimorphism in the energetics of reproduction and growth of North Sea plaice, Pleuronectes platessa L. Journal of Fish Biology, 35(3):401–415, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spanakis:1989:GDB

[1680] E. Spanakis, N. Tsimenides, and E. Zouros. Genetic differences between populations of sardine, Sardina pilchardus, and anchovy, Engraulis encrasicolus, in the Aegean and Ionian seas. Journal of Fish Biology, 35 (3):417–437, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bres:1989:EPR

[1681] M. Bres. The effects of prey relative abundance and chemical cues on prey selection in rainbow trout. Journal of Fish Biology, 35(3):439–445, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartwell:1989:TVA

[1682] S. I. Hartwell, J. H. Jin, D. S. Cherry, and J. Cairns, Jr. Toxicity versus avoidance response of golden shiner, Notemigonus crysoleucas, to five metals. Journal of Fish Biology, 35(3):447–456, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:1989:POE

[1683] E. A. Santos and E. P. Santos. Preliminary observations on effects of thyroxine on yolk-sac larvae of Odontesthes argentinensis Val., 1835 (Pisces, Atherinidae). Journal of Fish Biology, 35(3):457–458, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:1989:HMP

[1684] C. J. Robinson and T. J. Pitcher. Hunger motivation as a promoter of different behaviours within a shoal of herring: selection for homogeneity in fish shoal? Journal of Fish Biology, 35(3):459–460, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hindar:1989:FAS

[1685] K. Hindar and J. Nordland. A female Atlantic salmon, Salmo salar L., maturing sexually in the parr stage. Journal of Fish Biology, 35(3): 461–463, September 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:1989:FFR

[1686] M. R. Clark, K. J. King, and P. J. McMillan. The food and feeding relationships of black oreo, Allocyttus niger, smooth oreo, Pseudocyttus maculatus, and eight other fish species from the continental slope of the south-west Chatham Rise, New Zealand. Journal of Fish Biology, 35 (4):465–484, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radtke:1989:MJB

[1687] R. L. Radtke and B. Morales-Nin. Mediterranean juvenile bluefin tuna: life history patterns. Journal of Fish Biology, 35(4):485–496, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deJonge:1989:TTG

[1688] J. de Jonge, A. J. H. de Ruiter, and R. van den Hurk. Testis-testicular gland complex of two Tripterygion species (Blennioidei, Teleostei): differences between territorial and non-territorial males. Journal of Fish Biology, 35(4):497–508, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandes:1989:RRO

[1689] M. N. Fernandes and F. T. Rantin. Respiratory responses of Oreochromis niloticus (Pisces, Cichlidae) to environmental hypoxia under different thermal conditions. Journal of Fish Biology, 35(4):509–519, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thuvander:1989:CER

[1690] A. Thuvander. Cadmium exposure of rainbow trout, Salmo gairdneri Richardson: effects on immune functions. Journal of Fish Biology, 35 (4):521–529, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ladich:1989:SPR

[1691] F. Ladich. Sound production by the river bullhead, Cottus gobio L. (Cottidae, Teleostei). Journal of Fish Biology, 35(4):531–538, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrison:1989:GJA

[1692] B. R. S. Morrison. The growth of juvenile Atlantic salmon, Salmo salar L., and brown trout, Salmo trutta L., in a Scottish river system subject to cooling-water discharge. Journal of Fish Biology, 35(4):539–556, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spry:1989:IDW

[1693] D. J. Spry and C. M. Wood. The influence of dietary and waterborne zinc on heat-stable metal ligands in rainbow trout, Salmo gairdneri Richardson: quantification by^{109}Cd radioassay and evaluation of the assay. Journal of Fish Biology, 35(4):557–576, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lucas:1989:EID

[1694] M. C. Lucas. Effects of implanted dummy transmitters on mortality, growth and tissue reaction in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 35(4):577–587, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:1989:FSR

[1695] M. J. Phillips. The feeding sounds of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 35(4):589–592, October 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rasotto:1989:AHG

[1696] M. B. Rasotto and E. Zulian. Abnormal hermaphroditism in Gambusia affinis holbrooki from a hot spring of north-eastern Italy. Journal of Fish Biology, 35(4):593–595, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goolish:1989:COD

[1697] E. M. Goolish. A comparison of oxygen debt in small and large rainbow trout, Salmo gairdneri Richardson. Journal of Fish Biology, 35(4):597– 598, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchmann:1989:RBH

[1698] K. Buchmann. Relationship between host size of Anguilla anguilla and the infection level of the monogeneans Pseudodactylogyrus spp. Journal of Fish Biology, 35(4):599–601, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:BRh

[1699] Anonymous. Book reviews. Journal of Fish Biology, 35(4):603–606, October 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brule:1989:AEP

[1700] T. Brulé. Analysis of the enzyme polymorphism in a plaice, Pleuronectes platessa L., population from the north-west coast of Brittany, France. Journal of Fish Biology, 35(5):607–620, November 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorman:1989:SAB

[1701] J. A. Dorman. Some aspects of the biology of the garfish Belone belone (L.) from southern Ireland. Journal of Fish Biology, 35(5):621–629, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seiwald:1989:HFS

[1702] M. Seiwald and R. A. Patzner. Histological, fine-structural and histochemical differences in the testicular glands of gobiid and blenniid fishes. Journal of Fish Biology, 35(5):631–640, November 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spanier:1989:FAC

[1703] E. Spanier, S. Pisanty, M. Tom, and G. Almog-Shtayer. The fish assemblage on a coralligenous shallow shelf off the Mediterranean coast of northern Israel. Journal of Fish Biology, 35(5):641–649, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamilton:1989:PGC

[1704] K. E. Hamilton, A. Ferguson, J. B. Taggart, T. Tómasson, A. Walker, and E. Fahy. Post-glacial colonization of brown trout, Salmo trutta L.: Ldh-5 as a phylogeographic marker locus. Journal of Fish Biology, 35 (5):651–664, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crozier:1989:ADB

[1705] W. W. Crozier and I. J. J. Moffett. Amount and distribution of biochemical-genetic variation among wild populations and a hatchery stock of Atlantic salmon, Salmo salar L., from north-east Ireland. Journal of Fish Biology, 35(5):665–677, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beddington:1989:LTC

[1706] J. R. Beddington, C. A. Mills, F. Beards, M. J. Minski, and J. N. B. Bell. Long-term changes in strontium-90 concentrations within a freshwater predator–prey system. Journal of Fish Biology, 35(5):679–686, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hynes:1989:MDG

[1707] R. A. Hynes, E. J. Duke, and P. Joyce. Mitochondrial DNA as a genetic marker for brown trout, Salmo trutta L., populations. Journal of Fish Biology, 35(5):687–701, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moser:1989:RMJ

[1708] M. L. Moser and W. F. Hettler. Routine metabolism of juvenile spot, Leiostomus xanthums (Lacépède), as a function of temperature, salinity and weight. Journal of Fish Biology, 35(5):703–707, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hlnterleitner:1989:PEA

[1709] S. Hlnterleitner, J. Thurner-Flür, W. Wieser, and N. El-Flky. Profiles of enzyme activity in larvae of two cyprinid species with contrasting life styles (Cyprinidae; Teleostei). Journal of Fish Biology, 35(5):709–718, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:1989:PLM

[1710] M. A. Munro, P. J. Whitfield, and R. Diffley. Pomphorhynchus laevis (Müller) in the flounder, Platichthys flesus L., in the tidal River Thames: population structure, microhabitat utilization and reproductive status in the field and under conditions of controlled salinity. Journal of Fish Biology, 35(5):719–735, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Handy:1989:VEI

[1711] R. D. Handy, F. B. Eddy, and G. Romain. In vitro evidence for the ionoregulatory role of rainbow trout mucus in acid, acid/aluminium and zinc toxicity. Journal of Fish Biology, 35(5):737–747, November 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1989:Ab

[1712] Anonymous. Announcement. Journal of Fish Biology, 35(6):i, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1989:DSI

[1713] N. F. Smith, C. Talbot, and F. B. Eddy. Dietary salt intake and its relevance to ionic regulation in freshwater salmonids. Journal of Fish Biology, 35(6):749–753, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnhagen:1989:BBI

[1714] C. Magnhagen and L. Kvarnemo. Big is better: the importance of size for reproductive success in male Pomatoschistus minutus (Pallas) (Pisces, Gobiidae). Journal of Fish Biology, 35(6):755–763, December 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shuter:1989:ETP

[1715] B. J. Shuter, P. E. Ihssen, D. L. Wales, and E. J. Snucins. The effects of temperature, pH and water hardness on winter starvation of youngof-the-year smallmouth bass, Micropterus dolomieui Lacépède. Journal of Fish Biology, 35(6):765–780, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharp:1989:IRW

[1716] G. J. E. Sharp, A. W. Pike, and C. J. Secombes. The immune response of wild rainbow trout, Salmo gairdneri Richardson, to naturally acquired plerocercoid infections of Diphyllobothrium dendriticum (Nitzsch, 1824) and D. ditremum (Creplin, 1825). Journal of Fish Biology, 35(6):781– 794, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cui:1989:GFA

[1717] Y. Cui and R. J. Wootton. Grouping fails to affect the growth and energy budget of a cyprinid, Phoxinus phoxinus (L.). Journal of Fish Biology, 35 (6):795–798, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:1989:ILM

[1718] G. F. Turner, T. J. Pitcher, and A. S. Grimm. Identification of the Lake Malaŵi Oreochromis (Nyasalapia) spp. using multivariate morphometric techniques. Journal of Fish Biology, 35(6):799–812, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mugiya:1989:ORI

[1719] Y. Mugiya and T. Uchimura. Otolith resorption induced by anaerobic stress in the goldfish, Carassius auratus. Journal of Fish Biology, 35(6): 813–818, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arefjev:1989:KVS

[1720] V. A. Arefjev. Karyotype variability in successive generations after hybridization between the great sturgeon, Huso huso (L.), and the sterlet, Acipenser ruthenus L. Journal of Fish Biology, 35(6):819–828, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:1989:OCL

[1721] J. Davenport and V. Thorsteinsson. Observations on the colours of lumpsuckers, Cyclopterus lumpus L. Journal of Fish Biology, 35(6):829–838, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:1989:CMA

[1722] P. Mořan, A. M. Pendás, E. García-Vázquez, and A. R. Linde. Chromosomal and morphological analysis of two populations of Salmo trutta sbp. fario employed in repopulation. Journal of Fish Biology, 35(6):839–843, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:1989:IVE

[1723] D. Burton and M. P. M. Burton. Individual variability in the epidermis of prespawning inshore winter flounder, Pseudopleuronectes americanus. Journal of Fish Biology, 35(6):845–853, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McWilliams:1989:KCS

[1724] P. G. McWilliams and K. L. Shephard. Kinetic characteristics of the sodium uptake mechanism during the development of embryos and fry of Atlantic salmon, Salmo salar L., in an improved water quality. Journal of Fish Biology, 35(6):855–868, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polunin:1989:LED

[1725] N. V. C. Polunin and E. B. Brothers. Low efficiency of dietary carbon and nitrogen conversion to growth in an herbivorous coral-reef fish in the wild. Journal of Fish Biology, 35(6):869–879, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everall:1989:EWH

[1726] N. C. Everall, N. A. A. Macfarlane, and R. W. Sedgwick. The effects of water hardness upon the uptake, accumulation and excretion of zinc in the brown trout, Salmo trutta L. Journal of Fish Biology, 35(6):881–892, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gulland:1989:FPT

[1727] J. A. Gulland. Fish populations and their management. Journal of Fish Biology, 35(sA):1–9, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pauly:1989:FCT

[1728] D. Pauly. Food consumption by tropical and temperate fish populations: some generalizations. Journal of Fish Biology, 35(sA):11–20, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:1989:YFD

[1729] M. T. Scott and L. A. Nielsen. Young fish distribution in backwaters and main-channel borders of the Kanawha River, West Virginia. Journal of Fish Biology, 35(sA):21–27, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Houde:1989:SEE

[1730] E. D. Houde. Subtleties and episodes in the early life of fishes. Journal of Fish Biology, 35(sA):29–38, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laprise:1989:OCL

[1731] R. Laprise and J. J. Dodson. Ontogenetic changes in the longitudinal distribution of two species of larval fish in a turbid well-mixed estuary. Journal of Fish Biology, 35(sA):39–47, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heath:1989:CSE

[1732] M. Heath and K. Richardson. Comparative study of early-life survival variability of herring Clupea harengus, in the north-eastern Atlantic. Journal of Fish Biology, 35(sA):49–57, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsukamoto:1989:SDM

[1733] K. Tsukamoto, H. Kuwada, J. Hirokawa, M. Oya, S. Sekiya, H. Fujimoto, and K. Imaizuml. Size-dependent mortality of red sea bream, Pagrus major, juveniles released with fluorescent otolith-tags in News Bay, Japan. Journal of Fish Biology, 35(sA):59–69, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lough:1989:MCH

[1734] R. G. Lough and G. R. Bolz. The movement of cod and haddock larvae onto the shoals of Georges Bank. Journal of Fish Biology, 35(sA):71–79, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Al-Hossaini:1989:OMI

[1735] M. Al-Hossaini, Q. Liu, and T. J. Pitcher. Otolith microstructure indicating growth and mortality among plaice, Pleuronectes platessa L., post-larval sub-cohorts. Journal of Fish Biology, 35(sA):81–90, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:1989:CPC

[1736] J. M. Elliott. The critical-period concept for juvenile survival and its relevance for population regulation in young sea trout, Salmo trutta. Journal of Fish Biology, 35(sA):91–98, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Helle:1989:RBS

[1737] J. H. Helle. Relation between size-at-maturity and survival of progeny in chum salmon, Oncorhynchus keta (Walbaum). Journal of Fish Biology, 35(sA):99–107, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schneider:1989:ISS

[1738] D. C. Schneider. Identifying the spatial scale of density-dependent interaction of predators with schooling fish in the southern Labrador Current. Journal of Fish Biology, 35(sA):109–115, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromley:1989:EDD

[1739] P. J. Bromley. Evidence for density-dependent growth in North Sea gadoids. Journal of Fish Biology, 35(sA):117–123, December 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townsend:1989:PCF

[1740] C. R. Townsend. Population cycles in freshwater fish. Journal of Fish Biology, 35(sA):125–131, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Palomera:1989:FME

[1741] I. Palomera and J. Lleonart. Field mortality estimates of anchovy larvae, Engraulis encrasicolus, in the western Mediterranean. Journal of Fish Biology, 35(sA):133–138, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borisov:1989:LTV

[1742] V. M. Borisov and A. A. Elizarov. Long-term variations of biotic and abiotic conditions in the ecosystem of the Barents Sea. Journal of Fish Biology, 35(sA):139–144, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siddeek:1989:ENM

[1743] M. S. M. Siddeek. The estimation of natural mortality in Irish Sea plaice, Pleuronectes platessa L., using tagging methods. Journal of Fish Biology, 35(sA):145–154, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rose:1989:PVC

[1744] G. A. Rose and W. C. Leggett. Predicting variability in catch-per-effort in Atlantic cod, Gadus morhua, trap and gillnet fisheries. Journal of Fish Biology, 35(sA):155–161, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coutin:1989:ELT

[1745] P. C. Coutin and A. I. Payne. The effects of long-term exploitation of demersal fish populations off the coast of Sierra Leone, West Africa. Journal of Fish Biology, 35(sA):163–167, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:1989:AFL

[1746] Q. Liu, T. Pitcher, and M. Al-Hossaini. Ageing with fisheries lengthfrequency data, using information about growth. Journal of Fish Biology, 35(sA):169–177, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Livingston:1989:KFS

[1747] P. A. Livingston. Key fish species, northern fur seals, Callorhinus ursinus, and fisheries interactions involving walleye pollock, Theragra chalcogramma, in the eastern Bering Sea. Journal of Fish Biology, 35(sA):179–186, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

desClers:1989:MRD

[1748] S. des Clers. Modelling regional differences in ‘sealworm’, Pseudoterranova decipiens (Nematoda, Ascaridoidea), infections in some North Atlantic cod, Gadus morhua, stocks. Journal of Fish Biology, 35(sA): 187–192, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jamieson:1989:DHP

[1749] A. Jamieson and A. J. Birley. The demography of a haemoglobin polymorphism in the Atlantic cod, Gadus morhua L. Journal of Fish Biology, 35(sA):193–204, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:1989:GVM

[1750] E. Verspoor and W. C. Jordan. Genetic variation at the me-2 locus in the Atlantic salmon within and between rivers: evidence for its selective maintenance. Journal of Fish Biology, 35(sA):205–213, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chadwick:1989:RTP

[1751] E. M. P. Chadwick and R. R. Claytor. Run timing of pelagic fishes in Gulf of St Lawrence: area and species effects. Journal of Fish Biology, 35(sA):215–223, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skulason:1989:SSE

[1752] S. Skúlason, S. S. Snorrason, D. L. G. Noakes, M. M. Ferguson, and H. J. Malmquist. Segregation in spawning and early life history among polymorphic Arctic charr, Salvelinus alpinus, in Thingvallavatn, Iceland. Journal of Fish Biology, 35(sA):225–232, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernatchez:1989:PBI

[1753] L. Bernatchez, J. J. Dodson, and S. Boivin. Population bottlenecks: influence on mitochondrial DNA diversity and its effect in coregonine stock discrimination. Journal of Fish Biology, 35(sA):233–244, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jorstad:1989:GVP

[1754] K. E. Jørstad and G. Nævdal. Genetic variation and population structure of cod, Gadus morhua L., in some fjords in northern Norway. Journal of Fish Biology, 35(sA):245–252, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hovey:1989:MDA

[1755] S. J. Hovey, D. P. F. King, D. Thompson, and A. Scott. Mitochondrial DNA and allozyme analysis of Atlantic salmon, Salmo salar L., in England and Wales. Journal of Fish Biology, 35(sA):253–260, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deLeaniz:1989:GDC

[1756] C. Garcia de Leániz, E. Verspoor, and A. D. Hawkins. Genetic determination of the contribution of stocked and wild Atlantic salmon, Salmo salar L., to the angling fisheries in two Spanish rivers. Journal of Fish Biology, 35(sA):261–270, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:1989:PRU

[1757] P. J. B. Hart. Predicting resource utilization: the utility of optimal foraging models. Journal of Fish Biology, 35(sA):271–277, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:1989:IBR

[1758] I. G. Cowx. Interaction between the roach, Rutilus rutilus, and dace, Leuciscus leuciscus, populations in a river catchment in south-west England. Journal of Fish Biology, 35(sA):279–284, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beall:1989:IRB

[1759] E. Beall, M. Heland, and C. Marty. Interspecific relationships between emerging Atlantic salmon, Salmo salar, and Coho salmon, Oncorhynchus kisutch, juveniles. Journal of Fish Biology, 35(sA):285–293, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:1989:DVS

[1760] J. E. Thorpe. Developmental variation in salmonid populations. Journal of Fish Biology, 35(sA):295–303, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Agnalt:1989:LTC

[1761] A.-L. Agnalt. Long-term changes in growth and age at maturity of mackerel, Scomber scombrus L., from the North Sea. Journal of Fish Biology, 35(sA):305–311, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Danzmann:1989:GVC

[1762] R. G. Danzmann, M. M. Ferguson, and F. W. Allendorf. Genetic variability and components of fitness in hatchery strains of rainbow trout. Journal of Fish Biology, 35(sA):313–319, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:1989:NSA

[1763] T. J. Martin. Niche separation in Ambassis (Pisces: Perciformes) species co-existing in estuaries of south-east Africa. Journal of Fish Biology, 35(sA):321–328, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Judd:1989:RCS

[1764] A. G. Judd and M. Hovland. The role of chemosynthesis in supporting fish stocks in the North Sea. Journal of Fish Biology, 35(sA):329–330, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clemmesen:1989:RDR

[1765] C. M. Clemmesen. RNA/DNA ratios of laboratory-reared and wild herring larvae determined with a highly sensitive fluorescence method. Journal of Fish Biology, 35(sA):331–333, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowell:1989:DFG

[1766] C. Rowell, K. Stokes, and R. Law. Does fishing generate selection differentials? Journal of Fish Biology, 35(sA):335–337, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monaghan:1989:TSS

[1767] P. Monaghan, J. D. Uttley, and J. D. Okill. Terns and sandeels: seabirds as indicators of changes in marine fish populations. Journal of Fish Biology, 35(sA):339–340, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pierce:1989:DTC

[1768] G. J. Pierce, J. S. W. Diack, and P. R. Boyle. Digestive tract contents of seals in the Moray Firth area of Scotland. Journal of Fish Biology, 35(sA):341–343, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:1989:ESS

[1769] P. J. Smith, A. J. Birley, and A. Jamieson. Electrophoretic studies and the stock concept in marine fishes. Journal of Fish Biology, 35(sA): 345–346, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almatar:1989:ECC

[1770] S. M. Almatar. Energy content of Clyde spring-spawning herring, Clupea harengus L. Journal of Fish Biology, 35(sA):347–349, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Titus:1989:SAA

[1771] R. G. Titus and H. Mosegaard. Smolting at age 1 and its adaptive significance for migratory trout, Salmo trutta L., in a small Baltic-coast stream. Journal of Fish Biology, 35(sA):351–353, December 1989. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beverton:1989:CAS

[1772] R. J. H. Beverton. Closing address: The symposium in perspective. Journal of Fish Biology, 35(sA):355–363, December 1989. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).